A Bibliography of Publications of Nelson H. F. Beebe

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

             E-mail: beebe@math.utah.edu, beebe@acm.org,
                beebe@computer.org, beebe@ieee.org (Internet)
           WWW URL: http://www.math.utah.edu/~beebe/

22 October 2022 Version 4.368

Abstract

This bibliography records publications of Nelson H. F. Beebe.

Title word cross-reference

#1 [68]. #2 [99, 185]. #3 [100, 187]. #4 [115, 193]. #5 [117, 196]. #6 [202]. #7 [211]. #8 [224].

10 [49]. ^{+} [101, 109]. _{2} [57, 66, 59]. _{2}^{+} [66]. _{2}^{2+} [66]. A^{1}Π [101]. α [52]. Ax = λBx [61, 123]. F (m, t) [173]. expm1(x) = exp(x) − 1 [736]. log1p(x) = log(1 + x) [737]. He_{2} [105]. π [949]. X^{1}Σ^{+} [101].

1-UT [1789]. 118 [264]. 18 [187]. 18-Aug-80 [185, 193, 196, 202, 187]. 1964 [508]. 1969 [503, 1065, 509]. 1974 [511]. 1978 [1786]. 1979 [504, 1067, 1785, 512]. 1984 [513]. 1989 [505, 1068, 514]. 1994 [515]. 1999 [506, 1069, 510]. 2 [1789, 409, 408]. 2.0 [263]. 2.00 [216]. 2.3a [231]. 2.3c [235]. 20 [134, 233, 238]. 20/60 [250, 249]. ’2000 [658]. ’2001 [720]. ’2002 [760]. ’2003 [761, 778]. ’2004 [843, 625]. ’2005 [844]. 2007 [868, 1788]. 2009 [931, 1070]. ’2011 [999]. ’2012 [1064]. 2019 [1738, 930, 1066]. 2029 [1739, 1564]. 21st [1246]. 2311 [11, 26, 54]. 3 [138, 139, 219, 245]. 3.0 [263]. 3100 [255]. 3330 [39, 43, 54]. 360 [27]. 370 [39]. 4.4 [234]. 60 [250, 249]. 64 [652].

7 [711]. 754 [305, 310, 611, 690]. 77 [188, 227, 242]. 77/SFTRAN3 [188].

80 [187]. 80th [1790]. ’88 [528]. ’89 [527].

9} [416]. ’90 [530]. ’91 [529]. ’92 [526]. ’93 [524]. ’94 [525]. ’95 [497, 694, 706]. ’96 [518]. ’97 [517]. 978 [1789]. 978-3-319-64109-6 [1789]. 978-3-319-64110-2 [1789]. ’98 [537, 538]. ’99 [612].

= [604, 963, 964, 966, 970, 971, 969, 968, 967, 965, 1305, 1439, 1655, 1656, 1657, 1658, 1659, 1681, 1682, 1683, 1684, 1685, 879]. == [1355].

Aarhus [78, 60]. abstract [846]. Academy [1300]. Accepting [284]. Access [143, 151, 807, 811, 815, 830, 831, 832]. Accessible [876]. Accounts [802]. Accurate [270, 269]. ACM [277, 345, 336, 337, 338, 364, 363, 444, 446, 445, 391, 423, 584, 585, 586, 587, 593, 572, 574, 575, 576, 592, 628, 615, 643, 622, 669, 662, 663, 661, 685, 729, 712, 717, 746, 719, 721, 754, 755, 756, 772, 773, 775, 766, 767, 784, 783, 840, 841, 842, 833, 853, 855, 869, 876, 878, 877, 881, 882, 883, 884, 885, 890, 891, 926, 903, 935, 989, 1039, 1167, 1251, 1325, 1384, 1386, 1389, 1390, 1391, 1392, 1393, 1425, 1452, 1453, 1454, 1455, 1456, 1543, 1581, 1583, 1584, 1582, 1566, 1631, 1625, 1637, 1628, 1629, 377, 418, 419, 420, 421]. ACM} [422, 605, 770, 888, 1548, 315, 355, 563, 836, 341, 350]. ACM/IMS [1584]. Acta [606, 974]. Ada [338, 495, 1107, 1442]. Adaptive [881]. Additional [56]. Address [778, 845]. Administrators} [428, 429]. Adrien [952]. Advanced [311, 356]. Advances [648, 1186, 1289, 1406, 1436, 1491]. Affects [839]. Aids [542, 543]. Alan [837]. Albert [777]. Alchemy [1446]. Algebra [210, 329, 330, 372, 394, 395, 461, 459, 462, 464, 460, 620, 834, 996, 939, 1233, 1255, 1346, 1347, 1420, 1475, 703, 817, 819, 820, 821, 822, 824, 826, 1248, 1362]. Algebraic [437, 523, 659, 1786]. Algorithm [206, 864, 893]. Algorithmica [849]. Algorithmics [877]. Algorithms [126, 280, 315, 352, 355, 333, 480, 872, 1393, 1786, 481, 840, 799, 1087]. allgemeine [1327, 1326]. Alston [1472]. Ambix [1446]. American [486, 503, 504, 505, 506, 487, 550, 551, 552, 588, 589, 590, 591, 994, 993, 1021, 1022, 1023, 1024, 1027, 1026, 1025, 1020, 1012, 1015, 1013, 1014, 1016, 1017, 1018, 1019, 1029, 1028, 1318, 1319, 1320, 1300, 1522, 1523, 1524, 1525, 1526, 1574, 1621, 1687, 1688, 1689, 1690, 1691, 1692, 1693, 1694, 1695, 1696, 1697, 1698, 1699, 1700, 1701, 1702, 1703, 1704, 1705, 1706, 1707, 424, 637, 638, 639, 636, 1126, 1127, 1113, 1114, 1115, 1116, 1117, 1118, 1119, 1120, 1121, 1122, 1123, 1124, 1125]. Amsterdam [597, 627, 940]. Analysis [335, 458, 602, 607, 902, 906, 1173, 1390, 111, 1790, 609, 646, 678, 677, 941, 931, 930, 1290, 1291, 1292, 1293, 1536, 457, 573, 1211, 1174, 1175, 1289, 1480, 1252]. Animation [792, 490]. Anion [53]. Annales [1210]. Annals [683, 682, 745, 1063, 1198, 1203, 1204, 1205, 1206, 1207, 1208, 1209, 1178, 1179, 1180, 1181, 1182, 1283]. Annual [649, 650, 913]. ANSI [482]. Antony [1748]. ANZIAM [1495]. Appendix [1537, 1544]. Application [109]. Applications [380, 1406, 1376, 1377, 1378, 1379, 1380, 1491, 461, 459, 462, 464, 620, 996, 939, 1255, 1346, 1347, 380, 378, 473, 480, 458, 460, 565, 723, 733, 784, 872, 1037, 1173, 1309, 1282, 1444, 1619, 1734, 1735, 1736, 1737, 1727]. Applied [316, 352, 467, 468, 507, 566, 548, 569, 608, 647, 671, 672, 673, 674, 675, 670, 758, 763, 841, 873, 909, 905, 962, 961, 1071, 1072, 1073, 1075, 1074, 1186, 1187, 1208, 1209, 1253, 1170, 1323, 1420, 1369, 1370, 1371, 1372, 1373, 1374, 1375, 1382, 1489, 1528, 1530, 1495, 1553, 1566, 253]. Approach [119, 109]. Approximate [603]. Approximation [95, 181]. Aquaculture [1717, 1721, 1722, 1719]. Aquatic [1655, 1656, 1657, 1658, 1659]. aquatiques [1655, 1656, 1657, 1658, 1659]. Arbitrary [49, 998]. Archimedes [1450]. Architecture [328, 652, 772, 767, 853, 1510]. Archiv [1404]. Archive [933, 1464, 519]. Areas [1784]. Arithmetic [306, 305, 370, 1529, 1570, 1622, 346, 818, 866, 867, 1157, 1421]. Arnold [1410, 1405]. Array [110]. Arts [1300]. ASA [1171]. Asian [746, 1384]. Aspects [1783, 1136]. Assembler [6, 7]. Assignments [130]. Associates [366]. Association [1321, 493, 494, 743, 1085, 398, 487, 994, 993, 1012, 1015, 1013, 1014, 1016, 1017, 1018, 1019]. AStA [1289]. Astronomy [1249]. AT&T [416, 928, 922]. Atom [109]. Atomic [85, 171, 1108]. Atoms [65, 183, 107]. ATS [51]. Attributes [122]. Audio [1325]. Aug [185, 187, 193, 196, 202]. August [1786]. Augusta [1442]. Australian [1359, 1495]. authidx [520]. Author [520]. Author/Editor [520]. Automated [1356]. Automatic [13]. automating [857]. Automation [685]. automatisieren [857]. Autonomous [881]. aux [848]. Available [22, 814, 827]. Awards [836]. Axel [1324]. AXIOM [372]. Aylmer [1595].

B [390, 1210, 1495, 1262, 199]. B-Spline [199]. Babbage [1441]. Back [732, 778]. Background [264]. Backup [235, 641, 688, 221, 225]. Bandsystem [101]. Base [49, 148]. Based [360]. Bases [443, 619]. Basic [839]. Basis [42, 64, 174, 180]. Batch [15, 803]. Bauer [1617]. Bay [780]. Bayesian [1211]. Be [113, 109]. Beebe [1789, 130, 332]. Bell [928, 918, 920, 919, 921, 923, 924, 922, 929, 927, 1626]. Ben [879]. Benchmarking [934].

Benchmarks [268, 296, 351]. Benford [995, 998]. Benoît [1061]. Beresford [835]. Berichte [1148]. Bethe [1059]. BIAS [962]. Bibliographer [779]. Bibliography [53, 61, 123, 133, 174, 217, 223, 281, 292, 302, 290, 282, 301, 288, 284, 287, 295, 294, 285, 280, 283, 293, 277, 289, 291, 286, 303, 316, 315, 311, 309, 304, 307, 318, 324, 326, 327, 317, 325, 352, 345, 353, 354, 355, 341, 356, 329, 334, 335, 336, 337, 338, 348, 342, 349, 333, 340, 344, 343, 330, 331, 332, 339, 357, 347, 358, 350, 366, 368, 369, 370, 372, 361, 362, 371, 360, 364, 365, 363, 367, 444, 446, 445, 442, 443, 382, 387, 388, 389, 391, 394, 399, 402, 403, 404, 406, 415, 416, 423, 432, 435, 376, 383]. Bibliography [390, 397, 418, 419, 420, 421, 422, 434, 373, 374, 414, 441, 417, 375, 378, 380, 381, 413, 437, 438, 440, 439, 392, 393, 433, 427, 405, 410, 407, 411, 412, 409, 408, 447, 448, 384, 386, 379, 424, 425, 426, 385, 398, 431, 396, 400, 401, 430, 428, 429, 395, 495, 461, 453, 467, 468, 470, 482, 486, 492, 496, 503, 504, 505, 506, 449, 450, 452, 451, 465, 466, 472, 473, 474, 479, 478, 475, 487, 485, 489, 490, 459, 462, 464, 491, 454, 488, 463, 481, 477, 476, 480, 455, 456, 457, 458, 469, 484, 483, 507, 498]. Bibliography [499, 500, 502, 501, 493, 494, 497, 471, 460, 533, 516, 532, 549, 550, 551, 552, 560, 562, 555, 556, 557, 553, 508, 509, 511, 512, 513, 514, 515, 510, 535, 566, 568, 521, 531, 541, 540, 534, 522, 523, 567, 565, 563, 564, 528, 527, 530, 529, 526, 524, 525, 518, 517, 537, 538, 561, 539, 554, 558, 548, 559, 542, 544, 545, 546, 547, 543, 519, 569, 602, 584, 585, 586, 587, 588, 589, 590, 591, 604, 578, 598, 599, 597, 601, 600, 579, 580, 571, 593, 581, 595, 573, 572, 574, 575, 576, 570, 577, 592, 594, 596, 628]. Bibliography [617, 618, 607, 614, 615, 643, 605, 623, 634, 635, 633, 613, 626, 629, 630, 632, 625, 640, 609, 637, 638, 639, 636, 642, 645, 644, 631, 612, 622, 624, 606, 608, 616, 619, 620, 621, 627, 646, 647, 648, 649, 650, 679, 683, 681, 652, 665, 668, 667, 653, 658, 669, 671, 672, 673, 674, 675, 670, 660, 651, 655, 656, 659, 662, 663, 661, 666, 664, 678, 677, 680, 682, 684, 685, 1738, 1739, 654, 728, 734, 723, 722, 725, 724, 729, 730, 727, 744, 718, 715, 713, 716, 714, 742, 740, 741, 739, 749, 750, 743, 720, 712, 717]. Bibliography [746, 726, 748, 719, 721, 731, 733, 735, 745, 747, 751, 760, 761, 754, 755, 753, 756, 757, 758, 759, 762, 769, 771, 772, 773, 775, 774, 763, 764, 765, 766, 767, 770, 777, 785, 790, 834, 836, 837, 838, 782, 787, 843, 844, 784, 783, 840, 841, 842, 792, 786, 788, 791, 797, 796, 794, 795, 799, 798, 793, 800, 833, 835, 850, 851, 853, 855, 852, 854, 849, 862, 860, 861, 865, 870, 868, 869, 876, 878, 871, 872, 873, 877, 886, 881, 882, 883, 884, 885, 887, 894, 917, 902, 890, 891, 893, 910, 911, 888, 889, 892, 900]. Bibliography [895, 931, 909, 926, 928, 918, 920, 919, 921, 923, 924, 922, 929, 927, 896, 897, 898, 899, 901, 903, 904, 905, 906, 907, 908, 914, 912, 913, 915, 916, 925, 930, 955, 959, 986, 995, 954, 953, 1001, 1002, 1003, 1004, 1005, 948, 947, 943, 944, 945, 946, 949, 1000, 941, 936, 937, 999, 933, 935, 989, 962, 975, 957, 956, 958, 960, 932, 994, 993, 938, 990, 992, 991, 996, 1006, 939, 940, 951, 961, 963, 964, 966, 970, 971, 969, 968, 967, 965, 974, 977, 979, 980, 982, 981, 978, 984, 983, 985, 987, 988, 942, 950, 952]. Bibliography [972, 973, 976, 1039, 1041, 1094, 1095, 1096, 1097, 1098, 1099, 1100, 1101, 1102, 1103, 1104, 1105, 1112, 1065, 1089, 1090, 1091, 1092, 1093, 1085, 1064, 1052, 1021, 1022, 1023, 1024, 1027, 1026, 1025, 1020, 1012, 1015, 1013, 1014, 1016, 1017, 1018, 1019, 1007, 1078, 1087, 1088, 1071, 1072, 1073, 1075, 1074, 1081, 1083, 1086, 1084, 1082, 1036, 1008, 1009, 1010, 1011, 1029, 1028, 1030, 1031, 1033, 1034, 1035, 1032, 1037, 1038, 1040, 1063, 1067, 1068, 1069, 1070, 1066, 1076, 1077, 1079, 1080, 1042, 1043, 1047, 1050, 1051, 1046, 1045, 1048, 1049, 1044, 1062, 1053, 1054, 1055, 1056, 1057, 1058]. Bibliography [1059, 1060, 1061, 1108, 1107, 1109, 1136, 1154, 1106, 1111, 1126, 1127, 1151, 1113, 1144, 1145, 1146, 1147, 1148, 1152, 1153, 1155, 1156, 1114, 1115, 1116, 1117, 1118, 1119, 1120, 1121, 1122, 1123, 1124, 1125, 1129, 1130, 1131, 1132, 1134, 1135, 1137, 1138, 1139, 1140, 1141, 1142, 1143, 1128, 1133, 1150, 1233, 1196, 1168, 1235, 1167, 1251, 1243, 1244, 1245, 1248, 1236, 1237, 1238, 1239, 1240, 1241, 1242, 1249, 1188, 1189, 1190, 1191, 1192, 1193, 1194, 1195, 1158, 1186, 1187, 1197, 1198, 1199, 1200, 1201, 1202, 1203, 1204, 1205, 1206, 1207, 1208, 1209, 1210, 1211, 1212, 1213, 1214, 1215, 1216]. Bibliography [1217, 1218, 1219, 1220, 1159, 1221, 1222, 1223, 1224, 1225, 1226, 1227, 1228, 1229, 1230, 1231, 1234, 1250, 1253, 1294, 1160, 1161, 1162, 1163, 1164, 1165, 1166, 1169, 1170, 1171, 1172, 1173, 1174, 1175, 1176, 1177, 1178, 1179, 1180, 1181, 1182, 1183, 1184, 1185, 1296, 1323, 1308, 1309, 1254, 1303, 1304, 1325, 1256, 1327, 1257, 1258, 1259, 1260, 1261, 1262, 1263, 1264, 1265, 1266, 1267, 1268, 1269, 1270, 1271, 1272, 1273, 1274, 1275, 1276, 1277, 1278, 1279, 1280, 1281, 1302, 1255, 1315, 1316, 1317, 1318, 1319, 1320, 1321, 1322, 1333, 1282, 1283, 1284, 1285, 1286, 1287, 1288, 1289, 1290]. Bibliography [1291, 1292, 1293, 1295, 1297, 1298, 1300, 1301, 1305, 1306, 1307, 1310, 1312, 1313, 1314, 1326, 1299, 1311, 1331, 1332, 1351, 1334, 1360, 1361, 1362, 1364, 1365, 1366, 1367, 1335, 1336, 1337, 1338, 1339, 1340, 1341, 1342, 1343, 1344, 1345, 1346, 1347, 1348, 1349, 1350, 1352, 1353, 1354, 1355, 1356, 1357, 1358, 1359, 1423, 1395, 1408, 1368, 1396, 1397, 1398, 1399, 1401, 1402, 1403, 1404, 1406, 1407, 1409, 1410, 1412, 1413, 1414, 1415, 1416, 1417, 1418, 1419, 1420, 1422, 1424, 1369, 1370, 1371, 1372, 1373, 1374, 1375, 1376, 1377, 1378, 1379, 1380, 1381, 1382, 1383, 1384, 1386, 1387, 1388]. Bibliography [1389, 1390, 1391, 1392, 1393, 1394, 1405, 1437, 1438, 1441, 1442, 1457, 1472, 1468, 1444, 1425, 1461, 1462, 1463, 1464, 1470, 1471, 1426, 1427, 1428, 1429, 1430, 1431, 1432, 1433, 1434, 1435, 1436, 1439, 1440, 1443, 1446, 1450, 1451, 1452, 1453, 1454, 1455, 1456, 1458, 1459, 1460, 1465, 1445, 1447, 1449, 1491, 1473, 1527, 1541, 1543, 1531, 1532, 1533, 1534, 1535, 1488, 1490, 1477, 1480, 1481, 1482, 1484, 1485, 1486, 1487, 1489, 1546, 1474, 1521, 1522, 1523, 1524, 1525, 1526, 1528, 1530, 1536, 1538, 1539, 1540, 1542, 1545, 1475, 1495, 1496, 1497, 1498, 1499, 1500, 1501, 1502, 1503, 1504, 1505]. Bibliography [1506, 1507, 1508, 1509, 1510, 1512, 1513, 1514, 1515, 1516, 1517, 1519, 1520, 1483, 1478, 1493, 1494, 1596, 1597, 1598, 1599, 1600, 1601, 1602, 1603, 1604, 1552, 1569, 1557, 1620, 1617, 1618, 1565, 1573, 1595, 1574, 1581, 1583, 1584, 1588, 1579, 1551, 1614, 1568, 1576, 1582, 1587, 1607, 1608, 1609, 1610, 1611, 1612, 1613, 1616, 1605, 1606, 1548, 1580, 1615, 1621, 1550, 1553, 1554, 1555, 1556, 1558, 1559, 1560, 1561, 1562, 1563, 1564, 1566, 1567, 1571, 1572, 1575, 1577, 1578, 1590, 1591, 1592, 1593, 1594, 1619, 1589, 1631, 1627, 1715, 1625, 1637, 1626, 1623, 1624, 1655, 1656, 1657, 1658, 1659, 1660]. Bibliography [1661, 1662, 1663, 1664, 1665, 1666, 1667, 1668, 1669, 1670, 1671, 1672, 1673, 1674, 1675, 1676, 1677, 1678, 1679, 1680, 1628, 1681, 1682, 1683, 1684, 1685, 1686, 1687, 1688, 1689, 1690, 1691, 1692, 1693, 1694, 1695, 1696, 1697, 1698, 1699, 1700, 1701, 1702, 1703, 1704, 1705, 1706, 1629, 1707, 1708, 1709, 1710, 1711, 1712, 1713, 1714, 1630, 1633, 1634, 1635, 1636, 1638, 1639, 1640, 1641, 1642, 1643, 1644, 1645, 1646, 1647, 1648, 1649, 1650, 1651, 1652, 1653, 1654, 1731, 1717, 1746, 1748, 1740, 1743, 1716, 1734, 1735, 1736, 1737, 1732, 1733, 1718, 1719, 1757, 1758, 1759, 1760, 1761]. Bibliography [1762, 1763, 1764, 1765, 1766, 1767, 1768, 1769, 1770, 1771, 1772, 1773, 1774, 1775, 1776, 1777, 1778, 1720, 1721, 1722, 1723, 1724, 1725, 1726, 1727, 1728, 1729, 1730, 1741, 1742, 1744, 1745, 1747, 1749, 1750, 1751, 1752, 1753, 1754, 1755, 1756, 377, 359, 1252, 346, 789, 1324, 536]. bibsql [934]. Bids [251]. Biennial [679]. Big [1630]. Biochemistry [1773, 1774, 1775, 1776, 1777]. Bioinformatics [775]. Biological [913]. Biology [775, 1680, 1718, 1779, 442, 1348, 1397, 1496, 1497, 1498, 1499, 1500, 1501, 1502, 1503, 1504, 1505, 1506, 1507, 1508, 1509, 1519, 1571, 1661, 1662, 1663, 1664, 1665, 1666, 1667, 1767, 1768, 1769, 1770, 1771, 1772]. biomedical [253]. Biometrika [1213, 1214, 1215, 1216, 1217, 1218, 1219, 1220, 1221, 1222, 1223, 1224, 1575]. Biophysics [1396]. Biotechnology [1778]. birthday [1790]. BIT [558, 141, 144, 184, 558]. Bitcoin [1423]. BJHS [1364]. black [191]. black-box [191]. BLAS [695]. BMDP [362]. Board [1681, 1682, 1683, 1684, 1685]. Bohr [950]. book [1157, 1468]. Books [281, 318, 324, 353, 354, 533, 728, 734, 769, 894, 1107]. Born [986]. Borwein [1448, 1361]. Boston [1468]. box [191]. boxes [880]. Brief [886, 1469]. Brinch [1746]. British [914, 912, 915, 916, 1076, 1365, 1395]. Broglie [1399, 1403]. BSHM [1395]. budoucnosti [271]. Bugs [138]. Building [1232]. Bulletin [604, 962, 963, 1108, 1354, 1395, 1396, 1397, 1772, 1065, 1067, 1068, 1069, 1070, 1066, 1564, 337, 604, 963, 1723]. Bunge [1478]. Business [1587]. Byte [371, 449, 450, 452, 451].

C [286, 434, 430, 735, 991, 1157, 57, 695, 689, 707, 708, 710, 785]. C# [850]. C/C [430, 735]. Cahiers [307]. CALCOMP [50, 23]. Calculation [1, 182, 949, 105, 107]. Calculations [3, 66, 92, 110, 112, 59, 108]. calculator [711]. Calisthenics [373]. Canada [1681, 1682, 1683, 1684, 1685, 1681, 1682, 1683, 1684, 1685]. Canadian [604, 963, 964, 966, 970, 971, 969, 968, 967, 965, 1305, 1655, 1656, 1657, 1658, 1659]. canadien [604, 963, 964, 966, 970, 971, 969, 968, 967, 965, 1655, 1656, 1657, 1658, 1659]. canadienne [1305]. Card [37]. Carlo [1037]. Case [442, 1779]. Cash [1423]. catalog [705]. Cataloging [2]. Catalogued [19, 18]. Cation [53]. CD [101]. Cell [442, 1497, 1498, 1499, 1500, 1501, 1502, 1503, 1504, 1505, 1506, 1507, 1508, 1509, 1571, 1779]. Centaurus [1136]. Center [35]. Century [1246]. CH [101, 109]. Challenges [1246]. Cham [1789]. Chance [1296]. Chances [1783]. Changes [21, 32]. Changing [12]. chapter [1157, 1469]. \special [275]. Character [142, 145, 227, 259]. Charges [39]. Charles [1748, 1441]. checking [326, 327, 997]. Chemical [190]. Chemistry [120, 148, 174, 180, 1785, 1787, 1786, 375, 613, 913, 977, 979, 980, 982, 981, 978, 1542, 129, 179, 204, 648, 1078, 1077, 1079, 1080, 1354, 1446, 1572, 1787]. Chicago [577, 1108]. Chips [649, 650]. Choice [92]. Chymia [913]. CI [92]. Clang [1232]. Clarke [1056]. CLASS=T [39]. classes [863, 864]. Classic [789, 847, 857]. Claude [925]. Clerk [1451]. Cleve [390]. Clipping [206]. CO [113]. Codd [874]. Code [772, 767, 695, 752, 1411]. Codes [579, 796]. Coding [146]. Cody [347, 781]. Cody/Waite [781]. Cold [1358]. Collected [53, 315, 355]. Collection [295]. College [233, 1477]. Collnet [1322]. Color [246, 804]. Column [313]. Combinatorics [870, 975]. Combined [80, 188]. Command [51, 240]. Command-Line [240]. Commands [275]. Comment [46]. Comments [267, 1246]. Common [492]. Communication [1151]. Communications [341, 418, 419, 420, 421, 422, 605, 770, 784, 888, 1244, 1286, 1335, 1336, 1337, 1338, 1339, 1340, 1341, 1342, 1343, 1344, 1345, 1545, 1548, 1580, 1030, 1031, 1033, 1034, 1035, 1032, 1539]. Companion [1323]. compare [610]. Comparison [126]. Compilation [15]. Compiler [20, 694, 704, 706, 785]. Compilers [805, 702, 1232, 1421]. Complete [290, 277, 289, 291, 286, 304, 350, 363, 367, 438, 440, 439, 396, 401, 430, 428, 429, 471, 460, 561, 539, 554, 558, 548, 559, 542, 544, 545, 546, 547, 543, 572, 574, 575, 576, 570, 577, 592, 594, 596, 622, 624, 606, 608, 616, 619, 620, 621, 627, 646, 647, 648, 669, 671, 672, 673, 674, 675, 670, 660, 651, 655, 656, 659, 662, 663, 661, 666, 664, 678, 677, 680, 682, 684, 685, 1738, 1739, 712, 717, 746, 726, 748, 719, 721, 731, 733, 735, 745, 747, 754, 755, 753, 756, 757, 758, 759, 762, 772, 773, 775, 774, 763]. Complete [764, 765, 766, 767, 770, 784, 783, 840, 841, 842, 792, 786, 788, 791, 797, 796, 794, 795, 799, 798, 793, 800, 833, 835, 849, 865, 869, 876, 878, 871, 872, 873, 877, 881, 882, 883, 884, 885, 887, 909, 926, 928, 918, 920, 919, 921, 923, 924, 922, 929, 927, 896, 897, 898, 899, 901, 903, 904, 905, 906, 907, 908, 914, 912, 913, 915, 916, 925, 930, 933, 935, 989, 962, 975, 957, 956, 958, 960, 932, 994, 993, 938, 990, 992, 991, 996, 1006, 939, 940, 951, 961, 963, 964, 966, 970, 971, 969, 968, 967, 965, 974]. Complete [977, 979, 980, 982, 981, 978, 984, 983, 985, 987, 988, 1021, 1022, 1023, 1024, 1027, 1026, 1025, 1020, 1012, 1015, 1013, 1014, 1016, 1017, 1018, 1019, 1007, 1078, 1087, 1088, 1071, 1072, 1073, 1075, 1074, 1081, 1083, 1086, 1084, 1082, 1008, 1009, 1010, 1011, 1029, 1028, 1030, 1031, 1033, 1034, 1035, 1032, 1037, 1038, 1040, 1063, 1067, 1068, 1069, 1070, 1066, 1076, 1077, 1079, 1080, 1042, 1043, 1047, 1050, 1051, 1046, 1045, 1048, 1049, 1044, 1062, 1151, 1113, 1144, 1145, 1146, 1147, 1148, 1152, 1153, 1155, 1156, 1114, 1115, 1116, 1117, 1118, 1119, 1120, 1121, 1122, 1123, 1124, 1125]. Complete [1129, 1130, 1131, 1132, 1134, 1135, 1137, 1138, 1139, 1140, 1141, 1142, 1143, 1167, 1251, 1243, 1244, 1245, 1248, 1236, 1237, 1238, 1239, 1240, 1241, 1242, 1249, 1188, 1189, 1190, 1191, 1192, 1193, 1194, 1195, 1158, 1186, 1187, 1197, 1198, 1199, 1200, 1201, 1202, 1203, 1204, 1205, 1206, 1207, 1208, 1209, 1210, 1211, 1212, 1213, 1214, 1215, 1216, 1217, 1218, 1219, 1220, 1159, 1221, 1222, 1223, 1224, 1225, 1226, 1227, 1228, 1229, 1230, 1231, 1234, 1250, 1253, 1294, 1160, 1161, 1162, 1163, 1164, 1165, 1166, 1169, 1170, 1171, 1172, 1173, 1174, 1175, 1176, 1177, 1178, 1179, 1180, 1181, 1182]. Complete [1183, 1184, 1185, 1303, 1304, 1325, 1256, 1327, 1257, 1258, 1259, 1260, 1261, 1262, 1263, 1264, 1265, 1266, 1267, 1268, 1269, 1270, 1271, 1272, 1273, 1274, 1275, 1276, 1277, 1278, 1279, 1280, 1281, 1302, 1255, 1315, 1316, 1317, 1318, 1319, 1320, 1321, 1322, 1333, 1282, 1283, 1284, 1285, 1286, 1287, 1288, 1289, 1290, 1291, 1292, 1293, 1295, 1297, 1298, 1300, 1301, 1305, 1306, 1307, 1310, 1312, 1313, 1314, 1326, 1334, 1360, 1361, 1362, 1364, 1365, 1366, 1367, 1335, 1336, 1337, 1338, 1339, 1340, 1341, 1342, 1343, 1344, 1345, 1346, 1347, 1348, 1349, 1350, 1352, 1353, 1354, 1355, 1356, 1357, 1358]. Complete [1359, 1408, 1368, 1396, 1397, 1398, 1399, 1401, 1402, 1403, 1404, 1406, 1407, 1409, 1410, 1412, 1413, 1414, 1415, 1416, 1417, 1418, 1419, 1420, 1422, 1424, 1369, 1370, 1371, 1372, 1373, 1374, 1375, 1376, 1377, 1378, 1379, 1380, 1381, 1382, 1383, 1384, 1386, 1387, 1388, 1389, 1390, 1391, 1392, 1393, 1394, 1425, 1461, 1462, 1463, 1464, 1470, 1471, 1426, 1428, 1429, 1430, 1431, 1432, 1433, 1434, 1435, 1436, 1439, 1440, 1443, 1446, 1450, 1451, 1452, 1453, 1454, 1455, 1456, 1458, 1459, 1460, 1543, 1531, 1532, 1533, 1534, 1535, 1488, 1490, 1477, 1480, 1481, 1482, 1484, 1485, 1486, 1487, 1489]. Complete [1546, 1474, 1521, 1522, 1523, 1524, 1525, 1526, 1528, 1530, 1536, 1538, 1539, 1540, 1542, 1545, 1475, 1495, 1496, 1497, 1498, 1499, 1500, 1501, 1502, 1503, 1504, 1505, 1506, 1507, 1508, 1509, 1510, 1512, 1513, 1514, 1515, 1516, 1517, 1519, 1520, 1483, 1574, 1581, 1583, 1584, 1588, 1579, 1551, 1614, 1568, 1576, 1582, 1587, 1607, 1608, 1609, 1610, 1611, 1612, 1613, 1616, 1605, 1606, 1548, 1580, 1615, 1621, 1550, 1553, 1554, 1555, 1556, 1558, 1559, 1560, 1561, 1562, 1563, 1564, 1566, 1567, 1571, 1572, 1575, 1577, 1578, 1590, 1591, 1592, 1593, 1594, 1619, 1589, 1625, 1637, 1626, 1623, 1624, 1655]. Complete [1656, 1657, 1658, 1659, 1660, 1661, 1662, 1663, 1664, 1665, 1666, 1667, 1668, 1669, 1670, 1671, 1672, 1673, 1674, 1675, 1676, 1677, 1678, 1679, 1680, 1628, 1681, 1682, 1683, 1684, 1685, 1686, 1687, 1688, 1689, 1690, 1691, 1692, 1693, 1694, 1695, 1696, 1697, 1698, 1699, 1700, 1701, 1702, 1703, 1704, 1705, 1706, 1629, 1707, 1708, 1709, 1710, 1711, 1712, 1713, 1714, 1630, 1633, 1634, 1635, 1636, 1638, 1639, 1640, 1641, 1642, 1643, 1644, 1645, 1646, 1647, 1648, 1649, 1650, 1651, 1652, 1653, 1654, 1732, 1733, 1718, 1719, 1757, 1758, 1759, 1760, 1761, 1762, 1763, 1764, 1765, 1766]. Complete [1767, 1768, 1769, 1770, 1771, 1772, 1773, 1774, 1775, 1776, 1777, 1778, 1720, 1721, 1722, 1724, 1725, 1726, 1727, 1728, 1729, 1730, 1741, 1742, 1744, 1745, 1747, 1749, 1750, 1751, 1752, 1753, 1754, 1755, 1756, 377]. Complexity [1381]. Composites [1781, 1784]. Composition [414]. comprehensive [253]. Compressing [33]. Compression [132, 1109]. Computability [1243]. Computation [129, 173, 270, 269, 437, 453, 684, 1738, 1739, 748, 736, 737, 882, 1167, 1183, 1246, 1400, 1540, 1789, 1071, 1072, 1073, 1075, 1074, 1788, 393, 468, 454, 542, 544, 545, 546, 547, 543, 580, 581, 616, 763, 794, 793, 873, 897, 905, 1253, 1341, 1342, 1343, 1344, 1345, 1580, 1553]. Computational [3, 210, 485, 571, 622, 624, 723, 775, 1001, 1002, 1003, 1004, 1005, 1006, 1007, 1078, 1008, 1009, 1010, 1011, 1077, 1079, 1080, 1233, 1290, 1291, 1292, 1293, 1298, 1369, 1370, 1371, 1372, 1373, 1374, 1375, 1436, 1528, 1530, 1536, 1538, 1520, 1588, 1616, 1572, 1589, 1724, 1788]. Computations [190, 1106]. Computer [24, 30, 84, 147, 207, 217, 223, 248, 292, 302, 328, 360, 391, 376, 383, 397, 373, 378, 392, 405, 410, 407, 411, 412, 409, 408, 447, 448, 396, 453, 466, 479, 489, 490, 491, 498, 499, 500, 502, 501, 521, 531, 541, 540, 564, 578, 598, 599, 597, 600, 595, 572, 575, 577, 643, 631, 627, 681, 665, 651, 666, 664, 722, 725, 729, 727, 718, 715, 713, 716, 714, 742, 740, 741, 739, 749, 750, 755, 782, 792, 798, 800, 853, 852, 890, 892, 901, 948, 947, 943, 944, 945, 946, 936, 937, 940, 1041, 1094, 1095, 1096, 1097, 1098]. Computer [1099, 1100, 1101, 1102, 1103, 1104, 1105, 1052, 1081, 1083, 1086, 1084, 1082, 1030, 1031, 1033, 1034, 1035, 1032, 1040, 1152, 1153, 1155, 1168, 1235, 1254, 1282, 1357, 1444, 1541, 1474, 1539, 1510, 1529, 1570, 1567, 1623, 1628, 1629, 1731, 1732, 1733, 1728, 1729, 1730, 1157, 626, 629, 630, 632, 35, 413, 895, 1565]. Computer-Based [360]. Computer-Human [575]. Computers [27, 465, 1376, 1377, 1378, 1379, 1380, 1531, 1532, 1533, 1534, 1535, 955, 959, 957, 956, 958, 960, 1551]. Computing [46, 289, 308, 363, 431, 570, 617, 662, 663, 661, 719, 721, 754, 756, 762, 766, 784, 833, 806, 885, 903, 904, 932, 1062, 1251, 1390, 1392, 1543, 1583, 1566, 1625, 1630, 380, 385, 475, 477, 456, 484, 483, 563, 554, 601, 683, 655, 682, 724, 745, 764, 870, 876, 975, 1088, 1106, 1176, 1177, 1284, 1443, 1747]. Concepts [226]. Concurrency [594, 684, 1738, 1739, 553]. Conference [204, 437, 1788, 1785]. Conferences [443, 628]. Configuration [3, 66, 92, 80, 147, 172, 107]. conformance [776]. Conjecture [886]. Conservation [1767, 1768, 1769, 1770, 1771]. Consistent [59]. Constantine [860]. Constants [228, 101, 113, 112, 107, 59, 108]. Construction [40, 62]. contd [676]. Contemporary [1333, 1349]. Context [1135]. Control [567, 1172, 1475, 566, 1170]. Controller [9]. Conventions [134]. Convert [49]. Cool [649]. COPE [9]. COPY [102]. Copying [4, 33]. Core [16, 45, 212, 226]. Corfu [1788]. Cornelius [954]. Correction [185, 187, 193, 196, 202, 211, 224, 360]. Corrections [165, 166, 178, 214]. Cost [250, 209, 611]. Countess [1442]. Coupled [108]. Coupling [59, 108]. CPU [856]. Crandall [1112]. Cross [73, 116]. Cross-reference [116]. Cross-Referencing [73]. Cryptographic [1234]. Cryptography [562, 623, 839, 889, 900, 1557, 579, 796, 1036, 1382]. Cryptologia [560]. Cryptology [559, 1381, 1383]. Crystal [156, 197]. CSC [234]. Cultural [932, 1136]. Curves [87]. CVIU} [392, 1155]. Cyber [1391]. Cyber-Physical [1391].

D [1256, 1257, 1258, 1259, 1260, 1261].

Dædalus [1300]. Data

[8, 4, 10, 33, 31, 26, 36, 54, 122, 132, 443, 400, 619, 883, 887, 1109, 1290, 1291, 1292, 1293, 1521, 1536, 1584, 628]. Database [444, 445, 533, 628]. databases [874]. Datamation} [387]. date [389, 392, 396, 496, 613, 836, 913, 1301, 1477, 916]. Daubechies [1437]. DEC [233, 250, 249]. DEC-20 [233, 250, 249]. DEC-20/60 [250, 249]. decade [964, 966, 970, 971, 969, 968, 967, 965, 1213, 1214, 1215, 1216, 1217, 1218, 1219, 1220, 1221, 1222, 1223, 1224, 1225, 1226, 1227, 1228, 1229, 1230, 1231, 1575, 1578]. Decimal [1421]. Decimal-arithmetic [1421]. DECstation [255]. DECsystem [134]. DECsystem-20 [134]. Dedicated [874]. Deep [1749]. della [1308]. DEM67 [212]. Democracy [839]. Demonstration [175, 176]. Densities [44]. Density [40, 64, 62]. Department [641, 688, 693, 705, 707, 708, 709, 710, 657, 769, 1247, 1466, 1467, 1476, 1479, 1492, 1518, 1585, 1586]. Dependent [113, 112, 59]. Descendant [177]. Design [256, 685, 1151, 1622, 846, 845, 290, 291]. Designs [579, 796]. Detection [360]. Determinants [170]. Determination [91, 109]. Development} [433, 344, 1721]. Device [161]. devices [209]. d’Histoire [1309, 1439]. Dialectica [1401]. Dialup [807]. Diatomic [1, 60, 78, 70, 108]. Dictyostelium [402]. differences [610]. Different [119]. Digital [374, 765, 1423, 1579]. Dijkstra [1299]. Dirac [952]. directions [867]. Directory [699]. discoideum [402]. Discovery [883]. Discrete [522, 523, 659, 1036]. Discussions [146, 147, 148]. Diseases [1708, 1709, 1710, 1711, 1712, 1713]. Disk [11, 26, 43, 808]. Disks [39]. Dispersion [59]. Dissemination [290, 291]. Dissertations [1466, 1467, 1476, 1479, 1492, 1518, 1585, 1586, 1149]. Distributed [416, 473, 474, 475, 598, 744, 762, 871, 904, 1740, 1741, 1742, 1747]. Distribution [68, 86, 99, 100, 115, 117, 167, 751, 1000, 186]. Dobb [373, 634, 635, 633]. Doctrine [1783]. Document [114, 124, 200, 67, 83, 158, 200]. Documentation [2, 146, 643, 1318]. Documents [298, 312, 320, 438, 440, 439, 561]. domain [237]. Donald [342]. Dongarra [359]. downloadable [209]. Dr [373, 634, 635, 633]. Drawing [280, 191]. Driver [236, 239, 247, 275, 237]. Drumcard [34]. Dump [167]. Dumps [16, 14]. DVI [237, 236, 239, 247, 275]. Dynamic [59]. Dynamical [758]. Dyson [1150].

ECMA [660]. Ecological [1638, 1639, 1640, 1641, 1642, 1643]. Ecology [442, 1760, 1761, 1762, 1763, 1764, 1765, 1779]. Econometrics [1488, 1490, 1481, 1482, 1484, 1485, 1486, 1487, 1489, 1568]. Economic [1587]. Economics [1167, 1766]. Edgar [874]. EDIT [71]. Editing [71, 103, 125, 279, 278, 323]. Edition [203, 201, 1252]. Editor [520, 150]. Editorial [129]. Editors [810]. Edsger [1299]. Eduard [1620]. Education [885, 1394, 1409]. Educational [766, 1368]. Edward [976]. Efficient [863]. Eigenvalue [61, 96, 106, 123, 120, 131, 691]. Eigenvalues [118]. eigenvector [691]. Eigenvectors [118]. Einstein [777]. EISPACK [691]. Electron [38, 36, 110, 63, 107]. Electronic [57, 66, 290, 282, 291, 414, 685, 657, 809, 906, 955, 1244, 1245, 1248]. Elementary [453, 781]. Elements [170]. Elmer [838]. Elwood [925]. Embedded [754]. Emergent [871]. Emerging [833]. Empirical [1366]. Emulator [856]. Encapsulated [299, 320, 319]. encoding [259]. Endowment [951]. Energies [65, 183, 111, 109]. Energy [112]. Energy-Weighted [112]. Engineering [561, 669, 1407, 1452, 1453, 1454, 1455, 1456, 1788, 578, 571, 570, 624, 1153, 1234, 1366, 1431, 1432, 1433, 1434, 1435, 1550]. Enrico [1055]. Environment [322]. Epsilon [308]. equation [696]. Equipment [248, 246]. Ernest [1332]. Errata [194]. Error [360]. Erwin [1054]. ESAIM [1297]. Estimates [250, 58]. ETNA [906]. Ettore [1133]. Eugene [1112, 1058]. European [1333]. Evaluating [251]. Evaluation [87, 869, 1392]. Everyone [839]. Evolution [1428, 1430]. Evolutionary [1637]. Exact [933]. exceptional [611]. Excitation [111, 109]. Experience [1738, 1739, 339, 594, 684, 904, 1062]. Experimental [53, 877, 1038, 1246]. Explorer} [400]. Exponential [603]. Express [694]. Expressions [1111].

EXTENDED [28, 256, 321, 781, 25, 198, 1385]. Extending [866, 1784]. extensible [711]. Extensive [1624].

F [1789, 332]. facilities [806]. Fall [1466, 1467]. Family [236, 239, 247, 1250, 1328, 1408, 237, 1262, 1263, 1264, 1265, 1266, 1267, 1268, 1269, 1270, 1271, 1272, 1273, 1274, 1275, 1276, 1277, 1278, 1279, 1280, 1281, 1477, 1546]. FAQ [802, 803, 806, 804, 805, 807, 808, 809, 810, 811, 812, 813, 814, 815, 816, 817, 818, 819, 820, 821, 822, 823, 824, 825, 826, 827, 828, 829, 830, 831, 832, 856, 997, 1330]. Fast [35, 549, 603]. Feler [118]. Fermi [1055]. Festschrift [1790]. Feynman [168, 942]. FH [113]. Fibonacci [987, 768]. Fields [1184, 1185]. Fifth [1788]. File [132, 300]. Files [37, 73, 76, 812, 610]. Filesystem [641, 688]. Filled [97]. Financial [898]. Finding [118]. Fish [1660, 1661, 1662, 1663, 1664, 1665, 1666, 1667, 1680, 1708, 1709, 1710, 1711, 1712, 1713, 1717, 1773, 1774, 1775, 1776, 1777]. Fisher [1595]. Fisheries [1655, 1656, 1657, 1658, 1659, 1681, 1682, 1683, 1684, 1685, 1686, 1687, 1688, 1689, 1690, 1691, 1692, 1693, 1694, 1695, 1696, 1697, 1698, 1699, 1700, 1701, 1702, 1703, 1704, 1705, 1706, 1707, 1644, 1645, 1646, 1647, 1648, 1719, 1722, 1660, 1680, 1717, 1720]. FisheriesSciences.com [1716]. Fishery [1723]. Fixed [1622]. Fixed-Point [1622]. Floating [228, 305, 310, 346, 690, 866, 867]. Floating-Point [228, 305, 310, 346, 690, 866, 867]. Fock [60, 95, 78, 79, 181, 107, 59, 108]. FOCS [665, 666, 664, 1040]. Font [1328, 209]. Fonts [399, 813, 687]. Forces [59]. Fork [195]. Form [64]. FORMAT [261]. Formatting [67, 83, 114, 124, 158]. Formerly [373, 372]. Forsythe [838]. ForTec [717]. FORTGLIB [18].

FORTHLIB [18]. Fortran

[694, 706, 717, 13, 15, 12, 14, 20, 28, 19, 25, 71, 103, 104, 125, 141, 142, 144, 145, 159, 188, 201, 227, 388, 389, 406, 496, 184, 242, 253, 261, 272, 676, 695, 689, 708, 709, 710, 732]. Forum [717, 732, 292, 302, 717, 1731, 1732, 1733]. Forward [778]. Foundations [564, 665, 666, 664, 725, 1040, 1129, 1388]. Frank [874, 1790, 862, 1154, 1632]. Freedom [839]. Freeman [1150]. Frequency [113, 112]. Frequency-Dependent [113, 112]. Friedrich [1617]. Frisch [1447]. Frontiers [1422, 1715, 1714]. FTP [811]. Fujitsu [694]. Fun [768]. Function [781, 1400, 1789, 698]. Functional [1398]. Functions [453, 603, 863, 864]. Fund [46]. Fundamenta [1351, 1352, 1353]. Funds [246]. Future [727, 800, 1728, 1729, 1730, 186, 267]. FYI [438].

G [1537]. Gambling [1782, 1783]. Game [907]. Games [1782]. Gamow [1053]. Gauss [863, 864]. Gauss-related [863, 864]. Gaussian [81, 173]. Gautschi [1552]. Gazette [1302]. gcc [1421]. Gebiete [1185]. Gene [349]. General [38, 215, 260, 1327, 1326, 63]. Generalized [61, 123, 120, 361, 730, 911]. generating [864]. Generation [727, 751, 800, 1000, 1728, 1729, 1730]. Generations [1461]. Geo [1458]. Geo- [1458]. Geometry [723, 1420, 1520]. George [771, 838, 1053, 1494]. German [857]. Geschichte [1355]. Ghostscript [693]. Gilbert [1196]. GIVENS [29]. Glossary [226]. glue [880]. GMAKE [256, 257]. GMIP} [392]. GNU [441, 692, 1157]. Go [1537]. Golub [349]. Google [893]. Goudsmit [1493]. Government [1579]. GP [824]. GPU [907]. Grading [98]. Graeme [1331]. GRAPH [191, 260]. Graphical [392, 485, 1156, 1616]. Graphics [207, 217, 223, 248, 246, 244, 251, 292, 302, 288, 376, 378, 392, 471, 907, 1628, 1731, 1732, 1733, 253, 243, 345, 391, 465, 651, 729, 726, 890, 1623]. Graphing [260]. Gray [874]. Greece [1788]. Greetings [254]. Grid [764]. Group [702, 955]. Groups [1381]. Guest [129]. Guide [136, 164, 154, 163, 194, 203, 219, 230, 879]. GUTenberg [307, 536].

H [1789, 28, 25, 349, 332, 1128, 1333, 7, 20, 59]. halieutiques [1655, 1656, 1657, 1658, 1659]. Hamming [1457]. Handbook [1400, 1789]. Hans [1059]. Hansen [1746]. Hardware [814]. Harris [1154]. Hartree [79, 60, 95, 78, 181, 107, 59, 108]. Harwell [128]. Hash [582]. Hashing [333]. Header [167]. Headers [13, 300]. HEALTH [1583]. Healthcare [1583]. Heap [995]. Heaps [998]. Heavens [676]. Heinz [1618]. Heisenberg [1060]. Held [1785]. Hellmann [168]. Henri [1210]. Henrici [1473]. Heritage [932]. Hewlett [417, 49]. HF [59]. Hidden [50, 44, 47, 162]. Hidden-Line [47, 50, 44, 162]. HIDEPLOT [56, 86]. High [246, 265, 406, 380, 724, 711]. high-order [711]. High-Performance [246, 265, 406, 380]. Higher [580, 794]. Higher-Order [580, 794]. Highlights [1529]. Hilbert [89]. hints [1149]. Historia [1131, 1465]. Historical [913, 1333]. History [683, 682, 745, 914, 912, 913, 915, 933, 990, 992, 991, 1076, 1107, 1136, 1148, 1249, 1303, 1304, 1334, 1365, 1350, 1354, 1355, 1395, 1468, 1461, 1463, 1439, 1446, 1450, 1458, 1465, 1519, 1408]. Hoare [1748]. hoc [711]. Hodgman [206]. honor [1790]. HOPOS [1304]. Hot [650, 679]. Householder [1472]. Human [575, 1582, 1629]. Human-Computer [1629]. Human-Robot [1582]. Hydroxyl [53]. Hyperbolic [270]. Hyperfine [107].

I.R.E [955]. I/O [6, 134]. IA [652]. IA-64 [652]. IBM [344, 343, 401, 11, 26, 27, 39, 54, 232]. ICCMSE [1788, 868]. ICERM [1246]. ICES [1669, 1670, 1671, 1672, 1673, 1674, 1675, 1676, 1677, 1678, 1679]. Icon [516]. Ideas [137, 188, 189, 321]. IEC [617]. IEEE [378, 381, 413, 472, 473, 474, 553, 571, 593, 626, 629, 630, 632, 650, 679, 651, 656, 682, 744, 753, 775, 895, 955, 959, 957, 956, 958, 960, 1325, 1431, 1432, 1433, 1434, 1435, 1510, 1565, 1551, 1550, 1556, 1558, 1559, 1560, 1561, 1562, 1563, 1740, 1743, 1741, 1742, 1744, 1745, 228, 305, 310, 564, 618, 611, 665, 666, 664, 690, 1040, 1623]. IEEE/ACM [593, 775, 1325]. IEN [561]. ignoring [610]. Ihre [857]. II [1749]. III [1196]. IJPEDS [871]. IMA [646, 647]. IMACS [467]. Image [392, 726, 1155]. Imaging [899, 414]. Impact [328]. Implementation [256, 321]. Improvement [1427]. Improvements [98]. IMS [1584]. IMWUT [1389]. Inc. [702]. Incomplete [1465]. Incorporating [320]. Independent [143, 141, 142, 144, 145, 151, 184, 227, 322, 104]. INDEXER [200]. Indexing [298, 520]. Indirect [108]. Industrial [566, 548, 1170, 1495]. Infeld [1387]. Inference [1512, 1513, 1514, 1515, 1516, 1517]. Informatica [606]. Informaticae [1351, 1352, 1353]. Information [25, 340, 479, 555, 556, 557, 574, 576, 592, 621, 746, 887, 935, 938, 1110, 1319, 1320, 1321, 1322, 1384, 1426, 1727, 765, 1726]. Informational [1407]. Informationsbehandling [558]. Informatique [598]. Informetrics [1317]. INFORMS [1443]. Ingrid [1437]. Initiative [1247]. initio} [57]. Innovations [1422]. Input [143, 151]. Input/Output [143, 151]. Installation [164, 154, 203, 219, 230]. Installer [700]. Institut [1210]. Institute [311, 356, 733, 1283]. Instruction [653]. instructions [611]. Integer [306]. Integral [80, 94, 110, 119, 132, 173]. Integrals [119]. Intel [680, 652]. Intelligence [1724]. Intelligent [926, 989]. Interaction [3, 66, 92, 80, 172, 1582, 1629, 107, 575]. Interactive [989, 1389, 1628]. Interface [143, 151, 161, 240, 382, 1537, 1544]. International [443, 375, 380, 437, 479, 478, 578, 598, 613, 723, 722, 725, 724, 726, 774, 795, 871, 977, 979, 980, 982, 981, 978, 1136, 1153, 1212, 1304, 1306, 1409, 1382, 1444, 1462, 1426, 1465, 1542, 1614, 1788]. Internazionale [1308]. Internet [712, 1581, 318, 353, 438, 440, 439, 561, 734, 894]. Interpretation [58]. Interval [370, 1106]. Introduction [234, 1469, 848, 262, 276, 297]. Ions [36]. IRE [955]. ISBN [1789]. ISDN [599]. Isis} [1144, 1145, 1146, 1147, 1137, 1138, 1139, 1140, 1141, 1142, 1143]. ISO [470]. ISSAC [437]. Issues [251, 732].

J [347, 1150, 359]. J. [972]. J.UCS [489, 798]. Jack [359]. James [874, 1451]. Japan [1465]. Japanese [1463]. Java [368, 369, 433, 614, 910, 1569, 768]. JavaWorld [535]. JDIQ [887]. JERIC [766]. JETC [833]. JGT [471, 907]. JHRI [1582]. Jiao [879]. Jim [874]. Job [21, 32, 121]. jobs [803, 857]. JOCCH [932]. John [788, 865, 953, 1056]. Jonathan [1448, 1361]. Journal [286, 316, 304, 344, 350, 373, 375, 380, 393, 427, 431, 428, 429, 479, 478, 475, 487, 485, 489, 490, 456, 457, 458, 484, 483, 507, 471, 460, 508, 509, 511, 512, 513, 514, 515, 510, 566, 568, 522, 523, 567, 554, 548, 559, 569, 578, 598, 573, 577, 643, 634, 635, 633, 613, 625, 609, 608, 619, 646, 647, 659, 678, 677, 723, 722, 725, 724, 726, 748, 733, 747, 758, 759, 774, 764, 765, 766, 797, 795, 798, 833, 871, 877, 887, 931, 898, 899, 907, 908, 914, 912, 915, 916, 930, 1001, 1002, 1003, 1004, 1005, 941, 962]. Journal [932, 994, 993, 1006, 961, 964, 966, 970, 971, 969, 968, 967, 965, 977, 979, 980, 982, 981, 978, 984, 983, 985, 988, 1012, 1015, 1013, 1014, 1016, 1017, 1018, 1019, 1007, 1078, 1087, 1088, 1036, 1008, 1009, 1010, 1011, 1076, 1077, 1079, 1080, 1136, 1153, 1245, 1248, 1236, 1237, 1238, 1239, 1240, 1241, 1242, 1249, 1187, 1234, 1250, 1170, 1171, 1172, 1173, 1174, 1175, 1176, 1177, 1183, 1185, 1304, 1256, 1327, 1257, 1258, 1259, 1260, 1261, 1262, 1263, 1264, 1265, 1266, 1267, 1268, 1269, 1270, 1271, 1272, 1273, 1274, 1275, 1276, 1277, 1278, 1279, 1280, 1281, 1317, 1319, 1320]. Journal [1321, 1322, 1333, 1282, 1295, 1298, 1305, 1306, 1326, 1334, 1348, 1355, 1356, 1357, 1358, 1359, 1395, 1408, 1398, 1409, 1420, 1369, 1370, 1371, 1372, 1373, 1374, 1375, 1382, 1383, 1444, 1461, 1426, 1428, 1429, 1430, 1439, 1443, 1446, 1459, 1460, 1465, 1480, 1481, 1482, 1484, 1485, 1486, 1487, 1489, 1546, 1474, 1521, 1528, 1530, 1538, 1540, 1542, 1495, 1497, 1498, 1499, 1500, 1501, 1502, 1503, 1504, 1505, 1506, 1507, 1508, 1509, 1512, 1513, 1514, 1515, 1516, 1517, 1519, 1520, 1483, 1614, 1568, 1576, 1582, 1587, 1607, 1608, 1609, 1610, 1611, 1612, 1613, 1616, 1571, 1572, 1577, 1619, 1624, 1655]. Journal [1656, 1657, 1658, 1659, 1661, 1662, 1663, 1664, 1665, 1666, 1667, 1669, 1670, 1671, 1672, 1673, 1674, 1675, 1676, 1677, 1678, 1679, 1681, 1682, 1683, 1684, 1685, 1687, 1688, 1689, 1690, 1691, 1708, 1709, 1710, 1711, 1712, 1713, 1634, 1635, 1636, 1716, 1721, 901, 1477, 430, 1081, 1083, 1086, 1084, 1082, 1477, 289, 343, 397, 374, 417, 384, 401, 495, 539, 681, 680, 735, 917, 892, 928, 918, 920, 919, 921, 923, 924, 922, 929, 927, 1310, 1488, 1495, 1567, 1626, 1722, 1727]. Journals [284, 392]. Julia [1527]. July [1785].

Kahan [787]. Kepler [886]. Keynote [778, 845]. Keypunch [34]. Keyword [69, 127]. King [1442]. Klassische [857]. Klein [87]. Knowledge [883]. Knuth [342].

Komentár [271]. Komödie [317, 358].

L [171]. L-S [171]. Laboratories [922]. Laboratory [331]. Labs [928, 929, 927, 1626]. Lahey [694]. Lahey/Fujitsu [694]. Lanczos [954]. Language [6, 286, 329, 330, 372, 388, 389, 395, 496, 516, 532, 728, 746, 773, 790, 834, 850, 1325, 1384, 1527, 1627, 614, 910, 1569, 361, 730, 911, 368, 369, 383, 1107, 1624]. Languages [334, 394, 423, 453, 470, 482, 596, 617, 1425, 377, 466, 722]. LAPACK [786, 695]. Large [96, 106, 131, 443, 619]. Law [995]. lcc [785]. Learned [436, 1633]. Learning [1360, 1637]. least [698]. least-squares [698]. Lecture [207, 405, 410, 407, 411, 412, 409, 408, 447, 448, 396, 491, 521, 531, 541, 540, 595, 624, 718, 715, 713, 716, 714, 742, 740, 741, 739, 749, 750, 782, 852, 948, 947, 943, 944, 945, 946, 1041, 1094, 1095, 1096, 1097, 1098, 1099, 1100, 1101, 1102, 1103, 1104, 1105, 1052, 1168, 1235, 1188, 1189, 1190, 1191, 1192, 1193, 1194, 1195, 1787]. Leo [973]. Leopold [1387]. Lessons [436]. Letters [423, 519, 555, 556, 557, 621, 938, 338, 340, 476, 1199, 1200, 1201, 1202, 1388, 1510, 1750]. Leurs [1309]. Libraries [116, 695, 707, 708, 709, 710]. Library [5, 582, 583, 781, 1400, 1622, 691, 696, 698, 875, 128, 692, 697, 1789]. Lifetimes [101]. Limited [751]. Limnology [1757, 1758, 1750, 1751, 1752, 1753, 1754, 1755, 1756]. Line [47, 240, 260, 331, 50, 44, 162, 253, 752, 1411]. Linear [89, 210, 461, 459, 462, 464, 460, 620, 996, 939, 1248, 1255, 1362, 1346, 1347, 696, 817]. Linking [695]. LINPACK [696]. Linux [539, 857, 404, 694]. Linux-Tasks [857]. Lise [1449]. Lisp [581, 793, 492]. List [19]. Listing [37, 122]. Listings [160]. Lists [657]. Literate [287]. Literature [285]. LLVM [1232]. LMS [748]. logarithmic [863, 864]. Logic [622]. Login [815]. London [1440]. Looking [778]. LOPLAS [423]. Lord [1332]. Louis [1399, 1403]. Lovelace [1442]. Low [66, 209, 1384]. Low-cost [209]. Low-Lying [66]. Low-Resource [1384]. Löwdin [640]. lualatex [1549]. Lucent [697]. Ludwig [1617]. Lying [66].

M [1448, 787, 861]. Mach} [435]. Machine [104, 114, 124, 143, 141, 142, 144, 145, 147, 151, 184, 227, 308, 382]. Machine-Independent [143, 141, 142, 144, 145, 151, 184, 227, 104]. Machine-transportable [114, 124]. Machines [801, 851]. Macro [215]. Macros [319]. MACSYMA} [394]. Magazine [367, 387, 383, 386, 425, 535, 642, 645, 644, 449, 450, 452, 451, 371]. Magazines [385]. Magma} [1233]. mail [809]. Mailing [657]. Maintenance [1428, 1429]. Majorana [1133]. MAKE [256]. Management [628, 935, 1322, 1687, 1688, 1689, 1690, 1691]. Mandelbrot [1061]. Manipulation [141, 142, 144, 145, 184, 227]. Manual [152, 257, 153, 205]. Manuals [533, 704]. Manuscripts [284]. MAP [73]. Maple [488, 329, 819, 776]. Mapping [73]. Maps [116]. March [1128]. Marine [1715, 1669, 1670, 1671, 1672, 1673, 1674, 1675, 1676, 1677, 1678, 1679, 1714, 1649, 1650, 1651, 1652, 1653, 1654, 1718, 1760, 1761, 1762, 1763, 1764, 1765, 1766, 1778]. Mario [1478]. Markup [361, 596, 730, 911]. Marsaglia [771]. Master [216, 295]. Matching [189, 1111]. math [875]. MathCW [1789, 1400, 781]. Mathematica [917, 330, 820, 1131]. Mathematical [277, 442, 503, 504, 505, 506, 550, 551, 552, 508, 509, 511, 512, 513, 514, 515, 510, 542, 543, 588, 589, 590, 591, 604, 573, 625, 931, 930, 963, 984, 983, 985, 988, 1036, 1029, 1028, 1198, 1174, 1175, 1246, 1302, 1288, 1295, 1348, 1396, 1397, 1383, 1400, 1495, 1496, 1577, 1772, 1782, 1779, 1789]. Mathematical-Function [1400, 1789]. Mathematics [467, 468, 454, 566, 544, 545, 546, 547, 616, 641, 688, 693, 705, 707, 708, 709, 710, 657, 733, 769, 763, 873, 897, 901, 905, 964, 966, 970, 971, 969, 968, 967, 965, 1071, 1072, 1073, 1075, 1074, 1253, 1170, 1286, 1395, 1409, 1376, 1377, 1378, 1379, 1380, 1477, 1521, 1495, 1621, 1553, 1369, 1370, 1371, 1372, 1373, 1374, 1375, 1528, 1530, 608, 647, 961, 1495, 264, 1247, 1466, 1467, 1476, 1479, 1492, 1518, 1585, 1586, 507, 522, 558, 548, 748, 898, 1038, 1188, 1189, 1190, 1191, 1192, 1193, 1194, 1195, 1283, 1395, 1368, 1436, 1439, 1323]. Mathematik [463, 668, 667, 791, 896, 1554]. Mathématiques [1439, 604, 963, 964, 966, 970, 971, 969, 968, 967, 965]. Mathison [837]. Matlab [331, 817]. Matrices [40, 64, 62, 96, 106, 131]. Matrix [170, 265, 268, 296, 331, 458, 1173, 1252]. Matters [1329]. Maurice [952]. Max [986, 1311]. Maxima [821]. Maxwell [1451]. Measurement [1390]. mechanics [88]. Medicine [1496, 1355, 1590, 1591, 1592, 1593, 1594]. Medizin [1355]. Meeting [1469]. meets [874]. Meitner [1449]. Members [769]. Memory [328, 874]. Message [252, 382]. METAFONT [829]. Method [3, 52, 118, 172, 549]. Methodology [669, 1294]. Methods [1037, 1288, 1780, 1788, 1790, 1786, 523, 659, 1336, 1337, 1338, 1339, 1340, 1545]. Metrika [1287]. Metron [1596, 1597, 1598, 1599, 1600, 1601, 1602, 1603, 1604, 1605, 1606]. Metropolis [860]. Mexico [68, 99, 100, 115, 117]. u [271]. Michael [1361]. Micro [472]. Microprocessors [654]. Microsoft [384]. Milton [1331]. Minerva [1360]. Minimal [147]. minimization [698]. MINIX} [403]. MINPACK [698]. Miscellaneous [54]. Mobile [1389]. Modeling [442, 755, 757, 1392, 1779, 1668]. Modelling [1638, 1639, 1640, 1641, 1642, 1643]. Models [392, 1156]. Modern [990, 1042, 1043, 1047, 1050, 1051, 1046, 1045, 1048, 1049, 1044, 1787, 1788]. Modification [149]. MODNAME [72]. Mold [402]. Molecular [44, 47, 74, 80, 110, 182, 58, 156, 197]. Molecule [109, 82, 134, 157]. Molecules [1, 59, 108]. Moler [390]. MOLPRO [177]. Moments [111, 109]. Monographs [1365]. Monte [1037]. Monthly} [550, 551, 552, 588, 589, 590, 591, 1029, 1028, 503, 504, 505, 506]. Moore [854]. mostly [752, 1411]. Movies [816]. MPI [382]. Multi [250, 249]. Multi-Processor [250, 249]. Multilinear [1362]. multilingual [1547, 1549]. Multimedia [784, 656]. Multiple [818]. Multiple-precision [818]. Multiplet [65, 183]. Multiplication [265]. Multiply [268, 296]. Multipole [549]. Multiscale [757]. Multithreading [534]. MuPAD} [834, 822].

N [835]. N.A.Software [706]. NAG [699, 700, 701, 707, 708, 709, 710]. Name [12]. Naming [8]. NaN [879]. National [129]. NATO [311, 356]. Natural [913, 1668]. ndiff [610]. need [1363]. Nelson [1789, 332, 1332]. Netherlands [597, 627, 940]. Network [1247, 1282, 1424, 831, 832]. Networking [318, 353, 734, 894, 593]. Networks [598, 599, 597, 627, 940, 842]. Neumann [788]. Newcomb [998]. News} [1310, 853, 1039]. Newsletter} [1395, 488, 1631]. NHFB [295]. Nicholas [874, 860]. Niels [950]. NODELIST [48]. Nonorthogonal [64]. Nordic [554]. Nordisk [558]. Norman [1128]. North [1687, 1688, 1689, 1690, 1691]. Note [41, 199, 216, 150, 561]. Notes [56, 84, 68, 86, 78, 85, 70, 71, 81, 72, 73, 74, 75, 80, 69, 76, 77, 79, 82, 99, 100, 106, 115, 117, 131, 127, 179, 192, 206, 245, 314, 405, 410, 407, 411, 412, 409, 408, 447, 448, 396, 491, 521, 531, 541, 540, 595, 624, 700, 701, 718, 715, 713, 716, 714, 742, 740, 741, 739, 749, 750, 738, 782, 852, 948, 947, 943, 944, 945, 946, 1041, 1094, 1095, 1096, 1097, 1098, 1099, 1100, 1101, 1102, 1103, 1104, 1105, 1052, 1168, 1235, 1188, 1189, 1190, 1191, 1192, 1193, 1194, 1195, 1440, 1466, 1467, 1476, 1479, 1492, 1518, 1585]. Notes [1586, 1787, 786, 1452, 1453, 1454, 1455, 1456]. Notices [364, 584, 585, 586, 587, 615, 891]. NRCC [204]. NTM [1355]. NTS [321]. Nuclear [59, 108]. Number [751, 1000, 1614, 1607, 1608, 1609, 1610, 1611, 1612, 1613]. Numbers [21, 32, 49, 738, 752, 998, 1411]. numeric [610]. Numerica [974]. Numerical [314, 467, 481, 457, 460, 558, 602, 607, 609, 646, 678, 677, 799, 902, 906, 949, 941, 1464, 1475, 1790, 691, 696, 698, 817, 1786]. Numerische [463, 668, 667, 791, 896, 1554]. Nuncius [1130].

O [57, 6, 66, 134]. Object [73, 116]. Oceanography [1750, 1686, 1757, 1758, 1759, 1749, 1751, 1752, 1753, 1754, 1755, 1756]. Octave [817]. Office [574, 1681, 1682, 1683, 1684, 1685]. Official [1483]. Offline [16]. Olga [1438, 1438]. Olov [640]. On-Line [331]. One [1247]. Online} [744]. only [846]. Open [82, 386, 385, 1459, 1460]. operands [611]. Operating [403, 404, 416, 435, 679, 855]. Operations [1483]. Operator [42]. Operators [93]. Oppenheimer [972]. Opportunities [1246]. Optimal [1782]. Optimization [1637, 1780, 567, 759, 772, 767, 1172, 1475]. Oracle [533]. Orbital [42, 63]. Orbitals [38, 44, 92, 149]. Order [580, 794, 108, 109, 711]. O’Reilly [366]. Oriental [722]. Origin [290, 291]. ORSA [1443]. Orthodontia [373]. Osiris [1301, 1301]. Other [542, 543, 1107, 1784]. Otto [1447]. Outline [143, 151, 207]. Output [13, 143, 151]. Oxford [1615].

pêcheries [1681, 1682, 1683, 1684, 1685]. Pack [11, 26, 43]. Package [23, 47, 70, 81, 177, 195, 208, 260, 322, 348, 520, 703, 864]. Packard [417, 49]. PACMCGIT [1628]. PACMHCI [1629]. PACMPL [1425]. Page [13]. PageRank [893]. Panel [686, 687]. Paper [135, 275]. Papers [53, 437, 405, 410, 407, 411, 412, 409, 408, 447, 448, 491, 521, 531, 541, 540, 595, 718, 715, 713, 716, 714, 742, 740, 741, 739, 749, 750, 782, 852, 948, 947, 943, 944, 945, 946, 1041, 1094, 1095, 1096, 1097, 1098, 1099, 1100, 1101, 1102, 1103, 1104, 1105, 1168, 1235, 1285]. Parallel [382, 473, 474, 478, 475, 477, 476, 480, 762, 795, 871, 872, 904, 1251, 1740, 1741, 1742]. Parameters [1, 91, 212]. Pareto [998]. PARI [824]. PARI/GP [824]. Parlett [835]. Parsing [208, 261]. Part [388, 389, 409, 408, 447, 448, 496, 491, 521, 531, 541, 540, 595, 715, 713, 716, 714, 742, 740, 741, 739, 749, 750, 782, 852, 992, 991, 1749]. Partition [45]. Partitioned [10, 33]. Passing [382]. Past [769]. Paths [255]. Pattern [189, 1111]. Paul [952]. Pauli [1057]. PC [232]. pdflatex [1547]. pdfTEX [686]. Peierls [1445]. Per-Olov [640]. Perception [841]. Performance [246, 265, 328, 406, 380, 641, 688, 869, 1392]. Perils [1570]. Perspective [1132]. Perspectives [1134, 1333]. Perturbation [80]. Pete [1196]. Peter [1473]. PFORT [198]. Phillips [942]. Philosophical [1522, 1523, 1524, 1525, 1526]. Philosophy [916, 990, 992, 991, 1158, 1159, 1160, 1161, 1162, 1163, 1164, 1165, 1166, 1304, 1327, 1326, 1334, 1468, 1461, 1462, 1450]. Physical [913, 1333, 1391]. Physics [508, 509, 511, 512, 513, 514, 515, 510, 625, 1030, 1031, 1033, 1034, 1035, 1032, 1132, 1388, 1539, 990, 984, 983, 985, 988, 1295, 1577, 1001, 1002, 1003, 1004, 1005, 1006, 1007, 1008, 1009, 1010, 1011, 1042, 1043, 1047, 1050, 1051, 1046, 1045, 1048, 1049, 1044, 1129, 1237, 1238, 1239, 1240, 1241, 1242, 1333, 1298, 1349, 1538, 1576]. Physiology [442, 1773, 1774, 1775, 1776, 1777, 1779]. Physis [1308]. Picture [322]. PIP [1132]. Pitfalls [1570]. Plan [416]. Planck [1311]. Planning [1512, 1513, 1514, 1515, 1516, 1517]. Play [1782]. Plot [47, 68, 74, 99, 100, 115, 117, 776]. PLOT79 [216]. Plots [50, 162]. Plotter [23]. Plotting [44, 156, 197]. Plus [334, 825]. PLUTO [156, 197]. Poincaré [1210]. Point [228, 305, 310, 1622, 346, 690, 866, 867]. Polarizabilities [113, 112, 59]. Polarizability [58, 105]. Polarization [75, 160, 155, 105, 112, 108, 109]. Policy [213, 1089, 1090, 1091, 1092, 1093, 1360, 1573, 1649, 1650, 1651, 1652, 1653, 1654]. POLYATOM [177]. POMACS [1390]. PORT [697]. Portability [133]. Portable [136, 158, 159, 200, 201, 208, 1400, 1789, 253]. Portland [702]. Possible [29]. PostScript [776]. Postulates [88]. Potential [87]. Powerful [240]. powłoki [858]. pp [1789]. Practical [62, 304]. Practice [339, 594, 684, 1738, 1739, 904, 1062, 596, 1427, 1428, 1429, 1579]. Practices [1570, 762, 904]. precision [818]. predecessors [955]. Preface [179]. Preliminary [1128]. Preparation [34]. prepared [274, 1247]. Preprocessor [241]. Present [769]. President [252, 254, 262, 258, 266, 276, 297, 313]. PRETTY [166, 159, 186, 201, 214]. Prettyprinter [159, 188, 201, 242]. prettyprinting [326, 327]. Primes [583]. Primitives [141, 142, 144, 145, 184, 227]. Princeton [1323]. Principles [444, 628].

PRINT [102]. Printer [9]. Printing [823, 1408]. Privacy [1386, 753]. Probabilistic [1783]. Probability [565, 870, 975, 1199, 1200, 1201, 1202, 1210, 1184, 1185, 1297, 1307, 1407, 1734, 1735, 1736, 1737, 1244, 1245, 1236, 1186, 1187, 1208, 1178, 1179, 1180, 1181, 1182, 1306]. Problem [61, 96, 106, 123, 120, 131, 243]. Problems [29]. Procedure [76]. Procedures [19, 121, 18]. Proceedings [563, 564, 951, 1389, 1390, 1425, 1522, 1523, 1524, 1525, 1526, 1628, 1629, 1788, 1300, 311, 356, 1785]. Process [1427, 1430]. Processing [340, 392, 476, 555, 556, 557, 621, 722, 773, 938, 392, 746, 1325, 1384]. Processor [215, 250, 249]. processors [810]. Professional [373, 634, 635, 633, 955]. Program [12, 2, 5, 4, 49, 60, 84, 67, 83, 78, 85, 71, 73, 74, 75, 80, 69, 77, 79, 98, 127, 158, 159, 157, 201, 156, 197, 160, 155]. Programmer [136, 373, 634, 635, 633]. Programmers [152, 153, 205]. Programmierung [857]. Programming [1789, 104, 287, 324, 354, 334, 368, 369, 388, 389, 423, 453, 470, 482, 496, 516, 532, 617, 614, 850, 910, 1107, 1400, 1425, 1569, 377, 768, 1527, 1627, 478, 600, 731, 795, 1152, 1398, 1157]. Programowanie [858]. Programs [14, 22, 27, 71, 103, 125, 173, 175, 176]. Progress [1759, 1760, 1761, 1762, 1763, 1764, 1765]. Project [46, 536]. Projected [63]. Projection [42, 93]. Propagator [75, 160, 155, 105, 112, 108, 109]. Properties [70, 81, 182, 177]. Proposal [141, 142, 144, 145, 184, 227, 275]. Proposed [1157]. Pseudo [752, 1411, 738]. Pseudo-random [752, 1411, 738]. Pseudorandom [751, 1000]. Public [237, 1089, 1090, 1091, 1092, 1093, 1573]. Public-domain [237]. Publications [1448, 294, 273, 274, 309, 329, 334, 335, 336, 337, 338, 348, 342, 349, 333, 340, 344, 343, 330, 331, 332, 339, 357, 347, 368, 369, 370, 372, 361, 362, 371, 360, 364, 365, 382, 387, 388, 389, 391, 394, 399, 402, 403, 404, 406, 415, 416, 423, 432, 435, 376, 383, 390, 397, 418, 419, 420, 421, 422, 434, 373, 374, 414, 441, 417, 375, 378, 380, 381, 413, 392, 393, 433, 427, 384, 386, 379, 424, 425, 426, 385, 398, 400, 401, 430, 428, 429, 395, 461, 453, 467, 468, 470, 482, 486, 492, 496, 503, 504, 505, 506, 449]. Publications [450, 452, 451, 465, 466, 472, 473, 474, 479, 478, 475, 487, 485, 489, 490, 459, 462, 464, 454, 488, 463, 481, 477, 476, 480, 455, 456, 457, 458, 469, 484, 483, 507, 498, 499, 500, 502, 501, 493, 494, 460, 516, 532, 549, 550, 551, 552, 560, 562, 555, 556, 557, 553, 508, 509, 511, 512, 513, 514, 515, 510, 535, 566, 568, 534, 522, 523, 567, 565, 539, 554, 558, 548, 559, 542, 544, 545, 546, 547, 543, 602, 584, 585, 586, 587, 588, 589, 590, 591, 604, 578, 598, 599, 597, 601, 600, 579, 580, 571, 593, 581]. Publications [573, 570, 577, 592, 594, 596, 607, 614, 615, 643, 605, 623, 634, 635, 633, 613, 626, 629, 630, 632, 625, 640, 609, 637, 638, 639, 636, 642, 645, 644, 631, 606, 608, 616, 619, 620, 621, 627, 646, 647, 648, 683, 681, 652, 665, 668, 667, 653, 651, 655, 656, 659, 662, 663, 661, 666, 664, 678, 677, 680, 682, 684, 685, 1738, 1739, 654, 729, 730, 727, 744, 743, 719, 721, 731, 733, 735, 745, 747, 753, 756, 757, 758, 759, 762, 771, 763, 764, 765, 766, 767, 770, 777, 785, 790, 834, 836, 837, 838, 787, 788, 791]. Publications [797, 796, 794, 795, 799, 798, 793, 800, 833, 835, 850, 851, 853, 855, 854, 849, 862, 860, 861, 865, 871, 872, 873, 877, 886, 902, 890, 891, 893, 910, 911, 888, 889, 892, 900, 895, 931, 896, 897, 898, 899, 901, 903, 904, 905, 906, 908, 914, 912, 913, 915, 916, 925, 930, 986, 995, 954, 953, 1001, 1002, 1003, 1004, 1005, 949, 941, 936, 937, 938, 990, 992, 991, 996, 1006, 939, 940, 951, 961, 963, 964, 966, 970, 971, 969, 968, 967, 965, 974, 977, 979, 980, 982, 981, 978, 984, 983, 985, 987, 988, 942, 950, 952]. Publications [972, 973, 976, 1039, 1112, 1065, 1089, 1090, 1091, 1092, 1093, 1085, 1007, 1078, 1087, 1088, 1071, 1072, 1073, 1075, 1074, 1081, 1083, 1086, 1084, 1082, 1036, 1008, 1009, 1010, 1011, 1029, 1028, 1030, 1031, 1033, 1034, 1035, 1032, 1037, 1038, 1040, 1063, 1067, 1068, 1069, 1070, 1066, 1076, 1077, 1079, 1080, 1053, 1054, 1055, 1056, 1057, 1058, 1059, 1060, 1061, 1107, 1136, 1154, 1106, 1126, 1127, 1113, 1144, 1145, 1146, 1147, 1148, 1152, 1153, 1155, 1156, 1114, 1115, 1116, 1117, 1118, 1119, 1120, 1121, 1122, 1123, 1124, 1125, 1129, 1130, 1131, 1132, 1134, 1135, 1137, 1138, 1139, 1140, 1141]. Publications [1142, 1143, 1128, 1133, 1150, 1233, 1196, 1186, 1187, 1197, 1198, 1199, 1200, 1201, 1202, 1203, 1204, 1205, 1206, 1207, 1208, 1209, 1210, 1211, 1212, 1213, 1214, 1215, 1216, 1217, 1218, 1219, 1220, 1221, 1222, 1223, 1224, 1225, 1226, 1227, 1228, 1229, 1230, 1231, 1234, 1250, 1253, 1294, 1169, 1170, 1171, 1172, 1173, 1174, 1175, 1176, 1177, 1178, 1179, 1180, 1181, 1182, 1183, 1184, 1185, 1296, 1308, 1309, 1254, 1255, 1315, 1316, 1317, 1318, 1319, 1320, 1321, 1322, 1333, 1282, 1283, 1284, 1285, 1286, 1287, 1288, 1289, 1290, 1291, 1292, 1293, 1295, 1297, 1298, 1300, 1301, 1305, 1306]. Publications [1307, 1310, 1312, 1313, 1314, 1299, 1311, 1331, 1332, 1351, 1334, 1360, 1361, 1362, 1364, 1365, 1366, 1367, 1335, 1336, 1337, 1338, 1339, 1340, 1341, 1342, 1343, 1344, 1345, 1346, 1347, 1348, 1349, 1350, 1352, 1353, 1354, 1355, 1356, 1357, 1358, 1359, 1423, 1395, 1368, 1396, 1397, 1398, 1399, 1401, 1402, 1403, 1404, 1406, 1407, 1409, 1410, 1412, 1413, 1414, 1415, 1416, 1417, 1418, 1419, 1420, 1422, 1424, 1369, 1370, 1371, 1372, 1373, 1374, 1375, 1376, 1377, 1378, 1379, 1380, 1381, 1382, 1383, 1384, 1386, 1387, 1388, 1389, 1390, 1391, 1392, 1393, 1394, 1405, 1437, 1438, 1441, 1442, 1457]. Publications [1472, 1468, 1444, 1461, 1462, 1463, 1464, 1470, 1471, 1426, 1427, 1428, 1429, 1430, 1431, 1432, 1433, 1434, 1435, 1436, 1439, 1440, 1443, 1446, 1450, 1451, 1452, 1453, 1454, 1455, 1456, 1458, 1459, 1460, 1465, 1445, 1447, 1449, 1473, 1527, 1541, 1474, 1521, 1522, 1523, 1524, 1525, 1526, 1528, 1530, 1536, 1538, 1539, 1540, 1542, 1545, 1475, 1495, 1496, 1497, 1498, 1499, 1500, 1501, 1502, 1503, 1504, 1505, 1506, 1507, 1508, 1509, 1510, 1512, 1513, 1514, 1515, 1516, 1517, 1519, 1520, 1478, 1493, 1494, 1552, 1569, 1557, 1620, 1617, 1618, 1565, 1573, 1595, 1548, 1580, 1615, 1621, 1550, 1553, 1554]. Publications [1555, 1556, 1558, 1559, 1560, 1561, 1562, 1563, 1564, 1566, 1567, 1571, 1572, 1575, 1577, 1578, 1627, 1715, 1623, 1624, 1655, 1656, 1657, 1658, 1659, 1660, 1661, 1662, 1663, 1664, 1665, 1666, 1667, 1668, 1669, 1670, 1671, 1672, 1673, 1674, 1675, 1676, 1677, 1678, 1679, 1680, 1681, 1682, 1683, 1684, 1685, 1686, 1687, 1688, 1689, 1690, 1691, 1692, 1693, 1694, 1695, 1696, 1697, 1698, 1699, 1700, 1701, 1702, 1703, 1704, 1705, 1706, 1707, 1708, 1709, 1710, 1711, 1712, 1713, 1714, 1630, 1633, 1634, 1635, 1636, 1638, 1639, 1640, 1641, 1642, 1643, 1644, 1645, 1646, 1647, 1648, 1649, 1650]. Publications [1651, 1652, 1653, 1654, 1632, 1717, 1746, 1748, 1740, 1743, 1716, 1734, 1735, 1736, 1737, 1718, 1719, 1757, 1758, 1759, 1760, 1761, 1762, 1763, 1764, 1765, 1766, 1767, 1768, 1769, 1770, 1771, 1772, 1773, 1774, 1775, 1776, 1777, 1778, 1720, 1721, 1722, 1723, 1724, 1725, 1726, 1727, 1728, 1729, 1730, 1741, 1742, 1744, 1745, 1747, 1749, 1750, 1751, 1752, 1753, 1754, 1755, 1756, 377, 359, 346, 1324, 301, 283]. Publishers [366]. Publishing [290, 282, 291, 1633]. Punch [9, 17, 102]. Purchase [246]. putatively [610]. PVM [382]. Python [532].

QCPE [22]. QTP [43]. QTP999 [11]. QTPCOPY [17]. QUADLOG [864].

Quadrature [314, 863, 864]. Quality [887, 776]. Quantification [1171]. Quantum [120, 148, 174, 180, 179, 375, 613, 648, 977, 979, 980, 982, 981, 978, 1329, 1542, 1625, 1787, 88, 46]. Quarterly [1151, 987]. Quartet [107]. Query [728]. Queue [756].

R [865, 1310, 825, 1544]. Radical [53].

Ramon [854]. Random

[54, 143, 151, 738, 752, 998, 1411].

Random-Access [143, 151]. RDUMP [14]. Reader [9]. Reading [264]. Reasoning [1356]. recherches [1681, 1682, 1683, 1684, 1685]. Reconfigurable [878]. Record [167, 446]. Records [1440]. REDUCE [281, 395, 703, 826]. Reduced [40, 64, 62]. Rees [87]. Reference [152, 153, 205, 116]. Referencing [73]. Register [14]. Regular [1111]. Related [1184, 1185, 863, 864]. relational [874]. Release [124, 125, 165, 178, 185, 187, 193, 196, 202, 225, 245, 186, 230, 229, 231, 235]. Reliable [1106]. Remarks [128, 139, 177]. Replacements [250, 249]. Report [433, 1246, 1247, 1786, 434, 1726]. Reports [1148]. Request [46]. Research [344, 1428, 1429, 1459, 1579, 1681, 1682, 1683, 1684, 1685, 1721, 1749, 253, 1483, 1644, 1645, 1646, 1647, 1648, 1718, 1624]. Resource [129, 1384, 1668, 1766, 304]. Resources [766, 1511]. retrospective [846, 845]. Review [89, 1360, 1621, 379, 455, 855, 869, 1151, 1212, 1169, 1566]. Reviewed [1789]. Reviews [1042, 1043, 1047, 1050, 1051, 1046, 1045, 1048, 1049, 1044, 1680, 1719]. Revised [23]. Revision [47, 48]. Revue [1309, 1305, 1439]. REXX [790]. RFC [440]. Rice [865]. Richard [1748, 942, 1112, 1457, 1725]. Rivista [1308]. RJE [37]. Robert [972, 1447]. Robot [1582]. Ronald [1595]. Roos [80]. Root [269]. Roster [140, 204]. routines [864]. Royal [1256, 1257, 1258, 1259, 1260, 1261, 1262, 1263, 1264, 1265, 1266, 1267, 1268, 1269, 1270, 1271, 1272, 1273, 1274, 1275, 1276, 1277, 1278, 1279, 1280, 1281, 1440, 1546]. RSDUMP [14]. Rudolf [1445]. Ruhe [1324]. Rules [112]. Rust [1544, 1627]. Rutherford [1332, 1334]. Rutishauser [1618]. Rydberg [87].

S [39, 171, 334, 825]. S-Plus [334, 825]. S/370 [39]. sample [1385]. Sampler [1328]. Sampling [751, 1000]. Samuel [1493]. SAS [335, 828]. Saving [37]. Scalable [904, 1062]. SCAN [69, 127]. Scandinavian [1619]. SCF [85]. Schrödinger [1054]. Science [233, 578, 600, 571, 570, 624, 665, 666, 664, 839, 992, 991, 1089, 1090, 1091, 1092, 1093, 1040, 1136, 1148, 1152, 1153, 1135, 1327, 1320, 1321, 1334, 1360, 1365, 1355, 1409, 1394, 1470, 1450, 1465, 1573, 1584, 1719, 1784, 1788, 498, 499, 500, 502, 501, 631, 936, 937, 1158, 1159, 1160, 1161, 1162, 1163, 1164, 1165, 1166, 1254, 1468, 1541, 405, 410, 407, 411, 412, 409, 408, 916, 447, 448, 396, 489, 491, 521, 531, 541, 540, 564, 595, 577, 725, 718, 715, 713, 716, 714, 742, 740, 741, 739, 749, 750, 782, 798, 852, 901, 914, 912, 915, 948, 947]. Science [943, 944, 945, 946, 1041, 1094, 1095, 1096, 1097, 1098, 1099, 1100, 1101, 1102, 1103, 1104, 1105, 1052, 1063, 1076, 1134, 1168, 1235, 1197, 1303, 1304, 1319, 1326, 1468, 1461, 1462, 1463, 1471, 1521, 1715, 1669, 1670, 1671, 1672, 1673, 1674, 1675, 1676, 1677, 1678, 1679, 1714, 1630]. Sciences [348, 1036, 1309, 1655, 1656, 1657, 1658, 1659, 913, 479, 899, 913, 933, 1309, 1300, 1407, 1458, 1491, 1474]. Scientiarum [1465]. Scientific [424, 400, 484, 483, 637, 638, 639, 636, 692, 731, 1088, 1126, 1127, 1113, 1114, 1115, 1116, 1117, 1118, 1119, 1120, 1121, 1122, 1123, 1124, 1125, 1176, 1177, 253]. Scientists [1108, 1108]. Scientometrics [1322, 1315, 1316, 1312, 1313, 1314, 1555].

Scienza [1308]. Scott [1472]. Scratchpad [372]. Script [879]. Scripting [532, 790, 847, 857, 789]. scripts [848]. Sea [1749]. Search [69, 127]. Second [203, 201, 1252, 108, 109]. Security [31, 753, 1386, 1726, 1727, 592, 1357, 1424, 1426, 1531, 1532, 1533, 1534, 1535]. Selected [76, 217, 223, 654, 777, 942, 950, 952, 972, 973, 976, 1053, 1054, 1055, 1056, 1057, 1058, 1059, 1060, 1061, 1133, 1150, 1299, 1311, 1331, 1332, 1405, 1445, 1447, 1449, 1478, 1493, 1494]. Selection [751, 1000]. Selective [122]. Self [59]. Self-Consistent [59]. Sensor [842]. September [1787, 1788]. Sequence [32]. Sequences [21]. Series [566, 1170, 1256, 1257, 1258, 1259, 1260, 1261, 1262, 1263, 1264, 1265, 1266, 1267, 1268, 1269, 1270, 1271, 1272, 1273, 1274, 1275, 1276, 1277, 1278, 1279, 1280, 1281, 1301, 1480, 1487, 1546, 1495, 1468, 1760, 1761, 1762, 1763, 1764, 1765]. Serif [367]. Set [8, 31, 122, 141, 142, 144, 145, 184, 185, 187, 193, 196, 202, 211, 224, 227, 259]. Sets [4, 10, 33, 26, 54, 174, 180, 653]. SFTRAN [138, 137, 139, 219]. SFTRAN/3 [138, 139, 219]. SFTRAN3 [154, 160, 152, 188, 242, 241, 153, 205]. SGML [361, 730, 911]. Shōkai [859]. Shannon [925]. Shell [82, 789, 847, 857, 848, 879]. Shell-Programmierung [857]. Shells [97]. shieru [859]. Shift [190]. Short [218]. SIAM [455, 456, 457, 458, 484, 483, 507, 566, 522, 523, 567, 573, 609, 608, 659, 678, 677, 758, 759, 931, 898, 899, 930, 941, 961, 1169, 1170, 1171, 1172, 1173, 1174, 1175, 1176, 1177, 1420, 1469, 1521]. SIAM/ASA [1171]. Siegbahn [80]. SIGACT [444, 628, 1039]. SIGACT-SIGMOD [444, 628]. SIGAda [338]. SIGAPP [1566]. SIGARCH [853]. SIGCSE [1065, 1067, 1068, 1069, 1070, 1066, 1564]. SIGGRAPH [391, 729, 890]. SIGMETRICS [869]. SIGMICRO [1631].

SIGMOD [444, 446, 628]. Significance [1615]. SIGNUM [336]. SIGOPS [855]. SIGPLAN [364, 584, 585, 586, 587, 615, 891]. SIGSAM [337]. SIGSOFT [1452, 1453, 1454, 1455, 1456]. similar [610]. Simple [322]. Simplifications [110]. Simulation [1341, 1342, 1343, 1344, 1345, 1580, 755, 757, 1183, 1540]. Sinc [1790]. Single [121]. Sir [1445, 1595, 1748]. Sizes [45]. skryptów [858]. Slater [170, 1056]. Slime [402]. small [610]. Smithsonian} [425, 642, 645, 644]. Social [348, 1471, 1543]. Society [566, 548, 1170, 1304, 1256, 1257, 1258, 1259, 1260, 1261, 1262, 1263, 1264, 1265, 1266, 1267, 1268, 1269, 1270, 1271, 1272, 1273, 1274, 1275, 1276, 1277, 1278, 1279, 1280, 1281, 1319, 1320, 1365, 1395, 1440, 1446, 1465, 1546, 1495, 1483, 1522, 1523, 1524, 1525, 1526, 1692, 1693, 1694, 1695, 1696, 1697, 1698, 1699, 1700, 1701, 1702, 1703, 1704, 1705, 1706, 1707]. Software [133, 136, 140, 148, 135, 189, 204, 222, 244, 309, 339, 357, 373, 634, 635, 633, 669, 699, 700, 701, 705, 774, 827, 1366, 1400, 1427, 1428, 1429, 1430, 1431, 1432, 1433, 1434, 1435, 1452, 1453, 1454, 1455, 1456, 1550, 1630, 1785, 1789, 690, 1634, 1635, 1636, 277, 381, 568, 747, 908, 1464, 1459, 1460, 1743, 1744, 1745]. solution [696]. solutions [691, 698]. Some [27, 126]. Sommerfeld [1410, 1405]. Sorting [126]. Source [433, 1460]. Space [41, 89, 1458]. Spaces [89]. Spanish [1483]. Spanning [41]. Sparse [96, 106, 131]. Spatial [1393]. Specification [147, 275]. Specifications [90, 248]. Spectral [1, 36, 58]. Spectroscopic [91, 101]. Spectrum} [1556, 1558, 1559, 1560, 1561, 1562, 1563]. Speech [773, 1325]. Speed [724]. Spell [997]. Spelling [360]. spheres [192]. Spin [38, 59, 108, 63]. spin-orbital [63]. Spin-Spin [59, 108]. Spline [199, 86, 77].

Spontaneous [1461]. Spring

[1476, 1479, 1492]. Springer [1789, 1394]. SPSS [348]. SQL [728]. Square [269]. squares [698]. Standard [141, 142, 144, 145, 184, 227, 300, 361, 730, 911]. Standardized [8]. Standards [146, 135, 204, 309, 357, 618, 660, 1785, 470, 482, 617]. Stanisław [861]. State [52, 243]. Statement [213]. Statements [261]. States [66, 107]. Static [105]. Statistical [335, 348, 487, 484, 747, 994, 993, 1012, 1015, 1013, 1014, 1016, 1017, 1018, 1019, 1237, 1238, 1239, 1240, 1241, 1242, 1197, 1212, 1294, 1177, 1183, 1256, 1257, 1258, 1259, 1260, 1261, 1262, 1263, 1264, 1265, 1266, 1267, 1268, 1269, 1270, 1271, 1272, 1273, 1274, 1275, 1276, 1277, 1278, 1279, 1280, 1281, 1283, 1285, 1289, 1491, 1546, 1540, 1512, 1513, 1514, 1515, 1516, 1517, 1576]. Statistician [1256, 1257, 1258, 1259, 1260, 1261, 486, 1021, 1022, 1023, 1024, 1027, 1026, 1025, 1020, 1574]. Statistics [334, 962, 1199, 1200, 1201, 1202, 1284, 1290, 1291, 1292, 1293, 1305, 1306, 1336, 1337, 1338, 1339, 1340, 1341, 1342, 1343, 1344, 1345, 1536, 1545, 1483, 1580, 1590, 1591, 1592, 1593, 1594, 1619, 825, 828, 1203, 1204, 1205, 1206, 1207, 316, 352, 485, 569, 671, 672, 673, 674, 675, 670, 909, 962, 1198, 1209, 1210, 1250, 1286, 1288, 1297, 1335, 1359, 1406, 1490, 1588, 1587, 1616, 1621, 1589]. statistique [1305]. Status [247]. STD [439]. Stenger [1790, 862, 1632]. Step [121]. Steven [1725]. Stewart [1196]. Stiefel [1620]. STOC [662, 663, 661, 719, 721, 903]. Storage [783]. STORE [102]. Storia [1308]. String [1111]. Structural [1780]. Structure [57]. Structured [728]. Structures [156, 197]. STTT [774]. Student [1469]. Studies [442, 913, 990, 992, 991, 1368, 1468, 1462, 1463, 1471, 1450, 1749, 1782, 1779, 1358, 1470]. Study [57, 311, 356, 879]. Styles [1466, 1467, 1476, 1479, 1492, 1518, 1585, 1586, 1385].

Subroutine [128]. Sudhoffs [1404]. Suggested [146, 147, 148]. sukuriputo [859]. Sum [112]. Summary [51, 157, 272]. Summer [1518, 1585, 1586]. Sun [704, 255]. Supercomputer [380, 426]. Supercomputing [432, 379, 469, 1422, 497, 528, 527, 530, 529, 526, 524, 525, 518, 517, 537, 538, 612, 658, 720, 760, 761, 843, 844, 999, 1064, 427, 797]. Supplement [94]. Supplemental [395]. Support [279, 278, 310, 323, 1421]. Surfaces [199]. Survey [210, 218, 244]. Surveys [363, 1307]. Sutherland [206]. Symbolic [329, 330, 372, 394, 437, 393, 395, 580, 581, 834, 794, 793, 703, 819, 820, 821, 822, 824, 826]. Symmetry [97, 94]. Symposia [444, 437, 563, 564, 628, 662, 663, 661, 719, 721, 903, 665, 666, 664, 1040, 649, 650]. syntax [326, 327]. Synthese [1412, 1413, 1414, 1415, 1416, 1417, 1418, 1419]. SysAdmin [428, 429]. System [157, 200, 251, 245, 348, 403, 404, 416, 435, 441, 400, 592, 918, 920, 919, 921, 923, 924, 922, 1233, 1474, 253, 68, 99, 100, 115, 117, 226, 304, 335, 351, 431, 533]. Systems [38, 289, 343, 423, 384, 386, 401, 428, 429, 473, 568, 578, 617, 679, 744, 754, 833, 855, 871, 881, 926, 908, 935, 989, 1153, 1423, 1390, 1391, 1392, 1393, 1634, 1635, 1636, 377, 63, 444, 445, 474, 599, 572, 574, 576, 628, 685, 727, 758, 800, 878, 1740, 1728, 1729, 1730, 1741, 1742]. Szilard [973].

TAAS [881]. Table [171, 582]. Tables [542, 543]. talks [1448]. Tangent [270]. Tape [90, 130, 167, 216, 220, 229, 221, 225, 231, 235, 222]. tar [220, 229, 231]. Tasks [857]. Taussky [1438]. Taussky-Todd [1438]. TCPS [1391]. TDS [1584]. TEAC [1167]. Technical [374, 488, 928, 918, 920, 919, 921, 923, 924, 922, 929, 927, 1626, 1726]. Technik [1355]. Technique [42]. Techniques [1628]. Technologies [697, 833, 1389]. Technology [431, 473, 680, 774, 878, 926, 1334, 1355, 712, 1320, 1321, 1350, 1450]. Technometrics [1225, 1226, 1227, 1228, 1229, 1230, 1231, 1578]. Technovation [1367]. TECO [232]. TECS [754]. Ted [874]. Teller [976]. TELO [1637]. Temporary [216]. TEST [1483, 690]. Testing [1000]. TeX [829]. TEX-Package-Independent [322]. TEXniques} [301, 283]. TEXnische [317, 358]. Text [67, 83, 158, 215, 1547, 1549]. TEXu [271]. Themes [1364]. Theorem [168, 169]. Theoretical [53, 498, 499, 500, 502, 501, 577, 631, 936, 937, 1236, 1254, 1541, 101]. Theoria [1402]. Theory [46, 80, 565, 563, 596, 662, 663, 661, 719, 721, 882, 903, 1184, 1185, 1329, 1336, 1337, 1338, 1339, 1340, 1545, 1619, 1734, 1735, 1736, 1737, 1781, 1784, 1614, 1607, 1608, 1609, 1610, 1611, 1612, 1613]. Theory/Configuration [80]. Theses [1466, 1467, 1476, 1479, 1492, 1518, 1585, 1586, 1149]. Things [55, 1581]. Thoughts [54]. Three [38, 36, 107, 63]. Three-Electron [38, 36, 107, 63]. THRI [1582]. Tidskrift [558]. TIIS [989]. Time [1480, 1487, 59, 752, 1411]. Time-Dependent [59]. Timing [162]. TIOT [1581]. Tips [298]. TIST [926]. TKDD [883]. TMACRO [215]. TMIS [935]. TOCE [885]. ToCT [882]. Today [386, 756]. Todd [1438]. Tomorrow [756]. TOMPECS [1392]. Toolbox [779]. Tools [140, 204, 222, 373, 471, 634, 635, 633, 774, 907, 907]. TOPC [1251]. Topical [1749]. Topics [146, 147, 148, 679]. TOPLAS [377]. TOPS [1386, 238]. TOPS-20 [238]. TQC [1625]. Transactions [277, 345, 445, 467, 474, 593, 572, 574, 575, 576, 592, 622, 669, 651, 685, 712, 746, 754, 755, 772, 773, 775, 767, 784, 783, 840, 841, 842, 876, 878, 881, 882, 883, 884, 885, 926, 906, 955, 959, 935, 989, 957, 956, 958, 960, 1167, 1251, 1325, 1384, 1386, 1391, 1392, 1393, 1431, 1432, 1433, 1434, 1435, 1543, 1581, 1583, 1584, 1551, 1582, 1550, 1625, 1637, 1623, 1692, 1693, 1694, 1695, 1696, 1697, 1698, 1699, 1700, 1701, 1702, 1703, 1704, 1705, 1706, 1707, 1740, 1741, 1742, 377]. Transfer [774]. Transformation [80]. Transformations [119]. Transition [52, 111, 109]. Translation [286]. transportable [114, 124]. Tree [280]. TRT2629 [17]. TSAS [1393]. TSC [1543]. TSLP [773]. TUG [254, 780, 778]. Tukey [953]. Turing [836, 837]. Turnaround [35]. Tutorial [1110]. TWEB [884]. Two [110, 1477, 863, 864]. Two-Electron [110]. Two-Year [1477]. Type [367]. Typesetting [415, 1149, 1547, 1549]. Typographic [399, 1624]. typography [273, 367].

Ubiquitous [1389]. Uhlenbeck [1494]. UIBMOL [82]. UK [1615]. Ulam [861]. Uncertainty [1171]. Understanding [392, 1155, 392]. Unicode [365]. Uniform [240]. Universal [489, 798]. University [1466, 1467, 1476, 1479, 1492, 1518, 1585, 1586, 1785, 233, 641, 688, 693, 705, 707, 708, 709, 710, 657, 769, 1247, 1469]. UNIX [428, 429, 857, 229, 231, 238, 256, 324, 354, 220, 441]. Unix/Linux [857]. Unix/Linux-Tasks [857]. UnixWorld [385]. Unrestricted [79]. Upgrade [255]. Upkeep [10]. Usage [24, 30, 808]. Use [293]. USENIX [398, 493, 494, 743, 1085]. User [163, 194, 257, 495, 701, 495]. Users [238, 430, 735]. Using [233, 312, 689, 707, 708, 709, 710, 880, 1400, 1789, 284]. UT [1789]. Utah [233, 641, 688, 693, 705, 707, 708, 709, 710, 657, 769, 1247, 1469, 1785, 90, 179, 160, 155, 222, 1466, 1467, 1476, 1479, 1492, 1518, 1585, 1586, 1787]. utilities [17]. Utility [12, 4, 27, 102, 125, 191].

v5.5 [694]. Varga [1725]. Variation [42]. Variational [1780]. Various [55]. VAX [221, 225, 234, 216, 235]. VAXIMA} [394]. verbatim [1547]. Verdet [113, 112]. Verifier [198]. Version [82, 173, 166, 177, 190, 192, 214, 219, 220, 221, 223, 245, 176, 205, 156, 323]. verwandte [1185]. Very [443, 619]. Vibrational [58]. Viewing [212]. Virial [169]. Virtual [149, 382, 792, 801, 851]. Visible [1624]. Vision [392, 1155]. Visual [653]. Visualization [400, 490, 651, 1623]. VLDB [619, 951]. VMS [216, 221, 225, 234, 225, 235]. Volumes [563, 564, 1094, 1095, 1096, 1097, 1098, 1099, 1100, 1101, 1102, 1103, 1104, 1105, 1052, 1168, 1235, 175, 176].

W [953, 1299, 1331]. W. [347].

Wahrscheinlichkeitstheorie [1185]. Waite [781]. Walter [1552]. War [1358]. Wavefunctions [171, 63]. Wearable [1389]. Web [884, 436, 830]. weight [863, 864]. Weighted [112]. Werner [1060]. Wesley [1457]. WG2 [146]. WG3 [147]. WG4 [148]. Wigner [1058]. William [787]. Window [245, 304, 351, 431]. Wine [856, 856]. Wired [831]. Wireless [832]. WIREs [1589]. Wissenschaften [1355]. Wissenschaftsgeschichte} [1148, 1404]. Wissenschaftstheorie [1326, 1327]. Wolfgang [1057]. word [810]. Work [53]. Working [135, 786]. Workshop [179, 1246, 1787, 1786]. Workshops [679]. World [376]. Worlds [792]. Wright [1196]. Writing [161, 1149]. Written [284].

X [304, 431, 52, 245, 304, 351]. X- [52]. xelatex [1549]. Xi [879]. Xue [879]. xxxvi [1789].

Year [1477]. Years [778].

Zealand [1334, 1359, 1495]. Zeitschrift [1185, 1327, 1326, 1355, 1404]. ZFS [1363]. Zhi [879]. Zipf [995, 998]. zur [1148].

References

Beebe:1969:CSP

[1] Nelson H. F. Beebe. Calculation of spectral parameters for diatomic molecules. Technical report, Quantum Theory Project, Departments of Chemistry and Physics, University of Florida, Gainesville, FL 32601, USA, 1969. 25 pp.

Beebe:1970:PDC

[2] Nelson H. F. Beebe. Program documentation and cataloging. Technical report, Quantum Theory Project, Departments of Chemistry and Physics, University of Florida, Gainesville, FL 32601, USA, February 16, 1970. 2 pp.

Beebe:1970:CMC

[3] Nelson H. F. Beebe. A computational method for configuration interaction calculations. Technical report, Quantum Theory Project, Departments of Chemistry and Physics, University of Florida, Gainesville, FL 32601, USA, March 23, 1970. 29 pp.

Beebe:1970:UPC

[4] Nelson H. F. Beebe. Utility program for copying data sets. Technical report, Quantum Theory Project, Departments of Chemistry and Physics, University of Florida, Gainesville, FL 32601, USA, March 24, 1970. 2 pp.

Beebe:1970:PL

[5] Nelson H. F. Beebe. Program library. Technical report, Quantum Theory Project, Departments of Chemistry and Physics, University of Florida, Gainesville, FL 32601, USA, April 27, 1970. 9 pp.

Beebe:1970:AL

[6] Nelson H. F. Beebe. I/O from assembler language. Technical report, Quantum Theory Project, Departments of Chemistry and Physics, University of Florida, Gainesville, FL 32601, USA, August 12, 1970. 11 pp.

Beebe:1970:AH

[7] Nelson H. F. Beebe. Assembler H. Technical report, Quantum Theory Project, Departments of Chemistry and Physics, University of Florida, Gainesville, FL 32601, USA, August 18, 1970. 2 pp.

Beebe:1970:SDS

[8] Nelson H. F. Beebe. Standardized data set naming. Technical report, Quantum Theory Project, Departments of Chemistry and Physics, University of Florida, Gainesville, FL 32601, USA, August 22, 1970. 2 pp.

Beebe:1970:CCR

[9] Nelson H. F. Beebe. COPE controller with reader, punch, and printer. Technical report, Quantum Theory Project, Departments of Chemistry and Physics, University of Florida, Gainesville, FL 32601, USA, October 26, 1970. 2 pp.

Beebe:1970:UPD

[10] Nelson H. F. Beebe. Upkeep of partitioned data sets. Technical report, Quantum Theory Project, Departments of Chemistry and Physics, University of Florida, Gainesville, FL 32601, USA, November 6, 1970. 3 pp.

Beebe:1970:IDP

[11] Nelson H. F. Beebe. IBM 2311 disk pack — QTP999. Technical report, Quantum Theory Project, Departments of Chemistry and Physics, University of Florida, Gainesville, FL 32601, USA, November 22, 1970. 1 pp.

Beebe:1970:FNC

[12] Nelson H. F. Beebe. FORTRAN name changing utility program. Technical report, Quantum Theory Project, Departments of Chemistry and Physics, University of Florida, Gainesville, FL 32601, USA, November 30, 1970. 1 pp.

Beebe:1970:APH

[13] Nelson H. F. Beebe. Automatic page headers in FORTRAN output. Technical report, Quantum Theory Project, Departments of Chemistry and Physics, University of Florida, Gainesville, FL 32601, USA, 1970. 1 pp.

Beebe:1970:RRR

[14] Nelson H. F. Beebe. RDUMP and RSDUMP: Register dumps for FOR-TRAN programs. Technical report, Quantum Theory Project, Departments of Chemistry and Physics, University of Florida, Gainesville, FL 32601, USA, 1970. 1 pp.

Beebe:1970:BCF

[15] Nelson H. F. Beebe. Batch compilation of FORTRAN. Technical report, Quantum Theory Project, Departments of Chemistry and Physics, University of Florida, Gainesville, FL 32601, USA, 1970. 2 pp.

Beebe:1970:OCD

[16] Nelson H. F. Beebe. Offline core dumps. Technical report, Quantum Theory Project, Departments of Chemistry and Physics, University of Florida, Gainesville, FL 32601, USA, 1970. 2 pp.

Beebe:1970:PTQ

[17] Nelson H. F. Beebe. PUNCH, TRT2629, QTPCOPY utilities. Technical report, Quantum Theory Project, Departments of Chemistry and Physics, University of Florida, Gainesville, FL 32601, USA, 1970. 3 pp.

Purvis:1971:CPF

[18] George D. Purvis, III and Nelson H. F. Beebe. Catalogued procedures FORT-GLIB and FORTHLIB. Technical report, Quantum Theory Project, Departments of Chemistry and Physics, University of Florida, Gainesville, FL 32601, USA, January 1971. 3 pp.

Beebe:1971:LFC

[19] Nelson H. F. Beebe. List of FORTRAN catalogued procedures. Technical report, Quantum Theory Project, Departments of Chemistry and Physics, University of Florida, Gainesville, FL 32601, USA, February 1, 1971. 1 pp.

Beebe:1971:FHC

[20] Nelson H. F. Beebe. FORTRAN H compiler. Technical report, Quantum Theory Project, Departments of Chemistry and Physics, University of Florida, Gainesville, FL 32601, USA, March 2, 1971. 1 pp.

Beebe:1971:CJN

[21] Nelson H. F. Beebe and Erkki Brändas. Changes in job numbers and sequences numbers. Technical report, Quantum Theory Project, Departments of Chemistry and Physics, University of Florida, Gainesville, FL 32601, USA, March 16, 1971. 1 pp.

Beebe:1971:QPA

[22] Nelson H. F. Beebe. QCPE programs available. Technical report, Quantum Theory Project, Departments of Chemistry and Physics, University of Florida, Gainesville, FL 32601, USA, April 6, 1971. 2 pp.

Beebe:1971:RCP

[23] Nelson H. F. Beebe. Revised Cal-Comp plotter package. Technical report, Quantum Theory Project, Departments of Chemistry and Physics, University of Florida, Gainesville, FL 32601, USA, April 16, 1971. 13 pp.

Beebe:1971:CUa

[24] Nelson H. F. Beebe. Computer usage. Technical report, Quantum Theory Project, Departments of Chemistry and Physics, University of Florida, Gainesville, FL 32601, USA, May 11, 1971. 1 pp.

Beebe:1971:MIF

[25] Nelson H. F. Beebe. More information on FORTRAN H EXTENDED. Technical report, Quantum Theory Project, Departments of Chemistry and Physics, University of Florida, Gainesville, FL 32601, USA, May 11, 1971. 1 pp.

Beebe:1971:NID

[26] Nelson H. F. Beebe. New IBM 2311 disk pack and data sets. Technical report, Quantum Theory Project, Departments of Chemistry and Physics, University of Florida, Gainesville, FL 32601, USA, May 11, 1971. 2 pp.

Beebe:1971:SUP

[27] Nelson H. F. Beebe. Some utility programs for the IBM 360 and 370 computers. Technical Report QTP-TF-234, Quantum Theory Project, Departments of Chemistry and Physics, University of Florida, Gainesville, FL 32601, USA, May 15, 1971. 49 pp.

Beebe:1971:FHE

[28] Nelson H. F. Beebe. FORTRAN H (EXTENDED). Technical report, Quantum Theory Project, Departments of Chemistry and Physics, University of Florida, Gainesville, FL 32601, USA, June 15, 1971. 5 pp.

Beebe:1971:PPG

[29] Nelson H. F. Beebe. Possible problems with GIVENS. Technical report, Quantum Theory Project, Departments of Chemistry and Physics, University of Florida, Gainesville, FL 32601, USA, August 4, 1971. 1 pp.

Beebe:1971:CUb

[30] Nelson H. F. Beebe. Computer usage. Technical report, Quantum Theory Project, Departments of Chemistry and Physics, University of Florida, Gainesville, FL 32601, USA, September 11, 1971. 1 pp.

Beebe:1971:DSS

[31] Nelson H. F. Beebe. Data set security. Technical report, Quantum Theory Project, Departments of Chemistry and Physics, University of Florida, Gainesville, FL 32601, USA, September 27, 1971. 1 pp.

Beebe:1971:CJS

[32] Nelson H. F. Beebe. Changes in job sequence numbers. Technical report, Quantum Theory Project, Departments of Chemistry and Physics, University of Florida, Gainesville, FL 32601, USA, September 27, 1971. 2 pp.

Beebe:1971:CCP

[33] Nelson H. F. Beebe. Compressing and copying partitioned data sets. Technical report, Quantum Theory Project, Departments of Chemistry and Physics, University of Florida, Gainesville, FL 32601, USA, September 27, 1971. 3 pp.

Beebe:1971:KDP

[34] Nelson H. F. Beebe. Keypunch drumcard preparation. Technical report, Quantum Theory Project, Departments of Chemistry and Physics, University of Florida, Gainesville, FL 32601, USA, December 11, 1971. 2 pp.

Beebe:1972:FTC

[35] Nelson H. F. Beebe. Fast turnaround at the Computer Center. Technical report, Quantum Theory Project, Departments of Chemistry and Physics, University of Florida, Gainesville, FL 32601, USA, February 2, 1972. 2 pp.

Beebe:1972:SDT

[36] Nelson H. F. Beebe. Spectral data for three-electron ions. Technical report, Quantum Theory Project, Departments of Chemistry and Physics, University of Florida, Gainesville, FL 32601, USA, March 21, 1972. 13 pp.

Beebe:1972:SFR

[37] Nelson H. F. Beebe. Saving files on RJE and card listing. Technical report, Quantum Theory Project, Departments of Chemistry and Physics, University of Florida, Gainesville, FL 32601, USA, March 27, 1972. 1 pp.

Beebe:1972:GSO

[38] Nelson H. F. Beebe. General Spin Orbitals for Three-Electron Systems. PhD thesis, University of Florida, Gainesville, FL 32611, USA, June 1972. URL

http://www.archive.org/details/
generalspinorbit00beebrich.

Beebe:1972:ICI

[39] Nelson H. F. Beebe. IBM S/370 charges, IBM 3330 disks, CLASS=T. Technical report, Quantum Theory Project, Departments of Chemistry and Physics, University of Florida, Gainesville, FL 32601, USA, September 13, 1972. 2 pp.

Beebe:1972:CRD

[40] Nelson H. F. Beebe. Construction of reduced density matrices. Technical report, Quantum Theory Project, Departments of Chemistry and Physics, University of Florida, Gainesville, FL 32601, USA, 1972. 7 pp.

Beebe:1972:NSS

[41] Nelson H. F. Beebe. A note on space spanning. Int. J. Quant. Chem., 6S: 439–441, 1972. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Beebe:1972:POT

[42] Nelson H. F. Beebe and Sten Lunell. A projection operator technique for orbital basis variation. Int. J. Quant. Chem., 6:1149–1155, 1972. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Beebe:1973:QDP

[43] Nelson H. F. Beebe. QTP 3330 disk pack. Technical report, Quantum Theory Project, Departments of Chemistry and Physics, University of Florida, Gainesville, FL 32601, USA, February 14, 1973. 3 pp.

Beebe:1973:HLP

[44] Nelson H. F. Beebe. Hidden-line plotting of molecular orbitals and densities. Technical report, Quantum Theory Project, Departments of Chemistry and Physics, University of Florida, Gainesville, FL 32601, USA, May 8, 1973. 5 pp.

Beebe:1973:CPS

[45] Nelson H. F. Beebe. Core partition sizes. Technical report, Quantum Theory Project, Departments of Chemistry and Physics, University of Florida, Gainesville, FL 32601, USA, June 1, 1973. 1 pp.

Beebe:1973:CQT

[46] Nelson H. F. Beebe. Comment on the Quantum Theory Project computing fund request. Technical report, Quantum Theory Project, Departments of Chemistry and Physics, University of Florida, Gainesville, FL 32601, USA, July 10, 1973. 1 pp.

Beebe:1973:RHL

[47] Nelson H. F. Beebe. Revision of hidden-line molecular plot package. Technical report, Quantum Theory Project, Departments of Chemistry and Physics, University of Florida, Gainesville, FL 32601, USA, September 24, 1973. 1 pp.

Beebe:1973:RN

[48] Nelson H. F. Beebe. Revision of NODELIST. Technical report, Quantum Theory Project, Departments of Chemistry and Physics, University of Florida, Gainesville, FL 32601, USA, September 24, 1973. 1 pp.

Beebe:1973:PCN

[49] Nelson H. F. Beebe. Program to convert numbers of an arbitrary base to base 10 on a Hewlett–Packard. Technical report, Quantum Theory Project, Departments of Chemistry and Physics, University of Florida, Gainesville, FL 32601, USA, 1973. 1 pp.

Beebe:1973:HLC

[50] Nelson H. F. Beebe. Hidden-line CAL-COMP plots. Technical report, Quantum Theory Project, Departments of Chemistry and Physics, University of Florida, Gainesville, FL 32601, USA, 1973. 2 pp.

Beebe:1973:ACS

[51] Nelson H. F. Beebe. ATS command summary. Technical report, Quantum Theory Project, Departments of Chemistry and Physics, University of Florida, Gainesville, FL 32601, USA, 1973. 7 pp.

Beebe:1973:TSX

[52] Nelson H. F. Beebe. On the transition state in the X-α method. Chem. Phys. Letters, 19(2):290–294, March 15, 1973. CODEN CH-PLBC. ISSN 0009-2614 (print), 1873-4448 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0009261473850778.

Beebe:1973:CPH

[53] Nelson H. F. Beebe and Yngve Öhrn. Collected papers on the hydroxyl cation, radical, and anion: A bibliography of theoretical and experimental work. Technical report, Quantum Theory Project, Departments of Chemistry and Physics, University of Florida, Gainesville, FL 32601, USA, 1973.

Beebe:1974:IID

[54] Nelson H. F. Beebe. IBM 3330, IBM 2311, data sets, and miscellaneous random thoughts. Technical report, Quantum Theory Project, Departments of Chemistry and Physics, University of Florida, Gainesville, FL 32601, USA, July 4, 1974. 3 pp.

Beebe:1974:VT

[55] Nelson H. F. Beebe. Various things. Technical report, Quantum Theory Project, Departments of Chemistry and Physics, University of Florida, Gainesville, FL 32601, USA, September 24, 1974. 2 pp.

Beebe:1974:ANH

[56] Nelson H. F. Beebe. Additional notes to HIDEPLOT. Technical report, Kemisk Institut, Aarhus University, Århus, Denmark, 1974. 12 pp.

Beebe:1974:EIS

[57] Nelson H. F. Beebe and J. R. Sabin. An ab initio study of the electronic structure of C_{2}O_{2}. Chem. Phys. Letters, 24(3):389–394, February 1, 1974. CODEN CHPLBC. ISSN 0009-2614 (print), 1873-4448 (electronic).

Jao:1974:MPE

[58] T. C. Jao, Nelson H. F. Beebe, W. B. Person, and J. R. Sabin. Molecular polarizability estimates for vibrational spectral interpretation. Chem. Phys. Letters, 26(4): 474–478, June 15, 1974. CODEN CHPLBC. ISSN 0009-2614 (print), 1873-4448 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0009261474803945.

Oddershede:1975:SCT

[59] Jens Oddershede, Poul Jørgensen, and Nelson H. F. Beebe. Selfconsistent time-dependent Hartree– Fock calculations of dynamic polarizabilities, dispersion forces, and nuclear spin-spin coupling constants for the H_{2} and HF molecules. J. Chem. Phys., 63(7):2996–3007, October 1, 1975. CODEN JCPSA6. ISSN 0021-9606 (print), 1089-7690

(electronic). URL http://ojps.aip.
org/jcpo;   http://ojps.aip.org/

journal_cgi/search?KEY=JCPSA6;
http://www.math.utah.edu/~beebe/
publications/1975/j-chem-phys-63-
7-2996-1-oct-1975.pdf.

Beebe:1975:ADH

[60] Nelson H. F. Beebe. The Århus diatomic Hartree–Fock program. Technical report, Kemisk Institut, Aarhus University, Århus, Denmark, November 1975. 7 pp.

Beebe:1975:BGE

[61] Nelson H. F. Beebe. Bibliography on the generalized eigenvalue problem Ax = λBx. Technical report, Kemisk Institut, Aarhus University, Århus, Denmark, 1975. 3 pp.

Beebe:1975:PCR

[62] Nelson H. F. Beebe. On the practical construction of reduced density matrices. Technical report, Kemisk Institut, Aarhus University, Århus, Denmark, 1975. 12 pp.

Beebe:1975:GSO

[63] Nelson H. F. Beebe and Sten Lunell. Projected general spinorbital wavefunctions for three-electron systems. J. Phys. B: Atom. Molec. Phys., 8(14):2320–2326, October 1, 1975. CODEN JPAMA4. ISSN 0953-4075 (print), 1361-6455 (electronic). URL http://www.

iop.org/EJ/article/0022-3700/8/
14/013/jbv8i14p2320.pdf;   http:
//www.math.utah.edu/~beebe/publications/
1975/jbv8i14p2320.pdf.

Beebe:1975:FRD

[64] Nelson H. F. Beebe. On the form of reduced density matrices in a nonorthogonal basis. Technical report, Kemisk Institut, Aarhus University, Århus, Denmark, 1975.

Beebe:1976:MEA

[65] Nelson H. F. Beebe. Multiplet energies of atoms. Technical report, Kemisk Institut, Aarhus University, Århus, Denmark, February 1976. 5 pp.

Beebe:1976:CIC

[66] Nelson H. F. Beebe, Erik W. Thulstrup, and Andreas Andersen. Configuration interaction calculations of lowlying electronic states of O_{2}, O_{2}^{+}, and O_{2}^{2+}. J. Chem. Phys., 64(5):2080–2093, March 1, 1976. CODEN JCPSA6. ISSN 0021-9606 (print), 1089-7690

                                        (electronic). URL http://ojps.aip.
                                        org/jcpo;   http://ojps.aip.org/
                                        journal_cgi/search?KEY=JCPSA6;
                                        http://www.math.utah.edu/~beebe/
                                        publications/1976/j-chem-phys-64-
                                        5-2080-1-mar-1976.pdf.

Beebe:1976:DTFa

[67] Nelson H. F. Beebe. DOCUMENT: A text formatting program. Technical report, Quantum Theory Project, Departments of Chemistry and Physics, University of Florida, Gainesville, FL 32601, USA, May 12, 1976. 9 pp.

Beebe:1976:DNM

[68] Nelson H. F. Beebe. Distribution notes for the Mexico Plot System, #1. Technical report, Kemisk Institut, Aarhus University, Århus, Denmark, May 18, 1976. 6 pp.

Beebe:1976:NSK

[69] Nelson H. F. Beebe. Notes on SCAN: Keyword search program. Technical report, Kemisk Institut, Aarhus University, Århus, Denmark, August 8, 1976. 1 pp.

Beebe:1976:NDP

[70] Nelson H. F. Beebe. Notes on the diatomic properties package. Technical report, Kemisk Institut, Aarhus University, Århus, Denmark, August 19, 1976. 2 pp.

Beebe:1976:NEP

[71] Nelson H. F. Beebe. Notes on EDIT — A program for editing FORTRAN programs. Technical report, Kemisk Institut, Aarhus University, Århus, Denmark, August 19, 1976. 2 pp.

Beebe:1976:NM

[72] Nelson H. F. Beebe. Notes on MOD-NAME. Technical report, Kemisk Institut, Aarhus University, Århus, Denmark, August 19, 1976. 2 pp.

Beebe:1976:NMPa

[73] Nelson H. F. Beebe. Notes on MAP — A program for mapping and crossreferencing object files. Technical report, Kemisk Institut, Aarhus University, Århus, Denmark, August 19, 1976. 2 pp.

Beebe:1976:NMPb

[74] Nelson H. F. Beebe. Notes on the molecular plot program. Technical report, Kemisk Institut, Aarhus University, Århus, Denmark, August 19, 1976. 2 pp.

Beebe:1976:NPP

[75] Nelson H. F. Beebe. Notes on the polarization propagator program. Technical report, Kemisk Institut, Aarhus University, Århus, Denmark, August 19, 1976. 2 pp.

Beebe:1976:NSPa

[76] Nelson H. F. Beebe. Notes on selected procedure files. Technical report, Kemisk Institut, Aarhus University, Århus, Denmark, August 19, 1976. 2 pp.

Beebe:1976:NSPb

[77] Nelson H. F. Beebe. Notes on the SPLINE program. Technical report, Kemisk Institut, Aarhus University, Århus, Denmark, August 19, 1976. 2 pp.

Beebe:1976:NAD

[78] Nelson H. F. Beebe. Notes on the Aarhus diatomic Hartree–Fock program. Technical report, Kemisk Institut, Aarhus University, Århus, Denmark, August 19, 1976. 3 pp.

Beebe:1976:NUH

[79] Nelson H. F. Beebe. Notes on the Unrestricted Hartree–Fock program. Technical report, Kemisk Institut, Aarhus University, Århus, Denmark, August 19, 1976. 3 pp.

Beebe:1976:NRS

[80] Nelson H. F. Beebe. Notes on the Roos–Siegbahn molecular integral transformation program and combined perturbation theory/configuration interaction program. Technical report, Kemisk Institut, Aarhus University, Århus, Denmark, August 19, 1976. 4 pp.

Beebe:1976:NGP

[81] Nelson H. F. Beebe. Notes on the Gaussian properties package. Technical report, Kemisk Institut, Aarhus University, Århus, Denmark, August 19, 1976. 5 pp.

Beebe:1976:NUO

[82] Nelson H. F. Beebe. Notes on UIBMOL — an open shell version of MOLECULE. Technical report, Kemisk Institut, Aarhus University, Århus, Denmark, August 19, 1976. 9 pp.

Beebe:1976:DTFb

[83] Nelson H. F. Beebe. DOCUMENT: A text formatting program. Technical report, Quantum Theory Project, Departments of Chemistry and Physics, University of Florida, Gainesville, FL 32601, USA, August 19, 1976. 19 pp.

Beebe:1976:CPN

[84] Nelson H. F. Beebe. Computer program notes. Technical report, Kemisk Institut, Aarhus University, Århus, Denmark, August 19, 1976.

Beebe:1976:NAS

[85] Nelson H. F. Beebe. Notes on the atomic SCF program. Technical report, Kemisk Institut, Aarhus University, Århus, Denmark, August 19, 1976.

Beebe:1976:DNS

[86] Nelson H. F. Beebe. Distribution notes to SPLINE and HIDEPLOT. Technical report, Kemisk Institut, Aarhus University, Århus, Denmark, August 1976. 3 pp.

Beebe:1976:ERK

[87] Nelson H. F. Beebe. Evaluation of Rydberg–Klein–Rees potential curves. Technical report, Kemisk Institut, Aarhus University, Århus, Denmark, 1976. 4 pp.

Beebe:1976:PQM

[88] Nelson H. F. Beebe. Postulates of quantum mechanics. Technical report, Kemisk Institut, Aarhus University, Århus, Denmark, 1976. 4 pp.

Beebe:1976:RLS

[89] Nelson H. F. Beebe. Review of linear spaces and Hilbert space. Technical report, Kemisk Institut, Aarhus University, Århus, Denmark, 1976. 4 pp.

Beebe:1976:UTS

[90] Nelson H. F. Beebe. Utah tape specifications. Technical report, Kemisk Institut, Aarhus University, Århus, Denmark, 1976. 4 pp.

Beebe:1976:DSP

[91] Nelson H. F. Beebe. On the determination of spectroscopic parameters. Technical report, Kemisk Institut, Aarhus University, Århus, Denmark, 1976. 7 pp.

Beebe:1976:COC

[92] Nelson H. F. Beebe. On the choice of orbitals for CI [configuration interaction] calculations. Technical report, Kemisk Institut, Aarhus University, Århus, Denmark, 1976. 8 pp.

Beebe:1976:PO

[93] Nelson H. F. Beebe. Projection operators. Technical report, Kemisk Institut, Aarhus University, Århus, Denmark, 1976. 12 pp.

Beebe:1976:SIS

[94] Nelson H. F. Beebe. Supplement to Integral Symmetry. Technical report, Kemisk Institut, Aarhus University, Århus, Denmark, 1976. 13 pp.

Beebe:1976:HFA

[95] Nelson H. F. Beebe. The Hartree– Fock approximation. Technical report, Kemisk Institut, Aarhus University, Århus, Denmark, 1976. 21 pp.

Beebe:1976:EPL

[96] Nelson H. F. Beebe. The eigenvalue problem for large sparse matrices. Technical report, Quantum Theory Project, Departments of Chemistry and Physics, University of Florida, Gainesville, FL 32601, USA, 1976. 26 pp.

Beebe:1976:SFS

[97] Nelson H. F. Beebe. On the symmetry of filled shells. Int. J. Quant. Chem., 10:1007–1010, 1976. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Beebe:1977:GPI

[98] Nelson H. F. Beebe. Grading program improvements. Technical report, Quantum Theory Project, Departments of Chemistry and Physics, University of Florida, Gainesville, FL 32601, USA, February 27, 1977. 5 pp.

Beebe:1977:DNMa

[99] Nelson H. F. Beebe. Distribution notes for the Mexico Plot System, #2. Technical report, Kemisk Institut, Aarhus University, Århus, Denmark, March 15, 1977. 9 pp.

Beebe:1977:DNMb

[100] Nelson H. F. Beebe. Distribution notes for the Mexico Plot System, #3. Technical report, Kemisk Institut, Aarhus University, Århus, Denmark, March 29, 1977. 4 pp.

Elander:1977:PSB

[101] Nils Elander, Jens Oddershede, and Nelson H. F. Beebe. On the A^{1}Π— X^{1}Σ^{+} bandsystem in CH^{+} and CD^{+}: Theoretical spectroscopic constants and lifetimes. Astrophys. J., 216: 165–173, August 15, 1977. CO-DEN ASJOAB. ISSN 0004-637X (print), 1538-4357 (electronic). URL

http://www.math.utah.edu/~beebe/
publications/1977/astrophys-j-216-
165-1977.pdf.

Beebe:1977:CPP

[102] Nelson H. F. Beebe. COPY, PRINT, PUNCH, and STORE utility. Technical Report QTP.770921.01, Quantum Theory Project, Departments of Chemistry and Physics, University of Florida, Gainesville, FL 32601, USA, September 21, 1977. 9 pp.

Beebe:1977:EFP

[103] Nelson H. F. Beebe. Editing FOR-TRAN programs. Technical Report QTP.770927.02, Quantum Theory Project, Departments of Chemistry and Physics, University of Florida, Gainesville, FL 32601, USA, September 27, 1977. 6 pp.

Beebe:1977:MIF

[104] Nelson H. F. Beebe. Machineindependent FORTRAN programming. Technical Report QTP.771205.03, Quantum Theory Project, Departments of Chemistry and Physics, University of Florida, Gainesville, FL 32601, USA, December 5, 1977. 10 pp.

Jorgensen:1977:PPC

[105] Poul Jørgensen, Jack Simons, and Nelson H. F. Beebe. Polarization propagator calculation of the static polarizability of He_{2}. Technical report, Kemisk Institut, Aarhus University, Århus, Denmark, 1977. 16 pp.

Beebe:1977:NEP

[106] Nelson H. F. Beebe. Notes on the eigenvalue problem for large sparse matrices. Technical report, Quantum Theory Project, Departments of Chemistry and Physics, University of Florida, Gainesville, FL 32601, USA, 1977. 31 pp.

Lunell:1977:HFC

[107] Sten Lunell and Nelson H. F. Beebe. Hartree–Fock and configuration interaction calculation of hyperfine constants in quartet states of three-electron atoms. Physica Scripta, 15(4):268–272, 1977. CODEN PHSTBO. ISSN 0031-8949 (print), 1402-4896 (electronic).

URL http://www.math.utah.edu/~
beebe/publications/1977/physica-
scripta-15-4-268-1977.pdf.

Oddershede:1977:CHF

[108] Jens Oddershede, Poul Jørgensen, and Nelson H. F. Beebe. Coupled Hartree– Fock and second order polarization propagator calculations of indirect nuclear spin-spin coupling constants for diatomic molecules. Chem. Phys., 25: 451–458, 1977. CODEN CMPHC2. ISSN 0301-0104.

Oddershede:1977:DEE

[109] Jens Oddershede, Poul Jørgensen, and Nelson H. F. Beebe. Determination of excitation energies and transition moments in a second order polarization propagator approach. application to the Be atom and the CH^{+} molecule. Int. J. Quant. Chem., 12:655–670, 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Beebe:1977:STE

[110] Nelson H. F. Beebe and Jan Linderberg. Simplifications in the twoelectron integral array in molecular calculations. Int. J. Quant. Chem., 12 (4):683–705, October 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Jorgensen:1978:AEE

[111] Poul Jørgensen, Jens Oddershede, and Nelson H. F. Beebe. Analysis of excitation energies and transition moments. J. Phys. B: Atom. Molec. Phys., 11 (1):1–15, January 14, 1978. CODEN JPAMA4. ISSN 0953-4075 (print), 1361-6455 (electronic). URL http://

iopscience.iop.org/0022-3700/11/
1/007;   http://www.iop.org/EJ/
article/0022-3700/11/1/007/jbv11i1p1.
pdf;   http://www.math.utah.edu/
~beebe/publications/1978/jbv11i1p1.
pdf.

Jorgensen:1978:PPC

[112] Poul Jørgensen, Jens Oddershede, and Nelson H. F. Beebe. Polarization propagator calculations of frequencydependent polarizabilities, Verdet constants, and energy-weighted sum rules. J. Chem. Phys., 68(6):2527–2532, March 15, 1978. CODEN JCPSA6. ISSN 0021-9606 (print), 1089-7690

(electronic). URL http://ojps.aip.
org/jcpo;   http://ojps.aip.org/
journal_cgi/search?KEY=JCPSA6;
http://www.math.utah.edu/~beebe/
publications/1978/j-chem-phys-68-
6-2527-15-mar-1978.pdf.

Jorgensen:1978:FDP

[113] Poul Jørgensen, Jens Oddershede, Preben Albertsen, and Nelson H. F. Beebe. Frequency-dependent polarizabilities and Verdet constants for He, Be, CO, and FH. J. Chem. Phys., 68(6):2533–2543, March 15, 1978. CODEN JCPSA6. ISSN 0021-9606 (print), 1089-7690 (electronic).

URL http://ojps.aip.org/getpdf/
servlet/GetPDFServlet?filetype=
pdf&id=JCPSA6000068000006002533000001&
idtype=cvips;   http://ojps.aip.
org/jcpo; http://ojps.aip.org/journal_
cgi/search?KEY=JCPSA6;   http://
www.math.utah.edu/~beebe/publications/
1978/j-chem-phys-68-6-2533-15-mar-
1978.pdf.

Beebe:1978:MTDa

[114] Nelson H. F. Beebe. Machinetransportable document formatting. Technical Report QTP.780503.04, Quantum Theory Project, Departments of Chemistry and Physics, University of Florida, Gainesville, FL 32601, USA, May 3, 1978. 9 pp.

Beebe:1978:DNMa

[115] Nelson H. F. Beebe. Distribution notes for the Mexico Plot System, #4. Technical report, Quantum Theory Project, Departments of Chemistry and Physics, University of Florida, Gainesville, FL 32601, USA, May 17, 1978. 16 pp.

Beebe:1978:CRM

[116] Nelson H. F. Beebe. Cross-reference maps of object libraries. Technical Report QTP.780614.05, Quantum Theory Project, Departments of Chemistry and Physics, University of Florida, Gainesville, FL 32601, USA, June 14, 1978. 7 pp.

Beebe:1978:DNMb

[117] Nelson H. F. Beebe. Distribution notes for the Mexico Plot System, #5. Technical report, Quantum Theory Project, Departments of Chemistry and Physics, University of Florida, Gainesville, FL 32601, USA, June 14, 1978. 35 pp.

Beebe:1978:FMF

[118] Nelson H. F. Beebe. Feler’s method for finding eigenvalues and eigenvectors. In Moler and Shavitt [1786], pages 66–67. LCCN QD39.3.M3 W61 1978.

                                        URL http://escholarship.org/uc/
                                        item/3xb320bq.  UC-32, CONF-

780878.

Beebe:1978:DAI

[119] Nelson H. F. Beebe and Jan Linderberg. A different approach to integrals and integral transformations. In Moler and Shavitt [1786], pages 158– 162. LCCN QD39.3.M3 W61 1978.

                                        URL http://escholarship.org/uc/
                                        item/3xb320bq.  UC-32, CONF-

780878.

Beebe:1978:GEP

[120] Nelson H. F. Beebe. The generalized eigenvalue problem in quantum chemistry. In Moler and Shavitt [1786], pages 63–65. LCCN QD39.3.M3 W61

1978. URL http://escholarship.
org/uc/item/3xb320bq.  UC-32,

CONF-780878.

Beebe:1978:SSJ

[121] Nelson H. F. Beebe. Single step job procedures. Technical Report QTP.780703.06, Quantum Theory Project, Departments of Chemistry and Physics, University of Florida, Gainesville, FL 32601, USA, July 3, 1978. 5 pp.

Beebe:1978:SLD

[122] Nelson H. F. Beebe. Selective listing of data set attributes. Technical Report QTP.780725.07, Quantum Theory Project, Departments of Chemistry and Physics, University of Florida, Gainesville, FL 32601, USA, July 25, 1978. 3 pp.

Beebe:1978:BGE

[123] Nelson H. F. Beebe. Bibliography on the generalized eigenvalue problem, Ax = λBx. In Proceedings of the NRCC Workshop on Numerical Algorithms in Chemistry: Algebraic Methods, pages 71–72. National Resource for Computation in Chemistry, Lawrence Berkeley Laboratory, Berkeley, CA, USA, August 9–11, 1978.

Beebe:1978:MTDb

[124] Nelson H. F. Beebe. Machinetransportable document formatting [new release]. Technical Report QTP.780822.08, Quantum Theory Project, Departments of Chemistry and Physics, University of Florida, Gainesville, FL 32601, USA, August 22, 1978. 19 pp.

Beebe:1978:UEF

[125] Nelson H. F. Beebe. A utility for editing FORTRAN programs [new release]. Technical Report QTP.780825.09, Quantum Theory Project, Departments of Chemistry and Physics, University of Florida, Gainesville, FL 32601, USA, August 25, 1978. 18 pp.

Beebe:1978:CSS

[126] Nelson H. F. Beebe. A comparison of some sorting algorithms. Technical Report QTP-TF-234, Quantum Theory Project, Departments of Chemistry and Physics, University of Florida, Gainesville, FL 32601, USA, September 1, 1978. 70 pp.

Beebe:1978:NSK

[127] Nelson H. F. Beebe. Notes on SCAN: Keyword search program. Technical report, Kemisk Institut, Aarhus University, Århus, Denmark, October 8, 1978. 1 pp.

Beebe:1978:RHS

[128] Nelson H. F. Beebe. Remarks on the Harwell Subroutine Library. Technical report, College of Science Computer, University of Utah, Salt Lake City, UT 84112, USA, November 24, 1978. 6 pp.

Beebe:1978:GEA

[129] Nelson H. F. Beebe. Guest editorial about the National Resource for Computation in Chemistry. Quantum Chemistry Program Exchange Newsletter, 63:5–12, November 1978. CODEN QCBUEN. ISSN 0889-7514.

Beebe:1978:BTA

[130] Nelson H. F. Beebe. Beebe tape assignments. Technical report, Quantum Theory Project, Departments of Chemistry and Physics, University of Florida, Gainesville, FL 32601, USA, 1978. 4 pp.

Beebe:1978:NEP

[131] Nelson H. F. Beebe. Notes on the eigenvalue problem for large sparse matrices. Technical report, College of Science Computer, University of Utah, Salt Lake City, UT 84112, USA, 1978. 31 pp.

Beebe:1979:IFD

[132] Nelson H. F. Beebe. Integral file data compression. Technical report, College of Science Computer, University of Utah, Salt Lake City, UT 84112, USA, May 18, 1979. 39 pp.

Beebe:1979:BSP

[133] Nelson H. F. Beebe. Bibliography on software portability. In Harris and Beebe [1785], page 4.

URL http://escholarship.org/uc/
item/19b1d4sw.

Beebe:1979:DMC

[134] Nelson H. F. Beebe. DECsystem-20 MOLECULE I/O conventions. Technical report, College of Science Computer, University of Utah, Salt Lake City, UT 84112, USA, June 15, 1979. 6 pp.

Beebe:1979:WPS

[135] Nelson H. F. Beebe. Working paper on software standards. In Harris and Beebe [1785], page 13. URL

                                        http://escholarship.org/uc/item/
                                        19b1d4sw.

Beebe:1979:PGP

[136] Nelson H. F. Beebe. Programmer’s guide to portable software. Technical report, College of Science Computer, University of Utah, Salt Lake City, UT 84112, USA, June 30, 1979. 107 pp.

Beebe:1979:SI

[137] Nelson H. F. Beebe. SFTRAN ideas. Technical report, College of Science Computer, University of Utah, Salt Lake City, UT 84112, USA, July 2, 1979. 2 pp.

Beebe:1979:SB

[138] Nelson H. F. Beebe. SFTRAN/3 bugs. Technical report, College of Science Computer, University of Utah, Salt Lake City, UT 84112, USA, July 13, 1979. 2 pp.

Beebe:1979:SR

[139] Nelson H. F. Beebe. SFTRAN/3 remarks. Technical report, College of Science Computer, University of Utah, Salt Lake City, UT 84112, USA, July 13, 1979. 2 pp.

Beebe:1979:RST

[140] Nelson H. F. Beebe. A roster of software tools. In Proceedings of the NRCC Conference on Software Standards in Chemistry, pages 81–105. University of Utah, Salt Lake City, UT 84112, USA, July 25–27, 1979.

Beebe:1979:PSSa

[141] Nelson H. F. Beebe. Proposal for a standard set of primitives for machineindependent bit manipulation in FOR-TRAN. In Proceedings of the NRCC Conference on Software Standards in Chemistry, pages 106–115. University of Utah, Salt Lake City, UT 84112, USA, July 25–27, 1979.

Beebe:1979:PSSb

[142] Nelson H. F. Beebe. Proposal for a standard set of primitives for machineindependent character manipulation in FORTRAN. In Proceedings of the NRCC Conference on Software Standards in Chemistry, pages 116–128. University of Utah, Salt Lake City, UT 84112, USA, July 25–27, 1979.

Beebe:1979:OMI

[143] Nelson H. F. Beebe, J. S. Binkley, D. J. Duchamp, S. Elbert, A. D. McLean, G. D. Purvis, III, and R. C. Raffenetti. Outline for a machine-independent random-access input/output interface. In Proceedings of the NRCC Conference on Software Standards in Chemistry, pages 129–150. University of Utah, Salt Lake City, UT 84112, USA, July 25–27, 1979.

Beebe:1979:PSSc

[144] Nelson H. F. Beebe. Proposal for a standard set of primitives for machineindependent bit manipulation in FOR-TRAN. Technical report, College of Science Computer, University of Utah, Salt Lake City, UT 84112, USA, December 13, 1979. 10 pp.

Beebe:1979:PSSd

[145] Nelson H. F. Beebe. Proposal for a standard set of primitives for machineindependent character manipulation in FORTRAN. Technical report, College of Science Computer, University of Utah, Salt Lake City, UT 84112, USA, December 13, 1979. 13 pp.

Beebe:1979:STWa

[146] Nelson H. F. Beebe. Suggested topics for WG2 discussions: Documentation and coding standards. In Harris and Beebe [1785], page 3.

                                        URL http://escholarship.org/uc/
                                        item/19b1d4sw.

Beebe:1979:STWb

[147] Nelson H. F. Beebe. Suggested topics for WG3 discussions: Machine specification: What is a minimal computer configuration? In Harris and Beebe [1785], page 3.

                                        URL http://escholarship.org/uc/
                                        item/19b1d4sw.

Beebe:1979:STWc

[148] Nelson H. F. Beebe. Suggested topics for WG4 discussions: Quantum chemistry software base. In Harris and Beebe [1785], page 3.

                                        URL http://escholarship.org/uc/
                                        item/19b1d4sw.

Beebe:1979:MVO

[149] Nelson H. F. Beebe. Modification of virtual orbitals. Int. J. Quant. Chem., 15:589–600, 1979. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Harris:1979:EN

[150] Frank E. Harris and Nelson H. F. Beebe. Editors’ note. In Proceedings of the conference Software Standards in Chemistry, Held at the University of Utah, July 25–27, 1979 [1785], pages

ix–x. URL http://escholarship.
org/uc/item/19b1d4sw.

Beebe:1980:OMI

[151] Nelson H. F. Beebe, J. Stephen Binkley, David J. Duchamp, Stephen T. Elbert, A. Douglas McLean, George D. Purvis, III, and Richard C. Raffenetti. Outline for a machine-independent random-access input/output interface. Technical report, College of Science Computer, University of Utah, Salt Lake City, UT 84112, USA, January 26, 1980. 22 pp.

Beebe:1980:SPR

[152] Nelson H. F. Beebe, C. L. Lawson, and J. A. Flynn. SFTRAN3 programmers reference manual. Technical report, College of Science Computer, University of Utah, Salt Lake City, UT 84112, USA, March 3, 1980. 44 pp.

Beebe:sf3-refman

[153] C. L. Lawson, J. A. Flynn, and Nelson H. F. Beebe. SFTRAN3 programmers reference manual. Technical report, Jet Propulsion Laboratory and-College of Science Computer, University of Utah, California Institute of Technology, Pasadena, California, USA andSalt Lake City, UT 84112, USA, March 3, 1980. 44 pp.

Beebe:1980:SIG

[154] Nelson H. F. Beebe. SFTRAN3 installation guide. Technical report, College of Science Computer, University of Utah, Salt Lake City, UT 84112, USA, March 4, 1980. 69 pp.

Beebe:1980:UPP

[155] Nelson H. F. Beebe. The Utah Polarization Propagator Program. Technical report, College of Science Computer, University of Utah, Salt Lake City, UT 84112, USA, April 23, 1980. 46 pp.

Motherwell:1980:PPP

[156] Sam Motherwell, Joel L. Sussman, Zappora Shakked, David Barnett, and Nelson H. F. Beebe. PLUTO: A program for plotting molecular and crystal structures: Version 1.7. Technical report, University Chemical Laboratory, Lensfield Road, Cambridge CB2 1EW, England; The Weizmann Institute of Science, Rehovoth, Israel; Departments of Physics and Chemistry, University of Utah, Salt Lake City, UT 84112, USA, April 23, 1980. 57 pp.

Beebe:1980:SMP

[157] Nelson H. F. Beebe. Summary of MOLECULE program system. Technical report, College of Science Computer, University of Utah, Salt Lake City, UT 84112, USA, May 2, 1980. 8 pp.

Beebe:1980:DPT

[158] Nelson H. F. Beebe. DOCUMENT— A portable text formatting program. Technical report, College of Science Computer, University of Utah, Salt Lake City, UT 84112, USA, May 5, 1980. 138 pp.

Beebe:1980:PPF

[159] Nelson H. F. Beebe. PRETTY— A portable FORTRAN program prettyprinter. Technical report, College of Science Computer, University of Utah, Salt Lake City, UT 84112, USA, May 7, 1980. 33 pp.

Beebe:1980:SLU

[160] Nelson H. F. Beebe. SFTRAN3 listings of the Utah Polarization Propagator Program. Technical report, College of Science Computer, University of Utah, Salt Lake City, UT 84112, USA, May 13, 1980. 158 pp.

Beebe:1980:WDI

[161] Nelson H. F. Beebe. Writing a device interface for <PLOT79>. Technical report, College of Science Computer, University of Utah, Salt Lake City, UT 84112, USA, May 14, 1980. 28 pp.

Beebe:1980:THL

[162] Nelson H. F. Beebe. Timing of hiddenline plots. Technical report, College of Science Computer, University of Utah, Salt Lake City, UT 84112, USA, July 5, 1980. 2 pp.

Beebe:1980:UGP

[163] Nelson H. F. Beebe. A user’s guide to <PLOT79>. Technical report, College of Science Computer, University of Utah, Salt Lake City, UT 84112, USA, August 5, 1980. 314 pp.

Beebe:1980:IGP

[164] Nelson H. F. Beebe. Installation guide for <PLOT79>. Technical report, College of Science Computer, University of Utah, Salt Lake City, UT 84112, USA, August 15, 1980. 48 pp.

Beebe:1980:CPR

[165] Nelson H. F. Beebe. Corrections to <PLOT79> release 1.5. Technical report, College of Science Computer, University of Utah, Salt Lake City, UT 84112, USA, August 18, 1980. 17 pp.

Beebe:1980:CPV

[166] Nelson H. F. Beebe. Corrections to PRETTY: Version 3.0. Technical report, College of Science Computer, University of Utah, Salt Lake City, UT 84112, USA, August 18, 1980. 31 pp.

Beebe:1980:HRD

[167] Nelson H. F. Beebe. Header record dump of <PLOT79> distribution tape. Technical report, College of Science Computer, University of Utah, Salt Lake City, UT 84112, USA, August 18, 1980. 58 pp.

Beebe:1980:HFT

[168] Nelson H. F. Beebe. On the Hellmann– Feynman theorem. Technical report, College of Science Computer, University of Utah, Salt Lake City, UT 84112, USA, August 27, 1980. 3 pp.

Beebe:1980:VT

[169] Nelson H. F. Beebe. On the virial theorem. Technical report, College of Science Computer, University of Utah, Salt Lake City, UT 84112, USA, August 27, 1980. 7 pp.

Beebe:1980:MES

[170] Nelson H. F. Beebe. Matrix elements with Slater determinants. In Harris et al. [1787], page 10. Two volumes.

Beebe:1980:TAW

[171] Nelson H. F. Beebe. Table of atomic L-S wavefunctions. Technical report, College of Science Computer, University of Utah, Salt Lake City, UT 84112, USA, September 9, 1980. 14 pp.

Beebe:1980:MCI

[172] Nelson H. F. Beebe. The method of configuration interaction. Technical report, College of Science Computer, University of Utah, Salt Lake City, UT 84112, USA, September 21, 1980. 26 pp.

Beebe:1980:CGI

[173] Nelson H. F. Beebe. The computation of F (m, t) in Gaussian integral programs: Version 1.5. Technical report, College of Science Computer, University of Utah, Salt Lake City, UT 84112, USA, October 31, 1980. 65 pp.

Beebe:1980:BBS

[174] Nelson H. F. Beebe. Bibliography of basis sets in quantum chemistry. Technical report, College of Science Computer, University of Utah, Salt Lake City, UT 84112, USA, November 3, 1980. 6 pp.

Beebe:1980:PDP

[175] Nelson H. F. Beebe. <PLOT79> demonstration programs (2 volumes). Technical report, College of Science Computer, University of Utah, Salt Lake City, UT 84112, USA, November 3, 1980. 576 pp.

Beebe:plot79-demos

[176] Nelson H. F. Beebe. <PLOT79> demonstration programs (2 volumes):

Version 1.3. Technical report, College of Science Computer, University of Utah, Salt Lake City, UT 84112, USA, November 3, 1980. 576 pp.

Beebe:1980:RMD

[177] Nelson H. F. Beebe. Remarks on MOLPRO: A descendant of the POLY-ATOM properties package: Version 1.2. Technical report, College of Science Computer, University of Utah, Salt Lake City, UT 84112, USA, November 5, 1980. 12 pp.

Beebe:1980:CRP

[178] Nelson H. F. Beebe. Corrections to release 1.5 of <PLOT79>. Technical report, College of Science Computer, University of Utah, Salt Lake City, UT 84112, USA, December 3, 1980. 47 pp.

Beebe:1980:PNU

[179] Nelson H. F. Beebe. Preface to the notes for the Utah Workshop on Quantum Chemistry. In Harris et al. [1787], page 1. Two volumes.

Beebe:1980:BSQ

[180] Nelson H. F. Beebe. Basis sets in quantum chemistry. In Harris et al. [1787], pages 1–25. Two volumes.

Beebe:1980:HFA

[181] Nelson H. F. Beebe. The Hartree–Fock approximation. In Harris et al. [1787], pages 26–49. Two volumes.

Beebe:1980:CMP

[182] Nelson H. F. Beebe. The calculation of molecular properties. In Harris et al. [1787], pages 150–178. Two volumes.

Beebe:1980:MEA

[183] Nelson H. F. Beebe. Multiplet energies of atoms. In Harris et al. [1787], pages 561–583. Two volumes.

Beebe:1980:PSS

[184] Nelson H. F. Beebe. Proposal for a standard set of primitives for machineindependent bit manipulation in Fortran. Technical report, College of Science Computer, University of Utah, Salt Lake City, UT 84112, USA, 1980. This document is included in the <PLOT79> distribution for online access.

Beebe:1981:CSPa

[185] Nelson H. F. Beebe. Correction set #2 for <PLOT79> release 1.5 [18-Aug-80]. Technical report, College of Science Computer, University of Utah, Salt Lake City, UT 84112, USA, February 14, 1981. 8 pp.

Beebe:1981:NRP

[186] Nelson H. F. Beebe. New release of PRETTY, and future distribution. Technical report, College of Science Computer, University of Utah, Salt Lake City, UT 84112, USA, February 23, 1981. 2 pp.

Beebe:1981:CSPb

[187] Nelson H. F. Beebe. Correction set #3 for <PLOT79> release 1.5 [18-Aug-80]. Technical report, College of Science Computer, University of Utah, Salt Lake City, UT 84112, USA, March 27, 1981. 35 pp.

Beebe:1981:ICF

[188] Nelson H. F. Beebe. Ideas for a combined FORTRAN 77/SFTRAN3 prettyprinter. Technical report, College of Science Computer, University of Utah, Salt Lake City, UT 84112, USA, March 30, 1981. 2 pp.

Beebe:1981:IPM

[189] Nelson H. F. Beebe. Ideas for pattern matching software. Technical report, College of Science Computer, University of Utah, Salt Lake City, UT 84112, USA, May 12, 1981. 3 pp.

Beebe:1981:CSC

[190] Nelson H. F. Beebe. Chemical shift computations: Version 1.0. Technical report, College of Science Computer, University of Utah, Salt Lake City, UT 84112, USA, June 3, 1981. 9 pp.

Beebe:1981:GBB

[191] Nelson H. F. Beebe. GRAPH: a blackbox drawing utility. Technical report, College of Science Computer, University of Utah, Salt Lake City, UT 84112, USA, June 11, 1981. 18 pp.

Beebe:1981:NSV

[192] Nelson H. F. Beebe. Notes on spheres: Version 1.0. Technical report, College of Science Computer, University of Utah, Salt Lake City, UT 84112, USA, June 25, 1981. 3 pp.

Beebe:1981:CSPc

[193] Nelson H. F. Beebe. Correction set #4 for <PLOT79> release 1.5 [18-Aug-80]. Technical report, College of Science Computer, University of Utah, Salt Lake City, UT 84112, USA, July 9, 1981. 47 pp.

Beebe:1981:EUG

[194] Nelson H. F. Beebe. Errata in A User’s Guide to <PLOT79>. Technical report, College of Science Computer, University of Utah, Salt Lake City, UT 84112, USA, August 22, 1981. 4 pp.

Beebe:1982:FP

[195] Nelson H. F. Beebe. Fork package. Technical report, College of Science Computer, University of Utah, Salt Lake City, UT 84112, USA, January 16, 1982. 6 pp.

Beebe:1982:CSPa

[196] Nelson H. F. Beebe. Correction set #5 for <PLOT79> release 1.5 [18-Aug-80]. Technical report, College of Science Computer, University of Utah, Salt Lake City, UT 84112, USA, February 14, 1982. 38 pp.

Motherwell:1982:PPP

[197] Sam Motherwell, Joel L. Sussman, Zappora Shakked, David Barnett, and Nelson H. F. Beebe. PLUTO: A program for plotting molecular and crystal structures. Technical report, University Chemical Laboratory, Lensfield Road, Cambridge CB2 1EW, England; The Weizmann Institute of Science, Rehovoth, Israel; Departments of Physics and Chemistry, University of Utah, Salt Lake City, UT 84112, USA, February 16, 1982. 85 pp.

Beebe:1982:EPV

[198] Nelson H. F. Beebe. The Extended PFORT Verifier. Technical report, College of Science Computer, University of Utah, Salt Lake City, UT 84112, USA, February 1982. 92 pp. This is an authorized addition to Bell Laboratories Computing Science Technical Report #12, The PFORT Verifier, by A. D. Ryder and A. D. Hall.

Beebe:1982:NBS

[199] Nelson H. F. Beebe. A note on Bspline surfaces. Report, College of Science Computer, University of Utah, Salt Lake City, UT 84112, USA, March 23, 1982. 9 pp.

Beebe:1982:DIP

[200] Nelson H. F. Beebe. DOCUMENT and INDEXER: A portable document system. Technical report, College of Science Computer, University of Utah, Salt Lake City, UT 84112, USA, March 26, 1982.

Beebe:1982:PPF

[201] Nelson H. F. Beebe. PRETTY— A portable FORTRAN program prettyprinter, second edition. Technical report, College of Science Computer, University of Utah, Salt Lake City, UT 84112, USA, June 10, 1982. 46 pp.

Beebe:1982:CSPb

[202] Nelson H. F. Beebe. Correction set #6 for <PLOT79> release 1.5 [18-Aug-80]. Technical report, College of Science Computer, University of Utah, Salt Lake City, UT 84112, USA, June 30, 1982. 38 pp.

Beebe:1982:IGP

[203] Nelson H. F. Beebe. Installation guide for <PLOT79>, second edition. Technical report, College of Science Computer, University of Utah, Salt Lake City, UT 84112, USA, September 6, 1982. 79 pp.

Beebe:1982:NCS

[204] Nelson H. F. Beebe. NRCC Conference on Software Standards in Chemistry: A roster of software tools. Technical report, College of Science Computer, University of Utah, Salt Lake City, UT 84112, USA, October 23, 1982. 28 pp.

Lawson:1982:SPR

[205] C. L. Lawson, J. A. Flynn, and Nelson H. F. Beebe. SFTRAN3 programmers reference manual: Version 1.6. Technical report, Jet Propulsion Laboratory andCollege of Science Computer, University of Utah, California Institute of Technology, Pasadena, California, USA andSalt Lake City, UT 84112, USA, October 29, 1982. 57 pp.

Beebe:1982:NSH

[206] Nelson H. F. Beebe. Notes on the Sutherland–Hodgman clipping algorithm. Technical report, College of Science Computer, University of Utah, Salt Lake City, UT 84112, USA, December 2, 1982.

Beebe:1983:CGL

[207] Nelson H. F. Beebe. Computer graphics lecture outline. Technical report, College of Science Computer, University of Utah, Salt Lake City, UT 84112, USA, February 17, 1983. 6 pp.

Beebe:1983:PPP

[208] Nelson H. F. Beebe. A portable parsing package. Technical report, College of Science Computer, University of Utah, Salt Lake City, UT 84112, USA, April 1, 1983. 41 pp.

Beebe:1983:LCD

[209] Nelson H. F. Beebe. Low-cost downloadable font devices. TUGboat, 4(1): 11–12, April 1983. ISSN 0896-3207.

Beebe:1983:SCL

[210] Nelson H. F. Beebe. A survey of computational linear algebra. Technical report, College of Science Computer, University of Utah, Salt Lake City, UT 84112, USA, May 1983. 68 pp.

Beebe:1983:PCS

[211] Nelson H. F. Beebe. <PLOT79> correction set #7. Technical report, College of Science Computer, University of Utah, Salt Lake City, UT 84112, USA, June 28, 1983. 72 pp.

Beebe:1983:DCV

[212] Nelson H. F. Beebe. DEM67 CORE viewing parameters. Technical report, College of Science Computer, University of Utah, Salt Lake City, UT 84112, USA, September 29, 1983. 38 pp.

Beebe:1983:PPS

[213] Nelson H. F. Beebe. <PLOT79> policy statement. Technical report, College of Science Computer, University of Utah, Salt Lake City, UT 84112, USA, October 22, 1983. 4 pp.

Beebe:1983:CPV

[214] Nelson H. F. Beebe. Corrections to PRETTY: Version 3.0: Version 1.2. Technical report, College of Science Computer, University of Utah, Salt Lake City, UT 84112, USA, November 5, 1983. 31 pp.

Beebe:1984:TGT

[215] Nelson H. F. Beebe. TMACRO—A general text macro processor. Technical report, College of Science Computer, University of Utah, Salt Lake City, UT 84112, USA, May 9, 1984. 51 pp.

Beebe:1984:TNV

[216] Nelson H. F. Beebe. Temporary note on VAX VMS 2.00 [PLOT79 master] tape. Technical report, College of Science Computer, University of Utah, Salt Lake City, UT 84112, USA, May 17, 1984. 1 pp.

Beebe:1984:SCG

[217] Nelson H. F. Beebe. A selected computer graphics bibliography. Technical report, College of Science Computer, University of Utah, Salt Lake City, UT 84112, USA, June 20, 1984. 5 pp.

Beebe:1984:SSP

[218] Nelson H. F. Beebe. A short survey of <PLOT79>. Technical report, College of Science Computer, University of Utah, Salt Lake City, UT 84112, USA, June 23, 1984. 9 pp.

Beebe:1984:SIG

[219] Nelson H. F. Beebe. SFTRAN/3 installation guide: Version 1.8. Technical report, College of Science Computer, University of Utah, Salt Lake City, UT 84112, USA, October 31, 1984. 64 pp.

Beebe:1985:PUTa

[220] Nelson H. F. Beebe. The <PLOT79> Unix tar tape: Version 1.1. Technical report, College of Science Computer, University of Utah, Salt Lake City, UT 84112, USA, March 5, 1985. 29 pp.

Beebe:1985:PVVa

[221] Nelson H. F. Beebe. The <PLOT79> VAX VMS BACKUP tape: Version 2.5. Technical report, College of Science Computer, University of Utah, Salt Lake City, UT 84112, USA, March 5, 1985.

Beebe:1985:UST

[222] Nelson H. F. Beebe. The Utah Software Tools Tape. Technical report, College of Science Computer, University of Utah, Salt Lake City, UT 84112, USA, April 2, 1985. 17 pp.

Beebe:1985:SCG

[223] Nelson H. F. Beebe. A selected computer graphics bibliography: Version 1.10. Technical report, College of Science Computer, University of Utah, Salt Lake City, UT 84112, USA, May 3, 1985. 5 pp.

Beebe:1985:PCS

[224] Nelson H. F. Beebe. <PLOT79> correction set #8. Technical report, College of Science Computer, University of Utah, Salt Lake City, UT 84112, USA, May 5, 1985. 79 pp.

Beebe:1985:PVVb

[225] Nelson H. F. Beebe. The <PLOT79> VAX VMS BACKUP tape: VMS release 4.0 [release 2.2]. Technical report, College of Science Computer, University of Utah, Salt Lake City, UT 84112, USA, May 14, 1985.

Beebe:1985:GPC

[226] Nelson H. F. Beebe. Glossary of <PLOT79> and CORE System concepts. Technical report, College of Science Computer, University of Utah, Salt Lake City, UT 84112, USA, May 15, 1985. 17 pp.

Beebe:1985:PSS

[227] Nelson H. F. Beebe. Proposal for a standard set of primitives for machineindependent character manipulation in FORTRAN 77. Technical report, College of Science Computer, University of Utah, Salt Lake City, UT 84112, USA, May 23, 1985. 31 pp.

Beebe:1985:IFP

[228] Nelson H. F. Beebe. IEEE floatingpoint constants. Technical report, College of Science Computer, University of Utah, Salt Lake City, UT 84112, USA, June 17, 1985. 3 pp.

Beebe:1985:PUTb

[229] Nelson H. F. Beebe. The <PLOT79> UNIX tar tape [Release 2.3]. Technical report, College of Science Computer, University of Utah, Salt Lake City, UT 84112, USA, October 2, 1985. 50 pp.

Beebe:1985:IGP

[230] Nelson H. F. Beebe. Installation guide for <PLOT79> [Release 2.3]. Technical report, College of Science Computer, University of Utah, Salt Lake City, UT 84112, USA, October 28, 1985. 114 pp.

Beebe:1986:PUT

[231] Nelson H. F. Beebe. The <PLOT79> UNIX tar tape [Release 2.3a]. Technical report, College of Science Computer, University of Utah, Salt Lake City, UT 84112, USA, January 11, 1986. 58 pp.

Beebe:1986:TIP

[232] Nelson H. F. Beebe. TECO on the IBM PC. Technical report, College of Science Computer, University of Utah, Salt Lake City, UT 84112, USA, May 22, 1986. 62 pp.

Beebe:1986:ULA

[233] Nelson H. F. Beebe and Leslie Lamport. Using LaTEX at the University of Utah College of Science DEC-20. Technical report, College of Science Computer, University of Utah, Salt Lake City, UT 84112, USA, June 12, 1986. 75 pp.

Beebe:1986:ICV

[234] Nelson H. F. Beebe. Introduction to the CSC VAX 8600 VMS 4.4. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, October 15, 1986. 50 pp.

Beebe:1986:PVV

[235] Nelson H. F. Beebe. The <PLOT79> VAX VMS backup tape [Release 2.3c]. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, October 16, 1986. 52 pp.

Beebe:1987:TDDa

[236] Nelson H. F. Beebe. A TEX DVI driver family. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, April 15, 1987. 114 pp.

Beebe:1987:PDT

[237] Nelson H. F. Beebe. Public-domain TEX DVI driver family. TUGboat, 8(1): 41–42, April 1987. ISSN 0896-3207.

Beebe:1987:UTU

[238] Nelson H. F. Beebe. UNIX for TOPS-20 users. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, May 28, 1987. 29 pp. URL

http://www.math.utah.edu/~beebe/
reports/1987/t20unix.pdf.

Beebe:1987:TDDb

[239] Nelson H. F. Beebe. A TEX DVI driver family. TEXniques, 5:71–114, August 1987. Proceedings of the Eighth Annual Meeting of the TEX Users Group.

Beebe:1988:TPU

[240] Nelson H. F. Beebe. Towards a powerful uniform command-line interface. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, October 10, 1988. 11 pp.

Beebe:1988:NSP

[241] Nelson H. F. Beebe. A new SFTRAN3 preprocessor. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, November 1, 1988. 43 pp.

Beebe:1988:FSP

[242] Nelson H. F. Beebe. A Fortran 77 and SFTRAN3 prettyprinter. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, November 8, 1988. 18 pp.

Beebe:1989:STG

[243] N. Beebe. TEX and graphics: the state of the problem. Cahiers GUTenberg, 2:13–53, mai 1989. ISSN 1140-9304 (print), 2118-8254 (electronic). URL http://www.

gutenberg.eu.org/pub/GUTenberg/
publicationsPDF/2-beebe.pdf;   http:
//www.math.utah.edu/~beebe/publications/
1989/cahiers-gutenberg-2-13-mai-
1989.pdf.

Beebe:1989:GSS

[244] Nelson H. F. Beebe and Grant Gustafson. Graphics software survey. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, March 15, 1989. 81 pp.

Beebe:1989:XWS

[245] Nelson H. F. Beebe. X Window System Version 11 Release 3 notes. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, April 1, 1989. 12 pp.

Beebe:1989:FPH

[246] Nelson H. F. Beebe. Funds for purchase of high-performance color graphics equipment. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, May 9, 1989. 11 pp. This is the proposal that obtained funding for high-performance real-time graphics equipment [248].

Beebe:1989:TDD

[247] Nelson H. F. Beebe. TEX DVI driver family status. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, July 1, 1989. 20 pp.

Beebe:1989:CGE

[248] Nelson H. F. Beebe. Computer graphics equipment specifications. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, July 25, 1989. 5 pp. This is the Request for Proposals that eventually led to the selection of a Stardent 1520 system for the Center for Scientific Computing and the Department of Mathematics.

Beebe:1989:MPR

[249] Nelson H. F. Beebe. Multi-processor replacements for the DEC-20/60. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, July 26, 1989. 19 pp.

Beebe:1989:CEM

[250] Nelson H. F. Beebe. Cost estimates for multi-processor replacements for the DEC-20/60. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, August 31, 1989. 23 pp. This report is confidential for the Center for Scientific Computing and the Department of Mathematics.

Beebe:1989:IEG

[251] Nelson H. F. Beebe. Issues in evaluating graphics system bids. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, September 28, 1989. 11 pp. This report is confidential for the Center for Scientific Computing and the Department of Mathematics.

Beebe:1989:MNP

[252] Nelson H. F. Beebe. Message from the new President. TUGboat, 10(3):312– 314, November 1989. ISSN 0896-3207.

Beebe:1989:PCP

[253] Nelson H. F. Beebe and R. P. C. Rodgers. <PLOT79>: a comprehensive portable Fortran scientific line graphics system, as applied to biomedical research. Computers in Biology and Medicine, 19(6):385–402, 1989. CO-DEN CBMDAW. ISSN 0010-4825.

Beebe:1990:GTP

[254] Nelson H. F. Beebe. Greetings from the TUG President. Cahiers GUTenberg, 6:34–38, juillet 1990. ISSN 1140-9304 (print), 2118-8254 (electronic). URL http://www.

                                        gutenberg.eu.org/pub/GUTenberg/
                                        publicationsPDF/6-beebe.pdf;   http:
                                        //www.math.utah.edu/~beebe/publications/
                                        1990/cahiers-gutenberg-6-34-juillet-

1990.pdf. Presented to: Congrès GUTenberg, Toulouse, France, 15–17 May 1990.

Beebe:1990:SUP

[255] Nelson H. F. Beebe. Sun upgrade paths and the DECstation 3100. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, January 25, 1990. 8 pp.

Beebe:1990:DGE

[256] Nelson H. F. Beebe. The design of GMAKE—an extended implementation of UNIX MAKE. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, February 5, 1990. 32 pp.

Beebe:1990:UMG

[257] Nelson H. F. Beebe. A user’s manual for GMAKE. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, February 5, 1990. 78 pp.

Beebe:1990:Pa

[258] Nelson H. F. Beebe. From the President. TUGboat, 11(1):5–6, April 1990. ISSN 0896-3207.

Beebe:1990:CSE

[259] Nelson H. F. Beebe. Character set encoding. TUGboat, 11(2):171–175, June 1990. ISSN 0896-3207.

Beebe:1990:GGL

[260] Nelson H. F. Beebe. GRAPH: A general line graphing package. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, August 9, 1990. 100 pp.

Beebe:1990:PFF

[261] Nelson H. F. Beebe. Parsing Fortran FORMAT statements. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, August 20, 1990. 83 pp.

Beebe:1990:PI

[262] Nelson H. F. Beebe. President’s introduction. TUGboat, 11(3):335–336, September 1990. ISSN 0896-3207.

Beebe:1990:TM

[263] Nelson H. F. Beebe. TEX 3.0 and 2.0. TUGboat, 11(3):442– 443, September 1990. ISSN 0896-3207.

Beebe:1990:BRM

[264] Nelson H. F. Beebe. Background reading for Mathematics 118 and 119. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, November 7, 1990. 36 pp. Supplemental class notes prepared for Mathematics 118 and 119. This report is updated frequently.

Beebe:1990:HPM

[265] Nelson H. F. Beebe. High-performance matrix multiplication. Technical Report 3, Utah Supercomputing Institute, University of Utah, Salt Lake City, UT 84112, USA, November 1990. 27

pp. URL ftp://ftp.math.utah.edu/
pub/benchmarks/usirep.ps.z.

Beebe:1990:Pb

[266] Nelson H. F. Beebe. From the President. TUGboat, 11(4):485–487, November 1990. ISSN 0896-3207.

Beebe:1990:CFT

[267] Nelson H. F. Beebe. Comments on the future of TEX and . TUGboat, 11(4):490–494, November 1990. ISSN 0896-3207.

Beebe:1990:MMB

[268] Nelson H. F. Beebe. Matrix multiply benchmarks. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA,

1990. 317 pp. URL ftp://ftp.math.
utah.edu/pub/benchmarks/matrix-

multiply.*. This report is updated frequently.

Beebe:1991:ASR

[269] Nelson H. F. Beebe. Accurate square root computation. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, February 4, 1991. 23 pp. Supplemental class notes prepared for Mathematics 118.

Beebe:1991:AHT

[270] Nelson H. F. Beebe. Accurate hyperbolic tangent computation. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, February 7, 1991. 13 pp. Supplemental class notes prepared for Mathematics 119.

Beebe:1991:KKB

[271] Nelson H. F. Beebe. Komentár k budoucnosti TEXu a u. TEX Bulletin: Československé sdružení uživatelů TEXu: CS TUG, 1(1):3–

10, 1991. URL http://bulletin.
cstug.cz/pdf/bul911.pdf;   http:
//pavel.striz.cz/Zpravodaj/1991-

1/003.php. Czech translation of [267].

Beebe:1991:SF

[272] Nelson H. F. Beebe. A summary of Fortran. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, March 13, 1991. 61 pp. Supplemental class notes prepared for Mathematics 118 and 119.

Beebe:1991:PAT

[273] Nelson H. F. Beebe. Publications about TEX and typography. TUGboat, 12(12): 176–182, May 1991. ISSN 0896-3207.

Beebe:1991:PPT

[274] Nelson H. F. Beebe. Publications prepared with TEX. TUGboat, 12(12):183– 194, May 1991. ISSN 0896-3207.

Beebe:1991:PTT

[275] Nelson H. F. Beebe. A proposal for TEX \special commands and DVI driver paper specification. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, June 4, 1991. 12 pp. To be published in TUGboat (1992).

Beebe:1991:PIa

[276] Nelson H. F. Beebe. President’s introduction. TUGboat, 12(2):205–208, June 1991. ISSN 0896-3207.

Beebe:1991:CBA

[277] Nelson H. F. Beebe. A complete bibliography of ACM Transactions on Mathematical Software. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, August 17, 1991. 90 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-t.html#toms. This report is up-

dated frequently.

Beebe:1991:LESb

[278] Nelson H. F. Beebe. LaTEX editing support. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, October 7, 1991. 28 pp. To be published in TUGboat (1992).

Beebe:1991:LESa

[279] Nelson H. F. Beebe. LaTEX editing support. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, October 7, 1991. 69 pp. This is an on-line Emacs INFO version of part of [278].

Beebe:1991:BTD

[280] Nelson H. F. Beebe. A bibliography of tree drawing algorithms. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, October 27, 1991. 2 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-t.html#trees. This report is

updated frequently.

Beebe:1991:BBA

[281] Nelson H. F. Beebe. A bibliography of books about REDUCE. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, October 27, 1991. 4 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-r.html#redbooks. This report

is updated frequently.

Beebe:1991:BEP

[282] Nelson H. F. Beebe. A bibliography of electronic publishing. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, October 27, 1991. 6 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-

table-e.html#ep. This report is updated frequently.

Beebe:1991:BTP

[283] Nelson H. F. Beebe. A bibliography of TEXniques Publications. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, October 27, 1991. 6 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-t.html#texnique. This report

is updated frequently.

Beebe:1991:BJA

[284] Nelson H. F. Beebe. A bibliography of journals accepting manuscripts written using TEX. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, October 27, 1991. 8 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-t.html#texjourn. This report

is updated frequently.

Beebe:1991:BPL

[285] Nelson H. F. Beebe. A bibliography of PostScript literature. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, October 27, 1991. 10 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-p.html#postscri. This report

is updated frequently.

Beebe:1991:CBJ

[286] Nelson H. F. Beebe. A complete bibliography of the Journal of C Language Translation. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, October 27, 1991. 11 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-j.html#jclt. This journal is

no longer published.

Beebe:1991:BLP

[287] Nelson H. F. Beebe. A bibliography of literate programming. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, October 27, 1991. 12 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-l.html#litprog. This report is

updated frequently.

Beebe:1991:BGT

[288] Nelson H. F. Beebe. A bibliography of graphics in TEX. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, October 27, 1991. 14 pp. URL http:

//www.math.utah.edu/pub/tex/bib/

                                        index-table-t.html#texgraph. This

report is updated frequently.

Beebe:1991:CBC

[289] Nelson H. F. Beebe. A complete bibliography of Computing Systems Journal. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, October 27, 1991. 17 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-c.html#compsys. This report is

updated frequently.

Beebe:1991:BEE

[290] Nelson H. F. Beebe. A complete bibliography of Electronic Publishing— Origin, Dissemination, and Design. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, October 27, 1991. 20 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-e.html#epodd. This report is

updated frequently.

Beebe:1991:CBE

[291] Nelson H. F. Beebe. A complete bibliography of Electronic Publishing— Origin, Dissemination, and Design. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, October 27, 1991. 20 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-e.html#epodd. This report is

updated frequently.

Beebe:1991:BCG

[292] Nelson H. F. Beebe. A bibliography of Computer Graphics Forum. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, October 27, 1991. 53 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-c.html#cgf. This report is up-

dated frequently.

Beebe:1991:BUT

[293] Nelson H. F. Beebe. A bibliography of the use of TEX. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, October 27, 1991. 68 pp. URL http:

//www.math.utah.edu/pub/tex/bib/
index-table-t.html#texbook2. This

report is updated frequently.

Beebe:1991:BPA

[294] Nelson H. F. Beebe. A bibliography of publications about TEX. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, October 27, 1991. 69 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-t.html#texbook1;   http:/
/www.math.utah.edu/pub/tex/bib/
index-table-t.html#texbook3. This

report is updated frequently.

Beebe:1991:BNM

[295] Nelson H. F. Beebe. A bibliography of NHFB’s master collection. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, October 27, 1991. 101 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-m.html#master. This report is

updated frequently.

Beebe:1991:MMB

[296] Nelson H. F. Beebe. Matrix multiply benchmarks. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, November 4, 1991. 317

                                        pp. URL ftp://ftp.math.utah.edu/
                                        pub/benchmarks/matrix-multiply.*

. This report is updated frequently.

Beebe:1991:PIb

[297] Nelson H. F. Beebe. President’s introduction. TUGboat, 12(3+4):351–352, November 1991. ISSN 0896-3207.

Beebe:1991:TIL

[298] Nelson H. F. Beebe. Tips on indexing LaTEX documents. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, December 10, 1991. 12 pp.

Beebe:1991:TEP

[299] Nelson H. F. Beebe. TEX and Encapsulated PostScript. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, December 10, 1991. 44 pp.

Beebe:1991:SFH

[300] Nelson H. F. Beebe. Standard file headers. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, December 15, 1991. 58 pp.

Beebe:1991:BET

[301] Nelson H. F. Beebe. A bibliography of TEXniques Publications. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, 1991. 4

pp. URL http://www.math.utah.
edu/pub/tex/bib/texniques.*. This

report is updated frequently.

Beebe:1991:BEC

[302] Nelson H. F. Beebe. A bibliography of Computer Graphics Forum. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, 1991. 21

pp. URL http://www.math.utah.
edu/pub/tex/bib/cgf.*. This report

is updated frequently.

Beebe:1991:ETB

[303] Nelson H. F. Beebe. A TUGboat bibliography. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, 1991.

54 pp. URL http://www.math.utah.
edu/pub/tex/bib/tugboat.*. This

report is updated frequently.

Beebe:1992:CBX

[304] Nelson H. F. Beebe. A complete bibliography of The X Resource: A practical journal of the X Window System. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, January 14, 1992. 26 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-x.html#xres. This report is up-

dated frequently.

Beebe:1992:IFPa

[305] Nelson H. F. Beebe. IEEE 754 floatingpoint arithmetic. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, January 17, 1992. 7 pp. Supplemental class notes prepared for Mathematics 118 and 119.

Beebe:1992:IA

[306] Nelson H. F. Beebe. Integer arithmetic. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, January 17, 1992. 8 pp. Supplemental class notes prepared for Mathematics 118 and 119.

Beebe:1992:CGB

[307] Nelson H. F. Beebe. A Cahiers GUTenberg bibliography. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, February 12, 1992. 12 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-g.html#gut. This report is up-

dated frequently.

Beebe:1992:CME

[308] Nelson H. F. Beebe. Computing the machine epsilon. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, February 12, 1992. 20 pp. Supplemental class notes prepared for Mathematics 118 and 119.

Beebe:1992:BPS

[309] Nelson H. F. Beebe. A bibliography of publications on software standards. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, February 26, 1992. 5 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-s.html#standard. This report

is updated frequently.

Beebe:1992:IFPb

[310] Nelson H. F. Beebe. IEEE 754 floatingpoint support. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, February 1992. ii + 41 pp. Supplemental class notes prepared for Mathematics 118 and 119.

Beebe:1992:BNA

[311] Nelson H. F. Beebe. A bibliography of NATO Advanced Study Institute Proceedings. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, March 26, 1992. 10 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-a.html#asi. This report is up-

dated frequently.

Beebe:1992:UMT

[312] Nelson H. F. Beebe. Using make for TEX documents. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, April 30, 1992. 14 pp.

Beebe:1992:PC

[313] Nelson H. F. Beebe. President’s column. TUGboat, 13(1):6–10, April 1992. ISSN 0896-3207.

Beebe:1992:NNQ

[314] Nelson H. F. Beebe. Notes on numerical quadrature. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, May 28, 1992. ii + 19 pp. Supplemental class notes prepared for Mathematics 118 and 119.

Beebe:1992:BCA

[315] Nelson H. F. Beebe. A bibliography of the Collected Algorithms of the ACM. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, July 21, 1992. 59 pp. This report is updated frequently.

Beebe:1992:BAS

[316] Nelson H. F. Beebe. A bibliography of Applied Statistics and Journal of Applied Statistics. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, July 21, 1992. 60 pp. URL http:

                                        //www.math.utah.edu/pub/tex/bib/
                                        index-table-a.html#as. This report

is updated frequently.

Beebe:1993:BTK

[317] Nelson H. F. Beebe and Peter Schmitt. A bibliography of Die TEXnische Komödie. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, March 12, 1993. 4 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-k.html#komoedie. This report

is updated frequently.

Beebe:1993:BBAa

[318] Nelson H. F. Beebe. A bibliography of books about the Internet and networking: 1969–1999. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, April 5, 1993. 80 pp. URL http:

//www.math.utah.edu/pub/tex/bib/
index-table-i.html#internet. This

report is updated frequently.

Beebe:1993:TME

[319] Nelson H. F. Beebe. TEX macros for Encapsulated PostScript. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, July 21, 1993. 44 pp.

Beebe:1993:IEP

[320] Nelson H. F. Beebe. Incorporating Encapsulated PostScript in TEX and LaTEX documents. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, August 30, 1993. 5 pp.

Beebe:1993:INE

[321] Nelson H. F. Beebe. Ideas for NTS—an extended TEX implementation. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, August 30, 1993. 17 pp.

Beebe:1993:STP

[322] Nelson H. F. Beebe. A simple TEXpackage-independent picture environment. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, August 30, 1993. 34 pp.

Beebe:1993:LES

[323] Nelson H. F. Beebe. LaTEX editing support, version 1.10. Report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, September 14, 1993. iii + 65

pp. URL ftp://ftp.math.utah.edu/
pub/emacs/ltxmode/.

Beebe:1993:BBAb

[324] Nelson H. F. Beebe. A bibliography of books about UNIX and UNIX programming. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, September 30, 1993. 54 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-u.html#unix. This report is up-

dated frequently.

Beebe:1993:TB

[325] Nelson H. F. Beebe. A TUGboat bibliography. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, September 30, 1993. 102 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-t.html#tugboat. This report is

updated frequently.

Beebe:1993:BPSa

[326] Nelson H. F. Beebe. Bibliography prettyprinting and syntax checking. TUGboat, 14(3):222, October 1993. ISSN 0896-3207. URL

http://www.tug.org/TUGboat/tb14-
3/tb40beebe.pdf.

Beebe:1993:BPSb

[327] Nelson H. F. Beebe. Bibliography prettyprinting and syntax checking. TUGboat, 14(4):395–419, December 1993. ISSN 0896-3207. URL

http://www.tug.org/TUGboat/tb14-
4/tb41beebe.pdf.

Beebe:1994:IMA

[328] Nelson H. F. Beebe. The impact of memory and architecture on computer performance. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, February 23, 1994. viii + 62

pp. URL http://www.math.utah.
edu/~beebe/memperf.pdf. Supple-

mental class notes prepared for Mathematics 118 and 119.

Beebe:1994:BPAa

[329] Nelson H. F. Beebe. A bibliography of publications about the Maple symbolic algebra language. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, March 5, 1994. 39 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-m.html#maple-extract. This

report is updated frequently.

Beebe:1994:BPMa

[330] Nelson H. F. Beebe. A bibliography of publications about the Mathematica symbolic algebra language. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, June 17, 1994. 14 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-m.html#mathematica. This re-

port is updated frequently.

Beebe:1994:BPMb

[331] Nelson H. F. Beebe. A bibliography of publications about the Matlab online matrix laboratory. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, July 7, 1994. 18 pp. URL http:

//www.math.utah.edu/pub/tex/bib/
index-table-m.html#matlab. This

report is updated frequently.

Beebe:1994:BPN

[332] Nelson H. F. Beebe. A bibliography of publications of Nelson H. F. Beebe. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, July 11, 1994. 19

pp. URL http://www.math.utah.
edu/pub/bibnet/authors/b/beebe-

nelson-h-f.*. This report is updated frequently.

Beebe:1994:BPH

[333] Nelson H. F. Beebe. A bibliography of publications on hashing algorithms. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, July 22, 1994. 126 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-h.html#hash. This report is up-

dated frequently.

Beebe:1994:BPAb

[334] Nelson H. F. Beebe. A bibliography of publications about the S and S-Plus statistics programming languages. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, August 11, 1994. 2 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-s.html#s-plus. This report is

updated frequently.

Beebe:1994:BPAc

[335] Nelson H. F. Beebe. A bibliography of publications about the SAS Statistical Analysis System. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, August 11, 1994. 37 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-s.html#sas. This report is up-

dated frequently.

Beebe:1994:BPAd

[336] Nelson H. F. Beebe. A bibliography of publications in ACM SIGNUM. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, August 17, 1994. 14 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-

                                        table-s.html#signum. This report is

updated frequently.

Beebe:1994:BPAe

[337] Nelson H. F. Beebe. A bibliography of publications in ACM SIGSAM Bulletin. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, August 17, 1994. 28 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-s.html#sigsam. This report is

updated frequently.

Beebe:1994:BPAf

[338] Nelson H. F. Beebe. A bibliography of publications in ACM SIGAda Ada Letters. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, August 17, 1994. 44 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-s.html#sigada. This report is

updated frequently.

Beebe:1994:BPSa

[339] Nelson H. F. Beebe. A bibliography of publications in Software—Practice and Experience. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, August 17, 1994. 83 pp. URL http:

                                        //www.math.utah.edu/pub/tex/bib/

index.html#spe. This report is updated frequently.

Beebe:1994:BPIa

[340] Nelson H. F. Beebe. A bibliography of publications in Information Processing Letters. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, August 17, 1994. 172 pp. URL http:

//www.math.utah.edu/pub/tex/bib/
index-table-i.html#infoproc. This

report is updated frequently.

Beebe:1994:BCAb

[341] Nelson H. F. Beebe. A bibliography of the Communications of the ACM. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, August 17, 1994.

363 pp. URL http://www.math.utah.
edu/pub/tex/bib/index.html#cacm.

This report is updated frequently.

Beebe:1994:BPD

[342] Nelson H. F. Beebe. A bibliography of publications of Donald E. Knuth. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, August 18, 1994.

59 pp. URL http://www.math.utah.
edu/pub/bibnet/authors/k/knuth-

donald-e.*. This report is updated frequently.

Beebe:1994:BPIc

[343] Nelson H. F. Beebe. A bibliography of publications in the IBM Systems Journal. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, August 24,

1994. 38 pp. URL http://www.math.
utah.edu/pub/tex/bib/index.html#

ibmsysj. This report is updated frequently.

Beebe:1994:BPIb

[344] Nelson H. F. Beebe. A bibliography of publications in the IBM Journal of Research and Development. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, August 24, 1994. 83 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-i.html#ibmjrd. This report is

updated frequently.

Beebe:1994:BAT

[345] Nelson H. F. Beebe. A bibliography of ACM Transactions on Graphics. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, August 27, 1994. 30 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-t.html#tog. This report is up-

dated frequently.

Juffa:1994:BPF

[346] Norbert Juffa and Nelson H. F. Beebe. A bibliography of publications on floating-point arithmetic. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, September 6, 1994. 217 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-f.html#fparith. This report is

updated frequently.

Beebe:1994:BPW

[347] Nelson H. F. Beebe. A bibliography of publications of W. J. Cody. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, September 7, 1994. 13 pp. URL http:

//www.math.utah.edu/pub/bibnet/
authors/c/cody-william-j.*. This

report is updated frequently.

Beebe:1994:BPAi

[348] Nelson H. F. Beebe. A bibliography of publications about the SPSS (Statistical Package for the Social Sciences) system. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, October 9, 1994. 50 pp. URL http://www.

math.utah.edu/pub/tex/bib/index.

html#spss. This report is updated frequently.

Beebe:1994:BPG

[349] Nelson H. F. Beebe and Stefano Foresti. A bibliography of publications of Gene H. Golub. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, October 30, 1994. 30

pp. URL http://www.math.utah.
edu/pub/bibnet/authors/g/golub-

gene-h.*. This report is updated frequently.

Beebe:1994:CBJ

[350] Nelson H. F. Beebe. A complete bibliography of the Journal of the ACM. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, October 24, 1994. 214 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-

                                        table-j.html#jacm. This report is up-

dated frequently.

Beebe:1994:XWS

[351] Nelson H. F. Beebe. X Window System benchmarks. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, 1994. 13 pp. This report is updated frequently.

Beebe:1994:BAA

[352] Nelson H. F. Beebe. A bibliography of the algorithms from Applied Statistics. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA,

                                        1994. URL http://www.math.utah.
                                        edu/pub/tex/bib/as.*. This report is

updated frequently.

Beebe:1994:BBAa

[353] Nelson H. F. Beebe. A bibliography of books about the Internet and networking. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA,

                                        1994. URL http://www.math.utah.
                                        edu/pub/tex/bib/internet.*. This

report is updated frequently.

Beebe:1994:BBAb

[354] Nelson H. F. Beebe. A bibliography of books about UNIX and UNIX programming. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA,

                                        1994. URL http://www.math.utah.

edu/pub/tex/bib/unix.*. This re-

port is updated frequently.

Beebe:1994:BCAa

[355] Nelson H. F. Beebe. A bibliography of the Collected Algorithms of the ACM. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA,

1994. URL http://www.math.utah.
edu/pub/tex/bib/acm.*. This report

is updated frequently.

Beebe:1994:BNA

[356] Nelson H. F. Beebe. A bibliography of NATO Advanced Study Institute Proceedings. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA,

1994. URL http://www.math.utah.
edu/pub/tex/bib/asi.*. This report

is updated frequently.

Beebe:1994:BPSb

[357] Nelson H. F. Beebe. A bibliography of publications on software standards. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA,

1994. URL http://www.math.utah.
edu/pub/tex/bib/standard.*. This

report is updated frequently.

Beebe:1994:BTK

[358] Nelson H. F. Beebe and Peter Schmitt. A bibliography of Die TEXnische Komödie. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA,

1994. URL http://www.math.utah.
edu/pub/tex/bib/komoedie.*. This

report is updated frequently.

Dongarra:1995:BPJ

[359] Jack J. Dongarra, Nelson H. F. Beebe, and Stefano Foresti. A bibliography of publications of Jack J. Dongarra. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, January 15, 1995. 24 pp. URL http:

//www.math.utah.edu/pub/bibnet/
authors/d/dongarra-jack-j.*. This

report is updated frequently.

Beebe:1995:BPC

[360] Nelson H. F. Beebe. A bibliography of publications on computer-based spelling error detection and correction. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, July 23, 1995. 20 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-s.html#spell. This report is

updated frequently.

Beebe:1995:BPAe

[361] Nelson H. F. Beebe. A bibliography of publications about SGML, the Standard Generalized Markup Language: 1981–1999. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, August 11, 1995. 36 pp. URL http:

//www.math.utah.edu/pub/tex/bib/

index.html#sgml. This report is updated frequently.

Beebe:1995:BPBa

[362] Nelson H. F. Beebe. A bibliography of publications on BMDP. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, August 17, 1995. 9 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-b.html#bmdp. This report is up-

dated frequently.

Beebe:1995:CBA

[363] Nelson H. F. Beebe. A complete bibliography of ACM Computing Surveys. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, August 17, 1995. 51 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-c.html#compsurv. This report

is updated frequently.

Beebe:1995:BPS

[364] Nelson H. F. Beebe. A bibliography of publications in ACM SIGPLAN Notices. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, August 17, 1995. 171 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-s.html#sigplan. This report is

updated frequently.

Beebe:1995:BPU

[365] Nelson H. F. Beebe. A bibliography of publications on Unicode. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, August 24, 1995. 6 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-u.html#unicode. This report is

updated frequently.

Beebe:1995:BOA

[366] Nelson H. F. Beebe. A bibliography of O’Reilly and Associates Publishers. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, September 14, 1995. 35 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-o.html#ora. This report is up-

dated frequently.

Beebe:1995:CBS

[367] Nelson H. F. Beebe. A complete bibliography of Serif: The Magazine of Type and Typography. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, November 14, 1995. 6 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-s.html#serif. This report is

updated frequently.

Beebe:1995:BPAa

[368] Nelson H. F. Beebe. A bibliography of publications about the Java Programming Language. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, December 17, 1995. 12 pp. URL http:

//www.math.utah.edu/pub/tex/bib/

index.html#java. This report is updated frequently.

Beebe:1995:BPAb

[369] Nelson H. F. Beebe. A bibliography of publications about the Java Programming Language. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, December 17, 1995. 12 pp. URL http:

//www.math.utah.edu/pub/tex/bib/
index-table-j.html#java. This re-

port is updated frequently.

Beebe:1995:BPAc

[370] Nelson H. F. Beebe and Jon Rokne. A bibliography of publications about interval arithmetic. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, December 18, 1995. 222 pp.

URL http://www.math.utah.edu/
pub/tex/bib/index-table-i.html#

intarith. This report is updated frequently.

Beebe:1995:BPBb

[371] Nelson H. F. Beebe. A bibliography of publications in Byte Magazine. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, December 27, 1995.

322 pp. URL http://www.math.utah.
edu/pub/tex/bib/index.html#byte.

This report is updated frequently.

Beebe:1995:BPAd

[372] Nelson H. F. Beebe. A bibliography of publications about the AXIOM (formerly, Scratchpad) symbolic algebra language. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, December 30, 1995. 9 pp. URL http://www.

math.utah.edu/pub/tex/bib/index.

html#axiom. This report is updated frequently.

Beebe:1996:BPDa

[373] Nelson H. F. Beebe. A bibliography of publications in Dr. Dobb’s Journal: Software Tools for the Professional Programmer. formerly: Dr. Dobb’s Journal of Computer Calisthenics and Orthodontia. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, January 3, 1996. 77 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-d.html#dr-dobbs. This report

is updated frequently.

Beebe:1996:BPDb

[374] Nelson H. F. Beebe. A bibliography of publications in Digital Technical Journal. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, January 4, 1996. 28 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-d.html#dectechj. This report

is updated frequently.

Beebe:1996:BPIa

[375] Nelson H. F. Beebe. A bibliography of publications in The International Journal of Quantum Chemistry: 1967– 1999. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, January 4, 1996. 195 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-i.html#ijqc.

Beebe:1996:BPCa

[376] Nelson H. F. Beebe. A bibliography of publications in Computer Graphics World. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, January 5, 1996. 51 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-c.html#cgworld. This report is

updated frequently.

Briggs:1996:CBP

[377] Preston Briggs and Nelson H. F. Beebe. A complete bibliography of publications in ACM Transactions on Programming Languages and Systems (TOPLAS). Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, January 6, 1996. 68 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-t.html#toplas. This report is

updated frequently.

Beebe:1996:BPIb

[378] Nelson H. F. Beebe. A bibliography of publications in IEEE Computer Graphics and Applications. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, January 7, 1996. 91 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-i.html#ieeecga. This report is

updated frequently.

Beebe:1996:BPSa

[379] Nelson H. F. Beebe. A bibliography of publications in Supercomputing Review. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, January 8, 1996. 16 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-s.html#superrev. This report

is updated frequently.

Beebe:1996:BPIc

[380] Nelson H. F. Beebe. A bibliography of publications in International Journal of Supercomputer Applications and International Journal of Supercomputer Applications and High-Performance Computing. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, January 8, 1996. 20 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-i.html#ijsa. This report is up-

dated frequently.

Beebe:1996:BPId

[381] Nelson H. F. Beebe. A bibliography of publications in IEEE Software. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, January 10, 1996. 53 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-i.html#ieeesoft. This report

is updated frequently.

Beebe:1996:BPAa

[382] Nelson H. F. Beebe. A bibliography of publications about PVM (Parallel Virtual Machine) and MPI (Message Passing Interface). Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, January 16, 1996. 17 pp. URL http:

//www.math.utah.edu/pub/tex/bib/
index-table-p.html#pvm. This re-

port is updated frequently.

Beebe:1996:BPCb

[383] Nelson H. F. Beebe. A bibliography of publications in Computer Language magazine. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, January 23, 1996. 28 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-c.html#complang.

Beebe:1996:BPM

[384] Nelson H. F. Beebe. A bibliography of publications in The Microsoft Systems Journal. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, January 26, 1996. 27 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-m.html#microsys.

Beebe:1996:BPUa

[385] Nelson H. F. Beebe. A bibliography of publications in UnixWorld’s Open Computing and Open Computing magazines. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, January 26, 1996. 29 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-o.html#opencomp.

Beebe:1996:BPO

[386] Nelson H. F. Beebe. A bibliography of publications in Open Systems Today magazine. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, January 26, 1996. 44 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-o.html#opensyst.

Beebe:1996:BPAb

[387] Nelson H. F. Beebe. A bibliography of publications about Datamation magazine. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, January 27, 1996. 109 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-d.html#datamation.

Beebe:1996:BPAc

[388] Nelson H. F. Beebe. A bibliography of publications about the Fortran programming language: Part 1: 1956– 1980. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, February 3, 1996. 247 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-f.html#fortran1.

Beebe:1996:BPAd

[389] Nelson H. F. Beebe. A bibliography of publications about the Fortran programming language: Part 2: 1981– date. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, February 3, 1996. 303 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-f.html#fortran2.

Beebe:1996:BPCc

[390] Nelson H. F. Beebe. A bibliography of publications of Cleve B. Moler. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, February 4, 1996.

6 pp. URL http://www.math.utah.
edu/pub/bibnet/authors/m/moler-
cleve-b.*.

Beebe:1996:BPAe

[391] Nelson H. F. Beebe. A bibliography of publications in ACM SIG-GRAPH Computer Graphics: 1972– 1999. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, February 6, 1996. 127 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-s.html#siggraph.

Beebe:1996:BPJa

[392] Nelson H. F. Beebe. A bibliography of publications in the journals Computer Graphics and Image Processing (1972–1982), Computer Vision, Graphics, and Image Processing (1983–1990), Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing (1991– 1994), Computer Vision, Graphics, and Image Processing. Image Understanding (1991–1994), Computer Vision and Image Understanding: CVIU (1995– date) and Graphical Models and Image Processing: GMIP (1995–date). Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, February 7, 1996. 127 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-c.html#cvgip.

Beebe:1996:BPJb

[393] Nelson H. F. Beebe. A bibliography of publications in the Journal of Symbolic Computation. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, February 8, 1996. 46 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-j.html#jsymcomp.

Beebe:1996:BPAf

[394] Nelson H. F. Beebe. A bibliography of publications about the MACSYMA and VAXIMA symbolic algebra languages. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, February 9, 1996. 26 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-m.html#macsyma.

Beebe:1996:SBP

[395] Nelson H. F. Beebe. A supplemental bibliography of publications about the Reduce symbolic algebra language. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, February 12, 1996. 31 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-r.html#redextra.

Beebe:1996:CBL

[396] Nelson H. F. Beebe. A complete bibliography of Lecture Notes in Computer Science (1973–date). Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, February 14, 1996. 210 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-l.html#lncs.

Beebe:1996:BPCd

[397] Nelson H. F. Beebe. A bibliography of publications in The Computer Journal. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, February 19, 1996. 76 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-c.html#compj.

Beebe:1996:BPUb

[398] Nelson H. F. Beebe. A bibliography of publications of the USENIX Association. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, February 19, 1996. 294 pp.

Beebe:1996:BPAg

[399] Nelson H. F. Beebe. A bibliography of publications about typographic fonts. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, February 28, 1996. 66 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-f.html#font.

Beebe:1996:CBPa

[400] Nelson H. F. Beebe and Keith Sams. A bibliography of publications about the Data Explorer scientific visualization system. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, March 13, 1996. 2 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-d.html#data-explorer. This

report is updated frequently.

Beebe:1996:CBPb

[401] Nelson H. F. Beebe. A complete bibliography of publications in the IBM Systems Journal. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, March 20, 1996. 113 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-i.html#ibmsysj. This report is

updated frequently.

Beebe:1996:BPAh

[402] Nelson H. F. Beebe. A bibliography of publications about the slime mold Dictyostelium discoideum. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, April 20, 1996. 361 pp. This report is updated frequently.

Beebe:1996:BPAi

[403] Nelson H. F. Beebe. A bibliography of publications about the MINIX operating system. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, May 4, 1996. 6 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-

table-m.html#minix. This report is

updated frequently.

Beebe:1996:BPAj

[404] Nelson H. F. Beebe. A bibliography of publications about the Linux operating system. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, May 4, 1996. 11 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-l.html#linux. This report is

updated frequently.

Beebe:1996:BPLa

[405] Nelson H. F. Beebe. A bibliography of papers in Lecture Notes in Computer Science (1990). Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, May 11, 1996. 3 pp. URL http:

//www.math.utah.edu/pub/tex/bib/
index-table-l.html#lncs1990. This

report is updated frequently.

Beebe:1996:BPAk

[406] Nelson H. F. Beebe. A bibliography of publications about High-Performance Fortran. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, May 11, 1996. 16 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-h.html#hpfortran. This re-

port is updated frequently.

Beebe:1996:BPLc

[407] Nelson H. F. Beebe. A bibliography of papers in Lecture Notes in Computer Science (1992). Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, May 11, 1996. 199 pp. URL http:

//www.math.utah.edu/pub/tex/bib/
index-table-l.html#lncs1992. This

report is updated frequently.

Beebe:1996:BPLg

[408] Nelson H. F. Beebe. A bibliography of papers in Lecture Notes in Computer Science (1995), Part 2 of 2. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, May 11, 1996. 203 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-l.html#lncs1995b. This re-

port is updated frequently.

Beebe:1996:BPLf

[409] Nelson H. F. Beebe. A bibliography of papers in Lecture Notes in Computer Science (1995), Part 1 of 2. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, May 11, 1996. 206 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-l.html#lncs1995a. This re-

port is updated frequently.

Beebe:1996:BPLb

[410] Nelson H. F. Beebe. A bibliography of papers in Lecture Notes in Computer Science (1991). Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, May 11, 1996. 213 pp. URL http:

//www.math.utah.edu/pub/tex/bib/
index-table-l.html#lncs1991. This

report is updated frequently.

Beebe:1996:BPLd

[411] Nelson H. F. Beebe. A bibliography of papers in Lecture Notes in Computer Science (1993). Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, May 11, 1996. 248 pp. URL http:

//www.math.utah.edu/pub/tex/bib/
index-table-l.html#lncs1993. This

report is updated frequently.

Beebe:1996:BPLe

[412] Nelson H. F. Beebe. A bibliography of papers in Lecture Notes in Computer Science (1994). Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, May 11, 1996. 304 pp. URL http:

//www.math.utah.edu/pub/tex/bib/
index-table-l.html#lncs1994. This

report is updated frequently.

Beebe:1996:BPIe

[413] Nelson H. F. Beebe. A bibliography of publications in IEEE Computer. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, May 14, 1996. 124 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-c.html#computer. This report

is updated frequently.

Beebe:1996:BPE

[414] Nelson H. F. Beebe. A bibliography of publications in Electronic Composition and Imaging. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, May 21, 1996. 12 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-e.html#eleccomp. This report

is updated frequently.

Beebe:1996:BPAl

[415] Nelson H. F. Beebe. A bibliography of publications about typesetting. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, May 21, 1996. 44 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-t.html#typeset. This report is

updated frequently.

Beebe:1996:BPAm

[416] Nelson H. F. Beebe. A bibliography of publications about the AT&T Plan 9 distributed operating system. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, May 27, 1996. 5 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-p.html#plan9. This report is

updated frequently.

Beebe:1996:BPH

[417] Nelson H. F. Beebe. A bibliography of publications in the Hewlett–Packard Journal. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, May 27, 1996. 95 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-

table-h.html#hpj. This report is up-

dated frequently.

Beebe:1996:BPCe

[418] Nelson H. F. Beebe. A bibliography of publications in the Communications of the ACM : 1958–1959. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, May 29, 1996. 2 pp. URL http:

//www.math.utah.edu/pub/tex/bib/
index-table-c.html#cacm1950. This

report is updated frequently.

Beebe:1996:BPCf

[419] Nelson H. F. Beebe. A bibliography of publications in the Communications of the ACM : 1960–1969. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, May 29, 1996. 91 pp. URL http:

//www.math.utah.edu/pub/tex/bib/
index-table-c.html#cacm1960. This

report is updated frequently.

Beebe:1996:BPCg

[420] Nelson H. F. Beebe. A bibliography of publications in the Communications of the ACM : 1970–1979. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, May 29, 1996. 104 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-c.html#cacm1970. This report

is updated frequently.

Beebe:1996:BPCh

[421] Nelson H. F. Beebe. A bibliography of publications in the Communications of the ACM : 1980–1989. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, May 29, 1996. 129 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-c.html#cacm1980. This report

is updated frequently.

Beebe:1996:BPCi

[422] Nelson H. F. Beebe. A bibliography of publications in the Communications of the ACM : 1990–1999. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, May 29, 1996. 200 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-c.html#cacm1990. This report

is updated frequently.

Beebe:1996:BPAn

[423] Nelson H. F. Beebe. A bibliography of publications in ACM Letters on Programming Languages and Systems (LOPLAS). Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, May 30, 1996. 7 pp.

Beebe:1996:BPSb

[424] Nelson H. F. Beebe. A bibliography of publications in Scientific American. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, June 20, 1996. 108 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-s.html#sciam.

Beebe:1996:BPSc

[425] Nelson H. F. Beebe. A bibliography of publications in Smithsonian magazine. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, June 22, 1996. 55 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-s.html#smithsonian.

Beebe:1996:BPSd

[426] Nelson H. F. Beebe. A bibliography of publications in Supercomputer. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, August 22, 1996. 5 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-s.html#supercom.

Beebe:1996:BPJe

[427] Nelson H. F. Beebe. A bibliography of publications in The Journal of Supercomputing. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, August 26, 1996. 21 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-j.html#jsuper. This report is

updated frequently.

Beebe:1996:CBPd

[428] Nelson H. F. Beebe. A complete bibliography of publications in SysAdmin: The Journal for UNIX Systems Administrators. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, August 31, 1996. 27 pp. This report is updated frequently.

Beebe:1996:CBPe

[429] Nelson H. F. Beebe. A complete bibliography of publications in SysAdmin: The Journal for UNIX Systems Administrators. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, August 31, 1996. 27 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-s.html#sysadmin. This report

is updated frequently.

Beebe:1996:CBPc

[430] Nelson H. F. Beebe. A complete bibliography of publications in the C Users Journal and the C/C++ Users Journal. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, August 31, 1996. 161 pp. This report is updated frequently.

Beebe:1996:BXJ

[431] Nelson H. F. Beebe. A bibliography of The X Journal: Computing Technology with the X Window System. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, September 5, 1996. 12 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-x.html#xjournal. This report

is updated frequently.

Beebe:1996:BPAo

[432] Nelson H. F. Beebe. A bibliography of publications about supercomputing. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, September 6, 1996. 5 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-j.html#javarep. This report is

updated frequently.

Beebe:1996:BPJd

[433] Nelson H. F. Beebe. A bibliography of publications in Java Report: The Source for Java Development. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, September 6, 1996. 5 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-j.html#javarep. This report is

updated frequently.

Beebe:1996:BPCj

[434] Nelson H. F. Beebe. A bibliography of publications in the C++ Report. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, September 12, 1996. 14 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-c.html#cppreport. This re-

port is updated frequently.

Beebe:1996:BPAp

[435] Nelson H. F. Beebe. A bibliography of publications about the Mach operating system. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, September 28, 1996. 17 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-

                                        table-m.html#mach. This report is up-

dated frequently.

Beebe:1996:WLL

[436] Nelson H. F. Beebe. Web lessons learned. Byte Magazine, 21(9):20, 22, September 1996. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic). URL http://

                                        www.byte.com/art/9609/sec2/art7.
                                        htm;   http://www.math.utah.edu/
                                        ~beebe/publications/1996/byte-21-
                                        9-20-sep-1996.html.

Beebe:1996:BPIf

[437] Nelson H. F. Beebe. A bibliography of ISSAC (International Symposia on Symbolic and Algebraic Computation) conference papers. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, October 1, 1996. 56 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-i.html#issac. This report is

updated frequently.

Beebe:1996:BPIg

[438] Nelson H. F. Beebe. A complete bibliography of Internet FYI documents. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, October 5, 1996. 6 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-f.html#fyi. This report is up-

dated frequently.

Beebe:1996:BPIi

[439] Nelson H. F. Beebe. A complete bibliography of Internet STD documents. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, October 5, 1996. 11 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-s.html#std. This report is up-

dated frequently.

Beebe:1996:BPIh

[440] Nelson H. F. Beebe. A complete bibliography of Internet RFC documents. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, October 5, 1996. 327 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-r.html#rfc. This report is up-

dated frequently.

Beebe:1996:BPG

[441] Nelson H. F. Beebe. A bibliography of publications about the GNU (Gnu is Not Unix) system. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, October 31, 1996. 51 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-g.html#gnu. This report is up-

dated frequently.

Beebe:1996:BCS

[442] Nelson H. F. Beebe. A bibliography of Case Studies in Mathematical Modeling—Ecology, Physiology, and Cell Biology. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, November 26, 1996. 64 pp. URL http:

                                        //www.math.utah.edu/pub/tex/bib/
                                        index-table-c.html#csmmepcb.

Beebe:1996:BIC

[443] Nelson H. F. Beebe. A bibliography of International Conferences on Very Large Data Bases. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, December 7, 1996. 21 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-v.html#vldb. This report is up-

dated frequently.

Beebe:1996:BASa

[444] Nelson H. F. Beebe. A bibliography of ACM SIGACT-SIGMOD Symposia on Principles of Database Systems. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, December 7, 1996. 22 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-p.html#pods. This report is up-

dated frequently.

Beebe:1996:BAT

[445] Nelson H. F. Beebe. A bibliography of ACM Transactions on Database Systems. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, December 7, 1996. 68 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-t.html#tods. This report is up-

dated frequently.

Beebe:1996:BASb

[446] Nelson H. F. Beebe. A bibliography of ACM SIGMOD Record. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, December 10, 1996. 81 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-s.html#sigmod. This report is

updated frequently.

Beebe:1996:BPLh

[447] Nelson H. F. Beebe. A bibliography of papers in Lecture Notes in Computer Science (1996), part 1 of 2. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, December 21, 1996. 210 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-l.html#lncs1996a. This re-

port is updated frequently.

Beebe:1996:BPLi

[448] Nelson H. F. Beebe. A bibliography of papers in Lecture Notes in Computer Science (1996), part 2 of 2. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, December 21, 1996. 289 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-l.html#lncs1996b. This re-

port is updated frequently.

Beebe:1997:BPBa

[449] Nelson H. F. Beebe. A bibliography of publications in Byte Magazine: 1970– 1979. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, January 4, 1997. 3 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-b.html#byte1970. This report

is updated frequently.

Beebe:1997:BPBb

[450] Nelson H. F. Beebe. A bibliography of publications in Byte Magazine: 1980– 1989. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, January 4, 1997. 49 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-b.html#byte1980. This report

is updated frequently.

Beebe:1997:BPBd

[451] Nelson H. F. Beebe. A bibliography of publications in Byte Magazine: 1995– 1999. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, January 4, 1997. 169 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-b.html#byte1995. This report

is updated frequently.

Beebe:1997:BPBc

[452] Nelson H. F. Beebe. A bibliography of publications in Byte Magazine: 1990– 1994. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, January 4, 1997. 375 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-b.html#byte1990. This report

is updated frequently.

Beebe:1997:BPAa

[453] Nelson H. F. Beebe. A bibliography of publications about the computation of elementary functions in computer programming languages. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, January 11, 1997. 27 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-e.html#elefunt. This report is

updated frequently.

Beebe:1997:BPMa

[454] Nelson H. F. Beebe. A bibliography of publications in Mathematics of Computation. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, January 15, 1997. 276 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-m.html#mathcomp. This report

is updated frequently.

Beebe:1997:BPSa

[455] Nelson H. F. Beebe. A bibliography of publications in SIAM Review. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, January 20, 1997. 129 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-s.html#siamreview. This re-

port is updated frequently.

Beebe:1997:BPSb

[456] Nelson H. F. Beebe. A bibliography of publications in SIAM Journal on Computing. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, January 20, 1997. 159 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-s.html#siamjcomput. This re-

port is updated frequently.

Beebe:1997:BPSc

[457] Nelson H. F. Beebe. A bibliography of publications in SIAM Journal on Numerical Analysis. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, January 20, 1997. 222 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-s.html#siamjnumeranal. This

report is updated frequently.

Beebe:1997:BPSd

[458] Nelson H. F. Beebe. A bibliography of publications in SIAM Journal on Matrix Analysis and Applications. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, January 22, 1997. 86 pp.

URL http://www.math.utah.edu/
pub/tex/bib/index-table-s.html#

siamjmatanaappl. This report is updated frequently.

Beebe:1997:BPLa

[459] Nelson H. F. Beebe. A bibliography of publications in Linear Algebra and its Applications: 1968–1969. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, January 24, 1997. 2 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-l.html#linala1960. This re-

port is updated frequently.

Beebe:1997:CBP

[460] Nelson H. F. Beebe. A complete bibliography of publications in Journal of Numerical Linear Algebra with Applications and Numerical Linear Algebra with Applications. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, January 24, 1997. 16 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-n.html#numlinaa. This report

is updated frequently.

Beebe:1997:BLP

[461] Nelson H. F. Beebe. A bibliography of publications in Linear Algebra and its Applications: 1970–1979. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, January 24, 1997. 24 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-l.html#linala1970. This re-

port is updated frequently.

Beebe:1997:BPLb

[462] Nelson H. F. Beebe. A bibliography of publications in Linear Algebra and its Applications: 1980–1989. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, January 24, 1997. 203 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-

                                        table-l.html#linala1980. This re-

port is updated frequently.

Beebe:1997:BPNa

[463] Nelson H. F. Beebe. A bibliography of publications in Numerische Mathematik. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, January 24, 1997. 213 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-n.html#nummath. This report is

updated frequently.

Beebe:1997:BPLc

[464] Nelson H. F. Beebe. A bibliography of publications in Linear Algebra and its Applications: 1990–1999. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, January 24, 1997. 241 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-l.html#linala1990. This re-

port is updated frequently.

Beebe:1997:BPCa

[465] Nelson H. F. Beebe. A bibliography of publications in Computers and Graphics. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, February 4, 1997. 143 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-c.html#compgraph. This re-

port is updated frequently.

Beebe:1997:BPCb

[466] Nelson H. F. Beebe. A bibliography of publications in Computer Languages. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, February 11, 1997. 34 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-c.html#complngs. This report

is updated frequently.

Beebe:1997:BPAb

[467] Nelson H. F. Beebe. A bibliography of publications in Applied Numerical Mathematics: Transactions of IMACS. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, February 11, 1997. 102 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-a.html#applnummath. This re-

port is updated frequently.

Beebe:1997:BPAc

[468] Nelson H. F. Beebe. A bibliography of publications in Applied Mathematics and Computation. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, February 28, 1997. 191 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-a.html#applmathcomput. This

report is updated frequently.

Beebe:1997:BPSe

[469] Nelson H. F. Beebe. A bibliography of publications about supercomputing. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, February 28, 1997. 418 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-s.html#super. This report is

updated frequently.

Beebe:1997:BPAd

[470] Nelson H. F. Beebe. A bibliography of publications about ISO Standards for programming languages. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, March 3, 1997. 13 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-i.html#isostd. This report is

updated frequently.

Beebe:1997:CBJ

[471] Nelson H. F. Beebe. A complete bibliography of the Journal of Graphics Tools: JGT. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, March 4, 1997. 3 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-j.html#jgraphtools. This re-

port is updated frequently.

Beebe:1997:BPIa

[472] Nelson H. F. Beebe. A bibliography of publications in IEEE Micro. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, April 8, 1997. 90 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-i.html#ieeemicro. This re-

port is updated frequently.

Beebe:1997:BPIb

[473] Nelson H. F. Beebe. A bibliography of publications in IEEE Parallel and Distributed Technology: Systems and Applications. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, April 11, 1997. 25 pp.

URL http://www.math.utah.edu/
pub/tex/bib/index-table-i.html#

ieeepardisttech. This report is updated frequently.

Beebe:1997:BPIc

[474] Nelson H. F. Beebe. A bibliography of publications in IEEE Transactions on Parallel and Distributed Systems. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, April 11, 1997. 92 pp.

URL http://www.math.utah.edu/
pub/tex/bib/index-table-i.html#
ieeetranspardistsys. This report is

updated frequently.

Beebe:1997:BPJa

[475] Nelson H. F. Beebe. A bibliography of publications in Journal of Parallel and Distributed Computing. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, April 12, 1997. 150 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-j.html#jpardistcomp. This

report is updated frequently.

Beebe:1997:BPPb

[476] Nelson H. F. Beebe. A bibliography of publications in the Parallel Processing Letters. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, April 14, 1997. 34 pp.

URL http://www.math.utah.edu/
pub/tex/bib/index-table-p.html#
parallelprocesslett. This report is

updated frequently.

Beebe:1997:BPPa

[477] Nelson H. F. Beebe. A bibliography of publications in Parallel Computing. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, April 14, 1997. 140 pp.

URL http://www.math.utah.edu/
pub/tex/bib/index-table-p.html#
parallelcomputing. This report is up-

dated frequently.

Beebe:1997:BPIe

[478] Nelson H. F. Beebe. A bibliography of publications in the International Journal of Parallel Programming. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, April 26, 1997. 30 pp.

URL http://www.math.utah.edu/
pub/tex/bib/index-table-i.html#
intjparallelprogram. This report is

updated frequently.

Beebe:1997:BPId

[479] Nelson H. F. Beebe. A bibliography of publications in the International Journal of Computer and Information Sciences. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, April 26, 1997. 44 pp.

URL http://www.math.utah.edu/
pub/tex/bib/index-table-i.html#
intjcomputinfsci. This report is up-

dated frequently.

Beebe:1997:BPPc

[480] Nelson H. F. Beebe. A bibliography of publications in Parallel Algorithms and Applications. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, April 28, 1997. 12 pp.

URL http://www.math.utah.edu/
pub/tex/bib/index-table-p.html#
parallelalgorithmsappl. This re-

port is updated frequently.

Beebe:1997:BPNb

[481] Nelson H. F. Beebe. A bibliography of publications in Numerical Algorithms. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, April 29, 1997. 44 pp.

URL http://www.math.utah.edu/
pub/tex/bib/index-table-n.html#

numeralgorithms. This report is updated frequently.

Beebe:1997:BPAe

[482] Nelson H. F. Beebe. A bibliography of publications about ANSI Standards for programming languages. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, April 30, 1997. 21 pp. URL http:

//www.math.utah.edu/pub/tex/bib/
index-table-a.html#ansistd. This

report is updated frequently.

Beebe:1997:BPSg

[483] Nelson H. F. Beebe. A bibliography of publications in SIAM Journal on Scientific Computing. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, May 3, 1997. 56 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-s.html#siamjscicomput. This

report is updated frequently.

Beebe:1997:BPSf

[484] Nelson H. F. Beebe. A bibliography of publications in SIAM Journal on Scientific and Statistical Computing. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, May 3, 1997. 99 pp.

                                        URL http://www.math.utah.edu/
                                        pub/tex/bib/index-table-s.html#
                                        siamjscistatcomp. This report is up-

dated frequently.

Beebe:1997:BPJc

[485] Nelson H. F. Beebe. A bibliography of publications in Journal of Computational and Graphical Statistics. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, May 5, 1997. 9 pp.

                                        URL http://www.math.utah.edu/
                                        pub/tex/bib/index-table-j.html#
                                        jcomputgraphstat. This report is up-

dated frequently.

Beebe:1997:BPAf

[486] Nelson H. F. Beebe. A bibliography of publications in The American Statistician. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, May 5, 1997. 84 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-a.html#amstat. This report is

updated frequently.

Beebe:1997:BPJb

[487] Nelson H. F. Beebe. A bibliography of publications in The Journal of the American Statistical Association. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, May 5, 1997. 256 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-j.html#jamstatassoc. This

report is updated frequently.

Beebe:1997:BPMb

[488] Nelson H. F. Beebe. A bibliography of publications in the Maple Newsletter and the Maple Technical Newsletter. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, May 9, 1997. 25 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-m.html#maple-tech. This re-

port is updated frequently.

Beebe:1997:BPJd

[489] Nelson H. F. Beebe. A bibliography of publications in J.UCS: Journal of Universal Computer Science. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, May 28, 1997. 25 pp. URL http:

                                        //www.math.utah.edu/pub/tex/bib/
                                        index-table-j.html#jucs. This re-

port is updated frequently.

Beebe:1997:BPJe

[490] Nelson H. F. Beebe. A bibliography of publications in The Journal of Visualization and Computer Animation. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, June 3, 1997. 16 pp.

                                        URL http://www.math.utah.edu/
                                        pub/tex/bib/index-table-j.html#
                                        jviscompanimation. This report is up-

dated frequently.

Beebe:1997:BPLd

[491] Nelson H. F. Beebe. A bibliography of papers in Lecture Notes in Computer Science (1997), part 1 of 2. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, August 6, 1997. 164 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-l.html#lncs1997a. This re-

port is updated frequently.

Beebe:1997:BPAg

[492] Nelson H. F. Beebe. A bibliography of publications about Common Lisp. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, August 7, 1997. 33 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-c.html#common-lisp. This re-

port is updated frequently.

Beebe:1997:BPUa

[493] Nelson H. F. Beebe. A bibliography of publications of the USENIX Association: 1980–1989. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, August 14, 1997. 157 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-u.html#usenix1980. This re-

port is updated frequently.

Beebe:1997:BPUb

[494] Nelson H. F. Beebe. A bibliography of publications of the USENIX Association: 1990–1999. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, August 14, 1997. 296 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-u.html#usenix1990. This re-

port is updated frequently.

Beebe:1997:BAU

[495] Nelson H. F. Beebe. A bibliography of Ada User and Ada User Journal. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, September 9, 1997. 18 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-a.html#adauserj. This report

is updated frequently.

Beebe:1997:BPAh

[496] Nelson H. F. Beebe. A bibliography of publications about the Fortran programming language: Part 3: 1990– date. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, October 12, 1997. 190 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-f.html#fortran3. This report

is updated frequently.

Beebe:1997:BS

[497] Nelson H. F. Beebe. A bibliography of Supercomputing ’95. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, November 7, 1997. 28 pp.

URL http://www.math.utah.edu/
pub/tex/bib/index-table-s.html#
supercomputing95. This report is up-

dated frequently.

Beebe:1997:BPTa

[498] Nelson H. F. Beebe. A bibliography of publications in Theoretical Computer Science: 1975–1979. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, November 23, 1997. 21 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-t.html#tcs1975. This report is

updated frequently.

Beebe:1997:BPTb

[499] Nelson H. F. Beebe. A bibliography of publications in Theoretical Computer Science: 1980–1984. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, November 23, 1997. 46 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-

table-t.html#tcs1980. This report is

updated frequently.

Beebe:1997:BPTc

[500] Nelson H. F. Beebe. A bibliography of publications in Theoretical Computer Science: 1985–1989. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, November 23, 1997. 65 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-t.html#tcs1985. This report is

updated frequently.

Beebe:1997:BPTe

[501] Nelson H. F. Beebe. A bibliography of publications in Theoretical Computer Science: 1995–1999. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, November 23, 1997. 87 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-t.html#tcs1995. This report is

updated frequently.

Beebe:1997:BPTd

[502] Nelson H. F. Beebe. A bibliography of publications in Theoretical Computer Science: 1990–1994. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, November 23, 1997. 127 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-t.html#tcs1990. This report is

updated frequently.

Beebe:1997:BPAi

[503] Nelson H. F. Beebe. A bibliography of publications in American Mathematical Monthly: 1960–1969. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, December 4, 1997. 1 pp.

                                        URL http://www.math.utah.edu/
                                        pub/tex/bib/index-table-a.html#
                                        amermathmonthly1960. This report is

updated frequently.

Beebe:1997:BPAj

[504] Nelson H. F. Beebe. A bibliography of publications in American Mathematical Monthly: 1970–1979. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, December 4, 1997. 13 pp.

                                        URL http://www.math.utah.edu/
                                        pub/tex/bib/index-table-a.html#
                                        amermathmonthly1970. This report is

updated frequently.

Beebe:1997:BPAk

[505] Nelson H. F. Beebe. A bibliography of publications in American Mathematical Monthly: 1980–1989. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, December 4, 1997. 75 pp.

                                        URL http://www.math.utah.edu/
                                        pub/tex/bib/index-table-a.html#
                                        amermathmonthly1980. This report is

updated frequently.

Beebe:1997:BPAl

[506] Nelson H. F. Beebe. A bibliography of publications in American Mathematical Monthly: 1990–1999. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, December 4, 1997. 78 pp.

URL http://www.math.utah.edu/
pub/tex/bib/index-table-a.html#
amermathmonthly1990. This report is

updated frequently.

Beebe:1997:BPSh

[507] Nelson H. F. Beebe. A bibliography of publications in the SIAM Journal on Applied Mathematics. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, December 11, 1997. 247 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-s.html#siamjapplmath. This

report is updated frequently.

Beebe:1998:BPJa

[508] Nelson H. F. Beebe. A bibliography of publications in the Journal of Mathematical Physics: 1960– 1964. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, January 3, 1998. 2 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-j.html#jmathphys1960. This

report is updated frequently.

Beebe:1998:BPJb

[509] Nelson H. F. Beebe. A bibliography of publications in the Journal of Mathematical Physics: 1965–1969. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, January 3, 1998. 58 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-j.html#jmathphys1965. This

report is updated frequently.

Beebe:1998:BPJh

[510] Nelson H. F. Beebe. A bibliography of publications in the Journal of Mathematical Physics: 1995–1999. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, January 3, 1998. 147 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-j.html#jmathphys1995. This

report is updated frequently.

Beebe:1998:BPJc

[511] Nelson H. F. Beebe. A bibliography of publications in the Journal of Mathematical Physics: 1970–1974. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, January 3, 1998. 181 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-j.html#jmathphys1970. This

report is updated frequently.

Beebe:1998:BPJd

[512] Nelson H. F. Beebe. A bibliography of publications in the Journal of Mathematical Physics: 1975–1979. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, January 3, 1998. 201 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-j.html#jmathphys1975. This

report is updated frequently.

Beebe:1998:BPJe

[513] Nelson H. F. Beebe. A bibliography of publications in the Journal of Mathematical Physics: 1980–1984. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, January 3, 1998. 223 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-j.html#jmathphys1980. This

report is updated frequently.

Beebe:1998:BPJf

[514] Nelson H. F. Beebe. A bibliography of publications in the Journal of Mathematical Physics: 1985–1989. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, January 3, 1998. 232 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-j.html#jmathphys1985. This

report is updated frequently.

Beebe:1998:BPJg

[515] Nelson H. F. Beebe. A bibliography of publications in the Journal of Mathematical Physics: 1990–1994. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, January 3, 1998. 254 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-j.html#jmathphys1990. This

report is updated frequently.

Beebe:1998:BPAa

[516] Nelson H. F. Beebe. A bibliography of publications about the Icon programming language. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, January 12, 1998. 6 pp. URL http:

//www.math.utah.edu/pub/tex/bib/
index-table-i.html#icon.

Beebe:1998:BSj

[517] Nelson H. F. Beebe. A bibliography of Supercomputing ’97. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, March 21, 1998. 14 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-s.html#supercomputing97.

Beebe:1998:BSi

[518] Nelson H. F. Beebe. A bibliography of Supercomputing ’96. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, March 23, 1998. 13 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-s.html#supercomputing96.

Beebe:1998:LBA

[519] Nelson H. F. Beebe. Letters: Bibliography archive. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 32 (1):1–3, March 1998. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic). URL http://portal.

acm.org/citation.cfm?id=294833.
294835&dl=GUIDE&dl=ACM&idx=J707&
part=periodical&WantType=periodical&
title=ACM%20SIGSAM%20Bulletin&CFID=
11742174&CFTOKEN=31030343;   http:

//www.math.utah.edu/~beebe/publications/
1998/sigsam-32-1-1-march-1998.pdf.

Beebe:1998:AAE

[520] Nelson H. F. Beebe. authidx: An author/editor indexing package. TUGboat, 19(1):12–19, March

1998.  ISSN 0896-3207.  URL

http://www.math.utah.edu/~beebe/
publications/1998/tb58beeb.pdf.

Beebe:1998:BPLa

[521] Nelson H. F. Beebe. A bibliography of papers in Lecture Notes in Computer Science (1997), part 2 of

2. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, April 4, 1998. 238 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-l.html#lncs1997b. This re-

port is updated frequently.

Beebe:1998:BPSa

[522] Nelson H. F. Beebe. A bibliography of publications in the SIAM Journal on Discrete Mathematics. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, April 9, 1998. 66 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-s.html#siamjdiscrmath. This

report is updated frequently.

Beebe:1998:BPSb

[523] Nelson H. F. Beebe. A bibliography of publications in the SIAM Journal on Algebraic and Discrete Methods. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, April 11, 1998. 60 pp.

                                        URL http://www.math.utah.edu/
                                        pub/tex/bib/index-table-s.html#
                                        siamjalgdiscmeth. This report is up-

dated frequently.

Beebe:1998:BSf

[524] Nelson H. F. Beebe. A bibliography of Supercomputing ’93. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, April 15, 1998. 20 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-s.html#supercomputing93.

Beebe:1998:BSg

[525] Nelson H. F. Beebe. A bibliography of Supercomputing ’94. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, April 15, 1998. 21 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-s.html#supercomputing94.

Beebe:1998:BSe

[526] Nelson H. F. Beebe. A bibliography of Supercomputing ’92. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, April 15, 1998. 23 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-s.html#supercomputing92.

Beebe:1998:BSb

[527] Nelson H. F. Beebe. A bibliography of Supercomputing ’89. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, April 16, 1998. 14 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-s.html#supercomputing89.

Beebe:1998:BSa

[528] Nelson H. F. Beebe. A bibliography of Supercomputing ’88. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, April 16, 1998. 21 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-s.html#supercomputing88.

Beebe:1998:BSd

[529] Nelson H. F. Beebe. A bibliography of Supercomputing ’91. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, April 16, 1998. 21 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-s.html#supercomputing91.

Beebe:1998:BSc

[530] Nelson H. F. Beebe. A bibliography of Supercomputing ’90. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, April 16, 1998. 22 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-s.html#supercomputing90.

Beebe:1998:BPLb

[531] Nelson H. F. Beebe. A bibliography of papers in Lecture Notes in Computer Science (1998), part 1 of 2. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, April 29, 1998. 52 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-l.html#lncs1998a. This re-

port is updated frequently.

Beebe:1998:BPAb

[532] Nelson H. F. Beebe. A bibliography of publications about the Python scripting and programming language. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, May 22, 1998. 14 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-p.html#python. This report is

updated frequently.

Beebe:1998:BBM

[533] Nelson H. F. Beebe. A bibliography of books and manuals on the Oracle Database System. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, June 29, 1998. 14 pp. URL http:

                                        //www.math.utah.edu/pub/tex/bib/
                                        index-table-o.html#oracle-books.

This report is updated frequently.

Beebe:1998:BPM

[534] Nelson H. F. Beebe. A bibliography of publications about multithreading. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, August 7, 1998. 15 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-m.html#multithreading. This

report is updated frequently.

Beebe:1998:BPJi

[535] Nelson H. F. Beebe. A bibliography of publications in JavaWorld magazine. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, August 14, 1998. 101 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-j.html#javaworld. This re-

port is updated frequently.

Thomas:1998:BPG

[536] Steve Thomas and Nelson H. F. Beebe. A bibliography of Project Gutenberg. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, August 14, 1998. 224 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-g.html#gutenberg. This re-

port is updated frequently.

Beebe:1998:BSk

[537] Nelson H. F. Beebe. A bibliography of Supercomputing ’98. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, October 7, 1998. 9 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-s.html#supercomputing97.

Beebe:1998:BSl

[538] Nelson H. F. Beebe. A bibliography of Supercomputing ’98. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, October 7, 1998. 9 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-s.html#supercomputing98.

Beebe:1998:CBPa

[539] Nelson H. F. Beebe. A complete bibliography of publications in the Linux Journal. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, October 9, 1998. 130 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-l.html#linux-journal.

Beebe:1998:BPLd

[540] Nelson H. F. Beebe. A bibliography of papers in Lecture Notes in Computer Science (1999), part 1 of 2. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, October 11, 1998. 170 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-l.html#lncs1999a.

Beebe:1998:BPLc

[541] Nelson H. F. Beebe. A bibliography of papers in Lecture Notes in Computer Science (1998), part 2 of 2. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, October 11, 1998. 348 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-l.html#lncs1998b.

Beebe:1998:CBPf

[542] Nelson H. F. Beebe. A complete bibliography of publications in Mathematical Tables and [Other] Aids to Computation, 1940–1949. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, October 16, 1998. 11 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-m.html#mathcomp1940.

Beebe:1998:CBPk

[543] Nelson H. F. Beebe. A complete bibliography of publications in Mathematical Tables and Other Aids to Computation, 1950–1959. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, October 16, 1998. 29 pp. URL http:

//www.math.utah.edu/pub/tex/bib/
index-table-m.html#mathcomp1950.

Beebe:1998:CBPg

[544] Nelson H. F. Beebe. A complete bibliography of publications in Mathematics of Computation, 1960–1969. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, October 16, 1998. 81 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-m.html#mathcomp1960.

Beebe:1998:CBPh

[545] Nelson H. F. Beebe. A complete bibliography of publications in Mathematics of Computation, 1970–1979. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, October 16, 1998. 98 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-m.html#mathcomp1970.

Beebe:1998:CBPi

[546] Nelson H. F. Beebe. A complete bibliography of publications in Mathematics of Computation, 1980–1989. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, October 16, 1998. 116 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-m.html#mathcomp1980.

Beebe:1998:CBPj

[547] Nelson H. F. Beebe. A complete bibliography of publications in Mathematics of Computation, 1990–1999. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, October 16, 1998. 118 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-m.html#mathcomp1990.

Beebe:1998:CBPd

[548] Nelson H. F. Beebe. A complete bibliography of publications in the Journal of the Society for Industrial and Applied Mathematics. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, October 17, 1998. 62 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-s.html#siam.

Beebe:1998:BPAc

[549] Nelson H. F. Beebe. A bibliography of publications about the Fast Multipole Method. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, October 28, 1998. 52 pp.

URL http://www.math.utah.edu/
pub/bibnet/subjects/index.html#
fastmultipole.

Beebe:1998:BPAd

[550] Nelson H. F. Beebe. A bibliography of publications in American Mathematical Monthly: 1930–1939. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, October 31, 1998. 2 pp.

URL http://www.math.utah.edu/
pub/tex/bib/index-table-a.html#
amermathmonthly1930.

Beebe:1998:BPAe

[551] Nelson H. F. Beebe. A bibliography of publications in American Mathematical Monthly: 1940–1949. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, October 31, 1998. 2 pp.

URL http://www.math.utah.edu/
pub/tex/bib/index-table-a.html#
amermathmonthly1940.

Beebe:1998:BPAf

[552] Nelson H. F. Beebe. A bibliography of publications in American Mathematical Monthly: 1950–1959. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, October 31, 1998. 2 pp.

URL http://www.math.utah.edu/
pub/tex/bib/index-table-a.html#
amermathmonthly1950.

Beebe:1998:BPId

[553] Nelson H. F. Beebe. A bibliography of publications in IEEE Concurrency. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, November 5, 1998. 7 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-i.html#ieeeconcurrency.

Beebe:1998:CBPb

[554] Nelson H. F. Beebe. A complete bibliography of publications in Nordic Journal of Computing. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, November 13, 1998. 16 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-n.html#nordic-j-computing.

Beebe:1998:BPIa

[555] Nelson H. F. Beebe. A bibliography of publications in Information Processing Letters: 1970–1979. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, November 13, 1998. 77 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-i.html#infoproc1970.

Beebe:1998:BPIb

[556] Nelson H. F. Beebe. A bibliography of publications in Information Processing Letters: 1980–1989. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, November 13, 1998. 158 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-i.html#infoproc1980.

Beebe:1998:BPIc

[557] Nelson H. F. Beebe. A bibliography of publications in Information Processing Letters: 1990–1999. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, November 13, 1998. 239 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-i.html#infoproc1990.

Beebe:1998:CBPc

[558] Nelson H. F. Beebe. A complete bibliography of publications in Nordisk Tidskrift for Informationsbehandling, BIT, and BIT Numerical Mathematics. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, November 15, 1998. 215 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-b.html#bit.

Beebe:1998:CBPe

[559] Nelson H. F. Beebe. A complete bibliography of publications in the Journal of Cryptology. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, November 21, 1998. 27 pp. URL http:

                                        //www.math.utah.edu/pub/tex/bib/
                                        index-table-j.html#jcryptology.

Beebe:1998:BPCa

[560] Nelson H. F. Beebe. A bibliography of publications in Cryptologia. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, November 23, 1998. 67 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-c.html#cryptologia.

Beebe:1998:CBI

[561] Nelson H. F. Beebe. A complete bibliography of Internet Engineering Note (IEN) documents. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, November 24, 1998. 4 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-i.html#ien.

Beebe:1998:BPCb

[562] Nelson H. F. Beebe. A bibliography of publications on cryptography. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, November 25, 1998. 293 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-c.html#cryptography.

Beebe:1998:BPVa

[563] Nelson H. F. Beebe. A bibliography of the proceedings volumes of the ACM Symposia on the Theory of Computing (1970–1997). Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, December 3, 1998. 5 pp. URL http:

//www.math.utah.edu/pub/tex/bib/
index-table-s.html#stoc.

Beebe:1998:BPVb

[564] Nelson H. F. Beebe. A bibliography of the proceedings volumes of the IEEE symposia on the Foundations of Computer Science (1975– 1997). Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, December 3, 1998. 5 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-f.html#focs.

Beebe:1998:BPT

[565] Nelson H. F. Beebe. A bibliography of publications in Theory of Probability and its Applications. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, December 8, 1998. 150 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-t.html#theoryprobabappl.

Beebe:1998:BPJj

[566] Nelson H. F. Beebe. A bibliography of publications in the Journal of the Society for Industrial and Applied Mathematics. Series A, On control and SIAM Journal on Control. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, December 9, 1998. 51 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-s.html#siamjcontrol.

Beebe:1998:BPSc

[567] Nelson H. F. Beebe. A bibliography of publications in the SIAM Journal on Control and Optimization. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, December 9, 1998. 236 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-s.html#siamjcontroloptim.

Beebe:1998:BPJk

[568] Nelson H. F. Beebe. A bibliography of publications in the Journal of Systems and Software. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, December 18, 1998. 125 pp. URL http:

                                        //www.math.utah.edu/pub/tex/bib/
                                        index-table-j.html#jsystsoftw.

Beebe:1999:BJA

[569] Nelson H. F. Beebe. A bibliography of the Journal of Applied Statistics. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, January 2, 1999. 28 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-j.html#japplstat.

Beebe:1999:CBPa

[570] Nelson H. F. Beebe. A complete bibliography of publications in Computing in Science and Engineering.

Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, January 10, 1999. 4 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-c.html#computscieng.

Beebe:1999:BPIa

[571] Nelson H. F. Beebe. A bibliography of publications in IEEE Computational Science and Engineering. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, January 10, 1999. 62 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-i.html#ieeecomputscieng.

Beebe:1999:CBAa

[572] Nelson H. F. Beebe. A complete bibliography of ACM Transactions on Computer Systems. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, January 14, 1999. 42 pp. URL http:

//www.math.utah.edu/pub/tex/bib/
index-table-t.html#tocs.

Beebe:1999:BPS

[573] Nelson H. F. Beebe. A bibliography of publications in SIAM Journal on Mathematical Analysis. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, January 15, 1999. 295 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-s.html#siamjmathana.

Beebe:1999:CBAb

[574] Nelson H. F. Beebe. A complete bibliography of ACM Transactions on Office Information Systems. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, January 18, 1999. 21 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-t.html#toois.

Beebe:1999:CBAc

[575] Nelson H. F. Beebe. A complete bibliography of ACM Transactions on Computer-Human Interaction. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, January 19, 1999. 13 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-t.html#tochi.

Beebe:1999:CBAd

[576] Nelson H. F. Beebe. A complete bibliography of ACM Transactions on Information Systems. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, January 19, 1999. 28 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-t.html#tois.

Beebe:1999:CBPb

[577] Nelson H. F. Beebe. A complete bibliography of publications in Chicago Journal of Theoretical Computer Science. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, January 28, 1999. 5 pp.

URL http://www.math.utah.edu/
pub/tex/bib/index-table-c.html#
chicjtheorcomputsci.

Beebe:1999:BPCa

[578] Nelson H. F. Beebe. A bibliography of publications in Computer Systems Science and Engineering and International Journal of Computer Systems Science and Engineering. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, February 4, 1999. 50 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-c.html#computsystscieng.

Beebe:1999:BPD

[579] Nelson H. F. Beebe. A bibliography of publications in Designs, Codes, and Cryptography. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, February 10, 1999. 48 pp.

URL http://www.math.utah.edu/
pub/tex/bib/index-table-d.html#
designscodescryptogr.

Beebe:1999:BPH

[580] Nelson H. F. Beebe. A bibliography of publications in Higher-Order Symbolic Computation. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, February 11, 1999. 5 pp.

URL http://www.math.utah.edu/
pub/tex/bib/index-table-h.html#
higherordersymbcomput.

Beebe:1999:BPLa

[581] Nelson H. F. Beebe. A bibliography of publications in Lisp and Symbolic Computation. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, February 12, 1999. 21 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-l.html#lispsymbcomput.

Beebe:1999:HTL

[582] Nelson H. F. Beebe. The hash table library. Report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, March 12, 1999. 10 pp. URL http:/

                                        /www.math.utah.edu/pub/emacs/;
                                        http://www.math.utah.edu/pub/emacs/
                                        hash-1.00.tar.gz.

Beebe:1999:PL

[583] Nelson H. F. Beebe. The primes library. Report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, March 12, 1999. 8 pp. URL http:/

                                        /www.math.utah.edu/pub/emacs/;
                                        http://www.math.utah.edu/pub/emacs/
                                        primes-1.01.tar.gz.

Beebe:1999:BPAa

[584] Nelson H. F. Beebe. A bibliography of publications in ACM SIGPLAN Notices, 1960–1969. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, May 3, 1999. 2 pp. URL http:

                                        //www.math.utah.edu/pub/tex/bib/

index-table-s.html#sigplan1960.

This report is updated frequently.

Beebe:1999:BPAb

[585] Nelson H. F. Beebe. A bibliography of publications in ACM SIGPLAN Notices, 1970–1979. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, May 3, 1999. 71 pp. URL http:

//www.math.utah.edu/pub/tex/bib/
index-table-s.html#sigplan1970.

This report is updated frequently.

Beebe:1999:BPAc

[586] Nelson H. F. Beebe. A bibliography of publications in ACM SIGPLAN Notices, 1980–1989. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, May 3, 1999. 182 pp. URL http:

//www.math.utah.edu/pub/tex/bib/
index-table-s.html#sigplan1980.

This report is updated frequently.

Beebe:1999:BPAd

[587] Nelson H. F. Beebe. A bibliography of publications in ACM SIGPLAN Notices, 1990–1999. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, May 3, 1999. 250 pp. URL http:

//www.math.utah.edu/pub/tex/bib/
index-table-s.html#sigplan1990.

This report is updated frequently.

Beebe:1999:BPAe

[588] Nelson H. F. Beebe. A bibliography of publications in American Mathematical Monthly: 1894–1899. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, June 28, 1999.

URL http://www.math.utah.edu/
pub/tex/bib/index-table-a.html#
amermathmonthly1890.

Beebe:1999:BPAf

[589] Nelson H. F. Beebe. A bibliography of publications in American Mathematical Monthly: 1900–1910. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, June 28, 1999.

URL http://www.math.utah.edu/
pub/tex/bib/index-table-a.html#
amermathmonthly1900.

Beebe:1999:BPAg

[590] Nelson H. F. Beebe. A bibliography of publications in American Mathematical Monthly: 1910–1919. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, June 28, 1999.

URL http://www.math.utah.edu/
pub/tex/bib/index-table-a.html#
amermathmonthly1910.

Beebe:1999:BPAh

[591] Nelson H. F. Beebe. A bibliography of publications in American Mathematical Monthly: 1920–1929. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, June 28, 1999.

URL http://www.math.utah.edu/
pub/tex/bib/index-table-a.html#
amermathmonthly1920.

Beebe:1999:CBPc

[592] Nelson H. F. Beebe. A complete bibliography of publications in ACM Transactions on Information and System Security. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, July 27, 1999. 4 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-t.html#tissec.

Beebe:1999:BPIb

[593] Nelson H. F. Beebe. A bibliography of publications in IEEE/ACM Transactions on Networking. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, July 27, 1999. 89 pp.

URL http://www.math.utah.edu/
pub/tex/bib/index-table-i.html#
ieeetransnetworking.

Beebe:1999:CBPd

[594] Nelson H. F. Beebe. A complete bibliography of publications in Concurrency: Practice and Experience. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, September 7, 1999. 81 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-c.html#cpe. This report is up-

dated frequently.

Beebe:1999:BPLb

[595] Nelson H. F. Beebe. A bibliography of papers in Lecture Notes in Computer Science (1999), part 2 of 2. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, September 14, 1999. 184 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-l.html#lncs1999b.

Beebe:1999:CBPe

[596] Nelson H. F. Beebe. A complete bibliography of publications in Markup Languages: Theory & Practice. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, September 20, 1999. 3 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-m.html#markuplang.

Beebe:1999:BPCd

[597] Nelson H. F. Beebe. A bibliography of publications in Computer Networks (Amsterdam, Netherlands: 1999). Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, September 27, 1999. 21 pp.

URL http://www.math.utah.edu/
pub/tex/bib/index-table-c.html#
compnetamsterdam1990.

Beebe:1999:BPCb

[598] Nelson H. F. Beebe. A bibliography of publications in Computer Networks: The International Journal of Distributed Informatique. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, September 27, 1999. 27 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-c.html#compnet.

Beebe:1999:BPCc

[599] Nelson H. F. Beebe. A bibliography of publications in Computer Networks and ISDN Systems. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, September 27, 1999. 240 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-c.html#compnetisdn.

Beebe:1999:BPCf

[600] Nelson H. F. Beebe. A bibliography of publications in Science of Computer Programming. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, October 9, 1999. 45 pp.

URL http://www.math.utah.edu/
pub/tex/bib/index-table-c.html#
scicomputprogram. This report is up-

dated frequently.

Beebe:1999:BPCe

[601] Nelson H. F. Beebe. A bibliography of publications in Computing. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, October 18, 1999. 241 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-c.html#computing. This re-

port is updated frequently.

Beebe:1999:BNA

[602] Nelson H. F. Beebe. A bibliography of numerical analysis publications, 1990– 1999. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, November 3, 1999. 91 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-s.html#numana1990. This re-

port is updated frequently.

Beebe:1999:FAE

[603] Nelson H. F. Beebe. Fast approximate exponential functions. Report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, December 7, 1999.

Beebe:1999:BPB

[604] Nelson H. F. Beebe. A bibliography of publications in Bulletin canadien de mathématiques = Canadian Mathematical Bulletin. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, December 22, 1999. 8 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-c.html#canmathbull. This re-

port is updated frequently.

Beebe:2000:BPCa

[605] Nelson H. F. Beebe. A bibliography of publications in the Communications of the ACM : 2000–2009. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, January 6, 2000. 4 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-c.html#cacm2000. This report

is updated frequently.

Beebe:2000:CBPa

[606] Nelson H. F. Beebe. A complete bibliography of publications in Acta Informatica. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, January 6, 2000. 134 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-c.html#actainfo. This report

is updated frequently.

Beebe:2000:BNA

[607] Nelson H. F. Beebe. A bibliography of numerical analysis publications, 2000– 2009. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, January 22, 2000. 1 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-c.html#numana2000. This re-

port is updated frequently.

Beebe:2000:CBPb

[608] Nelson H. F. Beebe. A complete bibliography of publications in the SIAM Journal on Applied Mathematics: 2000–2009. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, January 22, 2000. 6 pp.

URL http://www.math.utah.edu/
pub/tex/bib/index-table-s.html#
siamjapplmath2000. This report is up-

dated frequently.

Beebe:2000:BPSa

[609] Nelson H. F. Beebe. A bibliography of publications in SIAM Journal on Numerical Analysis: 2000– 2009. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, January 22, 2000. 7 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-s.html#siamjnumeranal2000.

This report is updated frequently.

Beebe:2000:NCP

[610] Nelson H. F. Beebe. ndiff — compare putatively similar files, ignoring small numeric differences. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, January 28, 2000. 7

                                        pp. URL http://www.math.utah.
                                        edu/pub/ndiff.

Beebe:2000:CIE

[611] Nelson H. F. Beebe. The cost of IEEE 754 exceptional operands and instructions. World-Wide Web document, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, February 11, 2000. 6 pp. URL

                                        http://www.math.utah.edu/~beebe/
                                        software/ieee/timops.html.

Beebe:2000:BS

[612] Nelson H. F. Beebe. A bibliography of Supercomputing ’99. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, February 24, 2000. 11 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-s.html#supercomputing99.

Beebe:2000:BPIa

[613] Nelson H. F. Beebe. A bibliography of publications in The International Journal of Quantum Chemistry: 2000– date. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, February 25, 2000. 195 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-i.html#ijqc2000.

Beebe:2000:BPAa

[614] Nelson H. F. Beebe. A bibliography of publications about the Java Programming Language: 2000–2009. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, March 11, 2000. 13 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-j.html#java2000. This report

is updated frequently.

Beebe:2000:BPAb

[615] Nelson H. F. Beebe. A bibliography of publications in ACM SIG-PLAN Notices, 2000–2009. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, March 22, 2000. 24 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-s.html#sigplan2000. This re-

port is updated frequently.

Beebe:2000:CBPc

[616] Nelson H. F. Beebe. A complete bibliography of publications in Mathematics of Computation, 2000–2009.

Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, April 3, 2000. 30 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-m.html#mathcomp2000.

Beebe:2000:BISa

[617] Nelson H. F. Beebe. A bibliography of IEC Standards for computing systems and programming languages. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, April 24, 2000. 4 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-i.html#iecstd.

Beebe:2000:BISb

[618] Nelson H. F. Beebe. A bibliography of IEEE standards. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, April 24, 2000. 8 pp. URL http:

//www.math.utah.edu/pub/tex/bib/
index-table-i.html#ieeestd.

Beebe:2000:CBPd

[619] Nelson H. F. Beebe. A complete bibliography of publications in the VLDB Journal: Very Large Data Bases. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, September 27, 2000. 40 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-v.html#vldbj.

Beebe:2000:CBPe

[620] Nelson H. F. Beebe. A complete bibliography of publications in Linear Algebra and its Applications: 2000– 2009. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, October 9, 2000. 44 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-l.html#linala2000. This re-

port is updated frequently.

Beebe:2000:CBPf

[621] Nelson H. F. Beebe. A complete bibliography of publications in Information Processing Letters: 2000–2009. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, October 12, 2000. 24 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-i.html#infoproc2000.

Beebe:2000:CBA

[622] Nelson H. F. Beebe. A complete bibliography of ACM Transactions on Computational Logic. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, October 13, 2000. 3 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-t.html#tocl.

Beebe:2000:BPCb

[623] Nelson H. F. Beebe. A bibliography of publications on cryptography: 2000– 2009. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, October 17, 2000. 8 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-c.html#cryptography2000.

Beebe:2000:CBL

[624] Nelson H. F. Beebe. A complete bibliography of Lecture Notes in Computational Science and Engineering. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, October 18, 2000. 5 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-l.html#lncse.

Beebe:2000:BPJ

[625] Nelson H. F. Beebe. A bibliography of publications in the Journal of Mathematical Physics: 2000–2004. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, October 20, 2000. 48 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-j.html#jmathphys2000. This

report is updated frequently.

Beebe:2000:BPIb

[626] Nelson H. F. Beebe. A bibliography of publications in IEEE Computer: 2000– 2009. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, October 21, 2000. 39 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-c.html#computer2000. This

report is updated frequently.

Beebe:2000:CBPg

[627] Nelson H. F. Beebe. A complete bibliography of publications in Computer Networks (Amsterdam, Netherlands: 2000–2009). Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, October 22, 2000. 26 pp.

URL http://www.math.utah.edu/
pub/tex/bib/index-table-j.html#
compnetamsterdam2000. This report

is updated frequently.

Beebe:2000:BAS

[628] Nelson H. F. Beebe. A bibliography of ACM SIGACT-SIGMOD Symposia on Principles of Database Systems and SIGMOD Conferences on Management of Data. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, October 26, 2000. 273 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-p.html#pods. This report is up-

dated frequently.

Beebe:2000:BPIc

[629] Nelson H. F. Beebe. A bibliography of publications in IEEE Computer: 1972– 1979. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, October 31, 2000. 24 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-c.html#computer1970. This

report is updated frequently.

Beebe:2000:BPId

[630] Nelson H. F. Beebe. A bibliography of publications in IEEE Computer: 1980– 1989. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, October 31, 2000. 34 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-c.html#computer1980. This

report is updated frequently.

Beebe:2000:BPT

[631] Nelson H. F. Beebe. A bibliography of publications in Theoretical Computer Science: 2000–2004. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, October 31, 2000. 59 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-t.html#tcs2000. This report is

updated frequently.

Beebe:2000:BPIe

[632] Nelson H. F. Beebe. A bibliography of publications in IEEE Computer: 1990– 1999. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, October 31, 2000. 225 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-c.html#computer1990. This

report is updated frequently.

Beebe:2000:BPDc

[633] Nelson H. F. Beebe. A bibliography of publications in Dr. Dobb’s Journal: Software Tools for the Professional Programmer: 2000–2009. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, November 9, 2000. 38 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-d.html#dr-dobbs-2000. This

report is updated frequently.

Beebe:2000:BPDa

[634] Nelson H. F. Beebe. A bibliography of publications in Dr. Dobb’s Journal: Software Tools for the Professional Programmer: 1980–1989. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, November 9, 2000. 116 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-d.html#dr-dobbs-1980. This

report is updated frequently.

Beebe:2000:BPDb

[635] Nelson H. F. Beebe. A bibliography of publications in Dr. Dobb’s Journal: Software Tools for the Professional Programmer: 1990–1999. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, November 9, 2000. 285 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-d.html#dr-dobbs-1990. This

report is updated frequently.

Beebe:2000:BPSe

[636] Nelson H. F. Beebe. A bibliography of publications in Scientific American: 2000–2009. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, November 12, 2000. 18 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-s.html#sciam2000.

Beebe:2000:BPSb

[637] Nelson H. F. Beebe. A bibliography of publications in Scientific American: 1970–1979. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, November 12, 2000. 40 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-s.html#sciam1970.

Beebe:2000:BPSc

[638] Nelson H. F. Beebe. A bibliography of publications in Scientific American: 1980–1989. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, November 12, 2000. 72 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-s.html#sciam1980.

Beebe:2000:BPSd

[639] Nelson H. F. Beebe. A bibliography of publications in Scientific American: 1990–1999. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, November 12, 2000. 235 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-s.html#sciam1990.

Beebe:2000:BPP

[640] Nelson H. F. Beebe. A bibliography of publications of Per-Olov Löwdin. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, November 17, 2000.

                                        53 pp. URL http://www.math.utah.
                                        edu/pub/bibnet/authors/l/lowdin-
                                        per-olov.*.

Beebe:2000:FBP

[641] Nelson H. F. Beebe. Filesystem backup performance at the University of Utah Mathematics Department. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, December 8, 2000. 6 pp.

URL http://www.math.utah.edu/~
beebe/amanda/amanda-2000.pdf.

Beebe:2000:BPSf

[642] Nelson H. F. Beebe. A bibliography of publications in Smithsonian magazine: 1980–1989. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, December 9, 2000. 5 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-s.html#smithsonian1980.

Beebe:2000:BPAc

[643] Nelson H. F. Beebe. A bibliography of publications in The ACM Journal of Computer Documentation. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, December 9, 2000. 20 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-j.html#jcd.

Beebe:2000:BPSh

[644] Nelson H. F. Beebe. A bibliography of publications in Smithsonian magazine: 2000–2009. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, December 9, 2000. 26 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-s.html#smithsonian2000.

Beebe:2000:BPSg

[645] Nelson H. F. Beebe. A bibliography of publications in Smithsonian magazine: 1990–1999. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, December 9, 2000. 162 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-s.html#smithsonian1990.

Beebe:2000:CBPh

[646] Nelson H. F. Beebe. A complete bibliography of publications in The IMA Journal of Numerical Analysis. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, December 23, 2000. 88 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-i.html#imajnumeranal.

Beebe:2000:CBPi

[647] Nelson H. F. Beebe. A complete bibliography of publications in The IMA Journal of Applied Mathematics. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, December 26, 2000. 95 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-i.html#imajapplmath.

Beebe:2000:CBPj

[648] Nelson H. F. Beebe. A complete bibliography of publications in Advances in Quantum Chemistry. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, December 30, 2000. 73 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-a.html#advquantumchem.

Beebe:2001:BAC

[649] Nelson H. F. Beebe. A bibliography of the annual Cool Chips Symposia. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, January 8, 2001. 9 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-c.html#cool-chips.

Beebe:2001:BAI

[650] Nelson H. F. Beebe. A bibliography of the annual IEEE Hot Chips Symposia. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, January 8, 2001. 31 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-h.html#hot-chips.

Beebe:2001:CBPa

[651] Nelson H. F. Beebe. A complete bibliography of publications in IEEE Transactions on Visualization and Computer Graphics. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, January 12, 2001. 39 pp.

URL http://www.math.utah.edu/
pub/tex/bib/index-table-i.html#
ieeetransviscomputgraph.

Beebe:2001:BPIa

[652] Nelson H. F. Beebe. A bibliography of publications on the Intel IA-64 architecture. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, January 13, 2001. 10 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-i.html#intel-ia-64.

Beebe:2001:BPV

[653] Nelson H. F. Beebe. A bibliography of publications on visual instruction sets. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, January 13, 2001. 10 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-v.html#visual-instruction-
                                        set.

Beebe:2001:SBP

[654] Nelson H. F. Beebe. A selected bibliography of publications about microprocessors. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, January 17, 2001. 23 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-c.html#microchip. This re-

port is updated frequently.

Beebe:2001:CBPb

[655] Nelson H. F. Beebe. A complete bibliography of publications in Computing. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, January 17, 2001. 354 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-c.html#computing. This re-

port is updated frequently.

Beebe:2001:CBPc

[656] Nelson H. F. Beebe. A complete bibliography of publications in IEEE MultiMedia. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, January 30, 2001. 80 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-i.html#ieeemultimedia. This

report is updated frequently.

Beebe:2001:UUM

[657] Nelson H. F. Beebe. University of Utah Mathematics Department electronic mailing lists. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, February 1,

2001. 40 pp. URL file://suncore.
math.utah.edu/local/mail/lists/
reports/mailing-list-report.ps.

This report is updated semesterly.

Beebe:2001:BS

[658] Nelson H. F. Beebe. A bibliography of Supercomputing ’2000. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, February 11, 2001. 32 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-s.html#supercomputing2000.

Beebe:2001:CBPd

[659] Nelson H. F. Beebe. A complete bibliography of publications in the SIAM Journal on Algebraic and Discrete Methods. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, February 20, 2001. 62 pp.

URL http://www.math.utah.edu/
pub/tex/bib/index-table-s.html#
siamjalgdiscmeth. This report is up-

dated frequently.

Beebe:2001:CBE

[660] Nelson H. F. Beebe. A complete bibliography of ECMA standards. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, February 22, 2001. 45 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-e.html#ecmastd. This report is

updated frequently.

Beebe:2001:CBPg

[661] Nelson H. F. Beebe. A complete bibliography of publications in the ACM Symposia on Theory of Computing (STOC) for 2000–2009. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, April 4, 2001. 20 pp. URL http:

//www.math.utah.edu/pub/tex/bib/
index-table-e.html#stoc2000. This

report is updated frequently.

Beebe:2001:CBPe

[662] Nelson H. F. Beebe. A complete bibliography of publications in the ACM Symposia on Theory of Computing (STOC) for 1980–1989. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, April 4, 2001. 54 pp. URL http:

//www.math.utah.edu/pub/tex/bib/
index-table-e.html#stoc1980. This

report is updated frequently.

Beebe:2001:CBPf

[663] Nelson H. F. Beebe. A complete bibliography of publications in the ACM Symposia on Theory of Computing (STOC) for 1990–1999. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, April 4, 2001. 153 pp. URL http:

//www.math.utah.edu/pub/tex/bib/
index-table-e.html#stoc1990. This

report is updated frequently.

Beebe:2001:CBPi

[664] Nelson H. F. Beebe. A complete bibliography of publications in the IEEE symposia on Foundations of Computer Science (FOCS) for 2000– 2009. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, April 5, 2001. 12 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-e.html#focs2000. This report

is updated frequently.

Beebe:2001:BPIb

[665] Nelson H. F. Beebe. A bibliography of publications in the IEEE symposia on Foundations of Computer Science (FOCS) for 1980–1989. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, April 5, 2001. 40 pp. URL http:

                                        //www.math.utah.edu/pub/tex/bib/
                                        index-table-e.html#focs1980. This

report is updated frequently.

Beebe:2001:CBPh

[666] Nelson H. F. Beebe. A complete bibliography of publications in the IEEE symposia on Foundations of Computer Science (FOCS) for 1990– 1999. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, April 5, 2001. 133 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-e.html#focs1990. This report

is updated frequently.

Beebe:2001:BPNb

[667] Nelson H. F. Beebe. A bibliography of publications in Numerische Mathematik (2000–2009). Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, April 16, 2001. 27 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-n.html#nummath2000. This re-

port is updated frequently.

Beebe:2001:BPNa

[668] Nelson H. F. Beebe. A bibliography of publications in Numerische Mathematik (1959–1999). Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, April 16, 2001. 445 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-n.html#nummath. This report is

updated frequently.

Beebe:2001:CBAa

[669] Nelson H. F. Beebe. A complete bibliography of ACM Transactions on Software Engineering and Methodology. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, April 20, 2001. 30 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-n.html#tosem. This report is

updated frequently.

Beebe:2001:CBAg

[670] Nelson H. F. Beebe. A complete bibliography of Applied Statistics: 2000– 2009. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, April 23, 2001. 11 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-n.html#as2000. This report is

updated frequently.

Beebe:2001:CBAb

[671] Nelson H. F. Beebe. A complete bibliography of Applied Statistics: 1952– 1959. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, April 23, 2001. 34 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-n.html#as1950. This report is

updated frequently.

Beebe:2001:CBAc

[672] Nelson H. F. Beebe. A complete bibliography of Applied Statistics: 1960– 1969. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, April 23, 2001. 44 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-n.html#as1960. This report is

updated frequently.

Beebe:2001:CBAd

[673] Nelson H. F. Beebe. A complete bibliography of Applied Statistics: 1990– 1999. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, April 23, 2001. 78 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-n.html#as1990. This report is

updated frequently.

Beebe:2001:CBAe

[674] Nelson H. F. Beebe. A complete bibliography of Applied Statistics: 1980– 1989. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, April 23, 2001. 84 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-n.html#as1980. This report is

updated frequently.

Beebe:2001:CBAf

[675] Nelson H. F. Beebe. A complete bibliography of Applied Statistics: 1970– 1979. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, April 23, 2001. 93 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-n.html#as1970. This report is

updated frequently.

Beebe:2001:FHC

[676] Nelson H. F. Beebe. Fortran in the heavens (contd.). ACM Fortran Forum, 20(1):10, April 2001. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic). URL

http://www.math.utah.edu/~beebe/
publications/2001/fortran-forum-
20-1-10-april-2001.pdf.

Beebe:2001:CBPk

[677] Nelson H. F. Beebe. A complete bibliography of publications in SIAM Journal on Numerical Analysis: 2000– . Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, May 9, 2001. 38 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-s.html#siamjnumeranal2000.

This report is updated frequently.

Beebe:2001:CBPj

[678] Nelson H. F. Beebe. A complete bibliography of publications in SIAM Journal on Numerical Analysis: 1964– 1999. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, May 9, 2001. 457 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-s.html#siamjnumeranal.

Beebe:2001:BBI

[679] Nelson H. F. Beebe. A bibliography of the biennial IEEE Hot Topics in Operating Systems Workshops. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, May 28, 2001. 34 pp. URL http:

//www.math.utah.edu/pub/tex/bib/
index-table-h.html#hotos. This re-

port is updated frequently.

Beebe:2001:CBPl

[680] Nelson H. F. Beebe. A complete bibliography of publications in the Intel Technology Journal. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, June 1, 2001. 18 pp. URL http:

//www.math.utah.edu/pub/tex/bib/
index-table-i.html#intel-tech-j.

This report is updated frequently.

Beebe:2001:BPC

[681] Nelson H. F. Beebe. A bibliography of publications in The Computer Journal: 2000–2009. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, June 7, 2001. 16 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-i.html#compj2000. This re-

port is updated frequently.

Beebe:2001:CBPm

[682] Nelson H. F. Beebe. A complete bibliography of publications in the IEEE Annals of the History of Computing. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, July 12, 2001. 61 pp.

URL http://www.math.utah.edu/
pub/tex/bib/index-table-i.html#
ieeeannhistcomput. This report is up-

dated frequently.

Beebe:2001:BPA

[683] Nelson H. F. Beebe. A bibliography of publications in the Annals of the History of Computing. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, July 14, 2001. 15 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-i.html#annhistcomput. This

report is updated frequently.

Beebe:2001:CBPn

[684] Nelson H. F. Beebe. A complete bibliography of publications in Concurrency and Computation: Practice and Experience. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, July 25, 2001. 12 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-i.html#ccpe. This report is up-

dated frequently.

Beebe:2001:CBPo

[685] Nelson H. F. Beebe. A complete bibliography of publications in ACM Transactions on Design Automation of Electronic Systems. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, July 27, 2001. 30 pp. URL http:

//www.math.utah.edu/pub/tex/bib/
index-table-i.html#todaes. This

report is updated frequently.

Beebe:TB22-3-181

[686] Nelson Beebe. pdfTEX Panel. TUGboat, 22(3):181–187, September 2001. ISSN 0896-3207. URL

                                        http://www.math.utah.edu/~beebe/
                                        publications/2001/hyper-pdf-panel.
                                        pdf;   http://www.math.utah.edu/
                                        ~beebe/publications/2001/tb72beebe-
                                        pdf.pdf.

Beebe:TB22-3-220

[687] Nelson Beebe. Fonts Panel. TUGboat, 22(3):220–227, September 2001. ISSN

                                        0896-3207. URL http://www.math.
                                        utah.edu/~beebe/publications/2001/
                                        hyper-font-panel.pdf;   http://
                                        www.math.utah.edu/~beebe/publications/
                                        2001/tb72beebe-fonts.pdf.

Beebe:2001:FBP

[688] Nelson H. F. Beebe. Filesystem backup performance at the University of Utah Mathematics Department. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, October 12, 2001. 6 pp. URL http://

                                        www.math.utah.edu/~beebe/amanda/
                                        amanda-2001.pdf. This is an update

of [641].

Beebe:2001:UCC

[689] Nelson H. F. Beebe. Using C and C++ with Fortran. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, November 27, 2001. 23 pp. URL http:

                                        //www.math.utah.edu/software/c-
                                        with-fortran.html.

Beebe:2001:IFP

[690] Nelson H. F. Beebe. IEEE 754 floating-point test software. World-Wide Web document., December 1,

                                        2001. URL http://www.math.utah.
                                        edu/~beebe/software/ieee/.

Beebe:2001:ENL

[691] Nelson H. F. Beebe. EISPACK: numerical library for eigenvalue and eigenvector solutions. World-Wide Web docu-

ment., 2001. URL http://www.math.
utah.edu/software/eispack.html.

Beebe:2001:GSL

[692] Nelson H. F. Beebe. GNU Scientific Library. World-Wide Web document.,

2001. URL http://www.math.utah.
edu/software/gsl.html.

Beebe:2001:GUU

[693] Nelson H. F. Beebe. Ghostscript at the University of Utah Mathematics Department. World-Wide Web docu-

ment., 2001. URL http://www.math.
utah.edu/software/gs.html.

Beebe:2001:LFF

[694] Nelson H. F. Beebe. Lahey/Fujitsu Fortran 95 compiler: Linux Express v5.5. World-Wide Web document.,

2001. URL http://www.math.utah.
edu/software/lahey.html.

Beebe:2001:LLB

[695] Nelson H. F. Beebe. Linking LAPACK and BLAS libraries with Fortran and C code. World-Wide Web document.,

2001. URL http://www.math.utah.
edu/software/lapack.html.

Beebe:2001:LNL

[696] Nelson H. F. Beebe. LINPACK: numerical library for linear equation solution. World-Wide Web document.,

2001. URL http://www.math.utah.
edu/software/linpack.html.

Beebe:2001:LTI

[697] Nelson H. F. Beebe. The Lucent Technologies Inc. PORT Library. World-Wide Web document.,

                                        2001. URL http://www.math.utah.
                                        edu/software/port.html.

Beebe:2001:MNL

[698] Nelson H. F. Beebe. MINPACK: numerical library for function minimization and least-squares solutions. World-Wide Web document., 2001.

                                        URL http://www.math.utah.edu/
                                        software/minpack.html.

Beebe:2001:NSD

[699] Nelson H. F. Beebe. NAG software directory. World-Wide Web document.,

                                        2001. URL http://www.math.utah.
                                        edu/software/nag.html.

Beebe:2001:NSI

[700] Nelson H. F. Beebe. NAG software installer notes. World-Wide Web docu-

                                        ment., 2001. URL http://www.math.
                                        utah.edu/software/nag-ins.html.

Beebe:2001:NSU

[701] Nelson H. F. Beebe. NAG software user notes. World-Wide Web document.,

                                        2001. URL http://www.math.utah.
                                        edu/software/nag-uns.html.

Beebe:2001:PGI

[702] Nelson H. F. Beebe. Portland Group, Inc. compilers. World-Wide Web docu-

                                        ment., 2001. URL http://www.math.
                                        utah.edu/software/pgi.html.

Beebe:2001:RSA

[703] Nelson H. F. Beebe. Reduce symbolic algebra package. World-Wide Web doc-

ument., 2001. URL http://www.math.
utah.edu/software/reduce.html.

Beebe:2001:SCM

[704] Nelson H. F. Beebe. Sun compiler manuals. World-Wide Web document.,

2001. URL http://www.math.utah.
edu/software/sun-manuals.html.

Beebe:2001:SCU

[705] Nelson H. F. Beebe. Software catalog: University of Utah Mathematics Department. World-Wide Web docu-

ment., 2001. URL http://www.math.
utah.edu/software/index.html.

Beebe:2001:SFC

[706] Nelson H. F. Beebe. N.A.Software Fortran 95 compiler. World-Wide Web document., 2001. URL http://www.

math.utah.edu/software/nasoftware.
html.

Beebe:2001:UNC

[707] Nelson H. F. Beebe. Using the NAG C libraries at the University of Utah Mathematics Department. World-Wide Web document., 2001.

URL http://www.math.utah.edu/
software/nag-libraries-c.html.

Beebe:2001:UNFa

[708] Nelson H. F. Beebe. Using the NAG Fortran libraries at the University of Utah Mathematics Department from C and C++. World-Wide Web docu-

ment., 2001. URL http://www.math.
utah.edu/software/nag-libraries-
c-to-fortran.html.

Beebe:2001:UNFb

[709] Nelson H. F. Beebe. Using the NAG Fortran libraries at the University of Utah Mathematics Department. World-Wide Web docu-

ment., 2001. URL http://www.math.
utah.edu/software/nag-libraries-
fortran.html.

Beebe:2001:UNFc

[710] Nelson H. F. Beebe. Using the NAG Fortran and C libraries at the University of Utah Mathematics Department. World-Wide Web document.,

2001. URL http://www.math.utah.
edu/software/nag-libraries.html.

Beebe:2002:HEH

[711] Nelson H. F. Beebe, Brian W. Kernighan, and Rob Pike. hoc 7: An extensible high-order calculator. World-Wide Web document., January

4, 2002. URL http://www.math.utah.
edu/pub/hoc/.

Beebe:2002:CBAa

[712] Nelson H. F. Beebe. A complete bibliography of ACM Transactions on Internet Technology. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, January 30, 2002. 3 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-t.html#toit.

Beebe:2002:BPLc

[713] Nelson H. F. Beebe. A bibliography of papers in Lecture Notes in Computer Science (2001), part 2 of 4. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, February 1, 2002. 460 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-l.html#lncs2001b.

Beebe:2002:BPLe

[714] Nelson H. F. Beebe. A bibliography of papers in Lecture Notes in Computer Science (2001), part 4 of 4. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, February 1, 2002. 461 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-l.html#lncs2001d.

Beebe:2002:BPLb

[715] Nelson H. F. Beebe. A bibliography of papers in Lecture Notes in Computer Science (2001), part 1 of 4. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, February 1, 2002. 463 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-l.html#lncs2001a.

Beebe:2002:BPLd

[716] Nelson H. F. Beebe. A bibliography of papers in Lecture Notes in Computer Science (2001), part 3 of 4. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, February 1, 2002. 463 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-l.html#lncs2001c.

Beebe:2002:CBAb

[717] Nelson H. F. Beebe. A complete bibliography of ACM ForTec Forum and ACM Fortran Forum. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, February 7, 2002. 58 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-f.html#fortran-forum.

Beebe:2002:BPLa

[718] Nelson H. F. Beebe. A bibliography of papers in Lecture Notes in Computer Science (2000). Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, February 14, 2002. 326 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-l.html#lncs2000.

Beebe:2002:CBPa

[719] Nelson H. F. Beebe. A complete bibliography of publications in the ACM Symposia on Theory of Computing (STOC) for 1960–1969. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, February 20, 2002. 5 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-s.html#stoc1960.

Beebe:2002:BS

[720] Nelson H. F. Beebe. A bibliography of Supercomputing ’2001. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, February 20, 2002. 12 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-s.html#supercomputing2001.

Beebe:2002:CBPb

[721] Nelson H. F. Beebe. A complete bibliography of publications in the ACM Symposia on Theory of Computing (STOC) for 1970–1979. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, February 20, 2002. 35 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-s.html#stoc1970.

Beebe:2002:BIJb

[722] Nelson H. F. Beebe. A bibliography of the International Journal of Computer Processing of Oriental Languages. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, February 25, 2002. 9 pp. URL http:

//www.math.utah.edu/pub/tex/bib/
index-table-f.html#ijcpol.

Beebe:2002:BIJa

[723] Nelson H. F. Beebe. A bibliography of the International Journal of Computational Geometry and Applications. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, February 25, 2002. 17 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-f.html#ijcga.

Beebe:2002:BIJd

[724] Nelson H. F. Beebe. A bibliography of the International Journal of High Speed Computing. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, February 25, 2002. 30 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-f.html#ijhsc.

Beebe:2002:BIJc

[725] Nelson H. F. Beebe. A bibliography of the International Journal of Foundations of Computer Science. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, February 25, 2002. 48 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-f.html#ijfcs.

Beebe:2002:CBI

[726] Nelson H. F. Beebe. A complete bibliography of the International Journal of Image and Graphics. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, February 26, 2002. 5 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-f.html#ijig.

Beebe:2002:BPF

[727] Nelson H. F. Beebe. A bibliography of publications in Future Generation Computer Systems. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, February 27, 2002. 80 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-f.html#futgencompsys.

Beebe:2002:BBAa

[728] Nelson H. F. Beebe. A bibliography of books about the Structured Query Language (SQL). Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, March 9, 2002. 29 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-s.html#sqlbooks.

Beebe:2002:BPAa

[729] Nelson H. F. Beebe. A bibliography of publications in ACM SIGGRAPH Computer Graphics: 2000–2009. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, March 13, 2002. 19 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-s.html#siggraph2000.

Beebe:2002:BPAb

[730] Nelson H. F. Beebe. A bibliography of publications about SGML, the Standard Generalized Markup Language: 2000–2009. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, March 28, 2002. 21 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-s.html#sgml2000. This report

is updated frequently.

Beebe:2002:CBPc

[731] Nelson H. F. Beebe. A complete bibliography of publications in Scientific Programming. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, March 28, 2002. 39 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-s.html#sciprogram. This re-

port is updated frequently.

Beebe:2002:BIF

[732] Nelson H. F. Beebe. Back issues of Fortran Forum. ACM Fortran Forum, 21(1):6–7, April 2002. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Beebe:2002:CBPd

[733] Nelson H. F. Beebe. A complete bibliography of publications in Journal of the Institute of Mathematics and Its Applications. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, April 5, 2002. 89 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-j.html#jinstmathappl. This

report is updated frequently.

Beebe:2002:BBAb

[734] Nelson H. F. Beebe. A bibliography of books about the Internet and networking: 2000–2009. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, April 30, 2002. 115 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-i.html#internet2000.

Beebe:2002:CBPe

[735] Nelson H. F. Beebe. A complete bibliography of publications in the C/C++ Users Journal: 2000–2009. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, May 15, 2002. 38 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-c.html#cccuj2000. This re-

port is updated frequently.

Beebe:2002:Ca

[736] Nelson H. F. Beebe. Computation of expm1(x) = exp(x) − 1. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, June 13, 2002. iii +

17 pp. URL http://www.math.utah.
edu/~beebe/reports/expm1.pdf;
http://www.math.utah.edu/~beebe/
reports/expm1.ps.gz.

Beebe:2002:Cb

[737] Nelson H. F. Beebe. Computation of log1p(x) = log(1 + x). Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, June 18, 2002. 3 +

11 pp. URL http://www.math.utah.
edu/~beebe/reports/log1p.pdf;
http://www.math.utah.edu/~beebe/
reports/log1p.ps.gz.

Beebe:2002:NPR

[738] Nelson H. F. Beebe. Notes on pseudo-random numbers. Web site, July 20, 2002. URL http://www.

math.utah.edu/~beebe/talks/2002/
pseudo-random-numbers-2002. This

document is a handout for undergraduate colloquia that have been presented several times, each time with updates. The latest edition is [1411].

Beebe:2002:BPLi

[739] Nelson H. F. Beebe. A bibliography of papers in Lecture Notes in Computer Science (2002), part 4 of 5. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, September 11, 2002. 272 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-l.html#lncs2002d.

Beebe:2002:BPLg

[740] Nelson H. F. Beebe. A bibliography of papers in Lecture Notes in Computer Science (2002), part 2 of 5. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, September 11, 2002. 317 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-l.html#lncs2002b.

Beebe:2002:BPLh

[741] Nelson H. F. Beebe. A bibliography of papers in Lecture Notes in Computer Science (2002), part 3 of 5. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, September 11, 2002. 322 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-l.html#lncs2002c.

Beebe:2002:BPLf

[742] Nelson H. F. Beebe. A bibliography of papers in Lecture Notes in Computer Science (2002), part 1 of 5. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, September 11, 2002. 326 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-l.html#lncs2002a.

Beebe:2002:BPU

[743] Nelson H. F. Beebe. A bibliography of publications of the USENIX Association: 2000–2009. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, October 15, 2002. 72 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-u.html#usenix2000.

Beebe:2002:BPI

[744] Nelson H. F. Beebe. A bibliography of publications in IEEE Distributed Systems Online. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, October 23, 2002. 19 pp.

                                        URL http://www.math.utah.edu/
                                        pub/tex/bib/index-table-i.html#
                                        ieeedistribsystonline.

Beebe:2002:CBPf

[745] Nelson H. F. Beebe. A complete bibliography of publications in the Annals of the History of Computing. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, November 3, 2002. 184 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-a.html#annhistcomput. This

report is updated frequently.

Beebe:2002:CBAc

[746] Nelson H. F. Beebe. A complete bibliography of ACM Transactions on Asian Language Information Processing. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, November 5, 2002. 3 pp. URL http:

//www.math.utah.edu/pub/tex/bib/
index-table-t.html#talip.

Beebe:2002:CBPg

[747] Nelson H. F. Beebe. A complete bibliography of publications in Journal of Statistical Software. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, November 18, 2002. 17 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-j.html#jstatsoft. This re-

port is updated frequently.

Beebe:2002:CBL

[748] Nelson H. F. Beebe. A complete bibliography of the LMS Journal of Computation and Mathematics. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, November 29, 2002. 8 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-l.html#lms-j-comput-math.

Beebe:2002:BPLj

[749] Nelson H. F. Beebe. A bibliography of papers in Lecture Notes in Computer Science (2002), part 5 of 5. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, December 1, 2002. 601 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-l.html#lncs2002e.

Beebe:2002:BPLk

[750] Nelson H. F. Beebe. A bibliography of papers in Lecture Notes in Computer Science (2003), part 1 of 4. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, December 1, 2002. 601 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-l.html#lncs2003a.

Beebe:2002:LBP

[751] Nelson H. F. Beebe. A limited bibliography of pseudorandom number generation, sampling, selection, and distribution. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, December 4, 2002. 13 pp. URL http:

                                        //www.math.utah.edu/pub/tex/bib/
                                        index-table-p.html#prng.

Beebe:2003:PRN

[752] Nelson H. F. Beebe. Pseudo-random numbers: [mostly] a line [of code] at a time. Lecture notes, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, January 14, 2003. 22 pp.

Beebe:2003:CBPa

[753] Nelson H. F. Beebe. A complete bibliography of publications in IEEE Security & Privacy. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, April 4, 2003. 13 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-i.html#ieeesecpriv. This re-

port is updated frequently.

Beebe:2003:CBAa

[754] Nelson H. F. Beebe. A complete bibliography of ACM Transactions on Embedded Computing Systems (TECS). Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, August 7, 2003. 5 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-t.html#tecs.

Beebe:2003:CBAb

[755] Nelson H. F. Beebe. A complete bibliography of ACM Transactions on Modeling and Computer Simulation. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, August 7, 2003. 31 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-t.html#tomacs.

Beebe:2003:CBPb

[756] Nelson H. F. Beebe. A complete bibliography of publications in ACM Queue: Tomorrow’s Computing Today. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, August 23, 2003. 10 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-

                                        table-q.html#queue. This report is

updated frequently.

Beebe:2003:CBPc

[757] Nelson H. F. Beebe. A complete bibliography of publications in Multiscale Modeling & Simulation. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, September 30, 2003. 6 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-s.html#siammms.

Beebe:2003:CBPd

[758] Nelson H. F. Beebe. A complete bibliography of publications in SIAM Journal on Applied Dynamical Systems. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, September 30, 2003. 7 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-s.html#siamjappldynsyst.

Beebe:2003:CBPe

[759] Nelson H. F. Beebe. A complete bibliography of publications in the SIAM Journal on Optimization. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, September 30, 2003. 114 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-s.html#siamjopt.

Beebe:2003:BSa

[760] Nelson H. F. Beebe. A bibliography of Supercomputing ’2002. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, November 26, 2003. 14 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-s.html#supercomputing2002.

This report is updated frequently.

Beebe:2003:BSb

[761] Nelson H. F. Beebe. A bibliography of Supercomputing ’2003. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, November 26, 2003. 15 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-s.html#supercomputing2003.

This report is updated frequently.

Beebe:2003:CBPf

[762] Nelson H. F. Beebe. A complete bibliography of publications in Parallel and Distributed Computing Practices. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, December 19, 2003. 26 pp.

URL http://www.math.utah.edu/
pub/tex/bib/index-table-p.html#
paralleldistcomppract. This report

is updated frequently.

Beebe:2004:CBPa

[763] Nelson H. F. Beebe. A complete bibliography of publications in Applied Mathematics and Computation: 2000–2004. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, January 9, 2004. 23 pp.

URL http://www.math.utah.edu/
pub/tex/bib/index-table-a.html#
applmathcomput2000.

Beebe:2004:CBPb

[764] Nelson H. F. Beebe. A complete bibliography of publications in the Journal of Grid Computing. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, January 14, 2004. 3 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-j.html#jgridcomp.

Beebe:2004:CBPc

[765] Nelson H. F. Beebe. A complete bibliography of publications in the Journal of Digital Information. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, January 14, 2004. 16 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-j.html#jdigitalinfo.

Beebe:2004:CBPd

[766] Nelson H. F. Beebe. A complete bibliography of publications in the ACM Journal on Educational Resources in Computing (JERIC). Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, February 2, 2004. 8 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-j.html#jeric.

Beebe:2004:CBPe

[767] Nelson H. F. Beebe. A complete bibliography of publications in ACM Transactions on Architecture and Code Optimization. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, February 6, 2004. 8 pp. URL http:

//www.math.utah.edu/pub/tex/bib/
index-table-t.html#taco.

Beebe:2004:JPF

[768] Nelson H. F. Beebe. Java programming: Fun with Fibonacci. World-Wide Web document., March 2004.

URL http://www.math.utah.edu/
~beebe/software/java/fibonacci/

index.html. This report summarizes the origin of the Fibonacci sequence, giving the full Latin text from the original book written in 1202 (not previously available on the Web). Computation of the Fibonacci sequence, and its term ratios, is implemented in at least 45 different programming languages. The report comments on the relative difficulty of the task in some of those languages, and on their suitability for numerical computation. It also provides a complete floating-point formatted output package for Java.

Beebe:2004:BBP

[769] Nelson H. F. Beebe. A bibliography of books by past and present University of Utah Mathematics Department members. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, July 3, 2004. 10 pp.

URL http://www.math.utah.edu/
pub/tex/bib/index-table-u.html#
utah-math-dept-books.

Beebe:2004:CBPf

[770] Nelson H. F. Beebe. A complete bibliography of publications in the Communications of the ACM : 1958–1959. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, July 14, 2004. 15 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-

                                        table-c.html#cacm1950. This report

is updated frequently.

Beebe:2004:BPG

[771] Nelson H. F. Beebe. A bibliography of publications of George Marsaglia. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, August 5, 2004. 3 pp. URL http:

                                        //www.math.utah.edu/pub/bibnet/
                                        authors/m/marsaglia-george.bib.

This report is updated frequently.

Beebe:2004:CBAa

[772] Nelson H. F. Beebe. A complete bibliography of ACM Transactions on Architecture and Code Optimization. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, August 5, 2004. 3 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-t.html#taco. This report is up-

dated frequently.

Beebe:2004:CBAb

[773] Nelson H. F. Beebe. A complete bibliography of ACM Transactions on Speech and Language Processing (TSLP). Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, November 22, 2004. 1 pp. URL http:

                                        //www.math.utah.edu/pub/tex/bib/
                                        index-table-t.html#tslp. This re-

port is updated frequently.

Beebe:2004:CBIb

[774] Nelson H. F. Beebe. A complete bibliography of the International Journal on Software Tools for Technology Transfer (STTT). Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, November 22, 2004. 1 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-t.html#sttt. This report is up-

dated frequently.

Beebe:2004:CBIa

[775] Nelson H. F. Beebe. A complete bibliography of IEEE/ACM Transactions on Computational Biology and Bioinformatics. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, November 22, 2004. 3 pp. URL http:

//www.math.utah.edu/pub/tex/bib/
index-table-t.html#tcbb. This re-

port is updated frequently.

Beebe:2004:MPQ

[776] Nelson H. F. Beebe. Maple plot quality and PostScript conformance. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, December 11, 2004. 21 pp. URL

http://www.math.utah.edu/~beebe/
reports/maple-plot-bugs.pdf. This

report is updated frequently.

Beebe:2004:SBP

[777] Nelson H. F. Beebe. A selected bibliography of publications by, and about, Albert Einstein. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, December 17, 2004. 98 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-e.html#einstein. This report

is updated frequently.

Beebe:2004:YTM

[778] Nelson H. F. Beebe. 25 years of TEX and : Looking back and looking forward: TUG 2003 keynote address. TUGboat, 25(1):7–30, 2004. ISSN 0896-3207. URL http://www.

                                        math.utah.edu/~beebe/talks/2003/
                                        tug2003/.

Beebe:2004:BT

[779] Nelson H. F. Beebe. A bibliographer’s toolbox. TUGboat, 25(1):89–104, 2004. ISSN 0896-3207. Presented at Practical TEX ’2004, Training and Techniques Workshops and Presentations on LaTEX, TEX, ConTEXt, AMS-LaTEX July 19–22, 2004, Holiday Inn Fisherman’s Wharf San Francisco, California.

Beebe:2004:TB

[780] Nelson Beebe, Wendy McKay, and Ross Moore. TUG at Bay. TUGboat, 24(2):291–293, 2004. ISSN 0896-3207.

Beebe:2005:ECW

[781] Nelson H. F. Beebe. The extended Cody/Waite elementary function library: mathcw. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, March 1, 2005. 50 (est.)

                                        pp. URL http://www.math.utah.
                                        edu/~beebe/reports/mathcw.ltx.

Beebe:2005:BPL

[782] Nelson H. F. Beebe. A bibliography of papers in Lecture Notes in Computer Science (2004), part 1 of 1. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, March 10, 2005. 9 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-l.html#lncs2004a.

Beebe:2005:CBAb

[783] Nelson H. F. Beebe. A complete bibliography of ACM Transactions on Storage. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, March 10, 2005. 9 pp. URL http:

//www.math.utah.edu/pub/tex/bib/
index-table-t.html#tos.

Beebe:2005:CBAa

[784] Nelson H. F. Beebe. A complete bibliography of the ACM Transactions on Multimedia Computing, Communications and Applications. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, April 14, 2005. 1 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-t.html#tomccap. This report is

updated frequently.

Beebe:2005:BPAa

[785] Nelson H. F. Beebe. A bibliography of publications about the lcc C compiler. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, April 18, 2005. 6 pp. URL http:

//www.math.utah.edu/pub/tex/bib/
index-table-l.html#lcc.

Beebe:2005:CBL

[786] Nelson H. F. Beebe. A complete bibliography of the LAPACK Working Notes. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, April 22, 2005. 33 pp. URL http:

//www.math.utah.edu/pub/tex/bib/
index-table-l.html#lawn.

Beebe:2005:BPW

[787] Nelson H. F. Beebe. A bibliography of publications of William M. Kahan. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, April 25, 2005. 23 pp. URL http:

//www.math.utah.edu/pub/bibnet/
authors/k/kahan-william-m.bib.

Beebe:2005:CBPa

[788] Nelson H. F. Beebe. A complete bibliography of the publications of John von Neumann. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, June 9, 2005. 75

pp. URL http://www.math.utah.
edu/pub/bibnet/authors/v/index.
html#von-neumann-john.

Robbins:2005:BCS

[789] Arnold Robbins and Nelson H. F. Beebe. A bibliography of Classic Shell Scripting. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, June 11, 2005. 12 pp. URL http:

//www.math.utah.edu/pub/tex/bib/
index-table-c.html#css.

Beebe:2005:BPAb

[790] Nelson H. F. Beebe. A bibliography of publications about the REXX scripting language. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, July 1, 2005. 17 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-r.html#rexx.

Beebe:2005:CBPb

[791] Nelson H. F. Beebe. A complete bibliography of publications in Numerische Mathematik (2000–2009). Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, July 5, 2005. 90 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-n.html#nummath2000. This re-

port is updated frequently.

Beebe:2005:CBC

[792] Nelson H. F. Beebe. A complete bibliography of Computer Animation and Virtual Worlds. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, July 6, 2005. 9 pp.

URL http://www.math.utah.edu/
pub/tex/bib/index-table-c.html#
computanimvirtualworlds.

Beebe:2005:CBPi

[793] Nelson H. F. Beebe. A complete bibliography of publications in Lisp and Symbolic Computation. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, July 6, 2005. 18 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-l.html#lispsymbcomput.

Beebe:2005:CBPe

[794] Nelson H. F. Beebe. A complete bibliography of publications in Higher-Order Symbolic Computation. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, July 6, 2005. 22 pp.

                                        URL http://www.math.utah.edu/
                                        pub/tex/bib/index-table-h.html#
                                        higherordersymbcomput.

Beebe:2005:CBPf

[795] Nelson H. F. Beebe. A complete bibliography of publications in the International Journal of Parallel Programming. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, July 6, 2005. 83 pp.

                                        URL http://www.math.utah.edu/
                                        pub/tex/bib/index-table-i.html#
                                        intjparallelprogram. This report is

updated frequently.

Beebe:2005:CBPd

[796] Nelson H. F. Beebe. A complete bibliography of publications in Designs, Codes, and Cryptography. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, July 6, 2005. 107 pp.

                                        URL http://www.math.utah.edu/
                                        pub/tex/bib/index-table-d.html#
                                        designscodescryptogr.

Beebe:2005:CBPc

[797] Nelson H. F. Beebe. A complete bibliography of publications in The Journal of Supercomputing. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, July 6, 2005. 111 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-j.html#jsuper. This report is

updated frequently.

Beebe:2005:CBPh

[798] Nelson H. F. Beebe. A complete bibliography of publications in J.UCS:

Journal of Universal Computer Science. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, July 6, 2005. 141 pp. URL http:

//www.math.utah.edu/pub/tex/bib/
index-table-j.html#jucs. This re-

port is updated frequently.

Beebe:2005:CBPg

[799] Nelson H. F. Beebe. A complete bibliography of publications in Numerical Algorithms. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, July 6, 2005. 172 pp.

URL http://www.math.utah.edu/
pub/tex/bib/index-table-n.html#

numeralgorithms. This report is updated frequently.

Beebe:2005:CBPj

[800] Nelson H. F. Beebe. A complete bibliography of publications in Future Generation Computer Systems. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, July 15, 2005. 201 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-f.html#futgencompsys.

Beebe:2005:VM

[801] Nelson H. F. Beebe. Virtual machines. World-Wide Web document., August 1,

2005. URL http://www.math.utah.
edu/~beebe/vm.html.

Beebe:2005:AF

[802] Nelson H. F. Beebe. Accounts FAQ. World-Wide Web frequentlyasked question document., September

                                        2005. URL http://www.math.utah.
                                        edu/faq/accounts. This report is up-

dated frequently.

Beebe:2005:BJF

[803] Nelson H. F. Beebe. Batch jobs FAQ. World-Wide Web frequentlyasked question document., September

                                        2005. URL http://www.math.utah.

edu/faq/batch/. This report is updated frequently.

Beebe:2005:CFa

[804] Nelson H. F. Beebe. Color FAQ. World-Wide Web frequently-asked question document., September 2005.

                                        URL http://www.math.utah.edu/

faq/color/. This report is updated frequently.

Beebe:2005:CFb

[805] Nelson H. F. Beebe. Compilers FAQ. World-Wide Web frequentlyasked question document., September

                                        2005. URL http://www.math.utah.
                                        edu/faq/compilers/. This report is

updated frequently.

Beebe:2005:CFF

[806] Nelson H. F. Beebe. Computing facilities FAQ. World-Wide Web frequentlyasked question document., September

                                        2005. URL http://www.math.utah.

edu/faq/s. This report is updated frequently.

Beebe:2005:DAF

[807] Nelson H. F. Beebe. Dialup access FAQ. World-Wide Web frequentlyasked question document., September

                                        2005. URL http://www.math.utah.

edu/faq/dialup/. This report is updated frequently.

Beebe:2005:DUF

[808] Nelson H. F. Beebe. Disk usage FAQ. World-Wide Web frequentlyasked question document., September

2005. URL http://www.math.utah.
edu/faq/disk-usage/. This report is

updated frequently.

Beebe:2005:EMF

[809] Nelson H. F. Beebe. Electronic mail FAQ. World-Wide Web frequentlyasked question document., September

2005. URL http://www.math.utah.

edu/faq/email/. This report is updated frequently.

Beebe:2005:EWP

[810] Nelson H. F. Beebe. Editors (and word processors) FAQ. World-Wide Web frequently-asked question document., September 2005. URL http://www.

math.utah.edu/faq/editors/. This

report is updated frequently.

Beebe:2005:FAF

[811] Nelson H. F. Beebe. FTP access FAQ. World-Wide Web frequentlyasked question document., September

2005. URL http://www.math.utah.

edu/faq/ftp/. This report is updated frequently.

Beebe:2005:FFa

[812] Nelson H. F. Beebe. Files FAQ. World-Wide Web frequently-asked question document., September 2005.

URL http://www.math.utah.edu/

faq/files/. This report is updated frequently.

Beebe:2005:FFb

[813] Nelson H. F. Beebe. Fonts FAQ. World-Wide Web frequently-asked question document., September 2005.

URL http://www.math.utah.edu/

faq/fonts/. This report is updated frequently.

Beebe:2005:HAF

[814] Nelson H. F. Beebe. Hardware available FAQ. World-Wide Web frequentlyasked question document., September

2005. URL http://www.math.utah.
edu/faq/hardware/. This report is up-

dated frequently.

Beebe:2005:LAF

[815] Nelson H. F. Beebe. Login access FAQ. World-Wide Web frequentlyasked question document., September

2005. URL http://www.math.utah.

edu/faq/login/. This report is updated frequently.

Beebe:2005:MF

[816] Nelson H. F. Beebe. Movies FAQ. World-Wide Web frequently-asked question document., September 2005.

URL http://www.math.utah.edu/

faq/movies/. This report is updated frequently.

Beebe:2005:MON

[817] Nelson H. F. Beebe. Matlab and Octave (numerical linear algebra) FAQ. World-Wide Web frequentlyasked question document., September

2005. URL http://www.math.utah.

edu/faq/matlab/. This report is updated frequently.

Beebe:2005:MPA

[818] Nelson H. F. Beebe. Multiple-precision arithmetic FAQ. World-Wide Web frequently-asked question document., September 2005. URL http://www.

math.utah.edu/faq/mp/. This report

is updated frequently.

Beebe:2005:MSAa

[819] Nelson H. F. Beebe. Maple (symbolic algebra) FAQ. World-Wide Web frequently-asked question document., September 2005. URL http://www.

math.utah.edu/faq/maple/. This re-

port is updated frequently.

Beebe:2005:MSAb

[820] Nelson H. F. Beebe. Mathematica (symbolic algebra) FAQ. World-Wide Web frequently-asked question document., September 2005.

URL http://www.math.utah.edu/
faq/mathematica/. This report is up-

dated frequently.

Beebe:2005:MSAc

[821] Nelson H. F. Beebe. Maxima (symbolic algebra) FAQ. World-Wide Web frequently-asked question document., September 2005. URL http://www.

math.utah.edu/faq/maxima/. This

report is updated frequently.

Beebe:2005:MSAd

[822] Nelson H. F. Beebe. MuPAD (symbolic algebra) FAQ. World-Wide Web frequently-asked question document., September 2005. URL http://www.

math.utah.edu/faq/mupad/. This re-

port is updated frequently.

Beebe:2005:PF

[823] Nelson H. F. Beebe. Printing FAQ. World-Wide Web frequentlyasked question document., September

2005. URL http://www.math.utah.

                                        edu/faq/printing/. This report is up-

dated frequently.

Beebe:2005:PGS

[824] Nelson H. F. Beebe. PARI/GP (symbolic algebra) FAQ. World-Wide Web frequently-asked question document., September 2005. URL http://www.

                                        math.utah.edu/faq/pari/. This re-

port is updated frequently.

Beebe:2005:PRS

[825] Nelson H. F. Beebe. S-Plus and R (statistics) FAQ. World-Wide Web frequently-asked question document., September 2005. URL http://www.

                                        math.utah.edu/faq/s-plus/. This

report is updated frequently.

Beebe:2005:RSA

[826] Nelson H. F. Beebe. Reduce (symbolic algebra) FAQ. World-Wide Web frequently-asked question document., September 2005. URL http://www.

                                        math.utah.edu/faq/reduce/. This

report is updated frequently.

Beebe:2005:SAF

[827] Nelson H. F. Beebe. Software available FAQ. World-Wide Web frequentlyasked question document., September

                                        2005. URL http://www.math.utah.
                                        edu/faq/software/. This report is up-

dated frequently.

Beebe:2005:SSF

[828] Nelson H. F. Beebe. SAS (statistics) FAQ. World-Wide Web frequentlyasked question document., September

                                        2005. URL http://www.math.utah.

edu/faq/sas/. This report is updated frequently.

Beebe:2005:TMF

[829] Nelson H. F. Beebe. TeX and METAFONT FAQ. World-Wide Web frequently-asked question document., September 2005. URL http://www.

math.utah.edu/faq/texmf/. This re-

port is updated frequently.

Beebe:2005:WAF

[830] Nelson H. F. Beebe. Web access FAQ. World-Wide Web frequentlyasked question document., September

2005. URL http://www.math.utah.

edu/faq/web/. This report is updated frequently.

Beebe:2005:WNAa

[831] Nelson H. F. Beebe. Wired network access FAQ. World-Wide Web frequentlyasked question document., September

2005. URL http://www.math.utah.

edu/faq/wired/. This report is updated frequently.

Beebe:2005:WNAb

[832] Nelson H. F. Beebe. Wireless network access FAQ. World-Wide Web frequently-asked question document., September 2005. URL http://www.

math.utah.edu/faq/wireless/. This

report is updated frequently.

Beebe:2005:CBPk

[833] Nelson H. F. Beebe. A complete bibliography of publications in the ACM Journal on Emerging Technologies in Computing Systems (JETC). Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, September 17, 2005. 2 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-j.html#jetc.

Beebe:2005:BPAc

[834] Nelson H. F. Beebe. A bibliography of publications about the MuPAD symbolic algebra language. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, October 18, 2005. 26 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-m.html#mupad.

Beebe:2005:CBPl

[835] Nelson H. F. Beebe. A complete bibliography of publications of Beresford N. Parlett. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, November 12, 2005. 53

pp. URL http://www.math.utah.
edu/pub/bibnet/authors/p.

Beebe:2005:BPAd

[836] Nelson H. F. Beebe. A bibliography of publications of the ACM Turing Awards (1966–date). Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, November 19, 2005. 10 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-a.html#acm-turing-awards.

Beebe:2005:BPAe

[837] Nelson H. F. Beebe. A bibliography of publications of Alan Mathison Turing. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, November 20, 2005. 28

pp. URL http://www.math.utah.
edu/pub/bibnet/authors/t/index.
html#turing-alan-mathison.

Beebe:2005:BPAf

[838] Nelson H. F. Beebe. A bibliography of publications of George Elmer Forsythe. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, November 22, 2005. 17

pp. URL http://www.math.utah.
edu/pub/bibnet/authors/f/index.
html#forsythe-george-elmer.

Beebe:2005:CFD

[839] Nelson H. F. Beebe. Cryptography, freedom, and democracy: How basic science affects everyone. Web site, November 30, 2005. URL http://www.

math.utah.edu/~beebe/talks/2005/
cryptography-and-freedom/. This

document is from a talk that has been given multiple times in undergraduate colloquia at two universities, and in the Salt Lake City public library.

Beebe:2005:CBAc

[840] Nelson H. F. Beebe. A complete bibliography of ACM Transactions on Algorithms. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, December 13, 2005. 4 pp. URL http:

//www.math.utah.edu/pub/tex/bib/
index-table-r.html#talg.

Beebe:2005:CBAd

[841] Nelson H. F. Beebe. A complete bibliography of ACM Transactions on Applied Perception. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, December 17, 2005. 11 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-t.html#tap.

Beebe:2005:CBAe

[842] Nelson H. F. Beebe. A complete bibliography of ACM Transactions on Sensor Networks. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, December 27, 2005. 3 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-t.html#tosn.

Beebe:2005:BSa

[843] Nelson H. F. Beebe. A bibliography of Supercomputing ’2004. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, December 27, 2005. 11 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-s.html#supercomputing2004.

This report is updated frequently.

Beebe:2005:BSb

[844] Nelson H. F. Beebe. A bibliography of Supercomputing ’2005. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, December 27, 2005. 13 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-s.html#supercomputing2005.

This report is updated frequently.

Beebe:2005:KAD

[845] Nelson H. F. Beebe. Keynote address: The design of TEX and - : A retrospective. TUGboat, 26 (1):33–51, ???? 2005. ISSN 0896-3207.

                                        URL http://www.math.utah.edu/~
                                        beebe/talks/2005/pt2005/pt2005.
                                        pdf;   http://www.math.utah.edu/
                                        ~beebe/talks/2005/pt2005/pt2005.
                                        ps;   http://www.math.utah.edu/~

beebe/talks/2005/pt2005/pt2005.

ps.gz. Proceedings of the Practical TeX 2005 conference, Chapel Hill, NC, June 14–17, 2005.

Beebe:2005:DTM

[846] Nelson H. F. Beebe. The design of TEX and : A retrospective (abstract only). TUGboat, 26(2):169, ???? 2005. ISSN 0896-3207. URL http:/

/www.math.utah.edu/~beebe/talks/
2005/pt2005/pt2005.pdf;   http:/
/www.math.utah.edu/~beebe/talks/
2005/pt2005/pt2005.ps;   http://
www.math.utah.edu/~beebe/talks/
2005/pt2005/pt2005.ps.gz.  Pro-

ceedings of the TUG 2005 conference, Wuhan, China, August 23–25, 2005.

Robbins:2005:CSS

[847] Arnold Robbins and Nelson H. F. Beebe. Classic Shell Scripting. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2005. ISBN 0-596-00595-4. xxii + 534 pp. LCCN QA76.76.O63 R633

2005. URL http://www.oreilly.
com/catalog/shellsrptg/.  Also

available in Chinese [879], French [848], German [857], Japanese [859], and Polish [858] translations.

Robbins:2005:ISS

[848] Arnold Robbins and Nelson H. F. Beebe. Introduction aux scripts shell. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2005. ISBN 2-84177-375-2. xxii + 558 pp. LCCN

???? URL http://www.silicon.fr/

                                        getarticle.asp?id=14015. French

translation of [847] by Eric Jacoboni.

Beebe:2006:CBP

[849] Nelson H. F. Beebe. A complete bibliography of publications in Algorithmica. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, January 7, 2006. 216 pp. URL http:

                                        //www.math.utah.edu/pub/tex/bib/
                                        index-table-a.html#algorithmica.

Beebe:2006:BPAa

[850] Nelson H. F. Beebe. A bibliography of publications about the C# programming language. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, March 15, 2006. 19 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-c.html#csharp.

Beebe:2006:BPAb

[851] Nelson H. F. Beebe. A bibliography of publications about virtual machines. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, April 10, 2006. 55 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-v.html#virtual-machines.

Beebe:2006:BPL

[852] Nelson H. F. Beebe. A bibliography of papers in Lecture Notes in Computer Science (2005) (part 1 of 4). Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, April 3, 2006. 5 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-l.html#lncs2005a.

Beebe:2006:BPAc

[853] Nelson H. F. Beebe. A bibliography of publications in ACM SIGARCH Computer Architecture News. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, May 12, 2006. 303 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-s.html#sigarch.

Beebe:2006:BPR

[854] Nelson H. F. Beebe. A bibliography of publications of Ramon E. Moore. Report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, May 27, 2006.

14 pp. URL http://www.math.utah.
edu/pub/bibnet/authors/p.

Beebe:2006:BPAd

[855] Nelson H. F. Beebe. A bibliography of publications in ACM SIGOPS Operating Systems Review. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, August 26, 2006. 212 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-o.html#opersysrev.

Beebe:2006:WWC

[856] Nelson H. F. Beebe. WINE (Wine Is Not a (CPU) Emulator) FAQ. World-Wide Web frequently-asked question document., September 1,

2006. URL http://www.math.utah.

edu/faq/wine/. This report is updated frequently.

Robbins:2006:KSP

[857] Arnold Robbins and Nelson H. F. Beebe. Klassische Shell-Programmierung:

[automatisieren Sie Ihre Unix/Linux-Tasks]. (German) [Classic Shell Scripting: [automating your Unix/Linux jobs]]. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2006. ISBN 3-89721-441-5. xxiii + 572 pp. LCCN QA76.76.O63 R563 2005. EUR 44.00; EUR 45.65 (AT). URL http://www.

gbv.de/dms/hebis-darmstadt/toc/

17645067X.pdf. German translation of [847] by Kathrin Lichtenberg.

Robbins:2006:PSP

[858] Arnold Robbins and Nelson H. F. Beebe. Programowanie skryptów powłoki. Helion, Gliwice, Poland, 2006. ISBN 83-246-0131-7. 557 + 2 pp. LCCN ???? URL

http://www.empik.com/b/o/19/f1/
19f16b85e0d75ae1d3a1e7062569fbb0.
jpg;   http://www.empik.com/programowanie-
skryptow-powloki-ksiazka,360529,

p. Polish translation of [847] by Przemysław Szeremiota.

Robbins:2006:SSS

[859] Arnold Robbins and Nelson H. F. Beebe. Shōkai shieru sukuriputo. Orairī Japan, Tōkyō, Japan, 2006. ISBN 4-87311-267-2. 345 pp. LCCN ???? Japanese translation of [847] by Aoi Hyūga.

Beebe:2007:BPN

[860] Nelson H. F. Beebe. A bibliography of publications of Nicholas Constantine Metropolis. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, March 20, 2007. 34

pp. URL http://www.math.utah.
edu/pub/bibnet/authors/m/.

Beebe:2007:BPS

[861] Nelson H. F. Beebe. A bibliography of publications of Stanisław M. Ulam. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, March 31,

2007. 45 pp. URL http://www.math.
utah.edu/pub/bibnet/authors/u/.

Beebe:2007:BPF

[862] Nelson H. F. Beebe. A bibliography of publications of Frank Stenger. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, May 9, 2007.

27 pp. URL http://www.math.utah.
edu/pub/bibnet/authors/s/.

Ball:2007:EGR

[863] James S. Ball and Nelson H. F. Beebe. Efficient Gauss-related quadrature for two classes of logarithmic weight functions. ACM Transactions on Mathematical Software, 33(3):19:1–19:21, August 2007. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Beebe:2007:AQP

[864] Nelson H. F. Beebe and James S. Ball. Algorithm 867: QUADLOG—a package of routines for generating Gaussrelated quadrature for two classes of logarithmic weight functions. ACM Transactions on Mathematical Software, 33(3):20:1–20:30, August 2007. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Beebe:2007:CBP

[865] Nelson H. F. Beebe. A complete bibliography of the publications of John R. Rice. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, December 29, 2007. 118

                                        pp. URL http://www.math.utah.
                                        edu/pub/bibnet/authors/r/index.
                                        html#rice-john-r.

Beebe:2007:ETM

[866] Nelson H. F. Beebe. Extending TEX and with floating-point arithmetic. TUGboat, 28(3):319–328, 2007. ISSN 0896-3207.

Beebe:2007:NDF

[867] Nelson H. F. Beebe. New directions in floating-point arithmetic. In Simos and Maroulis [1788], pages 155– 158. ISBN 0-7354-0476-3 (set), 0-7354-0477-1 (vol. 1), 0-7354-0478-X (vol. 2). ISSN 0094-243X (print), 1551-7616 (electronic), 1935-0465. LCCN Q183.9 .I524 2007. URL http://www.

                                        springer.com/physics/atoms/book/
                                        978-0-7354-0478-6.

Beebe:2008:BI

[868] Nelson H. F. Beebe. A bibliography of ICCMSE 2007. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, February 21, 2008. 61 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-i.html#iccmse2007.

Beebe:2008:CBAa

[869] Nelson H. F. Beebe. A complete bibliography of ACM SIGMETRICS Performance Evaluation Review. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, June 27, 2008. 206 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-s.html#sigmetrics.

Beebe:2008:BCP

[870] Nelson H. F. Beebe. A bibliography of Combinatorics, Probability and Computing. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, July 7, 2008. 134 pp.

URL http://www.math.utah.edu/
pub/tex/bib/index-table-c.html#
combinprobabcomput.

Beebe:2008:CBPa

[871] Nelson H. F. Beebe. A complete bibliography of publications in the International Journal of Parallel, Emergent and Distributed Systems: IJPEDS. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, July 10, 2008. 20 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-p.html#intjparemerdistsys.

This report is updated frequently.

Beebe:2008:CBPb

[872] Nelson H. F. Beebe. A complete bibliography of publications in Parallel Algorithms and Applications. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, July 11, 2008. 61 pp.

URL http://www.math.utah.edu/
pub/tex/bib/index-table-p.html#
parallelalgorithmsappl. This re-

port is updated frequently.

Beebe:2008:CBPc

[873] Nelson H. F. Beebe. A complete bibliography of publications in Applied Mathematics and Computation: 2005–2009. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, July 12, 2008. 643 pp.

                                        URL http://www.math.utah.edu/
                                        pub/tex/bib/index-table-a.html#
                                        applmathcomput2005.

Beebe:2009:BMR

[874] Nelson H. F. Beebe. BibTEX meets relational databases: Dedicated to the memory of Edgar Frank “Ted” Codd (1923–2003) and James Nicholas “Jim” Gray (1944–2007). TUGboat, 30(2):252–271, 2009. ISSN 0896-3207.

                                        URL http://www.tug.org/TUGboat/
                                        tb30-2/tb95beebe.pdf. TUG 2009

Conference Proceedings volume.

Beebe:2009:NML

[875] Nelson H. F. Beebe. A new math library. Int. J. Quant. Chem., 109(13): 3008–3025, November 5, 2009. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Beebe:2008:CBAb

[876] Nelson H. F. Beebe. A complete bibliography of ACM Transactions on Accessible Computing. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, October 21, 2008. 6 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-t.html#taccess.

Beebe:2008:CBPd

[877] Nelson H. F. Beebe. A complete bibliography of publications in the ACM Journal of Experimental Algorithmics. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, October 06, 2008. 26 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-j.html#jea.

Beebe:2008:CBAc

[878] Nelson H. F. Beebe. A complete bibliography of ACM Transactions on Reconfigurable Technology and Systems. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, November 04, 2008. 5 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-t.html#trets.

Robbins:2008:SJB

[879] Arnold Robbins and Nelson H. F. Beebe. Shell Jiao Ben Xue Xi Zhi NaN = Shell Script Study Guide. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2008. ISBN 7-111-25504-6. vi + 494 pp. LCCN ???? 79.00 renminbi. Simplified Chinese translation of [847].

Beebe:2009:UBG

[880] Nelson H. F. Beebe. Using boxes and glue in TEX and LaTEX. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, March 28, 2009. 36 pp. URL

http://www.math.utah.edu/~beebe/
reports/2009/boxes.pdf.

Beebe:2009:CBAa

[881] Nelson H. F. Beebe. A complete bibliography of ACM Transactions on Autonomous and Adaptive Systems (TAAS). Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, April 24, 2009. 10 pp. URL http:

//www.math.utah.edu/pub/tex/bib/
index-table-t.html#taas.

Beebe:2009:CBAb

[882] Nelson H. F. Beebe. A complete bibliography of ACM Transactions on Computation Theory (ToCT). Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, April 24, 2009. 2 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-t.html#toct.

Beebe:2009:CBAc

[883] Nelson H. F. Beebe. A complete bibliography of ACM Transactions on Knowledge Discovery from Data (TKDD). Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, April 24, 2009. 9 pp. URL http:

//www.math.utah.edu/pub/tex/bib/
index-table-t.html#tkdd.

Beebe:2009:CBAd

[884] Nelson H. F. Beebe. A complete bibliography of ACM Transactions on the Web (TWEB). Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, April 24, 2009. 9 pp. URL http:

//www.math.utah.edu/pub/tex/bib/
index-table-t.html#tweb.

Beebe:2009:CBAe

[885] Nelson H. F. Beebe. A complete bibliography of the ACM Transactions on Computing Education (TOCE). Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, June 01, 2009. 2 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-t.html#toce.

Beebe:2009:BBP

[886] Nelson H. F. Beebe. A brief bibliography of publications about the Kepler Conjecture. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, August 14, 2009. 20 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-k.html#kepler.

Beebe:2009:CBJ

[887] Nelson H. F. Beebe. A complete bibliography of the Journal of Data and Information Quality (JDIQ). Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, September 18, 2009. 3 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-j.html#jdiq. This report is up-

dated frequently.

Beebe:2010:BPCa

[888] Nelson H. F. Beebe. A bibliography of publications in the Communications of the ACM : 2010–2019. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, February 2, 2010. 8 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-

                                        table-c.html#cacm2010. This report

is updated frequently.

Beebe:2010:BPCb

[889] Nelson H. F. Beebe. A bibliography of publications on cryptography: 2010–2019. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, February 22, 2010. 2 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-c.html#cryptography2010.

Beebe:2010:BPAa

[890] Nelson H. F. Beebe. A bibliography of publications in ACM SIGGRAPH Computer Graphics: 2010–2019. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, March 16, 2010. 2 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-s.html#siggraph2010.

Beebe:2010:BPAb

[891] Nelson H. F. Beebe. A bibliography of publications in ACM SIG-PLAN Notices, 2010–2019. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, March 16, 2010. 6 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-s.html#sigplan2010. This re-

port is updated frequently.

Beebe:2010:BPCc

[892] Nelson H. F. Beebe. A bibliography of publications in The Computer Journal: 2010–2019. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, April 28, 2010. 22 pp. URL http:

//www.math.utah.edu/pub/tex/bib/
index-table-c.html#compj2010.

Beebe:2010:BPAc

[893] Nelson H. F. Beebe. A bibliography of publications about the Google PageRank algorithm. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, May 6, 2010. 60 pp. URL http:

//www.math.utah.edu/pub/tex/bib/
index-table-p.html#pagerank.

Beebe:2010:BBA

[894] Nelson H. F. Beebe. A bibliography of books about the Internet and networking: 2010–2019. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, May 9, 2010. 2 pp. URL http:

//www.math.utah.edu/pub/tex/bib/
index-table-i.html#internet2010.

Beebe:2010:BPI

[895] Nelson H. F. Beebe. A bibliography of publications in IEEE Computer, 2010– 2019. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, May 12, 2010. 10 pp. URL http:

//www.math.utah.edu/pub/tex/bib/
index-table-c.html#computer2010.

Beebe:2010:CBPa

[896] Nelson H. F. Beebe. A complete bibliography of publications in Numerische Mathematik (2010–2019). Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, May 14, 2010. 10 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-

                                        table-n.html#nummath2010. This re-

port is updated frequently.

Beebe:2010:CBPb

[897] Nelson H. F. Beebe. A complete bibliography of publications in Mathematics of Computation, 2010–2019. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, May 14, 2010. 18 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-m.html#mathcomp2010.

Beebe:2010:CBPc

[898] Nelson H. F. Beebe. A complete bibliography of publications in SIAM Journal on Financial Mathematics. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, May 19, 2010. 3 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-s.html#siamjfinancialmath.

This report is updated frequently.

Beebe:2010:CBPd

[899] Nelson H. F. Beebe. A complete bibliography of publications in SIAM Journal on Imaging Sciences. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, May 19, 2010. 12 pp.

                                        URL http://www.math.utah.edu/
                                        pub/tex/bib/index-table-s.html#

siamjimagingsci. This report is updated frequently.

Beebe:2010:BPCd

[900] Nelson H. F. Beebe. A bibliography of publications on cryptography: 1990–1999. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, May 22, 2010. 293 pp.

URL http://www.math.utah.edu/
pub/tex/bib/index-table-c.html#
cryptography1990.

Beebe:2010:CBPe

[901] Nelson H. F. Beebe. A complete bibliography of publications in the journal Mathematics in Computer Science. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, July 9, 2010. 18 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-m.html#math-comput-sci.

Beebe:2010:BNA

[902] Nelson H. F. Beebe. A bibliography of numerical analysis publications, 2010–2019. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, August 23, 2010. 3 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-n.html#numana2010. This re-

port is updated frequently.

Beebe:2010:CBPf

[903] Nelson H. F. Beebe. A complete bibliography of publications in the ACM Symposia on Theory of Computing (STOC) for 2010–2019. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, September 1, 2010. 20 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-e.html#stoc2010. This report

is updated frequently.

Beebe:2010:CBPg

[904] Nelson H. F. Beebe. A complete bibliography of publications in Parallel and Distributed Computing Practices and Scalable Computing: Practice and Experience. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, September 2, 2010. 79 pp.

                                        URL http://www.math.utah.edu/
                                        pub/tex/bib/index-table-p.html#
                                        paralleldistcomppract. This report

is updated frequently.

Beebe:2010:CBPh

[905] Nelson H. F. Beebe. A complete bibliography of publications in Applied Mathematics and Computation: 2010– 2014. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, September 4, 2010. 102 pp.

                                        URL http://www.math.utah.edu/
                                        pub/tex/bib/index-table-a.html#
                                        applmathcomput2010.

Beebe:2010:CBPi

[906] Nelson H. F. Beebe. A complete bibliography of publications in Electronic Transactions on Numerical Analysis (ETNA). Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, September 6, 2010. 97 pp. URL http:

                                        //www.math.utah.edu/pub/tex/bib/
                                        index-table-e.html#etna. This re-

port is updated frequently.

Beebe:2010:CBPj

[907] Nelson H. F. Beebe. A complete bibliography of the Journal of Graphics Tools: JGT and Journal of Graphics, GPU, and Game Tools. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, September 7, 2010. 34 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-g.html#jgraphtools. This re-

port is updated frequently.

Beebe:2010:CBPk

[908] Nelson H. F. Beebe. A complete bibliography of publications in the Journal of Systems and Software. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, September 9, 2010. 405 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-j.html#jsystsoftw.

Beebe:2010:CBAa

[909] Nelson H. F. Beebe. A complete bibliography of Applied Statistics: 2000– 2009. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, September 11, 2010. 6 pp. URL http:

//www.math.utah.edu/pub/tex/bib/
index-table-n.html#as2010. This

report is updated frequently.

Beebe:2010:BPAd

[910] Nelson H. F. Beebe. A bibliography of publications about the Java Programming Language: 2010–2019. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, September 11, 2010. 7 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-j.html#java2010. This report

is updated frequently.

Beebe:2010:BPAe

[911] Nelson H. F. Beebe. A bibliography of publications about SGML, the Standard Generalized Markup Language: 2010–2019. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, September 12, 2010. 5 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-s.html#sgml2010. This report

is updated frequently.

Beebe:2010:CBPn

[912] Nelson H. F. Beebe. A complete bibliography of publications in the British Journal for the History of Science: 2010–2019. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, September 30, 2010. 32 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-b.html#bjhs2010. This report

is updated frequently.

Beebe:2010:CBPo

[913] Nelson H. F. Beebe. A complete bibliography of publications in Chymia: Annual Studies in the History of Chemistry: 1948–1967; Historical Studies in the Physical Sciences: 1969– 1971; Historical Studies in the Physical and Biological Sciences: 1986– 2007; and Historical Studies in the Natural Sciences: 2008–date. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, September 30, 2010. 101 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-h.html#hsns. This report is up-

dated frequently.

Beebe:2010:CBPm

[914] Nelson H. F. Beebe. A complete bibliography of publications in the British Journal for the History of Science: 2000–2009. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, September 30, 2010. 353 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-b.html#bjhs2000. This report

is updated frequently.

Beebe:2010:CBPq

[915] Nelson H. F. Beebe. A complete bibliography of publications in the British Journal for the History of Science: 1962–1999. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, September 30, 2010. 744 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-b.html#bjhs. At version 2.00

[27-Nov-2012], publications for 1990– 1999 were moved into a new bibliography file [1076] to reduce the entry count below the maximum of about 4000 entries permitted in traditional bibtex implementations. That split reduced the page count to 658. The title year range was changed to 1962–1989.

Beebe:2010:CBPr

[916] Nelson H. F. Beebe. A complete bibliography of publications in the British Journal for the Philosophy of Science: 1950–date. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, October 9, 2010. 515 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-

                                        table-b.html#bjps. This report is up-

dated frequently.

Beebe:2010:BMJ

[917] Nelson H. F. Beebe. A bibliography of The Mathematica Journal. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, November 6, 2010. 60 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-m.html#mathematicaj. This

report is updated frequently.

Beebe:2010:CBBa

[918] Nelson H. F. Beebe. A complete bibliography of The Bell System Technical Journal, 1920–1929. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, November 10, 2010. 26 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-b.html#bstj1920. This report

is updated frequently.

Beebe:2010:CBBc

[919] Nelson H. F. Beebe. A complete bibliography of The Bell System Technical Journal, 1940–1949. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, November 10, 2010. 35 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-b.html#bstj1940. This report

is updated frequently.

Beebe:2010:CBBb

[920] Nelson H. F. Beebe. A complete bibliography of The Bell System Technical Journal, 1930–1939. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, November 10, 2010. 56 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-b.html#bstj1930. This report

is updated frequently.

Beebe:2010:CBBd

[921] Nelson H. F. Beebe. A complete bibliography of The Bell System Technical Journal, 1950–1959. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, November 10, 2010. 83 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-b.html#bstj1950. This report

is updated frequently.

Beebe:2010:CBBg

[922] Nelson H. F. Beebe. A complete bibliography of The Bell System Technical Journal, 1980–1983, AT&T Bell Laboratories Technical Journal, 1984, and AT&T Technical Journal, 1985–1989. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, November 10, 2010. 128 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-b.html#bstj1980. This report

is updated frequently.

Beebe:2010:CBBe

[923] Nelson H. F. Beebe. A complete bibliography of The Bell System Technical Journal, 1960–1969. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, November 10, 2010. 151 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-

                                        table-b.html#bstj1960. This report

is updated frequently.

Beebe:2010:CBBf

[924] Nelson H. F. Beebe. A complete bibliography of The Bell System Technical Journal, 1970–1979. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, November 10, 2010. 190 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-b.html#bstj1970. This report

is updated frequently.

Beebe:2010:CBPs

[925] Nelson H. F. Beebe. A complete bibliography of publications of Claude Elwood Shannon. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, November 20, 2010. 26

                                        pp. URL http://www.math.utah.
                                        edu/pub/bibnet/authors/s/. This

report is updated frequently.

Beebe:2010:CBAb

[926] Nelson H. F. Beebe. A complete bibliography of ACM Transactions on Intelligent Systems and Technology (TIST). Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, November 23, 2010. 9 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-t.html#tist. This report is up-

dated frequently.

Beebe:2010:CBBi

[927] Nelson H. F. Beebe. A complete bibliography of Bell Labs Technical Journal, 2010–2019. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, November 26, 2010. 12 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-b.html#bstj2010. This report

is updated frequently.

Beebe:2010:CBAc

[928] Nelson H. F. Beebe. A complete bibliography of AT&T Technical Journal, 1990–1996, and Bell Labs Technical Journal, 1996–1999. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, November 26, 2010. 60 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-b.html#bstj1990. This report

is updated frequently.

Beebe:2010:CBBh

[929] Nelson H. F. Beebe. A complete bibliography of Bell Labs Technical Journal, 2000–2009. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, November 26, 2010. 83 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-b.html#bstj2000. This report

is updated frequently.

Beebe:2010:CBPt

[930] Nelson H. F. Beebe. A complete bibliography of publications in SIAM Journal on Mathematical Analysis for 2010– 2019. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, November 29, 2010. 28 pp.

URL http://www.math.utah.edu/
pub/tex/bib/index-table-s.html#

                                        siamjmathana2010. This report is up-

dated frequently.

Beebe:2010:BPS

[931] Nelson H. F. Beebe. A bibliography of publications in SIAM Journal on Mathematical Analysis for 2000– 2009. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, November 29, 2010. 138 pp.

                                        URL http://www.math.utah.edu/
                                        pub/tex/bib/index-table-s.html#
                                        siamjmathana2000. This report is up-

dated frequently.

Beebe:2011:CBJa

[932] Nelson H. F. Beebe. A complete bibliography of Journal on Computing and Cultural Heritage (JOCCH), 2010– 2019. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, January 15, 2011. 5 pp. URL http:

                                        //www.math.utah.edu/pub/tex/bib/
                                        index-table-j.html#jocch. This re-

port is updated frequently.

Beebe:2011:CBAa

[933] Nelson H. F. Beebe. A complete bibliography of Archive for History of Exact Sciences. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, February 5, 2011. 174 pp.

                                        URL http://www.math.utah.edu/
                                        pub/tex/bib/index-table-a.html#
                                        archhistexactsci. This report is up-

dated frequently.

Beebe:2011:BB

[934] Nelson H. F. Beebe. Benchmarking bibsql. Report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, March 17, 2011. 12 pp. URL http://www.

math.utah.edu/pub/bibsql/bibsql-
benchmark.pdf.

Beebe:2011:CBAb

[935] Nelson H. F. Beebe. A complete bibliography of ACM Transactions on Management Information Systems (TMIS). Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, March 28, 2011. 4 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-t.html#tmis. This report is up-

dated frequently.

Beebe:2011:BPTa

[936] Nelson H. F. Beebe. A bibliography of publications in Theoretical Computer Science: 2005–2009. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, March 29, 2011. 238 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-t.html#tcs2005. This report is

updated frequently.

Beebe:2011:BPTb

[937] Nelson H. F. Beebe. A bibliography of publications in Theoretical Computer Science: 2010–2014. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, March 29, 2011. 28 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-t.html#tcs2010. This report is

updated frequently.

Beebe:2011:CBPa

[938] Nelson H. F. Beebe. A complete bibliography of publications in Information Processing Letters: 2010–2019. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, March 31, 2011. 47 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-i.html#infoproc2010.

Beebe:2011:CBPb

[939] Nelson H. F. Beebe. A complete bibliography of publications in Linear Algebra and its Applications: 2010–2019. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, March 31, 2011. 92 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-l.html#linala2010. This re-

port is updated frequently.

Beebe:2011:CBPc

[940] Nelson H. F. Beebe. A complete bibliography of publications in Computer Networks (Amsterdam, Netherlands: 2010–2019). Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, April 2, 2011. 43 pp.

                                        URL http://www.math.utah.edu/
                                        pub/tex/bib/index-table-j.html#
                                        compnetamsterdam2010. This report

is updated frequently.

Beebe:2011:BPS

[941] Nelson H. F. Beebe. A bibliography of publications in SIAM Journal on Numerical Analysis: 2010–2019. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, April 6, 2011. 37 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-s.html#siamjnumeranal2010.

This report is updated frequently.

Beebe:2011:SBPa

[942] Nelson H. F. Beebe. A selected bibliography of publications by, and about, Richard Phillips Feynman. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, April 9, 2011.

46 pp. URL http://www.math.utah.
edu/pub/bibnet/authors/f/. This

report is updated frequently.

Beebe:2011:BPLc

[943] Nelson H. F. Beebe. A bibliography of papers in Lecture Notes in Computer Science (2008). Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, April 22, 2011. 3 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-l.html#lncs2008a. This re-

port is updated frequently.

Beebe:2011:BPLd

[944] Nelson H. F. Beebe. A bibliography of papers in Lecture Notes in Computer Science (2009). Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, April 22, 2011. 3 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-l.html#lncs2009a. This re-

port is updated frequently.

Beebe:2011:BPLe

[945] Nelson H. F. Beebe. A bibliography of papers in Lecture Notes in Computer Science (2010). Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, April 22, 2011. 3 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-l.html#lncs2010a. This re-

port is updated frequently.

Beebe:2011:BPLf

[946] Nelson H. F. Beebe. A bibliography of papers in Lecture Notes in Computer Science (2011). Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, April 22, 2011. 3 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-l.html#lncs2011a. This re-

port is updated frequently.

Beebe:2011:BPLb

[947] Nelson H. F. Beebe. A bibliography of papers in Lecture Notes in Computer Science (2007). Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, April 22, 2011. 5 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-l.html#lncs2007a. This re-

port is updated frequently.

Beebe:2011:BPLa

[948] Nelson H. F. Beebe. A bibliography of papers in Lecture Notes in Computer Science (2006). Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, April 22, 2011. 18 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-l.html#lncs2006a. This re-

port is updated frequently.

Beebe:2011:BPNa

[949] Nelson H. F. Beebe. A bibliography of publications on the numerical calculation of π. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, April 23, 2011. 18 pp. URL http:

//www.math.utah.edu/pub/tex/bib/
index-table-p.html#pi. This report

is updated frequently.

Beebe:2011:SBPb

[950] Nelson H. F. Beebe. A selected bibliography of publications by, and about, Niels Bohr. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, May 5, 2011. 104

pp. URL http://www.math.utah.
edu/pub/bibnet/authors/b/. This

report is updated frequently.

Beebe:2011:CBPd

[951] Nelson H. F. Beebe. A complete bibliography of publications in the Proceedings of the VLDB Endowment. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, May 13, 2011. 62 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-v.html#vldbe. This report is

updated frequently.

Beebe:2011:SBPc

[952] Nelson H. F. Beebe. A selected bibliography of publications by, and about, Paul Adrien Maurice Dirac. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, May 20, 2011.

125 pp. URL http://www.math.utah.
edu/pub/bibnet/authors/d/. This

report is updated frequently.

Beebe:2011:BPJa

[953] Nelson H. F. Beebe. A bibliography of publications of John W. Tukey. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, June 1, 2011.

67 pp. URL http://www.math.utah.
edu/pub/bibnet/authors/t/. This

report is updated frequently.

Beebe:2011:BPC

[954] Nelson H. F. Beebe. A bibliography of publications of Cornelius Lanczos. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, June 13, 2011. 131

pp. URL http://www.math.utah.
edu/pub/bibnet/authors/l/. This

report is updated frequently.

Beebe:2011:BITa

[955] Nelson H. F. Beebe. A bibliography of IEEE Transactions on Computers (1968–1969) and its predecessors: IEEE Transactions on Electronic Computers, IRE Transactions on Electronic Computers, and Transactions of the I.R.E. Professional Group on Electronic Computers. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, July 01, 2011. 321 pp.

URL http://www.math.utah.edu/

pub/tex/bib/index-table-i.html#

ieeetranscomput. This report is updated frequently.

Beebe:2011:CBIb

[956] Nelson H. F. Beebe. A complete bibliography of IEEE Transactions on Computers (2010–2019). Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, July 05, 2011. 43 pp.

URL http://www.math.utah.edu/
pub/tex/bib/index-table-i.html#
ieeetranscomput2010. This report is

updated frequently.

Beebe:2011:CBIa

[957] Nelson H. F. Beebe. A complete bibliography of IEEE Transactions on Computers (2000–2009). Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, July 05, 2011. 239 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-i.html#ieeetranscomput2000.

This report is updated frequently.

Beebe:2011:CBIc

[958] Nelson H. F. Beebe. A complete bibliography of IEEE Transactions on Computers (1990–1999). Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, July 07, 2011. 241 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-i.html#ieeetranscomput1990.

This report is updated frequently.

Beebe:2011:BITb

[959] Nelson H. F. Beebe. A bibliography of IEEE Transactions on Computers (1980–1989). Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, July 11, 2011. 272 pp.

URL http://www.math.utah.edu/
pub/tex/bib/index-table-i.html#
ieeetranscomput1980. This report is

updated frequently.

Beebe:2011:CBId

[960] Nelson H. F. Beebe. A complete bibliography of IEEE Transactions on Computers (1970–1979). Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, July 13, 2011. 340 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-i.html#ieeetranscomput1970.

This report is updated frequently.

Beebe:2011:CBPe

[961] Nelson H. F. Beebe. A complete bibliography of publications in the SIAM Journal on Applied Mathematics: 2010–2019. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, July 22, 2011. 23 pp.

URL http://www.math.utah.edu/
pub/tex/bib/index-table-s.html#
siamjapplmath2010. This report is up-

dated frequently.

Beebe:2011:CBB

[962] Nelson H. F. Beebe. A complete bibliography of the Bulletin in Applied Statistics (BIAS) and Journal of Applied Statistics. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, September 06, 2011. 268 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-j.html#japplstat. This re-

port is updated frequently.

Beebe:2011:CBPf

[963] Nelson H. F. Beebe. A complete bibliography of publications in Bulletin canadien de mathématiques = Canadian Mathematical Bulletin. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, September 09, 2011. 583 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-c.html#canmathbull. This re-

port is updated frequently.

Beebe:2011:CBPg

[964] Nelson H. F. Beebe. A complete bibliography of publications in Canadian Journal of Mathematics = Journal canadien de mathématiques for the decade 1940–1949. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, September 11, 2011. 6 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-c.html#canjmath1940. This

report is updated frequently.

Beebe:2011:CBPn

[965] Nelson H. F. Beebe. A complete bibliography of publications in Canadian Journal of Mathematics = Journal canadien de mathématiques for the decade 2010–2019. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, September 11, 2011. 17 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-

                                        table-c.html#canjmath2010. This

report is updated frequently.

Beebe:2011:CBPh

[966] Nelson H. F. Beebe. A complete bibliography of publications in Canadian Journal of Mathematics = Journal canadien de mathématiques for the decade 1950–1959. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, September 11, 2011. 68 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-c.html#canjmath1950. This

report is updated frequently.

Beebe:2011:CBPm

[967] Nelson H. F. Beebe. A complete bibliography of publications in Canadian Journal of Mathematics = Journal canadien de mathématiques for the decade 2000–2009. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, September 11, 2011. 70 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-c.html#canjmath2000. This

report is updated frequently.

Beebe:2011:CBPl

[968] Nelson H. F. Beebe. A complete bibliography of publications in Canadian Journal of Mathematics = Journal canadien de mathématiques for the decade 1990–1999. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, September 11, 2011. 87 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-

table-c.html#canjmath1990. This

report is updated frequently.

Beebe:2011:CBPk

[969] Nelson H. F. Beebe. A complete bibliography of publications in Canadian Journal of Mathematics = Journal canadien de mathématiques for the decade 1980–1989. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, September 11, 2011. 94 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-c.html#canjmath1980. This

report is updated frequently.

Beebe:2011:CBPi

[970] Nelson H. F. Beebe. A complete bibliography of publications in Canadian Journal of Mathematics = Journal canadien de mathématiques for the decade 1960–1969. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, September 11, 2011. 115 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-c.html#canjmath1960. This

report is updated frequently.

Beebe:2011:CBPj

[971] Nelson H. F. Beebe. A complete bibliography of publications in Canadian Journal of Mathematics = Journal canadien de mathématiques for the decade 1970–1979. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, September 11, 2011. 153 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-

                                        table-c.html#canjmath1970. This

report is updated frequently.

Beebe:2011:SBPd

[972] Nelson H. F. Beebe. A selected bibliography of publications by, and about, J. Robert Oppenheimer. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, September 14, 2011.

                                        26 pp. URL http://www.math.utah.
                                        edu/pub/bibnet/authors/s/. This

report is updated frequently.

Beebe:2011:SBPe

[973] Nelson H. F. Beebe. A selected bibliography of publications by, and about, Leo Szilard. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, September 15, 2011. 37

                                        pp. URL http://www.math.utah.
                                        edu/pub/bibnet/authors/s/. This

report is updated frequently.

Beebe:2011:CBPo

[974] Nelson H. F. Beebe. A complete bibliography of publications in Acta Numerica. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, September 24, 2011. 20 pp. URL http:

                                        //www.math.utah.edu/pub/tex/bib/
                                        index-table-a.html#actanumerica.

This report is updated frequently.

Beebe:2011:CBC

[975] Nelson H. F. Beebe. A complete bibliography of Combinatorics, Probability and Computing. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, September 29, 2011. 171 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-c.html#combinprobabcomput.

Beebe:2011:SBPf

[976] Nelson H. F. Beebe. A selected bibliography of publications by, and about, Edward Teller. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, October 9, 2011. 81

pp. URL http://www.math.utah.
edu/pub/bibnet/authors/t/. This

report is updated frequently.

Beebe:2011:CBPp

[977] Nelson H. F. Beebe. A complete bibliography of publications in The International Journal of Quantum Chemistry: 1960–1969. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, October 12, 2011. 94 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-i.html#ijqc1960.

Beebe:2011:CBPu

[978] Nelson H. F. Beebe. A complete bibliography of publications in The International Journal of Quantum Chemistry: 2010–2019. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, October 12, 2011. 177 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-i.html#ijqc2010.

Beebe:2011:CBPq

[979] Nelson H. F. Beebe. A complete bibliography of publications in The International Journal of Quantum Chemistry: 1970–1979. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, October 12, 2011. 392 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-i.html#ijqc1970.

Beebe:2011:CBPr

[980] Nelson H. F. Beebe. A complete bibliography of publications in The International Journal of Quantum Chemistry: 1980–1989. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, October 12, 2011. 599 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-i.html#ijqc1980.

Beebe:2011:CBPt

[981] Nelson H. F. Beebe. A complete bibliography of publications in The International Journal of Quantum Chemistry: 2000–2009. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, October 12, 2011. 768 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-i.html#ijqc2000.

Beebe:2011:CBPs

[982] Nelson H. F. Beebe. A complete bibliography of publications in The International Journal of Quantum Chemistry: 1990–1999. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, October 12, 2011. 777 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-i.html#ijqc1990.

Beebe:2011:CBPw

[983] Nelson H. F. Beebe. A complete bibliography of publications in the Journal of Mathematical Physics: 2000–2004. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, October 17, 2011. 307 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-j.html#jmathphys2000.

Beebe:2011:CBPv

[984] Nelson H. F. Beebe. A complete bibliography of publications in the Journal of Mathematical Physics: 1995–1999. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, October 17, 2011. 349 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-j.html#jmathphys1995.

Beebe:2011:CBPx

[985] Nelson H. F. Beebe. A complete bibliography of publications in the Journal of Mathematical Physics: 2005–2009. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, October 17, 2011. 386 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-j.html#jmathphys2005.

Beebe:2011:BPAa

[986] Nelson H. F. Beebe. A bibliography of publications by, and about, Max Born. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, October 19, 2011. 84

pp. URL http://www.math.utah.
edu/pub/bibnet/authors/b/.

Beebe:2011:CBPy

[987] Nelson H. F. Beebe. A complete bibliography of publications in The Fibonacci Quarterly. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, October 21, 2011. 550 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-f.html#fibquart.

Beebe:2011:CBPz

[988] Nelson H. F. Beebe. A complete bibliography of publications in the Journal of Mathematical Physics: 2010–2014. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, October 28, 2011. 184 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-j.html#jmathphys2010.

Beebe:2011:CBAc

[989] Nelson H. F. Beebe. A complete bibliography of ACM Transactions on Interactive Intelligent Systems (TIIS). Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, November 07, 2011. 3 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-t.html#tiis.

Beebe:2011:CBPaa

[990] Nelson H. F. Beebe. A complete bibliography of publications in the Studies in History and Philosophy of Modern Physics. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, November 07, 2011. 166 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-s.html#shps-b.

Beebe:2011:CBPac

[991] Nelson H. F. Beebe. A complete bibliography of publications in the Studies in History and Philosophy of Science Part C:. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, November 08, 2011. 135 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-s.html#shps-c.

Beebe:2011:CBPab

[992] Nelson H. F. Beebe. A complete bibliography of publications in the Studies in History and Philosophy of Science Part A. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, November 08, 2011. 301 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-s.html#shps-a.

Beebe:2011:CBJv

[993] Nelson H. F. Beebe. A complete bibliography of The Journal of the American Statistical Association: 2010–2019. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, November 10, 2011. 40 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-j.html#jamstatassoc2010.

Beebe:2011:CBJb

[994] Nelson H. F. Beebe. A complete bibliography of The Journal of the American Statistical Association: 2000–2009. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, November 10, 2011. 298 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-j.html#jamstatassoc2000.

Beebe:2011:BPAb

[995] Nelson H. F. Beebe. A bibliography of publications about Benford’s Law, Heaps’ Law, and Zipf’s Law. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, November 12, 2011. 62 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-b.html#benfords-law.

Beebe:2011:CBPad

[996] Nelson H. F. Beebe. A complete bibliography of publications in Linear Algebra and its Applications: 1980–1989. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, November 30, 2011. 385 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-l.html#linala1980. This re-

port is updated frequently.

Beebe:2011:SCF

[997] Nelson H. F. Beebe. Spell checking FAQ. World-Wide Web frequentlyasked question document., December 9,

                                        2011. URL http://www.math.utah.
                                        edu/faq/spell-check/. This report is

updated frequently.

Beebe:2011:NBP

[998] Nelson H. F. Beebe. Newcomb, Benford, Pareto, Heaps, and Zipf: Are arbitrary numbers random? Web site, December 10, 2011. URL http://www.

                                        math.utah.edu/~beebe/talks/2012/

benfords-law/. This document contains detailed slides for an undergraduate colloquium talk that has been given several times, with updates. The latest version of 27 January 2016 is 40 pages.

Beebe:2011:BS

[999] Nelson H. F. Beebe. A bibliography of Supercomputing ’2011. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, December 16, 2011. 18 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-s.html#supercomputing2011.

Beebe:2011:BPNb

[1000] Nelson H. F. Beebe. A bibliography of pseudorandom number generation, sampling, selection, distribution, and testing. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, December 20, 2011. 194 pp. URL http:

//www.math.utah.edu/pub/tex/bib/
index-table-p.html#prng.

Beebe:2011:BPJb

[1001] Nelson H. F. Beebe. A bibliography of publications in Journal of Computational Physics: 1960–1969. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, December 31, 2011. 3 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-j.html#jcomputphys1960.

Beebe:2011:BPJc

[1002] Nelson H. F. Beebe. A bibliography of publications in Journal of Computational Physics: 1970–1979. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, December 31, 2011. 7 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-j.html#jcomputphys1970.

Beebe:2011:BPJd

[1003] Nelson H. F. Beebe. A bibliography of publications in Journal of Computational Physics: 1980–1989. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, December 31, 2011. 12 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-j.html#jcomputphys1980.

Beebe:2011:BPJe

[1004] Nelson H. F. Beebe. A bibliography of publications in Journal of Computational Physics: 1990–1999. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, December 31, 2011. 14 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-j.html#jcomputphys1990.

Beebe:2011:BPJf

[1005] Nelson H. F. Beebe. A bibliography of publications in Journal of Computational Physics: 2000–2009. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, December 31, 2011. 34 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-j.html#jcomputphys2000.

Beebe:2011:CBPae

[1006] Nelson H. F. Beebe. A complete bibliography of publications in Journal of Computational Physics: 2010–2014.

Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, December 31, 2011. 256 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-j.html#jcomputphys2010.

Beebe:2012:CBPa

[1007] Nelson H. F. Beebe. A complete bibliography of publications in Journal of Computational Physics: 1960–1969. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, January 1, 2012. 34 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-j.html#jcomputphys1960.

Beebe:2012:CBPb

[1008] Nelson H. F. Beebe. A complete bibliography of publications in Journal of Computational Physics: 1970–1979. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, January 1, 2012. 197 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-j.html#jcomputphys1970.

Beebe:2012:CBPc

[1009] Nelson H. F. Beebe. A complete bibliography of publications in Journal of Computational Physics: 1980–1989. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, January 1, 2012. 314 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-j.html#jcomputphys1980.

Beebe:2012:CBPd

[1010] Nelson H. F. Beebe. A complete bibliography of publications in Journal of Computational Physics: 1990–1999. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, January 2, 2012. 497 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-j.html#jcomputphys1990.

Beebe:2012:CBPe

[1011] Nelson H. F. Beebe. A complete bibliography of publications in Journal of Computational Physics: 2000–2009. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, January 3, 2012. 796 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-j.html#jcomputphys2000.

Beebe:2012:CBJa

[1012] Nelson H. F. Beebe. A complete bibliography of The Journal of the American Statistical Association: 1920–1929. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, January 26, 2012. 155 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-j.html#jamstatassoc1920.

Beebe:2012:CBJc

[1013] Nelson H. F. Beebe. A complete bibliography of The Journal of the American Statistical Association: 1940–1949. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, January 26, 2012. 202 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-j.html#jamstatassoc1940.

Beebe:2012:CBJd

[1014] Nelson H. F. Beebe. A complete bibliography of The Journal of the American Statistical Association: 1950–1959. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, January 26, 2012. 236 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-j.html#jamstatassoc1950.

Beebe:2012:CBJb

[1015] Nelson H. F. Beebe. A complete bibliography of The Journal of the American Statistical Association: 1930–1939. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, January 26, 2012. 258 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-j.html#jamstatassoc1930.

Beebe:2012:CBJe

[1016] Nelson H. F. Beebe. A complete bibliography of The Journal of the American Statistical Association: 1960–1969. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, January 26, 2012. 350 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-j.html#jamstatassoc1960.

Beebe:2012:CBJf

[1017] Nelson H. F. Beebe. A complete bibliography of The Journal of the American Statistical Association: 1970–1979. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, January 26, 2012. 453 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-j.html#jamstatassoc1970.

Beebe:2012:CBJg

[1018] Nelson H. F. Beebe. A complete bibliography of The Journal of the American Statistical Association: 1980–1989. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, January 26, 2012. 482 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-j.html#jamstatassoc1980.

Beebe:2012:CBJh

[1019] Nelson H. F. Beebe. A complete bibliography of The Journal of the American Statistical Association: 1990–1999. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, January 26, 2012. 518 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-j.html#jamstatassoc1990.

Beebe:2012:CBAh

[1020] Nelson H. F. Beebe. A complete bibliography of The American Statistician: 2010–2019. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, January 28, 2012. 21 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-a.html#amstat2010.

Beebe:2012:CBAa

[1021] Nelson H. F. Beebe. A complete bibliography of The American Statistician: 1940–1949. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, January 28, 2012. 27 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-a.html#amstat1940.

Beebe:2012:CBAb

[1022] Nelson H. F. Beebe. A complete bibliography of The American Statistician: 1950–1959. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, January 28, 2012. 88 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-a.html#amstat1950.

Beebe:2012:CBAc

[1023] Nelson H. F. Beebe. A complete bibliography of The American Statistician: 1960–1969. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, January 28, 2012. 130 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-a.html#amstat1960.

Beebe:2012:CBAd

[1024] Nelson H. F. Beebe. A complete bibliography of The American Statistician: 1970–1979. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, January 28, 2012. 134 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-a.html#amstat1970.

Beebe:2012:CBAg

[1025] Nelson H. F. Beebe. A complete bibliography of The American Statistician: 2000–2009. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, January 28, 2012. 172 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-a.html#amstat2000.

Beebe:2012:CBAf

[1026] Nelson H. F. Beebe. A complete bibliography of The American Statistician: 1990–1999. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, January 28, 2012. 203 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-a.html#amstat1990.

Beebe:2012:CBAe

[1027] Nelson H. F. Beebe. A complete bibliography of The American Statistician: 1980–1989. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, January 28, 2012. 208 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-a.html#amstat1980.

Beebe:2012:CBPg

[1028] Nelson H. F. Beebe. A complete bibliography of publications in American Mathematical Monthly: 2010– 2019. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, January 30, 2012. 49 pp.

                                        URL http://www.math.utah.edu/
                                        pub/tex/bib/index-table-a.html#
                                        amermathmonthly2010.

Beebe:2012:CBPf

[1029] Nelson H. F. Beebe. A complete bibliography of publications in American Mathematical Monthly: 2000– 2009. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, January 30, 2012. 418 pp.

URL http://www.math.utah.edu/
pub/tex/bib/index-table-a.html#
amermathmonthly2000.

Beebe:2012:CBPh

[1030] Nelson H. F. Beebe. A complete bibliography of publications in Computer Physics Communications: 1960–1969. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, February 11, 2012. 6 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-a.html#compphyscomm1960.

Beebe:2012:CBPi

[1031] Nelson H. F. Beebe. A complete bibliography of publications in Computer Physics Communications: 1970–1979. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, February 13, 2012. 169 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-a.html#compphyscomm1970.

Beebe:2012:CBPm

[1032] Nelson H. F. Beebe. A complete bibliography of publications in Computer Physics Communications: 2010–2019. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, February 15, 2012. 183 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-a.html#compphyscomm2010.

Beebe:2012:CBPj

[1033] Nelson H. F. Beebe. A complete bibliography of publications in Computer Physics Communications: 1980–1989. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, February 15, 2012. 381 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-a.html#compphyscomm1980.

Beebe:2012:CBPk

[1034] Nelson H. F. Beebe. A complete bibliography of publications in Computer Physics Communications: 1990–1999. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, February 15, 2012. 510 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-a.html#compphyscomm1990.

Beebe:2012:CBPl

[1035] Nelson H. F. Beebe. A complete bibliography of publications in Computer Physics Communications: 2000–2009. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, February 15, 2012. 599 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-a.html#compphyscomm2000.

Beebe:2012:CBPap

[1036] Nelson H. F. Beebe. A bibliography of publications in Journal of Discrete Mathematical Sciences and Cryptography. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, February 16, 2012. 84 pp.

                                        URL http://www.math.utah.edu/
                                        pub/tex/bib/index-table-j.html#
                                        jdiscrmathscicrypto. This report is

updated frequently.

Beebe:2012:CBPn

[1037] Nelson H. F. Beebe. A complete bibliography of publications in Monte Carlo Methods and Applications. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, February 29, 2012. 96 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-m.html#mcma.

Beebe:2012:CBPo

[1038] Nelson H. F. Beebe. A complete bibliography of publications in Experimental Mathematics. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, March 5, 2012. 137 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-e.html#expmath.

Beebe:2012:BPA

[1039] Nelson H. F. Beebe. A bibliography of publications in ACM SIGACT News. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, March 24, 2012. 302 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-s.html#sigact.

Beebe:2012:CBPp

[1040] Nelson H. F. Beebe. A complete bibliography of publications in the IEEE symposia on Foundations of Computer Science (FOCS) for 2010–2019. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, April 12, 2012. 29 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-e.html#focs2010. This report

is updated frequently.

Beebe:2012:BPLa

[1041] Nelson H. F. Beebe. A bibliography of papers in Lecture Notes in Computer Science (2012). Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, May 5, 2012. 56 pp. URL http:

                                        //www.math.utah.edu/pub/tex/bib/
                                        index-table-l.html#lncs2012. This

report is updated frequently.

Beebe:2012:CBRa

[1042] Nelson H. F. Beebe. A complete bibliography of Reviews of Modern Physics, 1920–1929. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, May 23, 2012. 3 pp.

                                        URL http://www.math.utah.edu/
                                        pub/tex/bib/index-table-r.html#

revmodphys1920. This report is updated frequently.

Beebe:2012:CBRb

[1043] Nelson H. F. Beebe. A complete bibliography of Reviews of Modern Physics, 1930–1939. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, May 23, 2012. 29 pp.

                                        URL http://www.math.utah.edu/
                                        pub/tex/bib/index-table-r.html#

revmodphys1930. This report is updated frequently.

Beebe:2012:CBRj

[1044] Nelson H. F. Beebe. A complete bibliography of Reviews of Modern Physics, 2010–2019. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, May 23, 2012. 36 pp.

URL http://www.math.utah.edu/
pub/tex/bib/index-table-r.html#

revmodphys2010. This report is updated frequently.

Beebe:2012:CBRg

[1045] Nelson H. F. Beebe. A complete bibliography of Reviews of Modern Physics, 1980–1989. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, May 23, 2012. 54 pp.

URL http://www.math.utah.edu/
pub/tex/bib/index-table-r.html#

revmodphys1980. This report is updated frequently.

Beebe:2012:CBRf

[1046] Nelson H. F. Beebe. A complete bibliography of Reviews of Modern Physics, 1970–1979. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, May 23, 2012. 59 pp.

URL http://www.math.utah.edu/
pub/tex/bib/index-table-r.html#

revmodphys1970. This report is updated frequently.

Beebe:2012:CBRc

[1047] Nelson H. F. Beebe. A complete bibliography of Reviews of Modern Physics, 1940–1949. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, May 23, 2012. 61 pp.

URL http://www.math.utah.edu/
pub/tex/bib/index-table-r.html#

revmodphys1940. This report is updated frequently.

Beebe:2012:CBRh

[1048] Nelson H. F. Beebe. A complete bibliography of Reviews of Modern Physics, 1990–1999. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, May 23, 2012. 83 pp.

URL http://www.math.utah.edu/
pub/tex/bib/index-table-r.html#

revmodphys1990. This report is updated frequently.

Beebe:2012:CBRi

[1049] Nelson H. F. Beebe. A complete bibliography of Reviews of Modern Physics, 2000–2009. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, May 23, 2012. 85 pp.

URL http://www.math.utah.edu/
pub/tex/bib/index-table-r.html#

revmodphys2000. This report is updated frequently.

Beebe:2012:CBRd

[1050] Nelson H. F. Beebe. A complete bibliography of Reviews of Modern Physics, 1950–1959. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, May 23, 2012. 107 pp.

URL http://www.math.utah.edu/
pub/tex/bib/index-table-r.html#

revmodphys1950. This report is updated frequently.

Beebe:2012:CBRe

[1051] Nelson H. F. Beebe. A complete bibliography of Reviews of Modern Physics, 1960–1969. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, May 23, 2012. 171 pp.

URL http://www.math.utah.edu/
pub/tex/bib/index-table-r.html#

revmodphys1960. This report is updated frequently.

Beebe:2012:BVL

[1052] Nelson H. F. Beebe. A bibliography of volumes Lecture Notes in Computer Science (2011). Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, May 25, 2012. 23 pp. URL http:

//www.math.utah.edu/pub/tex/bib/
index-table-l.html#lncs2011. This

report is updated frequently.

Beebe:2012:SBPa

[1053] Nelson H. F. Beebe. A selected bibliography of publications by, and about, George Gamow. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, May 30, 2012. 75

pp. URL http://www.math.utah.
edu/pub/bibnet/authors/g/. This

report is updated frequently.

Beebe:2012:SBPb

[1054] Nelson H. F. Beebe. A selected bibliography of publications by, and about, Erwin Schrödinger. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, June 14, 2012.

126 pp. URL http://www.math.utah.
edu/pub/bibnet/authors/s/. This

report is updated frequently.

Beebe:2012:SBPc

[1055] Nelson H. F. Beebe. A selected bibliography of publications by, and about, Enrico Fermi. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, June 25, 2012. 136

                                        pp. URL http://www.math.utah.
                                        edu/pub/bibnet/authors/f/. This

report is updated frequently.

Beebe:2012:SBPd

[1056] Nelson H. F. Beebe. A selected bibliography of publications by, and about, John Clarke Slater. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, July 3, 2012. 55

                                        pp. URL http://www.math.utah.
                                        edu/pub/bibnet/authors/s/. This

report is updated frequently.

Beebe:2012:SBPe

[1057] Nelson H. F. Beebe. A selected bibliography of publications by, and about, Wolfgang Pauli. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, July 9, 2012. 91

                                        pp. URL http://www.math.utah.
                                        edu/pub/bibnet/authors/p/. This

report is updated frequently.

Beebe:2012:SBPf

[1058] Nelson H. F. Beebe. A selected bibliography of publications by, and about, Eugene Wigner. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, August 2, 2012. 216

                                        pp. URL http://www.math.utah.
                                        edu/pub/bibnet/authors/w/. This

report is updated frequently.

Beebe:2012:SBPg

[1059] Nelson H. F. Beebe. A selected bibliography of publications by, and about, Hans Bethe. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, August 6, 2012. 144

pp. URL http://www.math.utah.
edu/pub/bibnet/authors/b/. This

report is updated frequently.

Beebe:2012:SBPh

[1060] Nelson H. F. Beebe. A selected bibliography of publications by, and about, Werner Heisenberg. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, August 24, 2012.

217 pp. URL http://www.math.utah.
edu/pub/bibnet/authors/h/. This

report is updated frequently.

Beebe:2012:SBPi

[1061] Nelson H. F. Beebe. A selected bibliography of publications by, and about, Benoît Mandelbrot. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, September 11, 2012.

199 pp. URL http://www.math.utah.
edu/pub/bibnet/authors/m/. This

report is updated frequently.

Beebe:2012:CBS

[1062] Nelson H. F. Beebe. A complete bibliography of Scalable Computing: Practice and Experience. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, November 10, 2012. 63 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-s.html#scpe. This report is up-

dated frequently.

Beebe:2012:CBPq

[1063] Nelson H. F. Beebe. A complete bibliography of publications in Annals of Science. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, November 12, 2012. 313 pp.

                                        URL http://www.math.utah.edu/
                                        pub/tex/bib/index-table-a.html#

annscience. This report is updated frequently.

Beebe:2012:BS

[1064] Nelson H. F. Beebe. A bibliography of Supercomputing ’2012. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, November 15, 2012. 23 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-s.html#supercomputing2012.

Beebe:2012:BPSa

[1065] Nelson H. F. Beebe. A bibliography of publications in SIGCSE Bulletin: 1960–1969. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, November 19, 2012. 4 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-s.html#sigcse1960. This re-

port is updated frequently.

Beebe:2012:CBPv

[1066] Nelson H. F. Beebe. A complete bibliography of publications in SIGCSE Bulletin: 2010–2019. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, November 19, 2012. 8 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-s.html#sigcse2010. This re-

port is updated frequently.

Beebe:2012:CBPr

[1067] Nelson H. F. Beebe. A complete bibliography of publications in SIGCSE Bulletin: 1970–1979. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, November 19, 2012. 135 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-s.html#sigcse1970. This re-

port is updated frequently.

Beebe:2012:CBPs

[1068] Nelson H. F. Beebe. A complete bibliography of publications in SIGCSE Bulletin: 1980–1989. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, November 19, 2012. 175 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-s.html#sigcse1980. This re-

port is updated frequently.

Beebe:2012:CBPt

[1069] Nelson H. F. Beebe. A complete bibliography of publications in SIGCSE Bulletin: 1990–1999. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, November 19, 2012. 279 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-s.html#sigcse1990. This re-

port is updated frequently.

Beebe:2012:CBPu

[1070] Nelson H. F. Beebe. A complete bibliography of publications in SIGCSE Bulletin: 2000–2009. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, November 19, 2012. 593 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-s.html#sigcse2000. This re-

port is updated frequently.

Beebe:2012:CBPaf

[1071] Nelson H. F. Beebe. A complete bibliography of publications in Applied Mathematics and Computation: 1975– 1979. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, November 21, 2012. 19 pp.

URL http://www.math.utah.edu/
pub/tex/bib/index-table-a.html#
applmathcomput1975. This report is

updated frequently.

Beebe:2012:CBPag

[1072] Nelson H. F. Beebe. A complete bibliography of publications in Applied Mathematics and Computation: 1980– 1985. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, November 21, 2012. 36 pp.

URL http://www.math.utah.edu/
pub/tex/bib/index-table-a.html#
applmathcomput1980. This report is

updated frequently.

Beebe:2012:CBPah

[1073] Nelson H. F. Beebe. A complete bibliography of publications in Applied Mathematics and Computation: 1985– 1989. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, November 21, 2012. 55 pp.

URL http://www.math.utah.edu/
pub/tex/bib/index-table-a.html#
applmathcomput1985. This report is

updated frequently.

Beebe:2012:CBPaj

[1074] Nelson H. F. Beebe. A complete bibliography of publications in Applied Mathematics and Computation: 1995– 1999. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, November 21, 2012. 156 pp.

URL http://www.math.utah.edu/
pub/tex/bib/index-table-a.html#
applmathcomput1995. This report is

updated frequently.

Beebe:2012:CBPai

[1075] Nelson H. F. Beebe. A complete bibliography of publications in Applied Mathematics and Computation: 1990– 1995. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, November 22, 2012. 104 pp.

URL http://www.math.utah.edu/
pub/tex/bib/index-table-a.html#
applmathcomput1990. This report is

updated frequently.

Beebe:2012:CBPw

[1076] Nelson H. F. Beebe. A complete bibliography of publications in the British Journal for the History of Science: 1990–1999. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, November 27, 2012. 449 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-b.html#bjhs1990. This report

is updated frequently.

Beebe:2012:CBPx

[1077] Nelson H. F. Beebe. A complete bibliography of publications in Journal of Computational Chemistry: 1980–1989. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, November 30, 2012. 183 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-j.html#jcomputchem1980.

Beebe:2012:CBPaa

[1078] Nelson H. F. Beebe. A complete bibliography of publications in Journal of Computational Chemistry: 2010–2019. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, November 30, 2012. 219 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-j.html#jcomputchem2010.

Beebe:2012:CBPy

[1079] Nelson H. F. Beebe. A complete bibliography of publications in Journal of Computational Chemistry: 1990–1999. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, November 30, 2012. 322 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-j.html#jcomputchem1990.

Beebe:2012:CBPz

[1080] Nelson H. F. Beebe. A complete bibliography of publications in Journal of Computational Chemistry: 2000–2009. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, November 30, 2012. 456 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-j.html#jcomputchem2000.

Beebe:2012:CBPak

[1081] Nelson H. F. Beebe. A complete bibliography of publications in The Computer Journal: 1950–1959. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, December 06, 2012. 30 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-c.html#compj1950. This re-

port is updated frequently.

Beebe:2012:CBPao

[1082] Nelson H. F. Beebe. A complete bibliography of publications in The Computer Journal: 1990–1999. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, December 06, 2012. 290 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-c.html#compj1990. This re-

port is updated frequently.

Beebe:2012:CBPal

[1083] Nelson H. F. Beebe. A complete bibliography of publications in The Computer Journal: 1960–1969. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, December 06, 2012. 353 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-c.html#compj1960. This re-

port is updated frequently.

Beebe:2012:CBPan

[1084] Nelson H. F. Beebe. A complete bibliography of publications in The Computer Journal: 1980–1989. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, December 06, 2012. 359 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-c.html#compj1980. This re-

port is updated frequently.

Beebe:2012:BPU

[1085] Nelson H. F. Beebe. A bibliography of publications of the USENIX Association: 2010–2019. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, December 07, 2012. 40 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-u.html#usenix2010.

Beebe:2012:CBPam

[1086] Nelson H. F. Beebe. A complete bibliography of publications in The Computer Journal: 1970–1979. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, December 07, 2012. 419 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-c.html#compj1970. This re-

port is updated frequently.

Beebe:2012:CBPab

[1087] Nelson H. F. Beebe. A complete bibliography of publications in the Journal of Algorithms. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, December 11, 2012. 279 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-j.html#jalg. This report is up-

dated frequently.

Beebe:2012:CBPac

[1088] Nelson H. F. Beebe. A complete bibliography of publications in the Journal of Scientific Computing. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, December 22, 2012. 279 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-j.html#jscicomput. This re-

port is updated frequently.

Beebe:2012:BPSb

[1089] Nelson H. F. Beebe. A bibliography of publications in Science and Public Policy: 1970–1979. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, December 23, 2012. 62 pp.

URL http://www.math.utah.edu/
pub/tex/bib/index-table-s.html#
scipublicpol1970. This report is up-

dated frequently.

Beebe:2012:BPSc

[1090] Nelson H. F. Beebe. A bibliography of publications in Science and Public Policy: 1980–1989. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, December 23, 2012. 99 pp.

URL http://www.math.utah.edu/
pub/tex/bib/index-table-s.html#
scipublicpol1980. This report is up-

dated frequently.

Beebe:2012:BPSd

[1091] Nelson H. F. Beebe. A bibliography of publications in Science and Public Policy: 1990–1999. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, December 23, 2012. 68 pp.

URL http://www.math.utah.edu/
pub/tex/bib/index-table-s.html#

                                        scipublicpol1990. This report is up-

dated frequently.

Beebe:2012:BPSe

[1092] Nelson H. F. Beebe. A bibliography of publications in Science and Public Policy: 2000–2009. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, December 23, 2012. 109 pp.

                                        URL http://www.math.utah.edu/
                                        pub/tex/bib/index-table-s.html#
                                        scipublicpol2000. This report is up-

dated frequently.

Beebe:2012:BPSf

[1093] Nelson H. F. Beebe. A bibliography of publications in Science and Public Policy: 2010–2019. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, December 23, 2012. 43 pp.

                                        URL http://www.math.utah.edu/
                                        pub/tex/bib/index-table-s.html#
                                        scipublicpol2010. This report is up-

dated frequently.

Beebe:2012:BPLb

[1094] Nelson H. F. Beebe. A bibliography of papers in Lecture Notes in Computer Science (2012): Volumes 6121–7125. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, December 31, 2012. 360 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-l.html#lncs2012a. This re-

port is updated frequently.

Beebe:2012:BPLc

[1095] Nelson H. F. Beebe. A bibliography of papers in Lecture Notes in Computer Science (2012): Volumes 7125–7199.

Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, December 31, 2012. 267 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-l.html#lncs2012b. This re-

port is updated frequently.

Beebe:2012:BPLd

[1096] Nelson H. F. Beebe. A bibliography of papers in Lecture Notes in Computer Science (2012): Volumes 7200–7249. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, December 31, 2012. 181 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-l.html#lncs2012c. This re-

port is updated frequently.

Beebe:2012:BPLe

[1097] Nelson H. F. Beebe. A bibliography of papers in Lecture Notes in Computer Science (2012): Volumes 7250–7299. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, December 31, 2012. 193 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-l.html#lncs2012d. This re-

port is updated frequently.

Beebe:2012:BPLf

[1098] Nelson H. F. Beebe. A bibliography of papers in Lecture Notes in Computer Science (2012): Volumes 7300–7349. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, December 31, 2012. 192 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-l.html#lncs2012e. This re-

port is updated frequently.

Beebe:2012:BPLg

[1099] Nelson H. F. Beebe. A bibliography of papers in Lecture Notes in Computer Science (2012): Volumes 7350–7399. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, December 31, 2012. 200 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-l.html#lncs2012f. This re-

port is updated frequently.

Beebe:2012:BPLh

[1100] Nelson H. F. Beebe. A bibliography of papers in Lecture Notes in Computer Science (2012): Volumes 7400–7449. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, December 31, 2012. 186 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-l.html#lncs2012g. This re-

port is updated frequently.

Beebe:2012:BPLi

[1101] Nelson H. F. Beebe. A bibliography of papers in Lecture Notes in Computer Science (2012): Volumes 7450–7499. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, December 31, 2012. 189 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-l.html#lncs2012h. This re-

port is updated frequently.

Beebe:2012:BPLj

[1102] Nelson H. F. Beebe. A bibliography of papers in Lecture Notes in Computer Science (2012): Volumes 7500–7549. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, December 31, 2012. 149 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-l.html#lncs2012i. This re-

port is updated frequently.

Beebe:2012:BPLk

[1103] Nelson H. F. Beebe. A bibliography of papers in Lecture Notes in Computer Science (2012): Volumes 7550–7599. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, December 31, 2012. 126 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-l.html#lncs2012j. This re-

port is updated frequently.

Beebe:2012:BPLl

[1104] Nelson H. F. Beebe. A bibliography of papers in Lecture Notes in Computer Science (2012): Volumes 7600–7649. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, December 31, 2012. 139 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-l.html#lncs2012k. This re-

port is updated frequently.

Beebe:2012:BPLm

[1105] Nelson H. F. Beebe. A bibliography of papers in Lecture Notes in Computer Science (2012): Volumes 7650–7699. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, December 31, 2012. 102 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-l.html#lncs2012l. This re-

port is updated frequently.

Beebe:2013:BPI

[1106] Nelson H. F. Beebe. A bibliography of publications in Interval Computations and Reliable Computing. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, January 5, 2013. 191 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-r.html#reliablecomputing.

This report is updated frequently.

Beebe:2013:BBO

[1107] Nelson H. F. Beebe. A bibliography of books and other publications about the Ada Programming Language and its history. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, January 15, 2013. 275 pp. URL http:

                                        //www.math.utah.edu/pub/tex/bib/
                                        index-table-a.html#adabooks. This

report is updated frequently.

Beebe:2013:BBA

[1108] Nelson H. F. Beebe. A bibliography of the Bulletin of the Atomic Scientists of Chicago and the Bulletin of the Atomic Scientists. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, January 17, 2013. 416 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-b.html#bullatsci. This re-

port is updated frequently.

Beebe:2013:BDC

[1109] Nelson H. F. Beebe and David Salomon. A bibliography on data compression. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, February 06, 2013. 698 pp.

URL http://www.math.utah.edu/
pub/tex/bib/index-table-d.html#

datacompression. This report is updated frequently.

Beebe:2013:BIT

[1110] Nelson H. F. Beebe. BibTEX information and tutorial. Report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, February 16, 2013. 50

pp. URL http://www.math.utah.
edu/pub/bibnet/bibtex-info.html.

Beebe:2013:BPM

[1111] Nelson H. F. Beebe. A bibliography on pattern matching, regular expressions, and string matching. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, March 13, 2013. 292 pp.

URL http://www.math.utah.edu/
pub/tex/bib/index-table-s.html#

string-matching. This report is updated frequently.

Beebe:2012:BPR

[1112] Nelson H. F. Beebe. A bibliography of publications of Richard Eugene Crandall. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, March 20, 2012. 30

pp. URL http://www.math.utah.
edu/pub/bibnet/authors/c/. This

report is updated frequently.

Beebe:2013:CBPa

[1113] Nelson H. F. Beebe. A complete bibliography of publications in Scientific American: 1890–1899. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, May 24, 2013. 1806 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-s.html#sciam1890. This re-

port is updated frequently.

Beebe:2013:CBPb

[1114] Nelson H. F. Beebe. A complete bibliography of publications in Scientific American: 1900–1909. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, May 24, 2013. 2023 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-s.html#sciam1900. This re-

port is updated frequently.

Beebe:2013:CBPc

[1115] Nelson H. F. Beebe. A complete bibliography of publications in Scientific American: 1910–1919. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, May 24, 2013. 1235 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-s.html#sciam1910. This re-

port is updated frequently.

Beebe:2013:CBPd

[1116] Nelson H. F. Beebe. A complete bibliography of publications in Scientific American: 1920–1929. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, May 24, 2013. 828 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-

table-s.html#sciam1920. This re-

port is updated frequently.

Beebe:2013:CBPe

[1117] Nelson H. F. Beebe. A complete bibliography of publications in Scientific American: 1930–1939. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, May 24, 2013. 428 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-s.html#sciam1930. This re-

port is updated frequently.

Beebe:2013:CBPf

[1118] Nelson H. F. Beebe. A complete bibliography of publications in Scientific American: 1940–1949. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, May 24, 2013. 372 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-s.html#sciam1940. This re-

port is updated frequently.

Beebe:2013:CBPg

[1119] Nelson H. F. Beebe. A complete bibliography of publications in Scientific American: 1950–1959. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, May 24, 2013. 288 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-s.html#sciam1950. This re-

port is updated frequently.

Beebe:2013:CBPh

[1120] Nelson H. F. Beebe. A complete bibliography of publications in Scientific American: 1960–1969. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, May 24, 2013. 313 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-s.html#sciam1960. This re-

port is updated frequently.

Beebe:2013:CBPi

[1121] Nelson H. F. Beebe. A complete bibliography of publications in Scientific American: 1970–1979. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, May 24, 2013. 325 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-s.html#sciam1970. This re-

port is updated frequently.

Beebe:2013:CBPj

[1122] Nelson H. F. Beebe. A complete bibliography of publications in Scientific American: 1980–1989. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, May 24, 2013. 405 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-s.html#sciam1980. This re-

port is updated frequently.

Beebe:2013:CBPk

[1123] Nelson H. F. Beebe. A complete bibliography of publications in Scientific American: 1990–1999. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, May 24, 2013. 866 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-s.html#sciam1990. This re-

port is updated frequently.

Beebe:2013:CBPl

[1124] Nelson H. F. Beebe. A complete bibliography of publications in Scientific American: 2000–2009. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, May 24, 2013. 836 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-s.html#sciam2000. This re-

port is updated frequently.

Beebe:2013:CBPm

[1125] Nelson H. F. Beebe. A complete bibliography of publications in Scientific American: 2010–2019. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, May 24, 2013. 308 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-s.html#sciam2010. This re-

port is updated frequently.

Beebe:2013:BPSa

[1126] Nelson H. F. Beebe. A bibliography of publications in Scientific American: 1880–1889. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, May 27, 2013. 2074 pp.

URL http://www.math.utah.edu/
pub/tex/bib/index-table-s.html#

sciam1880. This report is updated frequently.

Beebe:2013:BPSb

[1127] Nelson H. F. Beebe. A bibliography of publications in Scientific American: 1870–1879. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, June 1, 2013. 2277 pp.

                                        URL http://www.math.utah.edu/
                                        pub/tex/bib/index-table-s.html#

sciam1870. This report is updated frequently.

Beebe:2013:PBP

[1128] Nelson H. F. Beebe. A preliminary bibliography of publications of Norman H. March. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, June 3, 2013. 287

                                        pp. URL http://www.math.utah.
                                        edu/pub/bibnet/authors/m/. This

report is updated frequently.

Beebe:2013:CBPn

[1129] Nelson H. F. Beebe. A complete bibliography of publications in Foundations of Physics. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, June 12, 2013. 623 pp.

                                        URL http://www.math.utah.edu/
                                        pub/tex/bib/index-table-f.html#

foundphys. This report is updated frequently.

Beebe:2013:CBPo

[1130] Nelson H. F. Beebe. A complete bibliography of publications in Nuncius. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, June 17, 2013. 609 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-n.html#nuncius. This report is

updated frequently.

Beebe:2013:CBPp

[1131] Nelson H. F. Beebe. A complete bibliography of publications in Historia Mathematica. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, June 25, 2013. 589 pp.

URL http://www.math.utah.edu/
pub/tex/bib/index-table-h.html#

histmath. This report is updated frequently.

Beebe:2013:CBPq

[1132] Nelson H. F. Beebe. A complete bibliography of publications in Physics in Perspective (PIP). Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, June 28, 2013. 189 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-p.html#physperspect. This

report is updated frequently.

Beebe:2013:SBPa

[1133] Nelson H. F. Beebe. A selected bibliography of publications by, and about, Ettore Majorana. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, July 9, 2013.

46 pp. URL http://www.math.utah.
edu/pub/bibnet/authors/m/. This

report is updated frequently.

Beebe:2013:CBPr

[1134] Nelson H. F. Beebe. A complete bibliography of publications in Perspectives on Science. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, July 17, 2013. 113 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-p.html#perspectsci. This re-

port is updated frequently.

Beebe:2013:CBPs

[1135] Nelson H. F. Beebe. A complete bibliography of publications in Science in Context. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, July 26, 2013. 122 pp.

                                        URL http://www.math.utah.edu/
                                        pub/tex/bib/index-table-s.html#

scicontext. This report is updated frequently.

Beebe:2013:BPC

[1136] Nelson H. F. Beebe. A bibliography of publications in Centaurus: an International Journal of the History of Science and its Cultural Aspects. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, July 29, 2013. 236 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-c.html#centaurus. This re-

port is updated frequently.

Beebe:2013:CBPt

[1137] Nelson H. F. Beebe. A complete bibliography of publications in Isis: 1910– 1919. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, August 2, 2013. 30 pp. URL http:

                                        //www.math.utah.edu/pub/tex/bib/
                                        index-table-i.html#isis1910. This

report is updated frequently.

Beebe:2013:CBPu

[1138] Nelson H. F. Beebe. A complete bibliography of publications in Isis: 1920– 1929. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, August 2, 2013. 141 pp. URL http:

//www.math.utah.edu/pub/tex/bib/
index-table-i.html#isis1920. This

report is updated frequently.

Beebe:2013:CBPv

[1139] Nelson H. F. Beebe. A complete bibliography of publications in Isis: 1930– 1939. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, August 2, 2013. 207 pp. URL http:

//www.math.utah.edu/pub/tex/bib/
index-table-i.html#isis1930. This

report is updated frequently.

Beebe:2013:CBPw

[1140] Nelson H. F. Beebe. A complete bibliography of publications in Isis: 1940– 1949. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, August 2, 2013. 157 pp. URL http:

//www.math.utah.edu/pub/tex/bib/
index-table-i.html#isis1940. This

report is updated frequently.

Beebe:2013:CBPx

[1141] Nelson H. F. Beebe. A complete bibliography of publications in Isis: 1950– 1959. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, August 2, 2013. 188 pp. URL http:

//www.math.utah.edu/pub/tex/bib/
index-table-i.html#isis1950. This

report is updated frequently.

Beebe:2013:CBPy

[1142] Nelson H. F. Beebe. A complete bibliography of publications in Isis: 1960– 1969. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, August 2, 2013. 231 pp. URL http:

                                        //www.math.utah.edu/pub/tex/bib/
                                        index-table-i.html#isis1960. This

report is updated frequently.

Beebe:2013:CBPz

[1143] Nelson H. F. Beebe. A complete bibliography of publications in Isis: 1970– 1979. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, August 2, 2013. 344 pp. URL http:

                                        //www.math.utah.edu/pub/tex/bib/
                                        index-table-i.html#isis1970. This

report is updated frequently.

Beebe:2013:CBPaa

[1144] Nelson H. F. Beebe. A complete bibliography of publications in Isis: 1980– 1989. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, August 2, 2013. 535 pp. URL http:

                                        //www.math.utah.edu/pub/tex/bib/
                                        index-table-i.html#isis1980. This

report is updated frequently.

Beebe:2013:CBPab

[1145] Nelson H. F. Beebe. A complete bibliography of publications in Isis: 1990– 1999. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, August 2, 2013. 671 pp. URL http:

                                        //www.math.utah.edu/pub/tex/bib/
                                        index-table-i.html#isis1990. This

report is updated frequently.

Beebe:2013:CBPac

[1146] Nelson H. F. Beebe. A complete bibliography of publications in Isis: 2000– 2009. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, August 2, 2013. 662 pp. URL http:

//www.math.utah.edu/pub/tex/bib/
index-table-i.html#isis2000. This

report is updated frequently.

Beebe:2013:CBPad

[1147] Nelson H. F. Beebe. A complete bibliography of publications in Isis: 2010– 2019. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, August 2, 2013. 177 pp. URL http:

//www.math.utah.edu/pub/tex/bib/
index-table-i.html#isis2010. This

report is updated frequently.

Beebe:2013:CBPae

[1148] Nelson H. F. Beebe. A complete bibliography of publications in Berichte zur Wissenschaftsgeschichte [Reports on the History of Science]: 2010– 2019. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, August 5, 2013. 399 pp.

URL http://www.math.utah.edu/
pub/tex/bib/index-table-b.html#
berwissenschaftgesch. This report

is updated frequently.

Beebe:2013:TWH

[1149] Nelson H. F. Beebe. Typesetting and writing hints for theses and dissertations. Web site, August 23, 2013.

URL http://www.math.utah.edu/
pub/uuthesis/writing-tips.html.

Despite the title focused at an original audience, this document is of general use for technical writer, with dozens of short sections about common problems and solutions. It is updated frequently, and at the last revision of 5 October 2019, is a document of 72 pages (the count depends on browser fonts and paper size).

Beebe:2013:SBPb

[1150] Nelson H. F. Beebe. A selected bibliography of publications by, and about, Freeman J. Dyson. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, August 28, 2013.

                                        183 pp. URL http://www.math.utah.
                                        edu/pub/bibnet/authors/d/. This

report is updated frequently.

Beebe:2013:CBC

[1151] Nelson H. F. Beebe. A complete bibliography of Communication Design Quarterly Review. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, September 5, 2013. 6 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-c.html#commundesignqreview.

This report is updated frequently.

Beebe:2013:CBPaf

[1152] Nelson H. F. Beebe. A complete bibliography of publications in Science of Computer Programming. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, December 3, 2013. 213 pp.

                                        URL http://www.math.utah.edu/
                                        pub/tex/bib/index-table-c.html#
                                        scicomputprogram. This report is up-

dated frequently.

Beebe:2013:CBPag

[1153] Nelson H. F. Beebe. A complete bibliography of publications in Computer Systems Science and Engineering and International Journal of Computer Systems Science and Engineering. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, December 3, 2013. 135 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-c.html#computsystscieng.

Beebe:2013:BPF

[1154] Nelson H. F. Beebe. A bibliography of publications of Frank E. Harris. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, December 5, 2013. 83

pp. URL http://www.math.utah.
edu/pub/bibnet/authors/h/. This

report is updated frequently.

Beebe:2013:CBPah

[1155] Nelson H. F. Beebe. A complete bibliography of publications in Computer Vision and Image Understanding: CVIU. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, December 10, 2013. 357 pp. URL http:

//www.math.utah.edu/pub/tex/bib/
index-table-c.html#cviu.

Beebe:2013:CBPai

[1156] Nelson H. F. Beebe. A complete bibliography of publications in Graphical Models. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, December 10, 2013. 68 pp. URL http:

                                        //www.math.utah.edu/pub/tex/bib/
                                        index-table-g.html#graph-models.

Beebe:2014:PCC

[1157] Nelson H. F. Beebe. Proposed chapter on computer arithmetic for the GNU C Programming book. Report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, February 14, 2014. 35 pp. Production of the book in which this chapter was to appear has been delayed, and is not yet in print. A revision appears in [1570].

Beebe:2014:CBPa

[1158] Nelson H. F. Beebe. A complete bibliography of Philosophy of Science: 1930–1939. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, March 1, 2014. 50 pp.

                                        URL http://www.math.utah.edu/
                                        pub/tex/bib/index-table-p.html#
                                        philossci1930.

Beebe:2014:CBPb

[1159] Nelson H. F. Beebe. A complete bibliography of Philosophy of Science: 1940–1949. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, March 1, 2014. 79 pp.

                                        URL http://www.math.utah.edu/
                                        pub/tex/bib/index-table-p.html#
                                        philossci1940.

Beebe:2014:CBPc

[1160] Nelson H. F. Beebe. A complete bibliography of Philosophy of Science: 1950–1959. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, March 1, 2014. 112 pp.

URL http://www.math.utah.edu/
pub/tex/bib/index-table-p.html#
philossci1950.

Beebe:2014:CBPd

[1161] Nelson H. F. Beebe. A complete bibliography of Philosophy of Science: 1960–1969. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, March 1, 2014. 100 pp.

URL http://www.math.utah.edu/
pub/tex/bib/index-table-p.html#
philossci1960.

Beebe:2014:CBPe

[1162] Nelson H. F. Beebe. A complete bibliography of Philosophy of Science: 1970–1979. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, March 1, 2014. 111 pp.

URL http://www.math.utah.edu/
pub/tex/bib/index-table-p.html#
philossci1970.

Beebe:2014:CBPf

[1163] Nelson H. F. Beebe. A complete bibliography of Philosophy of Science: 1980–1989. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, March 1, 2014. 113 pp.

URL http://www.math.utah.edu/
pub/tex/bib/index-table-p.html#
philossci1980.

Beebe:2014:CBPg

[1164] Nelson H. F. Beebe. A complete bibliography of Philosophy of Science: 1990–1999. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, March 1, 2014. 124 pp.

                                        URL http://www.math.utah.edu/
                                        pub/tex/bib/index-table-p.html#
                                        philossci1990.

Beebe:2014:CBPh

[1165] Nelson H. F. Beebe. A complete bibliography of Philosophy of Science: 2000–2009. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, March 1, 2014. 155 pp.

                                        URL http://www.math.utah.edu/
                                        pub/tex/bib/index-table-p.html#
                                        philossci2000.

Beebe:2014:CBPi

[1166] Nelson H. F. Beebe. A complete bibliography of Philosophy of Science: 2010–2019. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, March 1, 2014. 47 pp.

                                        URL http://www.math.utah.edu/
                                        pub/tex/bib/index-table-p.html#
                                        philossci2010.

Beebe:2014:CBAa

[1167] Nelson H. F. Beebe. A complete bibliography of ACM Transactions on Economics and Computation (TEAC). Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, March 14, 2014. 5 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-t.html#teac.

Beebe:2014:BPLa

[1168] Nelson H. F. Beebe. A bibliography of papers in Lecture Notes in Computer Science (2013): Volumes 7126–?? Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, March 21, 2014. 6 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-l.html#lncs2013a. This re-

port is updated frequently.

Beebe:2014:CBPj

[1169] Nelson H. F. Beebe. A complete bibliography of publications in SIAM Review. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, March 29, 2014. 769 pp.

URL http://www.math.utah.edu/
pub/tex/bib/index-table-s.html#

siamreview. This report is updated frequently.

Beebe:2014:CBPk

[1170] Nelson H. F. Beebe. A complete bibliography of publications in the Journal of the Society for Industrial and Applied Mathematics. Series A, On control and SIAM Journal on Control. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, March 30, 2014. 67 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-s.html#siamjcontrol.

Beebe:2014:CBPl

[1171] Nelson H. F. Beebe. A complete bibliography of publications in SIAM/ ASA Journal on Uncertainty Quantification. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, March 30, 2014. 5 pp.

URL http://www.math.utah.edu/
pub/tex/bib/index-table-s.html#
siamasajuncertainquantif.

Beebe:2014:CBPm

[1172] Nelson H. F. Beebe. A complete bibliography of publications in the SIAM Journal on Control and Optimization. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, March 30, 2014. 236 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-s.html#siamjcontroloptim.

Beebe:2014:CBPn

[1173] Nelson H. F. Beebe. A complete bibliography of publications in SIAM Journal on Matrix Analysis and Applications. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, March 30, 2014. 326 pp.

                                        URL http://www.math.utah.edu/
                                        pub/tex/bib/index-table-s.html#

siamjmatanaappl. This report is updated frequently.

Beebe:2014:CBPo

[1174] Nelson H. F. Beebe. A complete bibliography of publications in SIAM Journal on Mathematical Analysis for 2000–2009. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, March 31, 2014. 149 pp.

                                        URL http://www.math.utah.edu/
                                        pub/tex/bib/index-table-s.html#
                                        siamjmathana2000. This report is up-

dated frequently.

Beebe:2014:CBPp

[1175] Nelson H. F. Beebe. A complete bibliography of publications in SIAM Journal on Mathematical Analysis for 2010–2019. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, March 31, 2014. 109 pp.

URL http://www.math.utah.edu/
pub/tex/bib/index-table-s.html#
siamjmathana2010. This report is up-

dated frequently.

Beebe:2014:CBPq

[1176] Nelson H. F. Beebe. A complete bibliography of publications in SIAM Journal on Scientific Computing. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, March 31, 2014. 490 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-s.html#siamjscicomput. This

report is updated frequently.

Beebe:2014:CBPr

[1177] Nelson H. F. Beebe. A complete bibliography of publications in SIAM Journal on Scientific and Statistical Computing. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, March 31, 2014. 119 pp.

URL http://www.math.utah.edu/
pub/tex/bib/index-table-s.html#
siamjscistatcomp. This report is up-

dated frequently.

Beebe:2014:CBPs

[1178] Nelson H. F. Beebe. A complete bibliography of publications in Annals of Probability (1970–1979). Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, April 20, 2014. 137 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-

                                        table-a.html#annprobab1970. This

report is updated frequently.

Beebe:2014:CBPt

[1179] Nelson H. F. Beebe. A complete bibliography of publications in Annals of Probability (1980–1989). Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, April 20, 2014. 191 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-a.html#annprobab1980. This

report is updated frequently.

Beebe:2014:CBPu

[1180] Nelson H. F. Beebe. A complete bibliography of publications in Annals of Probability (1990–1999). Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, April 20, 2014. 176 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-a.html#annprobab1990. This

report is updated frequently.

Beebe:2014:CBPv

[1181] Nelson H. F. Beebe. A complete bibliography of publications in Annals of Probability (2000–2009). Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, April 20, 2014. 157 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-a.html#annprobab2000. This

report is updated frequently.

Beebe:2014:CBPw

[1182] Nelson H. F. Beebe. A complete bibliography of publications in Annals of Probability (2010–2019). Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, April 20, 2014. 78 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-a.html#annprobab2010. This

report is updated frequently.

Beebe:2014:CBPx

[1183] Nelson H. F. Beebe. A complete bibliography of publications in the Journal of Statistical Computation and Simulation. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, April 22, 2014. 632 pp.

URL http://www.math.utah.edu/
pub/tex/bib/index-table-j.html#
jstatcomputsimul. This report is up-

dated frequently.

Beebe:2014:CBPy

[1184] Nelson H. F. Beebe. A complete bibliography of publications in Probability Theory and Related Fields. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, April 28, 2014. 347 pp.

URL http://www.math.utah.edu/
pub/tex/bib/index-table-p.html#
probab-theory-relat-fields. This

report is updated frequently.

Beebe:2014:CBPz

[1185] Nelson H. F. Beebe. A complete bibliography of publications in Zeitschrift für Wahrscheinlichkeitstheorie und verwandte Gebiete [Journal for Probability Theory and Related Fields]. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, April 28, 2014. 447 pp.

                                        URL http://www.math.utah.edu/
                                        pub/tex/bib/index-table-p.html#
                                        z-wahrscheinlichkeits-theor-verw-

geb. This report is updated frequently.

Beebe:2014:CBPaa

[1186] Nelson H. F. Beebe. A complete bibliography of publications in Advances in Applied Probability. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, May 4, 2014. 578 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-a.html#advapplprob. This re-

port is updated frequently.

Beebe:2014:CBPab

[1187] Nelson H. F. Beebe. A complete bibliography of publications in the Journal of Applied Probability. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, May 4, 2014. 918 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-j.html#japplprobab. This re-

port is updated frequently.

Beebe:2014:CBLa

[1188] Nelson H. F. Beebe. A complete bibliography of Lecture Notes in Mathematics (1960–1969). Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, May 9, 2014. 253 pp. URL http:

                                        //www.math.utah.edu/pub/tex/bib/
                                        index-table-l.html#lnm1960. This

report is updated frequently.

Beebe:2014:CBLb

[1189] Nelson H. F. Beebe. A complete bibliography of Lecture Notes in Mathematics (1970–1974). Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, May 9, 2014. 730 pp. URL http:

//www.math.utah.edu/pub/tex/bib/
index-table-l.html#lnm1970. This

report is updated frequently.

Beebe:2014:CBLc

[1190] Nelson H. F. Beebe. A complete bibliography of Lecture Notes in Mathematics (1975–1979). Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, May 9, 2014. 854 pp. URL http:

//www.math.utah.edu/pub/tex/bib/
index-table-l.html#lnm1975. This

report is updated frequently.

Beebe:2014:CBLd

[1191] Nelson H. F. Beebe. A complete bibliography of Lecture Notes in Mathematics (1980–1984). Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, May 9, 2014. 926 pp. URL http:

//www.math.utah.edu/pub/tex/bib/
index-table-l.html#lnm1980. This

report is updated frequently.

Beebe:2014:CBLe

[1192] Nelson H. F. Beebe. A complete bibliography of Lecture Notes in Mathematics (1985–1989). Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, May 9, 2014. 778 pp. URL http:

//www.math.utah.edu/pub/tex/bib/
index-table-l.html#lnm1985. This

report is updated frequently.

Beebe:2014:CBLf

[1193] Nelson H. F. Beebe. A complete bibliography of Lecture Notes in Mathematics (1990–1999). Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, May 9, 2014. 650 pp. URL http:

                                        //www.math.utah.edu/pub/tex/bib/
                                        index-table-l.html#lnm1990. This

report is updated frequently.

Beebe:2014:CBLg

[1194] Nelson H. F. Beebe. A complete bibliography of Lecture Notes in Mathematics (2000–2009). Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, May 9, 2014. 618 pp. URL http:

                                        //www.math.utah.edu/pub/tex/bib/
                                        index-table-l.html#lnm2000. This

report is updated frequently.

Beebe:2014:CBLh

[1195] Nelson H. F. Beebe. A complete bibliography of Lecture Notes in Mathematics (2010–2019). Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, May 9, 2014. 304 pp. URL http:

                                        //www.math.utah.edu/pub/tex/bib/
                                        index-table-l.html#lnm2010. This

report is updated frequently.

Beebe:2014:BPG

[1196] Nelson H. F. Beebe. A bibliography of publications of Gilbert Wright ‘Pete’ Stewart III. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, May 30, 2014. 49

                                        pp. URL http://www.math.utah.
                                        edu/pub/bibnet/authors/s/. This

report is updated frequently.

Beebe:2014:CBPac

[1197] Nelson H. F. Beebe. A complete bibliography of publications in Statistical Science. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, May 30, 2014. 275 pp. URL http:

//www.math.utah.edu/pub/tex/bib/
index-table-s.html#statsci. This

report is updated frequently.

Beebe:2014:CBPad

[1198] Nelson H. F. Beebe. A complete bibliography of publications in Annals of Mathematical Statistics. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, May 31, 2014. 763 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-a.html#annmathstat. This re-

port is updated frequently.

Beebe:2014:CBPae

[1199] Nelson H. F. Beebe. A complete bibliography of publications in Statistics and Probability Letters: 1980– 1989. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, June 3, 2014. 118 pp.

URL http://www.math.utah.edu/
pub/tex/bib/index-table-s.html#
statproblett1980. This report is up-

dated frequently.

Beebe:2014:CBPaf

[1200] Nelson H. F. Beebe. A complete bibliography of publications in Statistics and Probability Letters: 1990– 1999. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, June 3, 2014. 420 pp.

URL http://www.math.utah.edu/
pub/tex/bib/index-table-s.html#
statproblett1990. This report is up-

dated frequently.

Beebe:2014:CBPag

[1201] Nelson H. F. Beebe. A complete bibliography of publications in Statistics and Probability Letters: 2000– 2009. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, June 3, 2014. 596 pp.

URL http://www.math.utah.edu/
pub/tex/bib/index-table-s.html#
statproblett2000. This report is up-

dated frequently.

Beebe:2014:CBPah

[1202] Nelson H. F. Beebe. A complete bibliography of publications in Statistics and Probability Letters: 2010– 2019. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, June 3, 2014. 315 pp.

URL http://www.math.utah.edu/
pub/tex/bib/index-table-s.html#
statproblett2010. This report is up-

dated frequently.

Beebe:2014:CBPai

[1203] Nelson H. F. Beebe. A complete bibliography of publications in Annals of Statistics (1970–1979). Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, June 5, 2014. 157 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-

table-a.html#annstat1970. This re-

port is updated frequently.

Beebe:2014:CBPaj

[1204] Nelson H. F. Beebe. A complete bibliography of publications in Annals of Statistics (1980–1989). Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, June 5, 2014. 235 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-a.html#annstat1980. This re-

port is updated frequently.

Beebe:2014:CBPak

[1205] Nelson H. F. Beebe. A complete bibliography of publications in Annals of Statistics (1990–1999). Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, June 5, 2014. 215 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-a.html#annstat1990. This re-

port is updated frequently.

Beebe:2014:CBPal

[1206] Nelson H. F. Beebe. A complete bibliography of publications in Annals of Statistics (2000–2009). Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, June 5, 2014. 176 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-a.html#annstat2000. This re-

port is updated frequently.

Beebe:2014:CBPam

[1207] Nelson H. F. Beebe. A complete bibliography of publications in Annals of Statistics (2010–2019). Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, June 5, 2014. 100 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-a.html#annstat2010. This re-

port is updated frequently.

Beebe:2014:CBPan

[1208] Nelson H. F. Beebe. A complete bibliography of publications in Annals of Applied Probability. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, June 6, 2014. 277 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-a.html#annapplprobab. This

report is updated frequently.

Beebe:2014:CBPao

[1209] Nelson H. F. Beebe. A complete bibliography of publications in Annals of Applied Statistics. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, June 6, 2014. 145 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-a.html#annapplstat. This re-

port is updated frequently.

Beebe:2014:CBPap

[1210] Nelson H. F. Beebe. A complete bibliography of publications in Annales de l’Institut Henri Poincaré (B) Probability and Statistics. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, June 9, 2014. 319 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-

table-a.html#annihpb. This report is

updated frequently.

Beebe:2014:CBPaq

[1211] Nelson H. F. Beebe. A complete bibliography of publications in Bayesian Analysis. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, June 17, 2014. 100 pp. URL http:

//www.math.utah.edu/pub/tex/bib/
index-table-b.html#bayesiananal.

This report is updated frequently.

Beebe:2014:CBPar

[1212] Nelson H. F. Beebe. A complete bibliography of publications in the International Statistical Review. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, June 19, 2014. 494 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-i.html#intstatrev. This re-

port is updated frequently.

Beebe:2014:CBPas

[1213] Nelson H. F. Beebe. A complete bibliography of publications in Biometrika for the decade 1900–1909. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, June 25, 2014. 45 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-b.html#biometrika1900. This

report is updated frequently.

Beebe:2014:CBPat

[1214] Nelson H. F. Beebe. A complete bibliography of publications in Biometrika for the decade 1910–1919. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, June 25, 2014. 35 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-b.html#biometrika1910. This

report is updated frequently.

Beebe:2014:CBPau

[1215] Nelson H. F. Beebe. A complete bibliography of publications in Biometrika for the decade 1920–1929. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, June 25, 2014. 42 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-b.html#biometrika1920. This

report is updated frequently.

Beebe:2014:CBPav

[1216] Nelson H. F. Beebe. A complete bibliography of publications in Biometrika for the decade 1930–1939. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, June 25, 2014. 52 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-b.html#biometrika1930. This

report is updated frequently.

Beebe:2014:CBPaw

[1217] Nelson H. F. Beebe. A complete bibliography of publications in Biometrika for the decade 1940–1949. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, June 25, 2014. 46 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-b.html#biometrika1940. This

report is updated frequently.

Beebe:2014:CBPax

[1218] Nelson H. F. Beebe. A complete bibliography of publications in Biometrika for the decade 1950–1959. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, June 25, 2014. 145 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-b.html#biometrika1950. This

report is updated frequently.

Beebe:2014:CBPay

[1219] Nelson H. F. Beebe. A complete bibliography of publications in Biometrika for the decade 1960–1969. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, June 25, 2014. 189 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-b.html#biometrika1960. This

report is updated frequently.

Beebe:2014:CBPaz

[1220] Nelson H. F. Beebe. A complete bibliography of publications in Biometrika for the decade 1970–1979. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, June 25, 2014. 201 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-b.html#biometrika1970. This

report is updated frequently.

Beebe:2014:CBPba

[1221] Nelson H. F. Beebe. A complete bibliography of publications in Biometrika for the decade 1980–1989. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, June 25, 2014. 200 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-b.html#biometrika1980. This

report is updated frequently.

Beebe:2014:CBPbb

[1222] Nelson H. F. Beebe. A complete bibliography of publications in Biometrika for the decade 1990–1999. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, June 25, 2014. 192 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-b.html#biometrika1990. This

report is updated frequently.

Beebe:2014:CBPbc

[1223] Nelson H. F. Beebe. A complete bibliography of publications in Biometrika for the decade 2000–2009. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, June 25, 2014. 174 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-b.html#biometrika2000. This

report is updated frequently.

Beebe:2014:CBPbd

[1224] Nelson H. F. Beebe. A complete bibliography of publications in Biometrika for the decade 2010–2019. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, June 25, 2014. 70 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-b.html#biometrika2010. This

report is updated frequently.

Beebe:2014:CBPbe

[1225] Nelson H. F. Beebe. A complete bibliography of publications in Technometrics for the decade 1950–1959. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, June 27, 2014. 8 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-t.html#technometrics1950.

This report is updated frequently.

Beebe:2014:CBPbf

[1226] Nelson H. F. Beebe. A complete bibliography of publications in Technometrics for the decade 1960–1969. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, June 27, 2014. 160 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-t.html#technometrics1960.

This report is updated frequently.

Beebe:2014:CBPbg

[1227] Nelson H. F. Beebe. A complete bibliography of publications in Technometrics for the decade 1970–1979. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, June 27, 2014. 221 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-t.html#technometrics1970.

This report is updated frequently.

Beebe:2014:CBPbh

[1228] Nelson H. F. Beebe. A complete bibliography of publications in Technometrics for the decade 1980–1989. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, June 27, 2014. 222 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-t.html#technometrics1980.

This report is updated frequently.

Beebe:2014:CBPbi

[1229] Nelson H. F. Beebe. A complete bibliography of publications in Technometrics for the decade 1990–1999. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, June 27, 2014. 315 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-t.html#technometrics1990.

This report is updated frequently.

Beebe:2014:CBPbj

[1230] Nelson H. F. Beebe. A complete bibliography of publications in Technometrics for the decade 2000–2009. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, June 27, 2014. 304 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-t.html#technometrics2000.

This report is updated frequently.

Beebe:2014:CBPbk

[1231] Nelson H. F. Beebe. A complete bibliography of publications in Technometrics for the decade 2010–2019. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, June 27, 2014. 69 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-t.html#technometrics2010.

This report is updated frequently.

Beebe:2014:BCL

[1232] Nelson H. F. Beebe. Building the Clang + LLVM compilers. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, July 21, 2014. 11 pp.

URL http://www.math.utah.edu/
pub/llvm/clang-notes.html. This

report is updated frequently.

Beebe:2014:BPA

[1233] Nelson H. F. Beebe. A bibliography of publications about the Magma computational algebra system. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, August 14, 2014. 21 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-m.html#magma. This report is

updated frequently.

Beebe:2014:CBPbl

[1234] Nelson H. F. Beebe. A complete bibliography of publications in the Journal of Cryptographic Engineering. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, August 20, 2014. 15 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-j.html#jcryptoeng. This re-

port is updated frequently.

Beebe:2014:BPLb

[1235] Nelson H. F. Beebe. A bibliography of papers in Lecture Notes in Computer Science (2014): Volumes 8349–?? Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, August 21, 2014. 6 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-l.html#lncs2014a. This re-

port is updated frequently.

Beebe:2014:CBJa

[1236] Nelson H. F. Beebe. A complete bibliography of the Journal of Theoretical Probability. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, August 28, 2014. 252 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-j.html#jtheorprobab. This

report is updated frequently.

Beebe:2014:CBJb

[1237] Nelson H. F. Beebe. A complete bibliography of the Journal of Statistical Physics: 1960–1969. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, August 28, 2014. 9 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-j.html#jstatphys1960. This

report is updated frequently.

Beebe:2014:CBJc

[1238] Nelson H. F. Beebe. A complete bibliography of the Journal of Statistical Physics: 1970–1979. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, August 28, 2014. 145 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-j.html#jstatphys1970. This

report is updated frequently.

Beebe:2014:CBJd

[1239] Nelson H. F. Beebe. A complete bibliography of the Journal of Statistical Physics: 1980–1989. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, August 28, 2014. 425 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-j.html#jstatphys1980. This

report is updated frequently.

Beebe:2014:CBJe

[1240] Nelson H. F. Beebe. A complete bibliography of the Journal of Statistical Physics: 1990–1999. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, August 28, 2014. 508 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-j.html#jstatphys1990. This

report is updated frequently.

Beebe:2014:CBJf

[1241] Nelson H. F. Beebe. A complete bibliography of the Journal of Statistical Physics: 2000–2009. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, August 28, 2014. 457 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-j.html#jstatphys2000. This

report is updated frequently.

Beebe:2014:CBJg

[1242] Nelson H. F. Beebe. A complete bibliography of the Journal of Statistical Physics: 2010–2019. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, August 28, 2014. 238 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-

                                        table-j.html#jstatphys2010. This

report is updated frequently.

Beebe:2014:CBC

[1243] Nelson H. F. Beebe. A complete bibliography of Computability. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, September 1, 2014. 6 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-c.html#computability. This

report is updated frequently.

Beebe:2014:CBEa

[1244] Nelson H. F. Beebe. A complete bibliography of Electronic Communications in Probability. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, September 3, 2014. 128 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-e.html#ecp. This report is up-

dated frequently.

Beebe:2014:CBEb

[1245] Nelson H. F. Beebe. A complete bibliography of Electronic Journal of Probability. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, September 3, 2014. 169 pp. URL http:

                                        //www.math.utah.edu/pub/tex/bib/
                                        index-table-e.html#ejp. This re-

port is updated frequently.

Beebe:2014:COC

[1246] Nelson H. F. Beebe. Comments on “Opportunities and Challenges in 21st Century Experimental Mathematical Computation: ICERM Workshop Report”. Report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, September 5, 2014. 10 pp. URL http:/

/www.math.utah.edu/~beebe/icerm/
ICERM-2014-NOTES.pdf;   http://
www.math.utah.edu/~beebe/icerm/
ICERM-2014.pdf.

Beebe:2014:MDN

[1247] Nelson H. F. Beebe. Mathematics Department network: A report prepared for the University of Utah ‘One Network’ Initiative. Report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, September 8, 2014. 15 pp. This report contains restricted financial and network data, and is not available outside the project for which it was written.

Beebe:2014:CBEc

[1248] Nelson H. F. Beebe. A complete bibliography of The Electronic Journal of Linear Algebra. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, September 3, 2014. 120 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-e.html#ela. This report is up-

dated frequently.

Beebe:2014:CBJh

[1249] Nelson H. F. Beebe. A complete bibliography of the Journal for the History of Astronomy. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, October 3, 2014. 373 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-j.html#jhistastron. This re-

port is updated frequently.

Beebe:2014:CBPbm

[1250] Nelson H. F. Beebe. A complete bibliography of publications in the journal family Statistics. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, October 10, 2014. 386 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-s.html#statistics. This re-

port is updated frequently.

Beebe:2014:CBAb

[1251] Nelson H. F. Beebe. A complete bibliography of ACM Transactions on Parallel Computing (TOPC). Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, October 17, 2014. 3 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-t.html#topc.

Horn:2014:BMA

[1252] Roger A. Horn and Nelson H. F. Beebe. A bibliography of Matrix Analysis, second edition. Report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, November 24, 2014. 43 pp.

                                        URL http://www.math.utah.edu/
                                        pub/bibnet/subjects/index.html#
                                        matrix-analysis-2ed.

Beebe:2014:CBPbn

[1253] Nelson H. F. Beebe. A complete bibliography of publications in Applied Mathematics and Computation: 2015–2019. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, January 7, 2015. 31 pp.

URL http://www.math.utah.edu/
pub/tex/bib/index-table-a.html#
applmathcomput2015. This report is

updated frequently.

Beebe:2015:BPT

[1254] Nelson H. F. Beebe. A bibliography of publications in Theoretical Computer Science: 2015–2019. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, January 7, 2015. 21 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-t.html#tcs2015. This report is

updated frequently.

Beebe:2015:CBPa

[1255] Nelson H. F. Beebe. A complete bibliography of publications in Linear Algebra and its Applications: 2015–2019. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, January 7, 2015. 31 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-l.html#linala2015. This re-

port is updated frequently.

Beebe:2015:CBJa

[1256] Nelson H. F. Beebe. A complete bibliography of Journal of the Royal Statistical Society, Series D (The Statistician): 1950–1959. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, January 23, 2015. 32 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-j.html#jrss-d-1950.

Beebe:2015:CBJb

[1257] Nelson H. F. Beebe. A complete bibliography of Journal of the Royal Statistical Society, Series D (The Statistician): 1960–1969. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, January 23, 2015. 75 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-j.html#jrss-d-1960.

Beebe:2015:CBJc

[1258] Nelson H. F. Beebe. A complete bibliography of Journal of the Royal Statistical Society, Series D (The Statistician): 1970–1979. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, January 23, 2015. 73 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-j.html#jrss-d-1970.

Beebe:2015:CBJd

[1259] Nelson H. F. Beebe. A complete bibliography of Journal of the Royal Statistical Society, Series D (The Statistician): 1980–1989. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, January 23, 2015. 113 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-j.html#jrss-d-1980.

Beebe:2015:CBJe

[1260] Nelson H. F. Beebe. A complete bibliography of Journal of the Royal Statistical Society, Series D (The Statistician): 1990–1999. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, January 23, 2015. 170 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-j.html#jrss-d-1990.

Beebe:2015:CBJf

[1261] Nelson H. F. Beebe. A complete bibliography of Journal of the Royal Statistical Society, Series D (The Statistician): 2000–2009. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, January 23, 2015. 78 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-j.html#jrss-d-2000.

Beebe:2015:CBJg

[1262] Nelson H. F. Beebe. A complete bibliography of the Journal of the Royal Statistical Society, Series B family. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, January 24, 2015. 423 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-j.html#jrss-b.

Beebe:2015:CBJh

[1263] Nelson H. F. Beebe. A complete bibliography of the Journal of the Royal Statistical Society, Series A family: 1830–1839. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, January 24, 2015. 20 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-j.html#jrss-a-1830.

Beebe:2015:CBJi

[1264] Nelson H. F. Beebe. A complete bibliography of the Journal of the Royal Statistical Society, Series A family: 1840–1849. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, January 24, 2015. 41 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-j.html#jrss-a-1840.

Beebe:2015:CBJj

[1265] Nelson H. F. Beebe. A complete bibliography of the Journal of the Royal Statistical Society, Series A family: 1850–1859. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, January 24, 2015. 36 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-j.html#jrss-a-1850.

Beebe:2015:CBJk

[1266] Nelson H. F. Beebe. A complete bibliography of the Journal of the Royal Statistical Society, Series A family: 1860–1869. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, January 24, 2015. 40 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-j.html#jrss-a-1860.

Beebe:2015:CBJl

[1267] Nelson H. F. Beebe. A complete bibliography of the Journal of the Royal Statistical Society, Series A family: 1870–1879. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, January 24, 2015. 46 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-j.html#jrss-a-1870.

Beebe:2015:CBJm

[1268] Nelson H. F. Beebe. A complete bibliography of the Journal of the Royal Statistical Society, Series A family: 1880–1889. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, January 24, 2015. 48 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-j.html#jrss-a-1880.

Beebe:2015:CBJn

[1269] Nelson H. F. Beebe. A complete bibliography of the Journal of the Royal Statistical Society, Series A family: 1890–1899. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, January 24, 2015. 47 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-j.html#jrss-a-1890.

Beebe:2015:CBJo

[1270] Nelson H. F. Beebe. A complete bibliography of the Journal of the Royal Statistical Society, Series A family: 1900–1909. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, January 24, 2015. 78 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-j.html#jrss-a-1900.

Beebe:2015:CBJp

[1271] Nelson H. F. Beebe. A complete bibliography of the Journal of the Royal Statistical Society, Series A family: 1910–1919. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, January 24, 2015. 126 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-j.html#jrss-a-1910.

Beebe:2015:CBJq

[1272] Nelson H. F. Beebe. A complete bibliography of the Journal of the Royal Statistical Society, Series A family: 1920–1929. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, January 24, 2015. 95 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-j.html#jrss-a-1920.

Beebe:2015:CBJr

[1273] Nelson H. F. Beebe. A complete bibliography of the Journal of the Royal Statistical Society, Series A family: 1930–1939. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, January 24, 2015. 116 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-j.html#jrss-a-1930.

Beebe:2015:CBJs

[1274] Nelson H. F. Beebe. A complete bibliography of the Journal of the Royal Statistical Society, Series A family: 1940–1949. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, January 24, 2015. 72 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-j.html#jrss-a-1940.

Beebe:2015:CBJt

[1275] Nelson H. F. Beebe. A complete bibliography of the Journal of the Royal Statistical Society, Series A family: 1950–1959. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, January 24, 2015. 120 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-j.html#jrss-a-1950.

Beebe:2015:CBJu

[1276] Nelson H. F. Beebe. A complete bibliography of the Journal of the Royal Statistical Society, Series A family: 1960–1969. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, January 24, 2015. 160 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-j.html#jrss-a-1960.

Beebe:2015:CBJv

[1277] Nelson H. F. Beebe. A complete bibliography of the Journal of the Royal Statistical Society, Series A family: 1970–1979. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, January 24, 2015. 173 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-j.html#jrss-a-1970.

Beebe:2015:CBJw

[1278] Nelson H. F. Beebe. A complete bibliography of the Journal of the Royal Statistical Society, Series A family: 1980–1989. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, January 24, 2015. 129 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-j.html#jrss-a-1980.

Beebe:2015:CBJx

[1279] Nelson H. F. Beebe. A complete bibliography of the Journal of the Royal Statistical Society, Series A family: 1990–1999. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, January 24, 2015. 108 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-j.html#jrss-a-1990.

Beebe:2015:CBJy

[1280] Nelson H. F. Beebe. A complete bibliography of the Journal of the Royal Statistical Society, Series A family: 2000–2009. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, January 24, 2015. 126 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-j.html#jrss-a-2000.

Beebe:2015:CBJz

[1281] Nelson H. F. Beebe. A complete bibliography of the Journal of the Royal Statistical Society, Series A family: 2010–2019. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, January 24, 2015. 70 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-j.html#jrss-a-2010.

Beebe:2015:CBPb

[1282] Nelson H. F. Beebe. A complete bibliography of publications in the Journal of Network and Computer Applications. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, January 28, 2015. 279 pp.

URL http://www.math.utah.edu/
pub/tex/bib/index-table-j.html#
jnetwcomputappl.

Beebe:2015:CBPc

[1283] Nelson H. F. Beebe. A complete bibliography of publications in Annals of the Institute of Statistical Mathematics. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, January 31, 2015. 422 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-a.html#anninststatmath.

Beebe:2015:CBPd

[1284] Nelson H. F. Beebe. A complete bibliography of publications in Statistics and Computing. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, January 31, 2015. 150 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-s.html#statcomput.

Beebe:2015:CBPe

[1285] Nelson H. F. Beebe. A complete bibliography of publications in Statistical Papers. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, February 1, 2015. 270 pp. URL http:

//www.math.utah.edu/pub/tex/bib/
index-table-s.html#statpapers.

Beebe:2015:CBPf

[1286] Nelson H. F. Beebe. A complete bibliography of publications in Communications in Mathematics and Statistics. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, February 2, 2015. 7 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-c.html#communmathstat. This

report is updated frequently.

Beebe:2015:CBPg

[1287] Nelson H. F. Beebe. A complete bibliography of publications in Metrika. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, February 2, 2015. 246 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-m.html#metrika. This report is

updated frequently.

Beebe:2015:CBPh

[1288] Nelson H. F. Beebe. A complete bibliography of publications in Mathematical Methods of Statistics. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, February 2, 2015. 65 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-m.html#mathmethodsstat.

Beebe:2015:CBPi

[1289] Nelson H. F. Beebe. A complete bibliography of publications in AStA. Advances in Statistical Analysis. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, February 5, 2015. 55 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-a.html#asta.

Beebe:2015:CBPj

[1290] Nelson H. F. Beebe. A complete bibliography of publications in Computational Statistics & Data Analysis (1980–1989). Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, February 6, 2015. 38 pp.

URL http://www.math.utah.edu/
pub/tex/bib/index-table-c.html#
computstatdataanal1980.

Beebe:2015:CBPk

[1291] Nelson H. F. Beebe. A complete bibliography of publications in Computational Statistics & Data Analysis (1990–1999). Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, February 6, 2015. 211 pp.

URL http://www.math.utah.edu/
pub/tex/bib/index-table-c.html#
computstatdataanal1990.

Beebe:2015:CBPl

[1292] Nelson H. F. Beebe. A complete bibliography of publications in Computational Statistics & Data Analysis (2000–2009). Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, February 6, 2015. 364 pp.

URL http://www.math.utah.edu/
pub/tex/bib/index-table-c.html#
computstatdataanal2000.

Beebe:2015:CBPm

[1293] Nelson H. F. Beebe. A complete bibliography of publications in Computational Statistics & Data Analysis (2010–2019). Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, February 6, 2015. 266 pp.

URL http://www.math.utah.edu/
pub/tex/bib/index-table-c.html#
computstatdataanal2010.

Beebe:2014:CBPbo

[1294] Nelson H. F. Beebe. A complete bibliography of publications in Statistical Methodology. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, February 6, 2015. 79 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-s.html#statmethodol. This

report is updated frequently.

Beebe:2015:CBPn

[1295] Nelson H. F. Beebe. A complete bibliography of publications in the Journal of Mathematical Physics: 2015–2019. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, February 7, 2015. 9 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-j.html#jmathphys2015.

Beebe:2015:BPCa

[1296] Nelson H. F. Beebe. A bibliography of publications in Chance. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, February 8, 2015. 150 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-c.html#chance.

Beebe:2015:CBPo

[1297] Nelson H. F. Beebe. A complete bibliography of publications in ESAIM:

Probability and Statistics. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, February 9, 2015. 45 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-e.html#esaim-probab-stat.

Beebe:2015:CBPp

[1298] Nelson H. F. Beebe. A complete bibliography of publications in Journal of Computational Physics: 2015–2019. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, February 23, 2015. 54 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-j.html#jcomputphys2015.

Beebe:2015:SBPa

[1299] Nelson H. F. Beebe. A selected bibliography of publications by, and about, Edsger W. Dijkstra. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, March 17, 2015.

281 pp. URL http://www.math.utah.
edu/pub/bibnet/authors/d/.

Beebe:2015:CBPq

[1300] Nelson H. F. Beebe. A complete bibliography of publications in Dædalus: proceedings of the American Academy of Arts and Sciences. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, March 20, 2015. 379 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-d.html#daedalus.

Beebe:2015:CBPr

[1301] Nelson H. F. Beebe. A complete bibliography of publications in Osiris (1936–1968) and Osiris (Series 2) (1985–date). Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, March 30, 2015. 93 pp.

                                        URL http://www.math.utah.edu/
                                        pub/tex/bib/index-table-o.html#
                                        osiris.

Beebe:2015:CBM

[1302] Nelson H. F. Beebe. A complete bibliography of The Mathematical Gazette (2010–2019). Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, May 12, 2015. 174 pp.

                                        URL http://www.math.utah.edu/
                                        pub/tex/bib/index-table-m.html#
                                        mathgaz2010.

Beebe:2015:CBHa

[1303] Nelson H. F. Beebe. A complete bibliography of History of Science. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, May 19, 2015. 231 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-h.html#histsciuk.

Beebe:2015:CBHb

[1304] Nelson H. F. Beebe. A complete bibliography of HOPOS: Journal of the International Society for the History of Philosophy of Science. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, May 25, 2015. 69 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-h.html#hopos.

Beebe:2015:CBPt

[1305] Nelson H. F. Beebe. A complete bibliography of publications in Canadian Journal of Statistics = Revue canadienne de statistique. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, June 9, 2015. 298 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-c.html#canjstat.

Beebe:2015:CBPu

[1306] Nelson H. F. Beebe. A complete bibliography of publications in International Journal of Statistics and Probability. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, June 9, 2015. 33 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-i.html#intjstatprobab.

Beebe:2015:CBPv

[1307] Nelson H. F. Beebe. A complete bibliography of publications in Probability Surveys. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, June 10, 2015. 15 pp. URL http:

//www.math.utah.edu/pub/tex/bib/
index-table-p.html#probabsurv.

Beebe:2015:BPP

[1308] Nelson H. F. Beebe. A bibliography of publications in Physis: Rivista Internazionale di Storia della Scienza. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, June 15, 2015. 137 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-p.html#physis.

Beebe:2015:BPR

[1309] Nelson H. F. Beebe. A bibliography of publications in Revue d’Histoire des Sciences et de Leurs Applications and Revue d’Histoire des Sciences. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, June 16, 2015. 857 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-r.html#revhistsci.

Beebe:2015:CBPw

[1310] Nelson H. F. Beebe. A complete bibliography of publications in R News and the R Journal. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, August 13, 2015. 61 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-r.html#rjournal.

Beebe:2015:SBPb

[1311] Nelson H. F. Beebe. A selected bibliography of publications by, and about, Max Planck. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, August 27, 2015. 153 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-p.html#planck-max.

Beebe:2015:CBPx

[1312] Nelson H. F. Beebe. A complete bibliography of publications in Scientometrics: 1970–1979. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, September 03, 2015. 8 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-s.html#scientometrics1970.

Beebe:2015:CBPy

[1313] Nelson H. F. Beebe. A complete bibliography of publications in Scientometrics: 1980–1989. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, September 03, 2015. 102 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-s.html#scientometrics1980.

Beebe:2015:CBPz

[1314] Nelson H. F. Beebe. A complete bibliography of publications in Scientometrics: 1990–1999. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, September 03, 2015. 178 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-s.html#scientometrics1990.

Beebe:2015:CBPaa

[1315] Nelson H. F. Beebe. A complete bibliography of publications in Scientometrics: 2000–2009. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, September 03, 2015. 268 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-s.html#scientometrics2000.

Beebe:2015:CBPab

[1316] Nelson H. F. Beebe. A complete bibliography of publications in Scientometrics: 2010–2019. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, September 03, 2015. 358 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-s.html#scientometrics2010.

Beebe:2015:CBPac

[1317] Nelson H. F. Beebe. A complete bibliography of publications in the Journal of Informetrics. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, September 09, 2015. 140 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-j.html#jinformetrics.

Beebe:2015:CBPad

[1318] Nelson H. F. Beebe. A complete bibliography of publications in American Documentation. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, September 14, 2015. 166 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-a.html#amdoc.

Beebe:2015:CBPae

[1319] Nelson H. F. Beebe. A complete bibliography of publications in Journal of the American Society for Information Science. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, September 14, 2015. 577 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-j.html#jasis.

Beebe:2015:CBPaf

[1320] Nelson H. F. Beebe. A complete bibliography of publications in Journal of the American Society for Information Science and Technology. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, September 14, 2015. 550 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-j.html#jasist.

Beebe:2015:CBPag

[1321] Nelson H. F. Beebe. A complete bibliography of publications in Journal of the Association for Information Science and Technology. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, September 14, 2015. 80 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-j.html#jaist.

Beebe:2015:CBPah

[1322] Nelson H. F. Beebe. A complete bibliography of publications in the Collnet Journal of Scientometrics and Information Management. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, September 14, 2015. 41 pp.

URL http://www.math.utah.edu/
pub/tex/bib/index-table-c.html#
collnet-j-scientometrics-inf-manage.

Beebe:2015:BPCb

[1323] Nelson H. F. Beebe. A bibliography of The Princeton Companion to Applied Mathematics. Report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, September 15, 2015.

URL http://www.math.utah.edu/
pub/bibnet/subjects/index.html#
companion-applied-mathematics.

Ruhe:2015:BPA

[1324] Axel Ruhe and Nelson H. F. Beebe. A bibliography of publications of Axel Ruhe (1942–2015). Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, September 29, 2015. 17 pp. URL http:

                                        //www.math.utah.edu/pub/bibnet/
                                        authors/r/ruhe-axel.bib.

Beebe:2015:CBI

[1325] Nelson H. F. Beebe. A complete bibliography of IEEE/ACM Transactions on Audio, Speech, and Language Processing. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, October 05, 2015. 98 pp. URL http:

                                        //www.math.utah.edu/pub/tex/bib/
                                        index-table-i.html#ieeeacmtaslp.

Beebe:2015:CBZ

[1326] Nelson H. F. Beebe. A complete bibliography of Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, October 27, 2015. 128 pp.

                                        URL http://www.math.utah.edu/
                                        pub/tex/bib/index-table-z.html#
                                        z-allg-wissenschaftstheor.

Beebe:2015:CBJaa

[1327] Nelson H. F. Beebe. A complete bibliography of the Journal for General Philosophy of Science / Zeitschrift für allgemeine Wissenschaftstheorie. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, October 27, 2015. 143 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-j.html#jgenphilossci.

Beebe:2015:FFS

[1328] Nelson H. F. Beebe. A font family sampler. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, November 4, 2015. 11 pp. URL http:/

/www.math.utah.edu/pub/uuthesis/
font-sampler/font-sampler.pdf.

Beebe:2015:QTM

[1329] Nelson H. F. Beebe. Quantum theory matters. Web site., November 11, 2015. URL http://www.

math.utah.edu/~beebe/talks/2015/
qtm;   http://www.math.utah.edu/
~beebe/talks/2018/qtm. This docu-

ment is a handout for a workshop on the history and significance of quantum theory, from the discovery of the electron, radioactivity, and X-rays in the 1890s, to the ‘new’ quantum theory of the 1920s, to the Manhattan Project in World War II, and on to the modern electronics and telecommunications industries, and the Internet and the World Wide Web.

Beebe:2015:BF

[1330] Nelson H. F. Beebe. BibTEX FAQ.

World-Wide Web frequently-asked question document., December 2015.

URL http://www.math.utah.edu/

faq/bibtex. This report is updated frequently.

Beebe:2015:SBPc

[1331] Nelson H. F. Beebe. A selected bibliography of publications by, and about, Graeme W. Milton. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, December 24, 2015.

                                        67 pp. URL http://www.math.utah.
                                        edu/pub/bibnet/authors/m/. This

report is updated frequently.

Beebe:2015:SBPd

[1332] Nelson H. F. Beebe. A selected bibliography of publications by, and about, Lord Ernest Rutherford of Nelson. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, December 30, 2015. 179

                                        pp. URL http://www.math.utah.
                                        edu/pub/bibnet/authors/r/.

Beebe:2015:CBPai

[1333] Nelson H. F. Beebe. A complete bibliography of publications in The European Physical Journal H: Historical Perspectives on Contemporary Physics. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, December 31, 2015. 29 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-e.html#eur-phys-j-h.

Beebe:2016:CBPa

[1334] Nelson H. F. Beebe. A complete bibliography of publications in The Rutherford Journal [the New Zealand Journal for the History and Philosophy of Science and Technology]. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, January 4, 2016. 8 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-r.html#rutherfordj.

Beebe:2016:CBPb

[1335] Nelson H. F. Beebe. A complete bibliography of publications in Communications in Statistics. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, January 28, 2016. 50 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-c.html#communstat.

Beebe:2016:CBPc

[1336] Nelson H. F. Beebe. A complete bibliography of publications in Communications in Statistics: Theory and Methods: 1970–1979. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, January 26, 2016. 89 pp.

URL http://www.math.utah.edu/
pub/tex/bib/index-table-c.html#
communstattheorymeth1970.

Beebe:2016:CBPd

[1337] Nelson H. F. Beebe. A complete bibliography of publications in Communications in Statistics: Theory and Methods: 1980–1989. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, January 28, 2016. 395 pp.

URL http://www.math.utah.edu/
pub/tex/bib/index-table-c.html#
communstattheorymeth1980.

Beebe:2016:CBPe

[1338] Nelson H. F. Beebe. A complete bibliography of publications in Communications in Statistics: Theory and Methods: 1990–1999. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, January 29, 2016. 402 pp.

                                        URL http://www.math.utah.edu/
                                        pub/tex/bib/index-table-c.html#
                                        communstattheorymeth1990.

Beebe:2016:CBPf

[1339] Nelson H. F. Beebe. A complete bibliography of publications in Communications in Statistics: Theory and Methods: 2000–2009. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, January 29, 2016. 368 pp.

                                        URL http://www.math.utah.edu/
                                        pub/tex/bib/index-table-c.html#
                                        communstattheorymeth2000.

Beebe:2016:CBPg

[1340] Nelson H. F. Beebe. A complete bibliography of publications in Communications in Statistics: Theory and Methods: 2010–2019. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, January 29, 2016. 377 pp.

                                        URL http://www.math.utah.edu/
                                        pub/tex/bib/index-table-c.html#
                                        communstattheorymeth2010.

Beebe:2016:CBPh

[1341] Nelson H. F. Beebe. A complete bibliography of publications in Communications in Statistics: Simulation and Computation: 1970– 1979. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, January 30, 2016. 22 pp.

                                        URL http://www.math.utah.edu/
                                        pub/tex/bib/index-table-c.html#
                                        communstatsimulcomput1970.

Beebe:2016:CBPi

[1342] Nelson H. F. Beebe. A complete bibliography of publications in Communications in Statistics: Simulation and Computation: 1980– 1989. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, January 30, 2016. 115 pp.

URL http://www.math.utah.edu/
pub/tex/bib/index-table-c.html#
communstatsimulcomput1980.

Beebe:2016:CBPj

[1343] Nelson H. F. Beebe. A complete bibliography of publications in Communications in Statistics: Simulation and Computation: 1990– 1999. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, January 30, 2016. 123 pp.

URL http://www.math.utah.edu/
pub/tex/bib/index-table-c.html#
communstatsimulcomput1990.

Beebe:2016:CBPk

[1344] Nelson H. F. Beebe. A complete bibliography of publications in Communications in Statistics: Simulation and Computation: 2000– 2009. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, January 30, 2016. 144 pp.

URL http://www.math.utah.edu/
pub/tex/bib/index-table-c.html#
communstatsimulcomput2000.

Beebe:2016:CBPl

[1345] Nelson H. F. Beebe. A complete bibliography of publications in Communications in Statistics: Simulation and Computation: 2010– 2019. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, January 30, 2016. 165 pp.

URL http://www.math.utah.edu/
pub/tex/bib/index-table-c.html#
communstatsimulcomput2010.

Beebe:2016:CBPm

[1346] Nelson H. F. Beebe. A complete bibliography of publications in Linear Algebra and its Applications: 1968–1969. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, February 17, 2016. 14 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-l.html#linala1960. This re-

port is updated frequently.

Beebe:2016:CBPn

[1347] Nelson H. F. Beebe. A complete bibliography of publications in Linear Algebra and its Applications: 1970–1979. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, February 17, 2016. 130 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-l.html#linala1970. This re-

port is updated frequently.

Beebe:2016:CBPo

[1348] Nelson H. F. Beebe. A complete bibliography of publications in the Journal of Mathematical Biology. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, February 18, 2016. 463 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-

table-j.html#jmathbiol. This re-

port is updated frequently.

Beebe:2016:CBPp

[1349] Nelson H. F. Beebe. A complete bibliography of publications in Contemporary Physics. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, February 22, 2016. 529 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-c.html#contempphys. This re-

port is updated frequently.

Beebe:2016:CBPq

[1350] Nelson H. F. Beebe. A complete bibliography of publications in History and Technology. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, March 2, 2016. 109 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-h.html#histtechnol. This re-

port is updated frequently.

Beebe:2016:BPF

[1351] Nelson H. F. Beebe. A bibliography of publications in Fundamenta Informaticae: 1990–1999. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, March 7, 2016. 82 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-f.html#fundinfo1990. This

report is updated frequently.

Beebe:2016:CBPs

[1352] Nelson H. F. Beebe. A complete bibliography of publications in Fundamenta Informaticae: 2000–2009. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, March 7, 2016. 222 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-f.html#fundinfo2000. This

report is updated frequently.

Beebe:2016:CBPt

[1353] Nelson H. F. Beebe. A complete bibliography of publications in Fundamenta Informaticae: 2010–2019. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, March 7, 2016. 175 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-f.html#fundinfo2010. This

report is updated frequently.

Beebe:2016:CBPu

[1354] Nelson H. F. Beebe. A complete bibliography of publications in the Bulletin for the History of Chemistry. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, March 13, 2016. 105 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-b.html#bullhistchem. This

report is updated frequently.

Beebe:2016:CBPv

[1355] Nelson H. F. Beebe. A complete bibliography of publications in NTM Zeitschrift für Geschichte der Wissenschaften, Technik und Medizin == NTM Journal of the History of Science, Technology and Medicine. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, March 15, 2016. 138 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-n.html#ntm. This report is up-

dated frequently.

Beebe:2016:CBPw

[1356] Nelson H. F. Beebe. A complete bibliography of publications in the Journal of Automated Reasoning. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, April 2, 2016. 142 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-j.html#jautomreason. This

report is updated frequently.

Beebe:2016:CBPx

[1357] Nelson H. F. Beebe. A complete bibliography of publications in the Journal of Computer Security. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, May 24, 2016. 96 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-j.html#jcompsecur. This re-

port is updated frequently.

Beebe:2016:CBPy

[1358] Nelson H. F. Beebe. A complete bibliography of publications in the Journal of Cold War Studies. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, June 23, 2016. 313 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-j.html#jcoldwarstud. This

report is updated frequently.

Beebe:2016:CBPz

[1359] Nelson H. F. Beebe. A complete bibliography of publications in the Australian Journal of Statistics and the Australian & New Zealand Journal of Statistics. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, July 16, 2016. 309 pp. URL http:

//www.math.utah.edu/pub/tex/bib/
index-table-a.html#anzjs. This re-

port is updated frequently.

Beebe:2016:CBPaa

[1360] Nelson H. F. Beebe. A complete bibliography of publications in Minerva: A Review of Science, Learning and Policy. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, July 21, 2016. 325 pp. URL http:

//www.math.utah.edu/pub/tex/bib/
index-table-m.html#minerva. This

report is updated frequently.

Beebe:2016:CBPab

[1361] Nelson H. F. Beebe. A complete bibliography of the publications of Jonathan Michael Borwein. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, August 17, 2016. 264 pp. URL http:

//www.math.utah.edu/pub/bibnet/
authors/b/borwein-jonathan-m.bib.

This report is updated frequently.

Beebe:2016:CBPac

[1362] Nelson H. F. Beebe. A complete bibliography of publications in Linear [and] Multilinear Algebra. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, September 21, 2016. 468 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-

table-l.html#linmultalgebra.bib.

This report is updated frequently.

Beebe:2016:ZWY

[1363] Nelson H. F. Beebe and Pieter J. Bowman. ZFS — and why you need it. Web site, November 10, 2016.

URL http://www.math.utah.edu/~
beebe/talks/2016/zfs;   http://
www.math.utah.edu/~beebe/talks/

2017/zfs. This document is the slides from a joint lecture on the importance and features of the Zettabyte File System (now just ZFS). The talk has been presented twice, with updates, to computer professionals and at the 2017 Rocky Mountain High-Performance Computing symposium.

Beebe:2016:CBPad

[1364] Nelson H. F. Beebe. A complete bibliography of publications in BJHS Themes. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, November 25, 2016. 4 pp. URL http:

//www.math.utah.edu/pub/tex/bib/
index-table-b.html#bjhs-themes.

This report is updated frequently.

Beebe:2016:CBPae

[1365] Nelson H. F. Beebe. A complete bibliography of publications in British Society for the History of Science Monographs. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, November 26, 2016. 5 pp.

URL http://www.math.utah.edu/
pub/tex/bib/index-table-b.html#

bshs-monographs. This report is updated frequently.

Beebe:2016:CBPaf

[1366] Nelson H. F. Beebe. A complete bibliography of publications in Empirical Software Engineering. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, December 2, 2016. 102 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-e.html#empir-software-eng.

This report is updated frequently.

Beebe:2016:CBPag

[1367] Nelson H. F. Beebe. A complete bibliography of publications in Technovation. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, December 24, 2016. 578 pp. URL http:

                                        //www.math.utah.edu/pub/tex/bib/
                                        index-table-t.html#technovation.

This report is updated frequently.

Beebe:2017:CBPa

[1368] Nelson H. F. Beebe. A complete bibliography of publications in Educational Studies in Mathematics. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, January 14, 2017. 448 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-e.html#educstudmath. This

report is updated frequently.

Beebe:2017:CBPb

[1369] Nelson H. F. Beebe. A complete bibliography of publications in Journal of Computational and Applied Mathematics (1970–1979). Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, January 14, 2017. 37 pp.

URL http://www.math.utah.edu/
pub/tex/bib/index-table-j.html#
jcomputapplmath1970. This report is

updated frequently.

Beebe:2017:CBPc

[1370] Nelson H. F. Beebe. A complete bibliography of publications in Journal of Computational and Applied Mathematics (1980–1989). Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, January 14, 2017. 162 pp.

URL http://www.math.utah.edu/
pub/tex/bib/index-table-j.html#
jcomputapplmath1980. This report is

updated frequently.

Beebe:2017:CBPd

[1371] Nelson H. F. Beebe. A complete bibliography of publications in Journal of Computational and Applied Mathematics (1990–1999). Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, January 14, 2017. 471 pp.

URL http://www.math.utah.edu/
pub/tex/bib/index-table-j.html#
jcomputapplmath1990. This report is

updated frequently.

Beebe:2017:CBPe

[1372] Nelson H. F. Beebe. A complete bibliography of publications in Journal of Computational and Applied Mathematics (2000–2005). Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, January 14, 2017. 358 pp.

URL http://www.math.utah.edu/
pub/tex/bib/index-table-j.html#
jcomputapplmath2000. This report is

updated frequently.

Beebe:2017:CBPf

[1373] Nelson H. F. Beebe. A complete bibliography of publications in Journal of Computational and Applied Mathematics (2005–2009). Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, January 14, 2017. 541 pp.

                                        URL http://www.math.utah.edu/
                                        pub/tex/bib/index-table-j.html#
                                        jcomputapplmath2005. This report is

updated frequently.

Beebe:2017:CBPg

[1374] Nelson H. F. Beebe. A complete bibliography of publications in Journal of Computational and Applied Mathematics (2010–2014). Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, January 14, 2017. 497 pp.

                                        URL http://www.math.utah.edu/
                                        pub/tex/bib/index-table-j.html#
                                        jcomputapplmath2010. This report is

updated frequently.

Beebe:2017:CBPh

[1375] Nelson H. F. Beebe. A complete bibliography of publications in Journal of Computational and Applied Mathematics (2015–2019). Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, January 14, 2017. 307 pp.

                                        URL http://www.math.utah.edu/
                                        pub/tex/bib/index-table-j.html#
                                        jcomputapplmath2015. This report is

updated frequently.

Beebe:2017:CBPi

[1376] Nelson H. F. Beebe. A complete bibliography of publications in Computers and Mathematics with Applications (1970–1979). Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, March 1, 2017. 41 pp.

URL http://www.math.utah.edu/
pub/tex/bib/index-table-c.html#
computmathappl1970. This report is

updated frequently.

Beebe:2017:CBPj

[1377] Nelson H. F. Beebe. A complete bibliography of publications in Computers and Mathematics with Applications (1980–1989). Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, March 1, 2017. 214 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-c.html#computmathappl1980.

This report is updated frequently.

Beebe:2017:CBPk

[1378] Nelson H. F. Beebe. A complete bibliography of publications in Computers and Mathematics with Applications (1990–1999). Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, March 1, 2017. 680 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-c.html#computmathappl1990.

This report is updated frequently.

Beebe:2017:CBPl

[1379] Nelson H. F. Beebe. A complete bibliography of publications in Computers and Mathematics with Applications (2000–2009). Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, March 1, 2017. 764 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-c.html#computmathappl2000.

This report is updated frequently.

Beebe:2017:CBPm

[1380] Nelson H. F. Beebe. A complete bibliography of publications in Computers and Mathematics with Applications (2010–2019). Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, March 1, 2017. 846 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-c.html#computmathappl2010.

This report is updated frequently.

Beebe:2017:CBPn

[1381] Nelson H. F. Beebe. A complete bibliography of publications in Groups Complexity Cryptology. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, March 17, 2017. 18 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-g.html#groups-complex-cryptol.

This report is updated frequently.

Beebe:2017:CBPo

[1382] Nelson H. F. Beebe. A complete bibliography of publications in International Journal of Applied Cryptography. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, March 17, 2017. 11 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-i.html#ijact. This report is

updated frequently.

Beebe:2017:CBPp

[1383] Nelson H. F. Beebe. A complete bibliography of publications in Journal of Mathematical Cryptology. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, March 17, 2017. 26 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-j.html#jmathcryptol. This

report is updated frequently.

Beebe:2017:CBPq

[1384] Nelson H. F. Beebe. A complete bibliography of publications in ACM Transactions on Asian and Low-Resource Language Information Processing. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, April 3, 2017. 15 pp. URL http:

//www.math.utah.edu/pub/tex/bib/
index-table-t.html#tallip. This

report is updated frequently.

Beebe:2017:ESB

[1385] Nelson H. F. Beebe. eXtended sample BibTEX styles. Report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, April 11, 2017. 25

pp. URL http://www.math.utah.
edu/pub/tex/x-bst/. This report has

four automatically-generated versions to illustrate different document font families, and four extended bibliography styles.

Beebe:2017:CBPr

[1386] Nelson H. F. Beebe. A complete bibliography of publications in ACM Transactions on Privacy and Security (TOPS). Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, April 3, 2017. 4 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-t.html#tops. This report is up-

dated frequently.

Beebe:2017:CBPs

[1387] Nelson H. F. Beebe. A complete bibliography of publications by, and about, Leopold Infeld. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, May 31, 2017. 67

pp. URL http://www.math.utah.
edu/pub/bibnet/authors/i/infeld-

leopold.bib. This report is updated frequently.

Beebe:2017:CBPt

[1388] Nelson H. F. Beebe. A complete bibliography of publications in Foundations of Physics Letters. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, June 3, 2017. 148 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-f.html#foundphyslett. This

report is updated frequently.

Beebe:2017:CBPu

[1389] Nelson H. F. Beebe. A complete bibliography of publications in Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT). Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, June 16, 2017. 2 pp. URL http:

//www.math.utah.edu/pub/tex/bib/

index-table-i.html#imwut. This re-

port is updated frequently.

Beebe:2017:CBPv

[1390] Nelson H. F. Beebe. A complete bibliography of publications in Proceedings of the ACM on Measurement and Analysis of Computing Systems (PO-MACS). Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, June 16, 2017. 7 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-p.html#pomacs. This report is

updated frequently.

Beebe:2017:CBPw

[1391] Nelson H. F. Beebe. A complete bibliography of publications in ACM Transactions on Cyber-Physical Systems (TCPS). Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, June 16, 2017. 4 pp. URL http:

//www.math.utah.edu/pub/tex/bib/
index-table-t.html#tcps. This re-

port is updated frequently.

Beebe:2017:CBPx

[1392] Nelson H. F. Beebe. A complete bibliography of publications in ACM Transactions on Modeling and Performance Evaluation of Computing Systems (TOMPECS). Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, June 16, 2017. 8 pp. URL http:

//www.math.utah.edu/pub/tex/bib/
index-table-t.html#tompecs. This

report is updated frequently.

Beebe:2017:CBPy

[1393] Nelson H. F. Beebe. A complete bibliography of publications in ACM Transactions on Spatial Algorithms and Systems (TSAS). Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, June 16, 2017. 7 pp. URL http:

                                        //www.math.utah.edu/pub/tex/bib/
                                        index-table-t.html#tsas. This re-

port is updated frequently.

Beebe:2017:CBPz

[1394] Nelson H. F. Beebe. A complete bibliography of publications in Science & Education (Springer). Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, June 19, 2017. 237 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-s.html#sci-educ-springer.

This report is updated frequently.

Beebe:2017:BPB

[1395] Nelson H. F. Beebe. A bibliography of publications in British Society for the History of Mathematics Newsletter and BSHM Bulletin: Journal of the British Society for the History of Mathematics. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, June 20, 2017. 61 pp. URL http:

                                        //www.math.utah.edu/pub/tex/bib/
                                        index-table-b.html#bshm. This re-

port is updated frequently.

Beebe:2017:CBPaa

[1396] Nelson H. F. Beebe. A complete bibliography of publications in the Bulletin of Mathematical Biophysics. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, June 29, 2017. 201 pp.

URL http://www.math.utah.edu/
pub/tex/bib/index-table-b.html#

bullmathbiophys. This report is updated frequently.

Beebe:2017:CBPab

[1397] Nelson H. F. Beebe. A complete bibliography of publications in the Bulletin of Mathematical Biology. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, June 29, 2017. 638 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-b.html#bullmathbiol. This

report is updated frequently.

Beebe:2017:CBPac

[1398] Nelson H. F. Beebe. A complete bibliography of publications in the Journal of Functional Programming. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, July 12, 2017. 176 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-j.html#jfunctprogram. This

report is updated frequently.

Beebe:2017:CBPad

[1399] Nelson H. F. Beebe. A complete bibliography of publications by, and about, Louis de Broglie. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, May 31, 2017.

67 pp. URL http://www.math.utah.
edu/pub/bibnet/authors/i/infeld-

leopold.bib. This report is updated frequently.

Beebe:2017:MFC

[1400] Nelson H. F. Beebe. The Mathematical-Function Computation Handbook: Programming Using the MathCW Portable Software Library. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2017. ISBN 3-319-64109-3 (hardcover), 3-319-64110-7 (e-book). xxxvi + 1114 pp. LCCN QA75.5-76.95. URL

                                        http://www.springer.com/us/book/
                                        9783319641096.

Beebe:2017:CBPae

[1401] Nelson H. F. Beebe. A complete bibliography of publications in Dialectica. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, July 20, 2017. 343 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-d.html#dialectica. This re-

port is updated frequently.

Beebe:2017:CBPaf

[1402] Nelson H. F. Beebe. A complete bibliography of publications in Theoria. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, July 20, 2017. 204 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-t.html#theoria. This report is

updated frequently.

Beebe:2017:CBPag

[1403] Nelson H. F. Beebe. A complete bibliography of publications by, and about, Louis de Broglie. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, July 29, 2017.

235 pp. URL http://www.math.utah.
edu/pub/bibnet/authors/d/. This

report is updated frequently.

Beebe:2017:CBPah

[1404] Nelson H. F. Beebe. A complete bibliography of publications in Sudhoffs Archiv: Zeitschrift für Wissenschaftsgeschichte. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, August 11, 2017. 557 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-s.html#sudhoffs-arch.

Beebe:2017:SBP

[1405] Nelson H. F. Beebe. A selected bibliography of publications by, and about, Arnold Sommerfeld. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, August 15, 2017.

143 pp. URL http://www.math.utah.
edu/pub/bibnet/authors/s/.

Beebe:2017:CBPai

[1406] Nelson H. F. Beebe. A complete bibliography of publications in Advances and Applications in Statistics. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, August 16, 2017. 136 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-a.html#advapplstat.

Beebe:2017:CBPaj

[1407] Nelson H. F. Beebe. A complete bibliography of publications in Probability in the Engineering and Informational Sciences. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, August 24, 2017. 267 pp. URL http:

                                        //www.math.utah.edu/pub/tex/bib/
                                        index-table-p.html#probab-engrg-
                                        inform-sci.

Beebe:2017:CBJ

[1408] Nelson H. F. Beebe. A complete bibliography of the journal family Printing History. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, August 28, 2017. 85 pp.

                                        URL http://www.math.utah.edu/
                                        pub/tex/bib/index-table-p.html#
                                        printing-history.

Beebe:2017:CBPak

[1409] Nelson H. F. Beebe. A complete bibliography of publications in International Journal of Science and Mathematics Education. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, September 9, 2017. 188 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-i.html#ijsme.

Beebe:2017:CBPal

[1410] Nelson H. F. Beebe. A complete bibliography of publications by, and about, Arnold Sommerfeld. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, September 12, 2017.

                                        199 pp. URL http://www.math.utah.
                                        edu/pub/bibnet/authors/s/.

Beebe:2017:PRN

[1411] Nelson H. F. Beebe. Pseudo-random numbers: [mostly] a line [of code] at a time. Web site, September 20, 2017.

URL http://www.math.utah.edu/~
beebe/talks/2017/pseudo-random-

numbers-2017. This document is a handout for undergraduate colloquia that have been presented several times, each time with updates. The original edition was [738].

Beebe:2017:CBPam

[1412] Nelson H. F. Beebe. A complete bibliography of publications in Synthese, 1930–1939. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, September 29, 2017. 61 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-s.html#synthese1930.

Beebe:2017:CBPan

[1413] Nelson H. F. Beebe. A complete bibliography of publications in Synthese, 1940–1949. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, September 29, 2017. 39 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-s.html#synthese1940.

Beebe:2017:CBPao

[1414] Nelson H. F. Beebe. A complete bibliography of publications in Synthese, 1950–1959. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, September 29, 2017. 28 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-s.html#synthese1950.

Beebe:2017:CBPap

[1415] Nelson H. F. Beebe. A complete bibliography of publications in Synthese, 1960–1969. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, September 29, 2017. 46 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-s.html#synthese1960.

Beebe:2017:CBPaq

[1416] Nelson H. F. Beebe. A complete bibliography of publications in Synthese, 1970–1979. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, September 29, 2017. 88 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-s.html#synthese1970.

Beebe:2017:CBPar

[1417] Nelson H. F. Beebe. A complete bibliography of publications in Synthese, 1990–1999. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, September 29, 2017. 128 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-s.html#synthese1990.

Beebe:2017:CBPas

[1418] Nelson H. F. Beebe. A complete bibliography of publications in Synthese, 2000–2009. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, September 29, 2017. 215 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-s.html#synthese2000.

Beebe:2017:CBPat

[1419] Nelson H. F. Beebe. A complete bibliography of publications in Synthese, 2010–2019. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, September 29, 2017. 295 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-s.html#synthese2010.

Beebe:2017:CBPau

[1420] Nelson H. F. Beebe. A complete bibliography of publications in the SIAM Journal on Applied Algebra and Geometry. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, October 01, 2017. 5 pp.

URL http://www.math.utah.edu/
pub/tex/bib/index-table-s.html#
siamjapplalgebrgeom.

Beebe:2017:DAS

[1421] Nelson H. F. Beebe. Decimalarithmetic support in gcc compilers. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, October 27, 2017. 13 pp. URL http://www.

math.utah.edu/pub/mathcw/dgcc/.

Beebe:2017:CBPav

[1422] Nelson H. F. Beebe. A complete bibliography of publications in Supercomputing Frontiers and Innovations. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, November 11, 2017. 19 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-s.html#superfri.

Beebe:2017:BPA

[1423] Nelson H. F. Beebe. A bibliography of publications about Bitcoin and digital cash systems. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, November 30, 2017. 104 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-b.html#bitcoin.

Beebe:2017:CBPaw

[1424] Nelson H. F. Beebe. A complete bibliography of publications in Network Security. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, December 4, 2017. 478 pp.

URL http://www.math.utah.edu/
pub/tex/bib/index-table-s.html#
network-security.

Beebe:2018:CBPa

[1425] Nelson H. F. Beebe. A complete bibliography of Proceedings of the ACM on Programming Languages (PACMPL). Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, January 10, 2018. 30 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-p.html#pacmpl.

Beebe:2018:CBPb

[1426] Nelson H. F. Beebe. A complete bibliography of publications in International Journal of Information Security. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, January 23, 2018. 96 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-i.html#intjinfosec.

Beebe:2018:CBPc

[1427] Nelson H. F. Beebe. A bibliography of publications in Software Process: Improvement and Practice. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, January 30, 2018. 64 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-s.html#spip.

Beebe:2018:CBPd

[1428] Nelson H. F. Beebe. A complete bibliography of publications in the Journal of Software Maintenance and Evolution: Research and Practice. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, January 31, 2018. 44 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-j.html#jsoftwmaintevol.

Beebe:2018:CBPe

[1429] Nelson H. F. Beebe. A complete bibliography of publications in the Journal of Software Maintenance: Research and Practice. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, January 31, 2018. 44 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-j.html#jsoftwmaint.

Beebe:2018:CBPf

[1430] Nelson H. F. Beebe. A complete bibliography of publications in the Journal of Software: Evolution and Process. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, January 31, 2018. 81 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-j.html#jsoftwevolproc.

Beebe:2018:CBPg

[1431] Nelson H. F. Beebe. A complete bibliography of publications in IEEE Transactions on Software Engineering (1970–1979). Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, February 1, 2018. 68 pp.

                                        URL http://www.math.utah.edu/
                                        pub/tex/bib/index-table-j.html#
                                        ieeetranssoftweng1970.

Beebe:2018:CBPh

[1432] Nelson H. F. Beebe. A complete bibliography of publications in IEEE Transactions on Software Engineering (1980–1989). Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, February 1, 2018. 244 pp.

                                        URL http://www.math.utah.edu/
                                        pub/tex/bib/index-table-j.html#
                                        ieeetranssoftweng1980.

Beebe:2018:CBPi

[1433] Nelson H. F. Beebe. A complete bibliography of publications in IEEE Transactions on Software Engineering (1990–1999). Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, February 1, 2018. 151 pp.

                                        URL http://www.math.utah.edu/
                                        pub/tex/bib/index-table-j.html#
                                        ieeetranssoftweng1990.

Beebe:2018:CBPj

[1434] Nelson H. F. Beebe. A complete bibliography of publications in IEEE Transactions on Software Engineering (2000–2009). Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, February 1, 2018. 180 pp.

URL http://www.math.utah.edu/
pub/tex/bib/index-table-j.html#
ieeetranssoftweng2000.

Beebe:2018:CBPk

[1435] Nelson H. F. Beebe. A complete bibliography of publications in IEEE Transactions on Software Engineering (2010–2019). Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, February 1, 2018. 148 pp.

URL http://www.math.utah.edu/
pub/tex/bib/index-table-j.html#
ieeetranssoftweng2010.

Beebe:2018:CBPl

[1436] Nelson H. F. Beebe. A complete bibliography of publications in Advances in Computational Mathematics. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, February 5, 2018. 200 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-a.html#advcomputmath.

Beebe:2018:BPAa

[1437] Nelson H. F. Beebe. A bibliography of publications by, and about, Ingrid Daubechies. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, March 13, 2018. 89

pp. URL http://www.math.utah.
edu/pub/bibnet/authors/d/.

Beebe:2018:BPAb

[1438] Nelson H. F. Beebe. A bibliography of publications by, and about, Olga Taussky [Olga Taussky-Todd]. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, March 29, 2018.

                                        78 pp. URL http://www.math.utah.
                                        edu/pub/bibnet/authors/t/.

Beebe:2018:CBPm

[1439] Nelson H. F. Beebe. A complete bibliography of publications in Revue d’Histoire Mathématiques = Journal for the History of Mathematics. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, April 3, 2018. 50 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-r.html#revhistmath.

Beebe:2018:CBPn

[1440] Nelson H. F. Beebe. A complete bibliography of publications in Notes and Records of the Royal Society of London. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, April 6, 2018. 341 pp. URL http:

                                        //www.math.utah.edu/pub/tex/bib/
                                        index-table-n.html#notes-rec-r-
                                        soc-lond.

Beebe:2018:BPAc

[1441] Nelson H. F. Beebe. A bibliography of publications by, and about, Charles Babbage. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, April 7, 2018. 106

                                        pp. URL http://www.math.utah.
                                        edu/pub/bibnet/authors/b/.

Beebe:2018:BPAd

[1442] Nelson H. F. Beebe. A bibliography of publications by, and about, Augusta Ada King, Countess of Lovelace. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, April 7,

2018. 25 pp. URL http://www.math.
utah.edu/pub/bibnet/authors/l/.

Beebe:2018:CBPo

[1443] Nelson H. F. Beebe. A complete bibliography of publications in the ORSA Journal on Computing and the IN-FORMS Journal on Computing. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, April 9, 2018. 252 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-i.html#informs-j-comput.

Beebe:2018:BPI

[1444] Nelson H. F. Beebe. A bibliography of publications in International Journal of Computer Applications. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, April 21, 2018. 109 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-i.html#ijca.

Beebe:2018:SBPa

[1445] Nelson H. F. Beebe. A selected bibliography of publications by, and about, Sir Rudolf Peierls. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, April 23, 2018.

110 pp. URL http://www.math.utah.
edu/pub/bibnet/authors/p/.

Beebe:2018:CBPp

[1446] Nelson H. F. Beebe. A complete bibliography of publications in Ambix: Journal of the Society for the History of Alchemy and Chemistry. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, April 25, 2018. 218 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-a.html#ambix.

Beebe:2018:SBPb

[1447] Nelson H. F. Beebe. A selected bibliography of publications by, and about, Otto Robert Frisch. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, May 2, 2018. 64

                                        pp. URL http://www.math.utah.
                                        edu/pub/bibnet/authors/f/.

Bailey:2018:PTA

[1448] David H. Bailey and Nelson H. F. Beebe. Publications and talks by (and about) Jonathan M. Borwein. Web site., 2018. URL https://www.

                                        jonborwein.org/jmbpapers/.

Beebe:2018:SBPc

[1449] Nelson H. F. Beebe. A selected bibliography of publications by, and about, Lise Meitner. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, May 19, 2018. 135

                                        pp. URL http://www.math.utah.
                                        edu/pub/bibnet/authors/m/.

Beebe:2018:CBPq

[1450] Nelson H. F. Beebe. A complete bibliography of publications in Archimedes:

New Studies in the History and Philosophy of Science and Technology. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, June 13,

2018. 179 pp. URL http://www.math.
utah.edu/pub/bibnet/authors/a/.

Beebe:2018:CBPr

[1451] Nelson H. F. Beebe. A complete bibliography of publications by, and about, James Clerk Maxwell. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, June 25, 2018.

166 pp. URL http://www.math.utah.
edu/pub/bibnet/authors/c/.

Beebe:2018:CBPs

[1452] Nelson H. F. Beebe. A complete bibliography of publications in ACM SIGSOFT Software Engineering Notes (1970–1979). Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, August 1, 2018. 17 pp.

URL http://www.math.utah.edu/
pub/tex/bib/index-table-s.html#
sigsoft1970.

Beebe:2018:CBPt

[1453] Nelson H. F. Beebe. A complete bibliography of publications in ACM SIGSOFT Software Engineering Notes (1980–1989). Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, August 1, 2018. 99 pp.

URL http://www.math.utah.edu/
pub/tex/bib/index-table-s.html#
sigsoft1980.

Beebe:2018:CBPu

[1454] Nelson H. F. Beebe. A complete bibliography of publications in ACM SIGSOFT Software Engineering Notes (1990–1999). Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, August 1, 2018. 150 pp.

                                        URL http://www.math.utah.edu/
                                        pub/tex/bib/index-table-s.html#
                                        sigsoft1990.

Beebe:2018:CBPv

[1455] Nelson H. F. Beebe. A complete bibliography of publications in ACM SIGSOFT Software Engineering Notes (2000–2009). Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, August 1, 2018. 246 pp.

                                        URL http://www.math.utah.edu/
                                        pub/tex/bib/index-table-s.html#
                                        sigsoft2000.

Beebe:2018:CBPw

[1456] Nelson H. F. Beebe. A complete bibliography of publications in ACM SIGSOFT Software Engineering Notes (2010–2019). Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, August 1, 2018. 109 pp.

                                        URL http://www.math.utah.edu/
                                        pub/tex/bib/index-table-s.html#
                                        sigsoft2010.

Beebe:2018:BPAe

[1457] Nelson H. F. Beebe. A bibliography of publications by, and about, Richard Wesley Hamming. Report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, August 8, 2018. URL http:

//www.math.utah.edu/pub/bibnet/
authors/h/.

Beebe:2018:CBPx

[1458] Nelson H. F. Beebe. A complete bibliography of publications in History of Geo- and Space Sciences. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, August 30, 2018. 22 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-h.html#hgss.

Beebe:2018:CBPy

[1459] Nelson H. F. Beebe. A complete bibliography of publications in the Journal of Open Research Software. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, September 8, 2018. 34 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-j.html#jors.

Beebe:2018:CBPz

[1460] Nelson H. F. Beebe. A complete bibliography of publications in the Journal of Open Source Software. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, September 13, 2018. 65 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-j.html#joss.

Beebe:2018:CBPaa

[1461] Nelson H. F. Beebe. A complete bibliography of publications in Spontaneous Generations: A Journal for the History and Philosophy of Science. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, September 25, 2018. 36 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-s.html#spontgen.

Beebe:2018:CBPab

[1462] Nelson H. F. Beebe. A complete bibliography of publications in International Studies in the Philosophy of Science. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, September 25, 2018. 137 pp.

                                        URL http://www.math.utah.edu/
                                        pub/tex/bib/index-table-s.html#
                                        intstudphilossci.

Beebe:2018:CBPac

[1463] Nelson H. F. Beebe. A complete bibliography of publications in Japanese Studies in the History of Science. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, October 8, 2018. 38 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-j.html#jshs.

Beebe:2018:CBPad

[1464] Nelson H. F. Beebe. A complete bibliography of publications in Archive of Numerical Software. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, October 15, 2018. 5 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-a.html#archnumsoft.

Beebe:2018:IBP

[1465] Nelson H. F. Beebe. An incomplete bibliography of publications in Historia Scientiarum (International Journal of the History of Science Society of Japan). Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, October 15, 2018. 99 pp. URL http:

//www.math.utah.edu/pub/tex/bib/
index-table-h.html#histscijpn.

Beebe:2018:NFUa

[1466] Nelson H. F. Beebe. Notes for the Fall 2018 University of Utah styles for Mathematics Department dissertations and theses in LaTEX. Report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, October 16, 2018. 20 pp. URL http:

//www.math.utah.edu/pub/uumath-
thesis/uumath-thesis-2018-10-16-

notes.pdf. These notes accompany an extensive set of LaTEX style files and sample Masters theses and doctoral dissertations that are intended to replace the vintage styles from the early 1990s.

Beebe:2018:NFUb

[1467] Nelson H. F. Beebe. Notes for the Fall 2018 University of Utah styles for Mathematics Department dissertations and theses in LaTEX. Report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, October 25, 2018. 23 pp. URL http:

//www.math.utah.edu/pub/uumath-
thesis/uumath-thesis-2018-10-25-

notes.pdf. These notes accompany an extensive set of LaTEX style files and sample Masters theses and doctoral dissertations that are intended to replace the vintage styles from the early 1990s.

Beebe:2018:BPB

[1468] Nelson H. F. Beebe. A bibliography of publications in the book series Boston Studies in the Philosophy of Science and Boston Studies in the Philosophy and History of Science. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, November 10, 2018. 910 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-b.html#bsphs.

Beebe:2018:BIT

[1469] Nelson H. F. Beebe. A brief introduction to TEX and LaTEX for the University of Utah SIAM Student Chapter meeting. Web site, November 13,

                                        2018. URL http://www.math.utah.
                                        edu/pub/tex/doc/latex-overview/.

This Web site contains the notes for a two-hour workshop that I have given several times to introduce novices to computer typesetting with TEX and LaTEX, up to the point of writing a simple book with bibliography and index.

Beebe:2018:CBPae

[1470] Nelson H. F. Beebe. A complete bibliography of publications in Science Studies. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, December 4, 2018. 22 pp. URL http:

                                        //www.math.utah.edu/pub/tex/bib/
                                        index-table-s.html#scistud.

Beebe:2018:CBPaf

[1471] Nelson H. F. Beebe. A complete bibliography of publications in Social Studies of Science. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, December 5, 2018. 396 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-s.html#socstudsci.

Beebe:2018:BPAf

[1472] Nelson H. F. Beebe. A bibliography of publications of Alston Scott Householder. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, December 6, 2018. 73

pp. URL http://www.math.utah.
edu/pub/bibnet/authors/h/.

Beebe:2019:BPAa

[1473] Nelson H. F. Beebe. A bibliography of publications by, and about, Peter Henrici. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, January 28, 2019. 43

pp. URL http://www.math.utah.
edu/pub/bibnet/authors/h/.

Beebe:2019:CBPa

[1474] Nelson H. F. Beebe. A complete bibliography of publications in the Journal of Computer and System Sciences. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, January 30, 2019. 496 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-j.html#jcompsyssci.

Beebe:2019:CBPb

[1475] Nelson H. F. Beebe. A complete bibliography of publications in Numerical Algebra, Control and Optimization. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, January 31, 2019. 53 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-n.html#naco.

Beebe:2019:NSUa

[1476] Nelson H. F. Beebe. Notes for the Spring 2019 University of Utah styles for Mathematics Department dissertations and theses in LaTEX. Report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, February 01, 2019. 28 pp. URL http://www.

                                        math.utah.edu/pub/uumath-thesis/
                                        uumath-thesis-2019-02-01-notes.

pdf. These notes accompany an extensive set of LaTEX style files and sample Masters theses and doctoral dissertations that are intended to replace the vintage styles from the early 1990s.

Beebe:2019:CBJa

[1477] Nelson H. F. Beebe. A complete bibliography of the journal family The Two-Year College Mathematics Journal (1970–1983) and The College Mathematics Journal (1984–date). Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, February 14, 2019. 625 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-c.html#collegemathj.

Beebe:2019:SBPa

[1478] Nelson H. F. Beebe. A selected bibliography of publications by, and about, Mario Bunge. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, February 24, 2019. 159

                                        pp. URL http://www.math.utah.
                                        edu/pub/bibnet/authors/b/.

Beebe:2019:NSUb

[1479] Nelson H. F. Beebe. Notes for the Spring 2019 University of Utah styles for Mathematics Department dissertations and theses in LaTEX. Report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, March 01, 2019. 34 pp. URL http://www.

math.utah.edu/pub/uumath-thesis/
uumath-thesis-2019-03-01-notes.

pdf. These notes accompany an extensive set of LaTEX style files and sample Masters theses and doctoral dissertations that are intended to replace the vintage styles from the early 1990s.

Beebe:2019:CBJb

[1480] Nelson H. F. Beebe. A complete bibliography of the Journal of Time Series Analysis. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, March 05, 2019. 274 pp. URL http:

//www.math.utah.edu/pub/tex/bib/
index-table-j.html#jtimeseranal.

Beebe:2019:CBJc

[1481] Nelson H. F. Beebe. A complete bibliography of the Journal of Econometrics (1970–1979). Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, March 06, 2019. 75 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-j.html#jeconometrics1970.

Beebe:2019:CBJd

[1482] Nelson H. F. Beebe. A complete bibliography of the Journal of Econometrics (1980–1989). Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, March 06, 2019. 172 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-j.html#jeconometrics1980.

Beebe:2019:CBT

[1483] Nelson H. F. Beebe. A complete bibliography of TEST: An Official Journal of the Spanish Society of Statistics and Operations Research. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, March 06, 2019. 183 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-t.html#test.

Beebe:2019:CBJe

[1484] Nelson H. F. Beebe. A complete bibliography of the Journal of Econometrics (1990–1999). Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, March 06, 2019. 213 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-j.html#jeconometrics1990.

Beebe:2019:CBJf

[1485] Nelson H. F. Beebe. A complete bibliography of the Journal of Econometrics (2000–2009). Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, March 06, 2019. 266 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-j.html#jeconometrics2000.

Beebe:2019:CBJg

[1486] Nelson H. F. Beebe. A complete bibliography of the Journal of Econometrics (2010–2019). Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, March 06, 2019. 317 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-j.html#jeconometrics2010.

Beebe:2019:CBJh

[1487] Nelson H. F. Beebe. A complete bibliography of the Journal of Time Series Econometrics. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, March 08, 2019. 26 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-j.html#jtimesereconom.

Beebe:2019:CBEa

[1488] Nelson H. F. Beebe. A complete bibliography of the The Econometrics Journal. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, March 09, 2019. 23 pp. URL http:

//www.math.utah.edu/pub/tex/bib/
index-table-e.html#economstat.

Beebe:2019:CBJi

[1489] Nelson H. F. Beebe. A complete bibliography of the Journal of Applied Econometrics. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, March 09, 2019. 26 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-j.html#jappleconometrics.

Beebe:2019:CBEb

[1490] Nelson H. F. Beebe. A complete bibliography of Econometrics and Statistics. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, March 09, 2019. 88 pp. URL http:

                                        //www.math.utah.edu/pub/tex/bib/
                                        index-table-e.html#economj.

Beebe:2019:BAA

[1491] Nelson H. F. Beebe. A bibliography of Advances and Applications in Statistical Sciences. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, March 20, 2019. 29 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-a.html#advapplstatsci.

Beebe:2019:NSUc

[1492] Nelson H. F. Beebe. Notes for the Spring 2019 University of Utah styles for Mathematics Department dissertations and theses in LaTEX. Report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, April 15, 2019. 34 pp. URL http:

                                        //www.math.utah.edu/pub/uumath-
                                        thesis/uumath-thesis-2019-04-15-

notes.pdf. These notes accompany an extensive set of LaTEX style files and sample Masters theses and doctoral dissertations that are intended to replace the vintage styles from the early 1990s.

Beebe:2019:SBPb

[1493] Nelson H. F. Beebe. A selected bibliography of publications by, and about, Samuel A. Goudsmit. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, April 16, 2019. 65

                                        pp. URL http://www.math.utah.
                                        edu/pub/bibnet/authors/g/.

Beebe:2019:SBPc

[1494] Nelson H. F. Beebe. A selected bibliography of publications by, and about, George E. Uhlenbeck. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, April 16, 2019. 96

pp. URL http://www.math.utah.
edu/pub/bibnet/authors/u/.

Beebe:2019:CBPc

[1495] Nelson H. F. Beebe. A complete bibliography of publications in the Journal of the Australian Mathematical Society: Series B, Applied Mathematics and the The ANZIAM Journal. The Australian & New Zealand Industrial and Applied Mathematics Journal. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, April 27, 2019. 379 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-a.html#anziamj.

Beebe:2019:CBPd

[1496] Nelson H. F. Beebe. A complete bibliography of publications in Mathematical Medicine and Biology. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, June 03, 2019. 162 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-m.html#mathmedbiol.

Beebe:2019:CBPe

[1497] Nelson H. F. Beebe. A complete bibliography of publications in the Journal of Cell Biology: 1955–1959. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, June 07, 2019. 126 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-j.html#jcellbiol1955.

Beebe:2019:CBPf

[1498] Nelson H. F. Beebe. A complete bibliography of publications in the Journal of Cell Biology: 1960–1964. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, June 07, 2019. 189 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-j.html#jcellbiol1960.

Beebe:2019:CBPg

[1499] Nelson H. F. Beebe. A complete bibliography of publications in the Journal of Cell Biology: 1965–1969. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, June 07, 2019. 246 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-j.html#jcellbiol1965.

Beebe:2019:CBPh

[1500] Nelson H. F. Beebe. A complete bibliography of publications in the Journal of Cell Biology: 1975–1979. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, June 07, 2019. 283 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-j.html#jcellbiol1975.

Beebe:2019:CBPi

[1501] Nelson H. F. Beebe. A complete bibliography of publications in the Journal of Cell Biology: 1970–1974. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, June 07, 2019. 299 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-j.html#jcellbiol1970.

Beebe:2019:CBPj

[1502] Nelson H. F. Beebe. A complete bibliography of publications in the Journal of Cell Biology: 2015–2019. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, June 07, 2019. 404 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-j.html#jcellbiol2015.

Beebe:2019:CBPk

[1503] Nelson H. F. Beebe. A complete bibliography of publications in the Journal of Cell Biology: 1980–1984. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, June 07, 2019. 484 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-j.html#jcellbiol1980.

Beebe:2019:CBPl

[1504] Nelson H. F. Beebe. A complete bibliography of publications in the Journal of Cell Biology: 2010–2014. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, June 07, 2019. 513 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-j.html#jcellbiol2010.

Beebe:2019:CBPm

[1505] Nelson H. F. Beebe. A complete bibliography of publications in the Journal of Cell Biology: 1985–1989. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, June 07, 2019. 569 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-j.html#jcellbiol1985.

Beebe:2019:CBPn

[1506] Nelson H. F. Beebe. A complete bibliography of publications in the Journal of Cell Biology: 1990–1994. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, June 07, 2019. 632 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-j.html#jcellbiol1990.

Beebe:2019:CBPo

[1507] Nelson H. F. Beebe. A complete bibliography of publications in the Journal of Cell Biology: 1995–1999. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, June 07, 2019. 645 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-j.html#jcellbiol1995.

Beebe:2019:CBPp

[1508] Nelson H. F. Beebe. A complete bibliography of publications in the Journal of Cell Biology: 2005–2009. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, June 07, 2019. 688 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-j.html#jcellbiol2005.

Beebe:2019:CBPq

[1509] Nelson H. F. Beebe. A complete bibliography of publications in the Journal of Cell Biology: 2000–2004. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, June 07, 2019. 712 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-j.html#jcellbiol2000.

Beebe:2019:CBPr

[1510] Nelson H. F. Beebe. A complete bibliography of publications in IEEE Computer Architecture Letters. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, June 25, 2019. 81 pp.

URL http://www.math.utah.edu/
pub/tex/bib/index-table-i.html#
ieeecomputarchitlett.

Beebe:2019:TLB

[1511] Nelson H. F. Beebe. TEX, LaTEX, and BibTEX resources. Web site., June 25,

2019. URL http://www.math.utah.

edu/pub/tex/doc/. This site supplies links to locally-written documents and journal articles about TeX, LaTeX, and BibTeX, as well as pointers to downloadable images and repository selection for the annual TeX Live releases, and archives of related free software.

Beebe:2019:CBPs

[1512] Nelson H. F. Beebe. A complete bibliography of publications in the Journal of Statistical Planning and Inference: 2020–2029. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, June 29, 2019. 6 pp.

URL http://www.math.utah.edu/
pub/tex/bib/index-table-j.html#
jstatplanninference2020.

Beebe:2019:CBPt

[1513] Nelson H. F. Beebe. A complete bibliography of publications in the Journal of Statistical Planning and Inference: 1970–1979. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, June 29, 2019. 22 pp.

                                        URL http://www.math.utah.edu/
                                        pub/tex/bib/index-table-j.html#
                                        jstatplanninference1970.

Beebe:2019:CBPu

[1514] Nelson H. F. Beebe. A complete bibliography of publications in the Journal of Statistical Planning and Inference: 1980–1989. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, June 29, 2019. 158 pp.

                                        URL http://www.math.utah.edu/
                                        pub/tex/bib/index-table-j.html#
                                        jstatplanninference1980.

Beebe:2019:CBPv

[1515] Nelson H. F. Beebe. A complete bibliography of publications in the Journal of Statistical Planning and Inference: 1990–1999. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, June 29, 2019. 375 pp.

                                        URL http://www.math.utah.edu/
                                        pub/tex/bib/index-table-j.html#
                                        jstatplanninference1990.

Beebe:2019:CBPw

[1516] Nelson H. F. Beebe. A complete bibliography of publications in the Journal of Statistical Planning and Inference: 2010–2019. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, June 29, 2019. 466 pp.

URL http://www.math.utah.edu/
pub/tex/bib/index-table-j.html#
jstatplanninference2010.

Beebe:2019:CBPx

[1517] Nelson H. F. Beebe. A complete bibliography of publications in the Journal of Statistical Planning and Inference: 2000–2009. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, June 29, 2019. 661 pp.

URL http://www.math.utah.edu/
pub/tex/bib/index-table-j.html#
jstatplanninference2000.

Beebe:2019:NSUd

[1518] Nelson H. F. Beebe. Notes for the Summer 2019 University of Utah styles for Mathematics Department dissertations and theses in LaTEX. Report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, July 01, 2019. 35 pp. URL http:

//www.math.utah.edu/pub/uumath-
thesis/uumath-thesis-2019-07-01-

notes.pdf. These notes accompany an extensive set of LaTEX style files and sample Bachelors and Masters theses and doctoral dissertations that are intended to replace the vintage styles from the early 1990s.

Beebe:2019:CBPy

[1519] Nelson H. F. Beebe. A complete bibliography of publications in Journal of the History of Biology. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, July 24, 2019. 213 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-j.html#jhistbiol.

Beebe:2019:CBPz

[1520] Nelson H. F. Beebe. A complete bibliography of publications in Journal of Computational Geometry. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, July 28, 2019. 27 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-j.html#jcomputgeom.

Beebe:2019:CBPaa

[1521] Nelson H. F. Beebe. A complete bibliography of publications in SIAM Journal on Mathematics of Data Science. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, August 23, 2019. 4 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-s.html#siamjmathdatasci.

Beebe:2019:CBPab

[1522] Nelson H. F. Beebe. A complete bibliography of publications in the Proceedings of the American Philosophical Society (1800– 1849). Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, August 27, 2019. 51 pp.

                                        URL http://www.math.utah.edu/
                                        pub/tex/bib/index-table-p.html#
                                        procamerphilsoc1800.

Beebe:2019:CBPac

[1523] Nelson H. F. Beebe. A complete bibliography of publications in the Proceedings of the American Philosophical Society (2000– 2049). Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, August 27, 2019. 186 pp.

URL http://www.math.utah.edu/
pub/tex/bib/index-table-p.html#
procamerphilsoc2000.

Beebe:2019:CBPad

[1524] Nelson H. F. Beebe. A complete bibliography of publications in the Proceedings of the American Philosophical Society (1850– 1899). Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, August 27, 2019. 368 pp.

URL http://www.math.utah.edu/
pub/tex/bib/index-table-p.html#
procamerphilsoc1850.

Beebe:2019:CBPae

[1525] Nelson H. F. Beebe. A complete bibliography of publications in the Proceedings of the American Philosophical Society (1900– 1949). Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, August 27, 2019. 380 pp.

URL http://www.math.utah.edu/
pub/tex/bib/index-table-p.html#
procamerphilsoc1900.

Beebe:2019:CBPaf

[1526] Nelson H. F. Beebe. A complete bibliography of publications in the Proceedings of the American Philosophical Society (1950– 1999). Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, August 27, 2019. 508 pp.

                                        URL http://www.math.utah.edu/
                                        pub/tex/bib/index-table-p.html#
                                        procamerphilsoc1950.

Beebe:2019:BPAb

[1527] Nelson H. F. Beebe. A bibliography of publications about the Julia programming language. Report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, September 7, 2019. 7 pp. URL http:

                                        //www.math.utah.edu/pub/tex/bib/
                                        index-table-j.html#julia. This re-

port is updated frequently.

Beebe:2019:CBPag

[1528] Nelson H. F. Beebe. A complete bibliography of publications in Journal of Computational and Applied Mathematics (2020–2024). Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, September 7, 2019. 3 pp.

                                        URL http://www.math.utah.edu/
                                        pub/tex/bib/index-table-j.html#
                                        jcomputapplmath2020. This report is

updated frequently.

Beebe:2019:HCA

[1529] Nelson H. F. Beebe. Highlights of computer arithmetic. Report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, September 14, 2019. 53 pp. This document supplements a book chapter on computer arithmetic, and is updated occasionally; the edition of 4 November 2019 has 69 pages. It is part of a set of handouts for workshops on computer arithmetic that I have given several times, and is supplemented by applications discussed in [1570].

Beebe:2019:CBPah

[1530] Nelson H. F. Beebe. A complete bibliography of publications in Journal of Computational and Applied Mathematics (2020–2024). Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, September 17, 2019. 3 pp.

URL http://www.math.utah.edu/
pub/tex/bib/index-table-j.html#
jcomputapplmath2020. This report is

updated frequently.

Beebe:2019:CBCa

[1531] Nelson H. F. Beebe. A complete bibliography of Computers and Security, 1980–1989. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, September 23, 2019. 186 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-c.html#computsecur1980.

Beebe:2019:CBCb

[1532] Nelson H. F. Beebe. A complete bibliography of Computers and Security, 2000–2009. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, September 23, 2019. 235 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-c.html#computsecur2000.

Beebe:2019:CBCc

[1533] Nelson H. F. Beebe. A complete bibliography of Computers and Security, 2010–2019. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, September 23, 2019. 315 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-c.html#computsecur2010.

Beebe:2019:CBCd

[1534] Nelson H. F. Beebe. A complete bibliography of Computers and Security, 1990–1999. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, September 23, 2019. 417 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-c.html#computsecur1990.

Beebe:2019:CBCe

[1535] Nelson H. F. Beebe. A complete bibliography of Computers and Security, 2020–2029. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, October 01, 2019. 8 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-c.html#computsecur2020.

Beebe:2019:CBPai

[1536] Nelson H. F. Beebe. A complete bibliography of publications in Computational Statistics & Data Analysis (2020–2029). Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, October 22, 2019. 3 pp.

                                        URL http://www.math.utah.edu/
                                        pub/tex/bib/index-table-c.html#
                                        computstatdataanal2020.

Beebe:2019:AGG

[1537] Nelson H. F. Beebe. Appendix G: Go interface. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, October 2, 2019. G-1–G-16 [946/947] pp. This is an online supplement to [1400].

Beebe:2019:CBPaj

[1538] Nelson H. F. Beebe. A complete bibliography of publications in Journal of Computational Physics: 2020–2024. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, November 2, 2019. 11 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-j.html#jcomputphys2020.

Beebe:2019:CBPak

[1539] Nelson H. F. Beebe. A complete bibliography of publications in Computer Physics Communications: 2020–2029. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, November 02, 2019. 25 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-a.html#compphyscomm2020.

Beebe:2019:CBPal

[1540] Nelson H. F. Beebe. A complete bibliography of publications in the Journal of Statistical Computation and Simulation: 2020–2029. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, November 16, 2019. 7 pp.

URL http://www.math.utah.edu/
pub/tex/bib/index-table-j.html#
jstatcomputsimul2020. This report

is updated frequently.

Beebe:2019:BPT

[1541] Nelson H. F. Beebe. A bibliography of publications in Theoretical Computer Science: 2020–2024. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, November 19, 2019. 28 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-t.html#tcs2020. This report is

updated frequently.

Beebe:2019:CBPam

[1542] Nelson H. F. Beebe. A complete bibliography of publications in The International Journal of Quantum Chemistry: 2020–2029. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, November 30, 2019. 6 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-i.html#ijqc2020.

Beebe:2019:CBA

[1543] Nelson H. F. Beebe. A complete bibliography of ACM Transactions on Social Computing (TSC). Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, December 6, 2019. 8 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-t.html#tsc.

Beebe:2019:ARR

[1544] Nelson H. F. Beebe. Appendix R: Rust interface. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, December 13, 2019. R-1–G-11 [994/ 995] pp. This is an online supplement to [1400].

Beebe:2019:CBPan

[1545] Nelson H. F. Beebe. A complete bibliography of publications in Communications in Statistics: Theory and Methods: 2020–2029. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, December 13, 2019. 12 pp.

URL http://www.math.utah.edu/
pub/tex/bib/index-table-c.html#
communstattheorymeth2020.

Beebe:2019:CBJj

[1546] Nelson H. F. Beebe. A complete bibliography of the Journal of the Royal Statistical Society, Series A family: 2020–2029. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, December 16, 2019. 5 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-j.html#jrss-a-2020.

Beebe:2019:TMV

[1547] Nelson H. F. Beebe. Typesetting multilingual verbatim text with pdflatex. Report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, December 18,

2019. 6 pp. URL http://www.math.
utah.edu/pub/tex/doc/mlverb.pdf.

Beebe:2020:CBPa

[1548] Nelson H. F. Beebe. A complete bibliography of publications in the Communications of the ACM : 2020–2029. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, January 02, 2020. 6 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-c.html#cacm2020. This report

is updated frequently.

Beebe:2020:TMT

[1549] Nelson H. F. Beebe. Typesetting multilingual text with lualatex and xelatex. Report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, January 08, 2020. 29 pp. URL http:

                                        //www.math.utah.edu/pub/tex/doc/
                                        mlverb.pdf.

Beebe:2020:CBPb

[1550] Nelson H. F. Beebe. A complete bibliography of publications in IEEE Transactions on Software Engineering (2020–2029). Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, January 9, 2020. 3 pp.

                                        URL http://www.math.utah.edu/
                                        pub/tex/bib/index-table-j.html#
                                        ieeetranssoftweng2020.

Beebe:2020:CBIa

[1551] Nelson H. F. Beebe. A complete bibliography of IEEE Transactions on Computers (2020–2029). Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, January 9, 2020. 4 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-i.html#ieeetranscomput2020.

This report is updated frequently.

Beebe:2020:BPAa

[1552] Nelson H. F. Beebe. A bibliography of publications by, and about, Walter Gautschi. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, January 10, 2020. 49

                                        pp. URL http://www.math.utah.
                                        edu/pub/bibnet/authors/g/.

Beebe:2020:CBPc

[1553] Nelson H. F. Beebe. A complete bibliography of publications in Applied Mathematics and Computation: 2020–2024. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, January 10, 2020. 56 pp.

URL http://www.math.utah.edu/
pub/tex/bib/index-table-a.html#
applmathcomput2020. This report is

updated frequently.

Beebe:2020:CBPd

[1554] Nelson H. F. Beebe. A complete bibliography of publications in Numerische Mathematik (2020–2029). Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, January 14, 2020. 3 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-n.html#nummath2020.

Beebe:2020:CBPe

[1555] Nelson H. F. Beebe. A complete bibliography of publications in Scientometrics: 2020–2029. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, January 14, 2020. 11 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-s.html#scientometrics2020.

Beebe:2020:CBPf

[1556] Nelson H. F. Beebe. A complete bibliography of publications in IEEE Spectrum (1960–1969). Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, January 14, 2020. 316 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-i.html#ieeespectrum1960.

Beebe:2020:BPC

[1557] Nelson H. F. Beebe. A bibliography of publications on cryptography: 2020–2029. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, January 15, 2020. 2 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-c.html#cryptography2020.

Beebe:2020:CBPg

[1558] Nelson H. F. Beebe. A complete bibliography of publications in IEEE Spectrum (1990–1999). Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, January 15, 2020. 195 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-i.html#ieeespectrum1990.

Beebe:2020:CBPh

[1559] Nelson H. F. Beebe. A complete bibliography of publications in IEEE Spectrum (2020–2029). Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, January 17, 2020. 5 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-i.html#ieeespectrum2020.

Beebe:2020:CBPi

[1560] Nelson H. F. Beebe. A complete bibliography of publications in IEEE Spectrum (2010–2019). Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, January 18, 2020. 256 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-i.html#ieeespectrum2010.

Beebe:2020:CBPj

[1561] Nelson H. F. Beebe. A complete bibliography of publications in IEEE Spectrum (2000–2009). Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, January 18, 2020. 285 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-i.html#ieeespectrum2000.

Beebe:2020:CBPk

[1562] Nelson H. F. Beebe. A complete bibliography of publications in IEEE Spectrum (1980–1989). Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, January 20, 2020. 377 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-i.html#ieeespectrum1980.

Beebe:2020:CBPl

[1563] Nelson H. F. Beebe. A complete bibliography of publications in IEEE Spectrum (1970–1979). Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, January 20, 2020. 480 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-i.html#ieeespectrum1970.

Beebe:2020:CBPm

[1564] Nelson H. F. Beebe. A complete bibliography of publications in SIGCSE Bulletin: 2020–2029. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, January 22, 2020. 3 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-s.html#sigcse2020. This re-

port is updated frequently.

Beebe:2020:BPI

[1565] Nelson H. F. Beebe. A bibliography of publications in IEEE Computer, 2020– 2029. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, January 22, 2020. 4 pp. URL http:

                                        //www.math.utah.edu/pub/tex/bib/
                                        index-table-c.html#computer2020.

Beebe:2020:CBPn

[1566] Nelson H. F. Beebe. A complete bibliography of publications in ACM SIGAPP Applied Computing Review. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, January 23, 2020. 60 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-s.html#sigapp.

Beebe:2020:CBPo

[1567] Nelson H. F. Beebe. A complete bibliography of publications in The Computer Journal: 2020–2029. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, January 28, 2020. 4 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-c.html#compj2020.

Beebe:2020:CBJa

[1568] Nelson H. F. Beebe. A complete bibliography of the Journal of Econometrics (2020–2029). Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, January 28, 2020. 8 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-j.html#jeconometrics2020.

Beebe:2020:BPAb

[1569] Nelson H. F. Beebe. A bibliography of publications about the Java Programming Language: 2020–2029. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, January 29, 2020. 1 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-j.html#java2020. This report

is updated frequently.

Beebe:2020:CAP

[1570] Nelson H. F. Beebe. Computer arithmetic — perils, pitfalls, and practices. Report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, January 31, 2020. 35 pp. Production of the book in which this chapter was to appear has been delayed, and the book is not yet in print. The original version appears in [1157]. This chapter is included in a set of handouts for a workshop on floatingpoint arithmetic, and is supplemented by applications in [1529]. The Texinfo markup system chosen by the book authors is severely limited in capabilities so that both online text and published PDF files can be produced; thus, the document design is inflexible, and use of color and alternate fonts is impossible.

Beebe:2020:CBPp

[1571] Nelson H. F. Beebe. A complete bibliography of publications in the Journal of Cell Biology: 2020–2024. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, February 11, 2020. 10 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-j.html#jcellbiol2020.

Beebe:2020:CBPq

[1572] Nelson H. F. Beebe. A complete bibliography of publications in Journal of Computational Chemistry: 2020–2029. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, February 11, 2020. 23 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-j.html#jcomputchem2020.

Beebe:2020:BPSa

[1573] Nelson H. F. Beebe. A bibliography of publications in Science and Public Policy: 2020–2029. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, February 29, 2020. 4 pp.

                                        URL http://www.math.utah.edu/
                                        pub/tex/bib/index-table-s.html#
                                        scipublicpol2020. This report is up-

dated frequently.

Beebe:2020:CBAa

[1574] Nelson H. F. Beebe. A complete bibliography of The American Statistician: 2020–2029. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, February 29, 2020. 5 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-a.html#amstat2020.

Beebe:2020:CBPr

[1575] Nelson H. F. Beebe. A complete bibliography of publications in Biometrika for the decade 2020–2029. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, February 29, 2020. 6 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-b.html#biometrika2020. This

report is updated frequently.

Beebe:2020:CBJb

[1576] Nelson H. F. Beebe. A complete bibliography of the Journal of Statistical Physics: 2020–2029. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, March 02, 2020. 10 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-j.html#jstatphys2020. This

report is updated frequently.

Beebe:2020:CBPs

[1577] Nelson H. F. Beebe. A complete bibliography of publications in the Journal of Mathematical Physics: 2020–2024. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, March 3, 2020. 15 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-j.html#jmathphys2020.

Beebe:2020:CBPt

[1578] Nelson H. F. Beebe. A complete bibliography of publications in Technometrics for the decade 2020–2029. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, March 4, 2020. 7 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-t.html#technometrics2020.

This report is updated frequently.

Beebe:2020:CBD

[1579] Nelson H. F. Beebe. A complete bibliography of Digital Government: Research and Practice. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, March 6, 2020. 3 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-d.html#dgov. This report is up-

dated frequently.

Beebe:2020:CBPah

[1580] Nelson H. F. Beebe. A complete bibliography of publications in Communications in Statistics: Simulation and Computation: 2020–2029. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, March 9, 2020. 6 pp.

                                        URL http://www.math.utah.edu/
                                        pub/tex/bib/index-table-c.html#
                                        communstatsimulcomput2020.

Beebe:2020:CBAb

[1581] Nelson H. F. Beebe. A complete bibliography of ACM Transactions on Internet of Things (TIOT). Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, April 7, 2020. 3 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-t.html#tiot.

Beebe:2020:CBJc

[1582] Nelson H. F. Beebe. A complete bibliography of Journal of Human–Robot Interaction (JHRI) and ACM Transactions on Human-Robot Interaction (THRI). Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, April 7, 2020. 33 pp. URL http:

//www.math.utah.edu/pub/tex/bib/
index-table-t.html#thri.

Beebe:2020:CBAc

[1583] Nelson H. F. Beebe. A complete bibliography of ACM Transactions on Computing for Healthcare (HEALTH). Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, April 3, 2020. 33 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-t.html#toch.

Beebe:2020:CBAd

[1584] Nelson H. F. Beebe. A complete bibliography of ACM/IMS Transactions on Data Science (TDS). Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, April 3, 2020. 33 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-t.html#tds.

Beebe:2020:NSUa

[1585] Nelson H. F. Beebe. Notes for the Summer 2020 University of Utah styles for Mathematics Department dissertations and theses in LaTEX. Report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, June 01, 2020. 35 pp. URL http:

//www.math.utah.edu/pub/uumath-
thesis/uumath-thesis-2020-06-01-

notes.pdf. These notes accompany an extensive set of LaTEX style files and sample Bachelors and Masters theses and doctoral dissertations that are intended to replace the vintage styles from the early 1990s.

Beebe:2020:NSUb

[1586] Nelson H. F. Beebe. Notes for the Summer 2020 University of Utah styles for Mathematics Department dissertations and theses in LaTEX. Report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, June 15, 2020. 36 pp. URL http:

                                        //www.math.utah.edu/pub/uumath-
                                        thesis/uumath-thesis-2020-06-15-

notes.pdf. These notes accompany an extensive set of LaTEX style files and sample Bachelors and Masters theses and doctoral dissertations that are intended to replace the vintage styles from the early 1990s.

Beebe:2020:CBJd

[1587] Nelson H. F. Beebe. A complete bibliography of Journal of Business & Economic Statistics. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, June 16, 2020. 386 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-j.html#jbuseconstat.

Beebe:2020:CBC

[1588] Nelson H. F. Beebe. A complete bibliography of Computational Statistics. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, June 18, 2020. 191 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-c.html#compstat.

Beebe:2020:CBW

[1589] Nelson H. F. Beebe. A complete bibliography of WIREs Computational Statistics. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, June 19, 2020. 73 pp. URL http:

//www.math.utah.edu/pub/tex/bib/
index-table-w.html#wires.

Beebe:2020:CBSa

[1590] Nelson H. F. Beebe. A complete bibliography Statistics in Medicine (1980– 1989). Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, June 19, 2020. 133 pp. URL http:

//www.math.utah.edu/pub/tex/bib/
index-table-s.html#statmed1980.

Beebe:2020:CBSb

[1591] Nelson H. F. Beebe. A complete bibliography Statistics in Medicine (1990– 1999). Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, June 19, 2020. 450 pp. URL http:

//www.math.utah.edu/pub/tex/bib/
index-table-s.html#statmed1990.

Beebe:2020:CBSc

[1592] Nelson H. F. Beebe. A complete bibliography Statistics in Medicine (2000– 2009). Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, June 19, 2020. 601 pp. URL http:

//www.math.utah.edu/pub/tex/bib/
index-table-s.html#statmed2000.

Beebe:2020:CBSd

[1593] Nelson H. F. Beebe. A complete bibliography Statistics in Medicine (2010– 2019). Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, June 19, 2020. 644 pp. URL http:

//www.math.utah.edu/pub/tex/bib/
index-table-s.html#statmed2010.

Beebe:2020:CBSe

[1594] Nelson H. F. Beebe. A complete bibliography Statistics in Medicine (2020– 2029). Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, June 19, 2020. 30 pp. URL http:

//www.math.utah.edu/pub/tex/bib/
index-table-s.html#statmed2020.

Beebe:2020:BPSb

[1595] Nelson H. F. Beebe. A bibliography of publications of Sir Ronald Aylmer Fisher. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, June 29, 2020. 119

pp. URL http://www.math.utah.
edu/pub/bibnet/authors/f/.

Beebe:2020:BMa

[1596] Nelson H. F. Beebe. A bibliography of Metron (1920–1929). Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, July 3, 2020. 2 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-m.html#metron1920.

Beebe:2020:BMb

[1597] Nelson H. F. Beebe. A bibliography of Metron (1930–1939). Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, July 3, 2020. 12 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-m.html#metron1930.

Beebe:2020:BMc

[1598] Nelson H. F. Beebe. A bibliography of Metron (1940–1949). Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, July 3, 2020. 5 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-m.html#metron1940.

Beebe:2020:BMd

[1599] Nelson H. F. Beebe. A bibliography of Metron (1950–1959). Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, July 3, 2020. 10 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-m.html#metron1950.

Beebe:2020:BMe

[1600] Nelson H. F. Beebe. A bibliography of Metron (1960–1969). Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, July 3, 2020. 13 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-m.html#metron1960.

Beebe:2020:BMf

[1601] Nelson H. F. Beebe. A bibliography of Metron (1970–1979). Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, July 3, 2020. 38 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-m.html#metron1970.

Beebe:2020:BMg

[1602] Nelson H. F. Beebe. A bibliography of Metron (1980–1989). Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, July 3, 2020. 40 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-m.html#metron1980.

Beebe:2020:BMh

[1603] Nelson H. F. Beebe. A bibliography of Metron (1990–1999). Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, July 3, 2020. 42 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-m.html#metron1990.

Beebe:2020:BMi

[1604] Nelson H. F. Beebe. A bibliography of Metron (2000–2009). Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, July 3, 2020. 44 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-m.html#metron2000.

Beebe:2020:CBMa

[1605] Nelson H. F. Beebe. A complete bibliography of Metron (2010–2019). Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, July 3, 2020. 44 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-m.html#metron2010.

Beebe:2020:CBMb

[1606] Nelson H. F. Beebe. A complete bibliography of Metron (2020–2029). Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, July 3, 2020. 3 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-m.html#metron2020.

Beebe:2020:CBJe

[1607] Nelson H. F. Beebe. A complete bibliography of the Journal of Number Theory (1960–1969). Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, July 16, 2020. 11 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-j.html#jnumbertheory1960.

Beebe:2020:CBJf

[1608] Nelson H. F. Beebe. A complete bibliography of the Journal of Number Theory (1970–1979). Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, July 16, 2020. 89 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-j.html#jnumbertheory1970.

Beebe:2020:CBJg

[1609] Nelson H. F. Beebe. A complete bibliography of the Journal of Number Theory (1980–1989). Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, July 16, 2020. 132 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-j.html#jnumbertheory1980.

Beebe:2020:CBJh

[1610] Nelson H. F. Beebe. A complete bibliography of the Journal of Number Theory (1990–1999). Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, July 16, 2020. 210 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-j.html#jnumbertheory1990.

Beebe:2020:CBJi

[1611] Nelson H. F. Beebe. A complete bibliography of the Journal of Number Theory (2000–2009). Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, July 16, 2020. 291 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-j.html#jnumbertheory2000.

Beebe:2020:CBJj

[1612] Nelson H. F. Beebe. A complete bibliography of the Journal of Number Theory (2010–2019). Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, July 16, 2020. 460 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-j.html#jnumbertheory2010.

Beebe:2020:CBJk

[1613] Nelson H. F. Beebe. A complete bibliography of the Journal of Number Theory (2020–2029). Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, July 16, 2020. 38 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-j.html#jnumbertheory2020.

Beebe:2020:CBIb

[1614] Nelson H. F. Beebe. A complete bibliography of the International Journal of Number Theory. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, July 21, 2020. 277 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-i.html#ijnt.

Beebe:2020:CBPai

[1615] Nelson H. F. Beebe. A complete bibliography of publications in Significance (Oxford, UK). Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, August 13, 2020. 190 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-s.html#significance.

Beebe:2020:CBJl

[1616] Nelson H. F. Beebe. A complete bibliography of the Journal of Computational and Graphical Statistics. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, August 14, 2020. 300 pp.

URL http://www.math.utah.edu/
pub/tex/bib/index-table-j.html#
jcomputgraphstat. Major update of

[471], completing coverage of all articles from 1992 to date, including DOI values. This report is updated frequently.

Beebe:2020:BPF

[1617] Nelson H. F. Beebe. A bibliography of publications of Friedrich Ludwig Bauer. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, August 20, 2020. 86

                                        pp. URL http://www.math.utah.
                                        edu/pub/bibnet/authors/b/bauer-
                                        friedrich-ludwig.bib.

Beebe:2020:BPH

[1618] Nelson H. F. Beebe. A bibliography of publications of Heinz Rutishauser. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, August 24, 2020. 48 pp. URL http:

                                        //www.math.utah.edu/pub/bibnet/
                                        authors/r/rutihauser-heinz.bib.

Beebe:2020:CBSf

[1619] Nelson H. F. Beebe. A complete bibliography of Scandinavian Journal of Statistics. Theory and Applications. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, August 31, 2020. 377 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-s.html#scandjstat.

Beebe:2020:BPE

[1620] Nelson H. F. Beebe. A bibliography of publications of Eduard Stiefel. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, September 05,

                                        2020. 60 pp. URL http://www.math.
                                        utah.edu/pub/bibnet/authors/s/.

Beebe:2020:CBPaj

[1621] Nelson H. F. Beebe. A complete bibliography of publications in the American Review of Mathematics and Statistics. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, September 30, 2020. 16 pp. URL http:

//www.math.utah.edu/pub/tex/bib/
index-table-a.html#arm.

Beebe:2020:DFP

[1622] Nelson H. F. Beebe. Design of a fixedpoint arithmetic library. Report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, September 2020. vii + 106 pp. This document is still under development. It describes the design of, and algorithms in, a large portable mathematical software library (already more than 38 000 lines of code) that the author developed in Fall 2020/Winter 2021, and that will eventually be released under an open-source license for free public access. A formal journal article in ACM Transactions on Mathematical Software is contemplated, although it can only be a supplement to this long document.

Beebe:2021:CBPa

[1623] Nelson H. F. Beebe. A complete bibliography of publications in IEEE Transactions on Visualization and Computer Graphics (2020–2029). Report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, February 5, 2021. 91 pp.

URL http://www.math.utah.edu/
pub/tex/bib/index-table-i.html#
ieeetransviscomputgraph2020. This

report is updated frequently.

Beebe:2021:CBPb

[1624] Nelson H. F. Beebe. A complete bibliography of publications in The Journal of Typographic Research and an extensive bibliography of publications in Visible Language. Report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, February 9, 2021. 186 pp.

                                        URL http://www.math.utah.edu/
                                        pub/tex/bib/index-table-v.html#
                                        visible-language. This report is up-

dated frequently.

Beebe:2021:CBAa

[1625] Nelson H. F. Beebe. A complete bibliography of ACM Transactions on Quantum Computing (TQC). Report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, March 10, 2021. 3 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-t.html#tqc. This report is up-

dated frequently.

Beebe:2021:CBB

[1626] Nelson H. F. Beebe. A complete bibliography of Bell Labs Technical Journal, 2020–2029. Report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, March 10, 2021. 1 pp. URL http:

                                        //www.math.utah.edu/pub/tex/bib/
                                        index-table-b.html#bstj2020. This

report is updated frequently.

Beebe:2021:BPA

[1627] Nelson H. F. Beebe. A bibliography of publications about the Rust programming language. Report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, March 23, 2021. 5 pp. URL http:

                                        //www.math.utah.edu/pub/tex/bib/
                                        index-table-r.html#rust. This re-

port is updated frequently.

Beebe:2021:CBPc

[1628] Nelson H. F. Beebe. A complete bibliography of Proceedings of the ACM on Computer Graphics and Interactive Techniques (PACMCGIT). Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, March 30, 2021. 16 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-p.html#pacmcgit.

Beebe:2021:CBPd

[1629] Nelson H. F. Beebe. A complete bibliography of Proceedings of the ACM on Human-Computer Interaction (PACMHCI). Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, March 31, 2021. 65 pp.

URL http://www.math.utah.edu/
pub/tex/bib/index-table-p.html#
pacmhci.

Beebe:2021:CBPe

[1630] Nelson H. F. Beebe. A complete bibliography of publications in Computing and Software for Big Science. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, April 9, 2021. 9 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-c.html#computsoftwbigsci.

Beebe:2021:BAS

[1631] Nelson H. F. Beebe. A bibliography of ACM SIGMICRO Newsletter. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, April 16, 2021. 65 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-s.html#sigmicro.

Beebe:2021:PAF

[1632] Nelson H. F. Beebe. Publications by, and about, Frank Stenger. In Baumann [1790], pages 385–399. ISBN 3-030-49715-1 (hardcover), 3-030-49716-X (e-book). ISSN 2297-0215. LCCN QA372.

Beebe:2021:CBPf

[1633] Nelson H. F. Beebe. A complete bibliography of publications in Learned Publishing. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, May 15, 2021. 310 pp.

                                        URL http://www.math.utah.edu/
                                        pub/tex/bib/index-table-l.html#
                                        learnpubl.

Beebe:2021:CBPg

[1634] Nelson H. F. Beebe. A complete bibliography of publications in the Journal of Systems and Software (2000–2009). Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, May 25, 2021. 257 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-j.html#jsystsoftw2000.

Beebe:2021:CBPh

[1635] Nelson H. F. Beebe. A complete bibliography of publications in the Journal of Systems and Software (2010–2019). Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, May 25, 2021. 487 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-j.html#jsystsoftw2010.

Beebe:2021:CBPi

[1636] Nelson H. F. Beebe. A complete bibliography of publications in the Journal of Systems and Software (2020–2029). Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, May 25, 2021. 71 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-j.html#jsystsoftw2020.

Beebe:2021:CBAb

[1637] Nelson H. F. Beebe. A complete bibliography of ACM Transactions on Evolutionary Learning and Optimization (TELO). Report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, August 21, 2021. 4 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-t.html#telo. This report is up-

dated frequently.

Beebe:2021:CBPj

[1638] Nelson H. F. Beebe. A complete bibliography of publications in Ecological Modelling (1970–1979). Report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, November 12, 2021. 47 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-e.html#ecolmodell1970. This

report is updated frequently.

Beebe:2021:CBPk

[1639] Nelson H. F. Beebe. A complete bibliography of publications in Ecological Modelling (1980–1989). Report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, November 12, 2021. 223 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-e.html#ecolmodell1980. This

report is updated frequently.

Beebe:2021:CBPl

[1640] Nelson H. F. Beebe. A complete bibliography of publications in Ecological Modelling (1990–1999). Report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, November 12, 2021. 394 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-e.html#ecolmodell1990. This

report is updated frequently.

Beebe:2021:CBPm

[1641] Nelson H. F. Beebe. A complete bibliography of publications in Ecological Modelling (2000–2009). Report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, November 12, 2021. 845 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-e.html#ecolmodell2000. This

report is updated frequently.

Beebe:2021:CBPn

[1642] Nelson H. F. Beebe. A complete bibliography of publications in Ecological Modelling (2010–2019). Report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, November 12, 2021. 894 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-e.html#ecolmodell2010. This

report is updated frequently.

Beebe:2021:CBPo

[1643] Nelson H. F. Beebe. A complete bibliography of publications in Ecological Modelling (2020–2029). Report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, November 12, 2021. 162 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-e.html#ecolmodell2020. This

report is updated frequently.

Beebe:2021:CBPp

[1644] Nelson H. F. Beebe. A complete bibliography of publications in Fisheries Research (1980–1989). Report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, November 12, 2021. 84 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-f.html#fishres1980. This re-

port is updated frequently.

Beebe:2021:CBPq

[1645] Nelson H. F. Beebe. A complete bibliography of publications in Fisheries Research (1990–1999). Report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, November 12, 2021. 270 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-f.html#fishres1990. This re-

port is updated frequently.

Beebe:2021:CBPr

[1646] Nelson H. F. Beebe. A complete bibliography of publications in Fisheries Research (2000–2009). Report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, November 12, 2021. 529 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-

                                        table-f.html#fishres2000. This re-

port is updated frequently.

Beebe:2021:CBPs

[1647] Nelson H. F. Beebe. A complete bibliography of publications in Fisheries Research (2010–2019). Report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, November 12, 2021. 697 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-f.html#fishres2010. This re-

port is updated frequently.

Beebe:2021:CBPt

[1648] Nelson H. F. Beebe. A complete bibliography of publications in Fisheries Research (2020–2029). Report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, November 12, 2021. 183 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-f.html#fishres2020. This re-

port is updated frequently.

Beebe:2021:CBPu

[1649] Nelson H. F. Beebe. A complete bibliography of publications in Marine Policy (1970–1979). Report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, November 12, 2021. 47 pp.

                                        URL http://www.math.utah.edu/
                                        pub/tex/bib/index-table-m.html#

marpolicy1970. This report is updated frequently.

Beebe:2021:CBPv

[1650] Nelson H. F. Beebe. A complete bibliography of publications in Marine Policy (1980–1989). Report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, November 12, 2021. 151 pp.

URL http://www.math.utah.edu/
pub/tex/bib/index-table-m.html#

marpolicy1980. This report is updated frequently.

Beebe:2021:CBPw

[1651] Nelson H. F. Beebe. A complete bibliography of publications in Marine Policy (1990–1999). Report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, November 12, 2021. 140 pp.

URL http://www.math.utah.edu/
pub/tex/bib/index-table-m.html#

marpolicy1990. This report is updated frequently.

Beebe:2021:CBPx

[1652] Nelson H. F. Beebe. A complete bibliography of publications in Marine Policy (2000–2009). Report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, November 12, 2021. 165 pp.

URL http://www.math.utah.edu/
pub/tex/bib/index-table-m.html#

marpolicy2000. This report is updated frequently.

Beebe:2021:CBPy

[1653] Nelson H. F. Beebe. A complete bibliography of publications in Marine Policy (2010–2019). Report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, November 12, 2021. 686 pp.

URL http://www.math.utah.edu/
pub/tex/bib/index-table-m.html#

marpolicy2010. This report is updated frequently.

Beebe:2021:CBPz

[1654] Nelson H. F. Beebe. A complete bibliography of publications in Marine Policy (2020–2029). Report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, November 12, 2021. 264 pp.

                                        URL http://www.math.utah.edu/
                                        pub/tex/bib/index-table-m.html#

marpolicy2020. This report is updated frequently.

Beebe:2021:CBPba

[1655] Nelson H. F. Beebe. A complete bibliography of publications in the Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques (1980–1989). Report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, November 15, 2021. 581 pp.

                                        URL http://www.math.utah.edu/
                                        pub/tex/bib/index-table-c.html#
                                        canjfishaquatsci1980. This report

is updated frequently.

Beebe:2021:CBPbb

[1656] Nelson H. F. Beebe. A complete bibliography of publications in the Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques (1990–1999). Report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, November 15, 2021. 660 pp.

                                        URL http://www.math.utah.edu/
                                        pub/tex/bib/index-table-c.html#
                                        canjfishaquatsci1990. This report

is updated frequently.

Beebe:2021:CBPbc

[1657] Nelson H. F. Beebe. A complete bibliography of publications in the Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques (2000–2009). Report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, November 15, 2021. 513 pp.

URL http://www.math.utah.edu/
pub/tex/bib/index-table-c.html#
canjfishaquatsci2000. This report

is updated frequently.

Beebe:2021:CBPbd

[1658] Nelson H. F. Beebe. A complete bibliography of publications in the Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques (2010–2019). Report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, November 15, 2021. 459 pp.

URL http://www.math.utah.edu/
pub/tex/bib/index-table-c.html#
canjfishaquatsci2010. This report

is updated frequently.

Beebe:2021:CBPbe

[1659] Nelson H. F. Beebe. A complete bibliography of publications in the Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques (2020–2029). Report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, November 15, 2021. 89 pp.

URL http://www.math.utah.edu/
pub/tex/bib/index-table-c.html#

                                        canjfishaquatsci2020. This report

is updated frequently.

Beebe:2021:CBPbf

[1660] Nelson H. F. Beebe. A complete bibliography of publications in Fish and Fisheries. Report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, November 15, 2021. 197 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-f.html#fishfish. This report

is updated frequently.

Beebe:2021:CBPbg

[1661] Nelson H. F. Beebe. A complete bibliography of publications in the Journal of Fish Biology (1960–1969). Report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, November 15, 2021. 8 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-j.html#jfishbiol1960. This

report is updated frequently.

Beebe:2021:CBPbh

[1662] Nelson H. F. Beebe. A complete bibliography of publications in the Journal of Fish Biology (1970–1979). Report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, November 15, 2021. 158 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-j.html#jfishbiol1970. This

report is updated frequently.

Beebe:2021:CBPbi

[1663] Nelson H. F. Beebe. A complete bibliography of publications in the Journal of Fish Biology (1980–1989).

Report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, November 15, 2021. 308 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-j.html#jfishbiol1980. This

report is updated frequently.

Beebe:2021:CBPbj

[1664] Nelson H. F. Beebe. A complete bibliography of publications in the Journal of Fish Biology (1990–1999). Report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, November 15, 2021. 430 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-j.html#jfishbiol1990. This

report is updated frequently.

Beebe:2021:CBPbk

[1665] Nelson H. F. Beebe. A complete bibliography of publications in the Journal of Fish Biology (2010–2019). Report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, November 15, 2021. 581 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-j.html#jfishbiol2010. This

report is updated frequently.

Beebe:2021:CBPbl

[1666] Nelson H. F. Beebe. A complete bibliography of publications in the Journal of Fish Biology (2000–2009). Report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, November 15, 2021. 621 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-j.html#jfishbiol2000. This

report is updated frequently.

Beebe:2021:CBPbm

[1667] Nelson H. F. Beebe. A complete bibliography of publications in the Journal of Fish Biology (2020–2029). Report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, November 15, 2021. 149 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-j.html#jfishbiol2020. This

report is updated frequently.

Beebe:2021:CBPbn

[1668] Nelson H. F. Beebe. A complete bibliography of publications in Natural Resource Modeling. Report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, November 15, 2021. 160 pp.

                                        URL http://www.math.utah.edu/
                                        pub/tex/bib/index-table-n.html#

natresourmodel. This report is updated frequently.

Beebe:2021:CBPbo

[1669] Nelson H. F. Beebe. A complete bibliography of publications in the ICES Journal of Marine Science (1920–1929). Report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, November 23, 2021. 48 pp.

                                        URL http://www.math.utah.edu/
                                        pub/tex/bib/index-table-i.html#

icesjmarsci1920. This report is updated frequently.

Beebe:2021:CBPbp

[1670] Nelson H. F. Beebe. A complete bibliography of publications in the ICES Journal of Marine Science (1930–1939). Report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, November 23, 2021. 178 pp.

URL http://www.math.utah.edu/
pub/tex/bib/index-table-i.html#

icesjmarsci1930. This report is updated frequently.

Beebe:2021:CBPbq

[1671] Nelson H. F. Beebe. A complete bibliography of publications in the ICES Journal of Marine Science (1940–1949). Report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, November 23, 2021. 16 pp.

URL http://www.math.utah.edu/
pub/tex/bib/index-table-i.html#

icesjmarsci1940. This report is updated frequently.

Beebe:2021:CBPbr

[1672] Nelson H. F. Beebe. A complete bibliography of publications in the ICES Journal of Marine Science (1950–1959). Report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, November 23, 2021. 103 pp.

URL http://www.math.utah.edu/
pub/tex/bib/index-table-i.html#

icesjmarsci1950. This report is updated frequently.

Beebe:2021:CBPbs

[1673] Nelson H. F. Beebe. A complete bibliography of publications in the ICES Journal of Marine Science (1960–1969). Report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, November 23, 2021. 65 pp.

URL http://www.math.utah.edu/

                                        pub/tex/bib/index-table-i.html#

icesjmarsci1960. This report is updated frequently.

Beebe:2021:CBPbt

[1674] Nelson H. F. Beebe. A complete bibliography of publications in the ICES Journal of Marine Science (1970–1979). Report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, November 23, 2021. 75 pp.

                                        URL http://www.math.utah.edu/
                                        pub/tex/bib/index-table-i.html#

icesjmarsci1970. This report is updated frequently.

Beebe:2021:CBPbu

[1675] Nelson H. F. Beebe. A complete bibliography of publications in the ICES Journal of Marine Science (1980–1989). Report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, November 23, 2021. 66 pp.

                                        URL http://www.math.utah.edu/
                                        pub/tex/bib/index-table-i.html#

icesjmarsci1980. This report is updated frequently.

Beebe:2021:CBPbv

[1676] Nelson H. F. Beebe. A complete bibliography of publications in the ICES Journal of Marine Science (1990–1999). Report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, November 23, 2021. 197 pp.

                                        URL http://www.math.utah.edu/
                                        pub/tex/bib/index-table-i.html#

icesjmarsci1990. This report is updated frequently.

Beebe:2021:CBPbw

[1677] Nelson H. F. Beebe. A complete bibliography of publications in the ICES Journal of Marine Science (2000–2009). Report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, November 23, 2021. 439 pp.

URL http://www.math.utah.edu/
pub/tex/bib/index-table-i.html#

icesjmarsci2000. This report is updated frequently.

Beebe:2021:CBPbx

[1678] Nelson H. F. Beebe. A complete bibliography of publications in the ICES Journal of Marine Science (2010–2019). Report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, November 23, 2021. 621 pp.

URL http://www.math.utah.edu/
pub/tex/bib/index-table-i.html#

icesjmarsci2010. This report is updated frequently.

Beebe:2021:CBPby

[1679] Nelson H. F. Beebe. A complete bibliography of publications in the ICES Journal of Marine Science (2020–2029). Report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, November 23, 2021. 144 pp.

URL http://www.math.utah.edu/
pub/tex/bib/index-table-i.html#

icesjmarsci2020. This report is updated frequently.

Beebe:2021:CBPbz

[1680] Nelson H. F. Beebe. A complete bibliography of publications in the Reviews in Fish Biology and Fisheries. Report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, November 23, 2021. 256 pp.

URL http://www.math.utah.edu/
pub/tex/bib/index-table-r.html#

revfishbiolfish. This report is updated frequently.

Beebe:2021:CBPca

[1681] Nelson H. F. Beebe. A complete bibliography of publications in the Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada (1930–1939). Report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, November 25, 2021. 12 pp.

URL http://www.math.utah.edu/
pub/tex/bib/index-table-j.html#
jfishresboardcan1930. This report

is updated frequently.

Beebe:2021:CBPcb

[1682] Nelson H. F. Beebe. A complete bibliography of publications in the Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada (1940–1949). Report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, November 25, 2021. 32 pp.

URL http://www.math.utah.edu/
pub/tex/bib/index-table-j.html#
jfishresboardcan1940. This report

is updated frequently.

Beebe:2021:CBPcc

[1683] Nelson H. F. Beebe. A complete bibliography of publications in the Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada (1950–1959). Report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, November 25, 2021. 89 pp.

URL http://www.math.utah.edu/
pub/tex/bib/index-table-j.html#
jfishresboardcan1950. This report

is updated frequently.

Beebe:2021:CBPcd

[1684] Nelson H. F. Beebe. A complete bibliography of publications in the Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada (1960–1969). Report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, November 25, 2021. 324 pp.

URL http://www.math.utah.edu/
pub/tex/bib/index-table-j.html#
jfishresboardcan1960. This report

is updated frequently.

Beebe:2021:CBPce

[1685] Nelson H. F. Beebe. A complete bibliography of publications in the Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada (1970–1979). Report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, November 25, 2021. 618 pp.

URL http://www.math.utah.edu/
pub/tex/bib/index-table-j.html#
jfishresboardcan1970. This report

is updated frequently.

Beebe:2021:CBPcf

[1686] Nelson H. F. Beebe. A complete bibliography of publications in Fisheries Oceanography. Report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, November 27, 2021. 288 pp.

                                        URL http://www.math.utah.edu/
                                        pub/tex/bib/index-table-f.html#

fishoceanogr. This report is updated frequently.

Beebe:2021:CBPcg

[1687] Nelson H. F. Beebe. A complete bibliography of publications in North American Journal of Fisheries Management: 1980–1989. Report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, November 27, 2021. 114 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-n.html#namjfishmanag1980.

This report is updated frequently.

Beebe:2021:CBPch

[1688] Nelson H. F. Beebe. A complete bibliography of publications in North American Journal of Fisheries Management: 1990–1999. Report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, November 27, 2021. 215 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-n.html#namjfishmanag1990.

This report is updated frequently.

Beebe:2021:CBPci

[1689] Nelson H. F. Beebe. A complete bibliography of publications in North American Journal of Fisheries Management: 2000–2009. Report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, November 27, 2021. 321 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-n.html#namjfishmanag2000.

This report is updated frequently.

Beebe:2021:CBPcj

[1690] Nelson H. F. Beebe. A complete bibliography of publications in North American Journal of Fisheries Management: 2010–2019. Report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, November 27, 2021. 300 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-n.html#namjfishmanag2010.

This report is updated frequently.

Beebe:2021:CBPck

[1691] Nelson H. F. Beebe. A complete bibliography of publications in North American Journal of Fisheries Management: 2020–2029. Report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, November 27, 2021. 77 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-n.html#namjfishmanag2020.

This report is updated frequently.

Beebe:2021:CBPcl

[1692] Nelson H. F. Beebe. A complete bibliography of publications in Transactions of the American Fisheries Society: 1870–1879. Report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, November 30, 2021. 14 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-t.html#transamfishsoc1870.

This report is updated frequently.

Beebe:2021:CBPcm

[1693] Nelson H. F. Beebe. A complete bibliography of publications in Transactions of the American Fisheries Society: 1880–1889. Report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, November 30, 2021. 26 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-t.html#transamfishsoc1880.

This report is updated frequently.

Beebe:2021:CBPcn

[1694] Nelson H. F. Beebe. A complete bibliography of publications in Transactions of the American Fisheries Society: 1890–1899. Report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, November 30, 2021. 23 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-t.html#transamfishsoc1890.

This report is updated frequently.

Beebe:2021:CBPco

[1695] Nelson H. F. Beebe. A complete bibliography of publications in Transactions of the American Fisheries Society: 1900–1909. Report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, November 30, 2021. 32 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-t.html#transamfishsoc1900.

This report is updated frequently.

Beebe:2021:CBPcp

[1696] Nelson H. F. Beebe. A complete bibliography of publications in Transactions of the American Fisheries Society: 1910–1919. Report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, November 30, 2021. 30 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-t.html#transamfishsoc1910.

This report is updated frequently.

Beebe:2021:CBPcq

[1697] Nelson H. F. Beebe. A complete bibliography of publications in Transactions of the American Fisheries Society: 1920–1929. Report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, November 30, 2021. 52 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-t.html#transamfishsoc1920.

This report is updated frequently.

Beebe:2021:CBPcr

[1698] Nelson H. F. Beebe. A complete bibliography of publications in Transactions of the American Fisheries Society: 1930–1939. Report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, November 30, 2021. 75 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-t.html#transamfishsoc1930.

This report is updated frequently.

Beebe:2021:CBPcs

[1699] Nelson H. F. Beebe. A complete bibliography of publications in Transactions of the American Fisheries Society: 1940–1949. Report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, November 30, 2021. 64 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-t.html#transamfishsoc1940.

This report is updated frequently.

Beebe:2021:CBPct

[1700] Nelson H. F. Beebe. A complete bibliography of publications in Transactions of the American Fisheries Society: 1950–1959. Report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, November 30, 2021. 81 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-t.html#transamfishsoc1950.

This report is updated frequently.

Beebe:2021:CBPcu

[1701] Nelson H. F. Beebe. A complete bibliography of publications in Transactions of the American Fisheries Society: 1960–1969. Report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, November 30, 2021. 155 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-t.html#transamfishsoc1960.

This report is updated frequently.

Beebe:2021:CBPcv

[1702] Nelson H. F. Beebe. A complete bibliography of publications in Transactions of the American Fisheries Society: 1970–1979. Report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, November 30, 2021. 231 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-t.html#transamfishsoc1970.

This report is updated frequently.

Beebe:2021:CBPcw

[1703] Nelson H. F. Beebe. A complete bibliography of publications in Transactions of the American Fisheries Society: 1980–1989. Report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, November 30, 2021. 182 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-t.html#transamfishsoc1980.

This report is updated frequently.

Beebe:2021:CBPcx

[1704] Nelson H. F. Beebe. A complete bibliography of publications in Transactions of the American Fisheries Society: 1990–1999. Report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, November 30, 2021. 213 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-t.html#transamfishsoc1990.

This report is updated frequently.

Beebe:2021:CBPcy

[1705] Nelson H. F. Beebe. A complete bibliography of publications in Transactions of the American Fisheries Society: 2000–2009. Report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, November 30, 2021. 288 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-t.html#transamfishsoc2000.

This report is updated frequently.

Beebe:2021:CBPcz

[1706] Nelson H. F. Beebe. A complete bibliography of publications in Transactions of the American Fisheries Society: 2010–2019. Report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, November 30, 2021. 301 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-t.html#transamfishsoc2010.

This report is updated frequently.

Beebe:2021:CBPda

[1707] Nelson H. F. Beebe. A complete bibliography of publications in Transactions of the American Fisheries Society: 2020–2029. Report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, November 30, 2021. 31 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-t.html#transamfishsoc2020.

This report is updated frequently.

Beebe:2021:CBPdb

[1708] Nelson H. F. Beebe. A complete bibliography of publications in Journal of Fish Diseases (1970–1979). Report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, December 4, 2021. 25 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-j.html#jfishdis1970. This

report is updated frequently.

Beebe:2021:CBPdc

[1709] Nelson H. F. Beebe. A complete bibliography of publications in Journal of Fish Diseases (1980–1989). Report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, December 4, 2021. 140 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-j.html#jfishdis1980. This

report is updated frequently.

Beebe:2021:CBPdd

[1710] Nelson H. F. Beebe. A complete bibliography of publications in Journal of Fish Diseases (1990–1999). Report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, December 4, 2021. 156 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-j.html#jfishdis1990. This

report is updated frequently.

Beebe:2021:CBPde

[1711] Nelson H. F. Beebe. A complete bibliography of publications in Journal of Fish Diseases (2000–2009). Report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, December 4, 2021. 210 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-j.html#jfishdis2000. This

report is updated frequently.

Beebe:2021:CBPdf

[1712] Nelson H. F. Beebe. A complete bibliography of publications in Journal of Fish Diseases (2010–2019). Report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, December 4, 2021. 357 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-j.html#jfishdis2010. This

report is updated frequently.

Beebe:2021:CBPdg

[1713] Nelson H. F. Beebe. A complete bibliography of publications in Journal of Fish Diseases (2020–2029). Report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, December 4, 2021. 104 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-j.html#jfishdis2020. This

report is updated frequently.

Beebe:2021:CBPdh

[1714] Nelson H. F. Beebe. A complete bibliography of publications in Frontiers in Marine Science (2010– 2019). Report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, December 22, 2021. 422 pp.

                                        URL http://www.math.utah.edu/
                                        pub/tex/bib/index-table-f.html#

frontmarsci2010. This report is updated frequently.

Beebe:2021:BPF

[1715] Nelson H. F. Beebe. A bibliography of publications in Frontiers in Marine Science (2020–2029). Report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, December 22, 2021. 611 pp.

                                        URL http://www.math.utah.edu/
                                        pub/tex/bib/index-table-f.html#

frontmarsci2020. This report is updated frequently.

Beebe:2022:BPJ

[1716] Nelson H. F. Beebe. A bibliography of publications in the Journal of FisheriesSciences.com. Report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, January 22, 2022. 131 pp.

                                        URL http://www.math.utah.edu/
                                        pub/tex/bib/index-table-j.html#
                                        jfisheriessciencescom. This report

is updated frequently.

Beebe:2022:BPAa

[1717] Nelson H. F. Beebe. A bibliography of publications in Aquaculture, Fish and Fisheries. Report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, February 21, 2022. 5 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-a.html#aquacfishfish. This

report is updated frequently.

Beebe:2022:CBPa

[1718] Nelson H. F. Beebe. A complete bibliography of publications in Marine Biology Research. Report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, March 3, 2022. 276 pp.

URL http://www.math.utah.edu/
pub/tex/bib/index-table-m.html#

marbiolres. This report is updated frequently.

Beebe:2022:CBPb

[1719] Nelson H. F. Beebe. A complete bibliography of publications in Reviews in Fisheries Science & Aquaculture. Report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, March 3, 2022. 124 pp.

URL http://www.math.utah.edu/
pub/tex/bib/index-table-r.html#

revfishsciaquac. This report is updated frequently.

Beebe:2022:CBPc

[1720] Nelson H. F. Beebe. A complete bibliography of publications in Fisheries. Report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, March 4, 2022. 490 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-f.html#fisheries. This re-

port is updated frequently.

Beebe:2022:CBPd

[1721] Nelson H. F. Beebe. A complete bibliography of publications in the Journal of Aquaculture Research & Development. Report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, March 8, 2022. 151 pp.

                                        URL http://www.math.utah.edu/
                                        pub/tex/bib/index-table-j.html#

jaquacultresdev. This report is updated frequently.

Beebe:2022:CBPe

[1722] Nelson H. F. Beebe. A complete bibliography of publications in the Fisheries and Aquaculture Journal. Report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, March 9, 2022. 75 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-f.html#fishaquacultj. This

report is updated frequently.

Beebe:2022:CBPf

[1723] Nelson H. F. Beebe. A bibliography of publications in Fishery Bulletin. Report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, March 18, 2022. 1272 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-f.html#fishbull. This report

is updated frequently.

Beebe:2022:CBPg

[1724] Nelson H. F. Beebe. A complete bibliography of publications in Computational Intelligence. Report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, April 1, 2022. 209 pp. URL http:

//www.math.utah.edu/pub/tex/bib/
index-table-c.html#computintell.

This report is updated frequently.

Beebe:2022:CBPh

[1725] Nelson H. F. Beebe. A complete bibliography of publications by, and about, Richard Steven Varga. Report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, April 20, 2022. 73

pp. URL http://www.math.utah.
edu/pub/bibnet/authors/v/index.
html#varga-richard-steven. This

report is updated frequently.

Beebe:2022:CBPi

[1726] Nelson H. F. Beebe. A complete bibliography of publications in the Information Security Technical Report. Report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, May 8, 2022. 122 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-i.html#infosectechrep. This

report is updated frequently.

Beebe:2022:CBPj

[1727] Nelson H. F. Beebe. A complete bibliography of publications in the Journal of Information Security and Applications. Report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, May 8, 2022. 237 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-j.html#jinfosecappl. This

report is updated frequently.

Beebe:2022:CBPk

[1728] Nelson H. F. Beebe. A complete bibliography of publications in Future Generation Computer Systems: 1984–2009. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, June 10, 2022. 304 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-f.html#futgencompsys2010.

Beebe:2022:CBPl

[1729] Nelson H. F. Beebe. A complete bibliography of publications in Future Generation Computer Systems: 2010–2019. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, June 10, 2022. 735 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-f.html#futgencompsys2010.

Beebe:2022:CBPm

[1730] Nelson H. F. Beebe. A complete bibliography of publications in Future Generation Computer Systems: 2020–2029. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, June 10, 2022. 382 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-f.html#futgencompsys2010.

Beebe:2022:BCG

[1731] Nelson H. F. Beebe. A bibliography of Computer Graphics Forum: 1982– 2009. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, June 10, 2022. 412 pp. URL http:

//www.math.utah.edu/pub/tex/bib/
index-table-c.html#cgf. This re-

port is updated frequently.

Beebe:2022:CBCa

[1732] Nelson H. F. Beebe. A complete bibliography of Computer Graphics Forum: 2010–2019. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, June 10, 2022. 493 pp.

URL http://www.math.utah.edu/
pub/tex/bib/index-table-c.html#

cgf. This report is updated frequently.

Beebe:2022:CBCb

[1733] Nelson H. F. Beebe. A complete bibliography of Computer Graphics Forum: 2020–2029. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, June 10, 2022. 112 pp.

URL http://www.math.utah.edu/
pub/tex/bib/index-table-c.html#

cgf. This report is updated frequently.

Beebe:2022:BPTa

[1734] Nelson H. F. Beebe. A bibliography of publications in Theory of Probability and its Applications: (1956–1999). Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, June 10, 2022. 524 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-t.html#theoryprobabappl.

Beebe:2022:BPTb

[1735] Nelson H. F. Beebe. A bibliography of publications in Theory of Probability and its Applications: (2000– 2009). Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, June 10, 2022. 107 pp.

URL http://www.math.utah.edu/

                                        pub/tex/bib/index-table-t.html#
                                        theoryprobabappl2000.

Beebe:2022:BPTc

[1736] Nelson H. F. Beebe. A bibliography of publications in Theory of Probability and its Applications: (2010– 2019). Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, June 10, 2022. 86 pp.

                                        URL http://www.math.utah.edu/
                                        pub/tex/bib/index-table-t.html#
                                        theoryprobabappl2010.

Beebe:2022:BPTd

[1737] Nelson H. F. Beebe. A bibliography of publications in Theory of Probability and its Applications: (2020– 2029). Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, June 10, 2022. 17 pp.

                                        URL http://www.math.utah.edu/
                                        pub/tex/bib/index-table-t.html#
                                        theoryprobabappl2020.

Beebe:2001:CBPp

[1738] Nelson H. F. Beebe. A complete bibliography of publications in Concurrency and Computation: Practice and Experience: 2000–2019. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, June 10, 2001. 762 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-i.html#ccpe. This report is up-

dated frequently.

Beebe:2001:CBPq

[1739] Nelson H. F. Beebe. A complete bibliography of publications in Concurrency and Computation: Practice and Experience: 2020–2029. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, June 10, 2001. 331 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-i.html#ccpe2020. This report

is updated frequently.

Beebe:2022:BPIa

[1740] Nelson H. F. Beebe. A bibliography of publications in IEEE Transactions on Parallel and Distributed Systems: 1990–2009. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, June 10, 2022. 387 pp.

URL http://www.math.utah.edu/
pub/tex/bib/index-table-i.html#
ieeetranspardistsys. This report is

updated frequently.

Beebe:2022:CBPn

[1741] Nelson H. F. Beebe. A complete bibliography of publications in IEEE Transactions on Parallel and Distributed Systems: 2010– 2019. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, June 10, 2022. 473 pp.

URL http://www.math.utah.edu/
pub/tex/bib/index-table-i.html#
ieeetranspardistsys2010. This re-

port is updated frequently.

Beebe:2022:CBPo

[1742] Nelson H. F. Beebe. A complete bibliography of publications in IEEE Transactions on Parallel and Distributed Systems: 2020– 2029. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, June 10, 2022. 110 pp.

                                        URL http://www.math.utah.edu/
                                        pub/tex/bib/index-table-i.html#
                                        ieeetranspardistsys2010. This re-

port is updated frequently.

Beebe:2022:BPIb

[1743] Nelson H. F. Beebe. A bibliography of publications in IEEE Software: 1984– 2009. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, June 10, 2022. 435 pp. URL http:

                                        //www.math.utah.edu/pub/tex/bib/
                                        index-table-i.html#ieeesoft. This

report is updated frequently.

Beebe:2022:CBPp

[1744] Nelson H. F. Beebe. A complete bibliography of publications in IEEE Software: 2010–2019. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, June 10, 2022. 221 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-i.html#ieeesoft2010. This

report is updated frequently.

Beebe:2022:CBPq

[1745] Nelson H. F. Beebe. A complete bibliography of publications in IEEE Software: 2020–2029. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, June 10, 2022. 66 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-i.html#ieeesoft2020. This

report is updated frequently.

Beebe:2022:BPAb

[1746] Nelson H. F. Beebe. A bibliography of publications by, and about, Per Brinch Hansen. Report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, July 21,

2022. 19 pp. URL http://www.math.
utah.edu/pub/bibnet/authors/h/.

Beebe:2022:CBPr

[1747] Nelson H. F. Beebe. A complete bibliography of publications in Distributed Computing. Report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, August 1, 2022. 148 pp.

URL http://www.math.utah.edu/
pub/tex/bib/index-table-d.html#
distribcomput.

Beebe:2022:BPAc

[1748] Nelson H. F. Beebe. A bibliography of publications by, and about, Sir Charles Antony Richard Hoare. Report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, July 30, 2022. 106

pp. URL http://www.math.utah.
edu/pub/bibnet/authors/h/.

Beebe:2022:CBPs

[1749] Nelson H. F. Beebe. A complete bibliography of publications in Deep Sea Research Part II: Topical Studies in Oceanography. Report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, August 11, 2022. 1382 pp. URL http:

//www.math.utah.edu/pub/tex/bib/
index-table-d.html#deep-sea-res-
part-2-top-stud-oceanogr.

Beebe:2022:CBPt

[1750] Nelson H. F. Beebe. A complete bibliography of publications in Limnology and Oceanography Letters. Report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, August 11, 2022. 51 pp.

URL http://www.math.utah.edu/
pub/tex/bib/index-table-l.html#
limnol-oceanogr-lett.

Beebe:2022:CBPu

[1751] Nelson H. F. Beebe. A complete bibliography of publications in Limnology and Oceanography: 1950–1959. Report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, August 11, 2022. 35 pp.

URL http://www.math.utah.edu/
pub/tex/bib/index-table-l.html#
limnol-oceanogr1950.

Beebe:2022:CBPv

[1752] Nelson H. F. Beebe. A complete bibliography of publications in Limnology and Oceanography: 1960–1969. Report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, August 11, 2022. 176 pp.

URL http://www.math.utah.edu/
pub/tex/bib/index-table-l.html#
limnol-oceanogr1960.

Beebe:2022:CBPw

[1753] Nelson H. F. Beebe. A complete bibliography of publications in Limnology and Oceanography: 1970–1979. Report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, August 11, 2022. 270 pp.

URL http://www.math.utah.edu/
pub/tex/bib/index-table-l.html#
limnol-oceanogr1970.

Beebe:2022:CBPx

[1754] Nelson H. F. Beebe. A complete bibliography of publications in Limnology and Oceanography: 1980–1989. Report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, August 11, 2022. 277 pp.

URL http://www.math.utah.edu/
pub/tex/bib/index-table-l.html#
limnol-oceanogr1980.

Beebe:2022:CBPy

[1755] Nelson H. F. Beebe. A complete bibliography of publications in Limnology and Oceanography: 1990–1999. Report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, August 11, 2022. 358 pp.

URL http://www.math.utah.edu/
pub/tex/bib/index-table-l.html#
limnol-oceanogr1990.

Beebe:2022:CBPz

[1756] Nelson H. F. Beebe. A complete bibliography of publications in Limnology and Oceanography: 2000–2009. Report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, August 11, 2022. 455 pp.

URL http://www.math.utah.edu/
pub/tex/bib/index-table-l.html#
limnol-oceanogr2000.

Beebe:2022:CBPba

[1757] Nelson H. F. Beebe. A complete bibliography of publications in Limnology and Oceanography: 2010–2019. Report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, August 11, 2022. 383 pp.

                                        URL http://www.math.utah.edu/
                                        pub/tex/bib/index-table-l.html#
                                        limnol-oceanogr2010.

Beebe:2022:CBPbb

[1758] Nelson H. F. Beebe. A complete bibliography of publications in Limnology and Oceanography: 2020–2029. Report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, August 11, 2022. 186 pp.

                                        URL http://www.math.utah.edu/
                                        pub/tex/bib/index-table-l.html#
                                        limnol-oceanogr2020.

Beebe:2022:CBPbc

[1759] Nelson H. F. Beebe. A complete bibliography of publications in Progress in Oceanography. Report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, August 11, 2022. 846 pp.

                                        URL http://www.math.utah.edu/
                                        pub/tex/bib/index-table-p.html#
                                        progr-oceanogr.

Beebe:2022:CBPbd

[1760] Nelson H. F. Beebe. A complete bibliography of publications in Marine Ecology Progress Series: 1970–1979. Report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, August 14, 2022. 12 pp.

URL http://www.math.utah.edu/
pub/tex/bib/index-table-m.html#
mar-ecol-prog-ser1970.

Beebe:2022:CBPbe

[1761] Nelson H. F. Beebe. A complete bibliography of publications in Marine Ecology Progress Series: 1980–1989. Report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, August 14, 2022. 395 pp.

URL http://www.math.utah.edu/
pub/tex/bib/index-table-m.html#
mar-ecol-prog-ser1980.

Beebe:2022:CBPbf

[1762] Nelson H. F. Beebe. A complete bibliography of publications in Marine Ecology Progress Series: 1990–1999. Report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, August 14, 2022. 979 pp.

URL http://www.math.utah.edu/
pub/tex/bib/index-table-m.html#
mar-ecol-prog-ser1990.

Beebe:2022:CBPbg

[1763] Nelson H. F. Beebe. A complete bibliography of publications in Marine Ecology Progress Series: 2000–2009. Report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, August 14, 2022. 1273 pp.

URL http://www.math.utah.edu/
pub/tex/bib/index-table-m.html#
mar-ecol-prog-ser2000.

Beebe:2022:CBPbh

[1764] Nelson H. F. Beebe. A complete bibliography of publications in Marine Ecology Progress Series: 2010–2019. Report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, August 14, 2022. 1149 pp.

                                        URL http://www.math.utah.edu/
                                        pub/tex/bib/index-table-m.html#
                                        mar-ecol-prog-ser2010.

Beebe:2022:CBPbi

[1765] Nelson H. F. Beebe. A complete bibliography of publications in Marine Ecology Progress Series: 2020–2029. Report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, August 14, 2022. 248 pp.

                                        URL http://www.math.utah.edu/
                                        pub/tex/bib/index-table-m.html#
                                        mar-ecol-prog-ser2020.

Beebe:2022:CBPbj

[1766] Nelson H. F. Beebe. A complete bibliography of publications in Marine Resource Economics: 2020–2029. Report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, September 29, 2022. 220 pp. URL http://www.

                                        math.utah.edu/pub/tex/bib/index-
                                        table-m.html#marinresourcecon.

Beebe:2022:CBPbk

[1767] Nelson H. F. Beebe. A complete bibliography of publications in Conservation Biology: 1980–1989. Report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, September 29, 2022. 36 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-c.html#conservbiol1980.

Beebe:2022:CBPbl

[1768] Nelson H. F. Beebe. A complete bibliography of publications in Conservation Biology: 1990–1999. Report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, September 29, 2022. 351 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-c.html#conservbiol1990.

Beebe:2022:CBPbm

[1769] Nelson H. F. Beebe. A complete bibliography of publications in Conservation Biology: 2000–2009. Report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, September 29, 2022. 593 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-c.html#conservbiol2000.

Beebe:2022:CBPbn

[1770] Nelson H. F. Beebe. A complete bibliography of publications in Conservation Biology: 2010–2019. Report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, September 29, 2022. 424 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-c.html#conservbiol2010.

Beebe:2022:CBPbo

[1771] Nelson H. F. Beebe. A complete bibliography of publications in Conservation Biology: 2020–2029. Report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, September 29, 2022. 121 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-c.html#conservbiol2020.

Beebe:2022:CBPbp

[1772] Nelson H. F. Beebe. A complete bibliography of publications in the Bulletin of Mathematical Biology: 2020– 2029. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, October 17, 2022. 19 pp.

URL http://www.math.utah.edu/
pub/tex/bib/index-table-b.html#
bullmathbiol2020. This report is up-

dated frequently.

Beebe:2022:CBPbq

[1773] Nelson H. F. Beebe. A complete bibliography of publications in Fish Physiology and Biochemistry: 1980– 1989. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, October 18, 2022. 44 pp.

URL http://www.math.utah.edu/
pub/tex/bib/index-table-f.html#
fishphysiolbiochem1980. This re-

port is updated frequently.

Beebe:2022:CBPbr

[1774] Nelson H. F. Beebe. A complete bibliography of publications in Fish Physiology and Biochemistry: 1990– 1999. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, October 18, 2022. 164 pp.

URL http://www.math.utah.edu/
pub/tex/bib/index-table-f.html#
fishphysiolbiochem1990. This re-

port is updated frequently.

Beebe:2022:CBPbs

[1775] Nelson H. F. Beebe. A complete bibliography of publications in Fish Physiology and Biochemistry: 2000– 2009. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, October 18, 2022. 126 pp.

URL http://www.math.utah.edu/
pub/tex/bib/index-table-f.html#
fishphysiolbiochem2000. This re-

port is updated frequently.

Beebe:2022:CBPbt

[1776] Nelson H. F. Beebe. A complete bibliography of publications in Fish Physiology and Biochemistry: 2010– 2019. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, October 18, 2022. 312 pp.

URL http://www.math.utah.edu/
pub/tex/bib/index-table-f.html#
fishphysiolbiochem2010. This re-

port is updated frequently.

Beebe:2022:CBPbu

[1777] Nelson H. F. Beebe. A complete bibliography of publications in Fish Physiology and Biochemistry: 2020– 2029. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, October 18, 2022. 97 pp.

URL http://www.math.utah.edu/
pub/tex/bib/index-table-f.html#
fishphysiolbiochem2020. This re-

port is updated frequently.

Beebe:2022:CBPbv

[1778] Nelson H. F. Beebe. A complete bibliography of publications in Marine Biotechnology. Technical report, University of Utah, Department of Mathematics, Salt Lake City, UT 84112-0090, USA, October 21, 2022. 412 pp. URL http://www.

math.utah.edu/pub/tex/bib/index-
table-m.html#marbiotechnol. This

report is updated frequently.

Othmer:1997:CSM

[1779] H. G. Othmer, F. R. Adler, M. A. Lewis, and J. C. Dallon, editors. Case Studies in Mathematical Modeling— Ecology, Physiology, and Cell Biology. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1997. ISBN 0-13-574039-8. viii + 411 pp. LCCN QH541.15.M3C37 1997. URL

http://www.math.utah.edu/books/
csmmepcb/.

Cherkaev:2000:VMS

[1780] Andrej Cherkaev. Variational Methods for Structural Optimization, volume 140 of Applied mathematical sciences. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2000. ISBN 0-387-98462-3 (hardcover). xxvi + 545 pp. LCCN QA1 .A647 v.140; TA658.8

.C53 2000. URL http://www.math.
utah.edu/book/vmso. NHFB was the

chief technical editor for this book, and wrote the author/editor indexing software, the TEX DVI-to-PostScript driver and the initial Web site for the

book, at http://www.math.utah.edu/
books/vmso/index.html.

Milton:2002:TC

[1781] Graeme W. Milton. The Theory of Composites, volume 6 of Cambridge Monographs on Applied and Computational Mathematics. Cambridge University Press, Cambridge, UK, 2002. ISBN 0-521-78125-6. xxviii + 719 pp. LCCN TA418.9.C6 M58 2001.

US$80. URL http://www.math.utah.

edu/books/tcbook. Series editors: P. G. Ciarlet, A. Iserles, Robert V. Kohn, and M. H. Wright.

Ethier:2007:OPM

[1782] Stewart N. Ethier and William R. Eadington, editors. Optimal Play: Mathematical Studies of Games and Gambling. Institute for the Study of Gambling and Commercial Gaming, Reno, NV, USA, 2007. ISBN 0-9796873-0-6. xxvi + 550 pp. LCCN QA273 .O685 2007.

Ethier:2010:DCP

[1783] Stewart N. Ethier. The Doctrine of Chances: Probabilistic Aspects of Gambling. Probability and its applications. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2010. ISBN 3-540-78782-8 (hardcover), 3-540-78783-6. xiv + 816 pp. LCCN QA271 .E84 2010.

Milton:2016:ETC

[1784] Graeme W. Milton, editor. Extending the Theory of Composites to Other Areas of Science. Milton–Patton Publishers, P.O. Box 581077, Salt Lake City, UT 85148, USA, 2016. ISBN 1-4835-6919-5 (print), 1-4835-6920-9 (e-book). xx + 422 pp. LCCN ???? URL http:/

/www.math.utah.edu/books/milton.

Harris:1979:PCS

[1785] Frank E. Harris and Nelson H. F.

Beebe, editors. Proceedings of the conference Software Standards in Chemistry, Held at the University of Utah, July 25–27, 1979, volume 7. Lawrence Berkeley Laboratory, University of California, Berkeley, CA, USA, July 1979.

URL http://escholarship.org/uc/
item/19b1d4sw.

Moler:1978:NAC

[1786] Cleve B. Moler and Isaiah Shavitt, editors. Numerical algorithms in chemistry: algebraic methods: report on the workshop, August 9– 11, 1978, LBL-8158. Lawrence Berkeley Laboratory, University of California, Berkeley, CA, USA, 1978. LCCN QD39.3.M3 W61 1978. URL

http://escholarship.org/uc/item/

3xb320bq. UC-32, CONF-780878.

Harris:1980:LNU

[1787] F. E. Harris, J. Michl, J. P. Simons, and R. P. Steiner, editors. Lecture Notes: Utah Chemistry Workshop, September 22–27, 1980, Modern Quantum Chemistry. Department of Chemistry, University of Utah, Salt Lake City, UT 84112, USA, 1980. Two volumes.

Simos:2007:CMS

[1788] Theodore E. Simos and George Maroulis, editors. Computation in Modern Science and Engineering: Proceedings of the [Fifth] International Conference on Computational Methods in Science and Engineering 2007 (ICCMSE 2007), Corfu, Greece, 25– 30 September 2007, volume 2A, 2B of AIP Conference Proceedings (#963). American Institute of Physics, Woodbury, NY, USA, 2007. ISBN 0-7354-0476-3 (set), 0-7354-0477-1 (vol. 1), 0-7354-0478-X (vol. 2). ISSN 0094-243X (print), 1551-7616 (electronic), 1935-0465. LCCN Q183.9 .I524 2007. URL

http://www.springer.com/physics/
atoms/book/978-0-7354-0478-6.

Anonymous:2017:RBN

[1789] Anonymous. Reviewed: Beebe, Nelson H. F. (1-UT), The mathematicalfunction computation handbook. Programming using the MathCW portable software library. Springer, Cham, 2017. xxxvi + 1115 pp. ISBN: 978-3-319-64109-6; 978-3-319-64110-2. Mathematical Reviews, December 2017. CO-DEN MAREAR. ISSN 1069-8221 (print), 2167-5163 (electronic). URL

http://ams.rice.edu/mathscinet/.

Baumann:2021:NSM

[1790] Gerd Baumann, editor. New Sinc methods of numerical analysis: Festschrift in honor of Frank Stenger’s 80th birthday. Trends in mathematics. Birkhäuser, Cham, Switzerland, 2021. ISBN 3-030-49715-1 (hardcover), 3-030-49716-X (e-book). ISSN 2297-0215. xvi + 404 pp. LCCN QA372.